A Complete Bibliography of The Journal of the American Statistical Association: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.11

Title word cross-reference

2 [Rue04]. 2^{n} [MOJ05]. $79.97 [McC02a]. _{2.5} [KSZL02]. C_{p} [KE02]. δ [FS06b]. ℓ_{1} [YZ09]. k [ADW07, EM05, FS06b]. L [PC00]. L^{2} [CS04]. L_{1} [WS07]. L_{2} [BY03]. L_{p} [LL05b]. M [KL06b, LP06, VH02]. P [BB00a, BB00b, Boo00, Eva00a, Mar00a, Pic00, RvdVV00a, RvdVV00b, Ste00c, HDW07, HDS06, Mar00b, MMB09, SSY07]. ψ [LS06, STZW03]. t [FHY07, MH04, McN06, UCTW07].

-Equivalence [FS06b]. -Estimation [LP06]. -Learning [LS06, STZW03]. -Norm [WS07]. -Regularized [YZ09]. -Sample [ADW07, EM05, FS06b]. -Statistics [PC00]. -Values [MMB09]. -Year [UCTW07].

0-387-95234-9 [McC02a].

1 [HR00]. 1012 [NVP06].

2000 [Han03]. 2003 [HI06]. 21st [Nol03a, SRRW00, Nol03b, Sco02]. 2A [Ann06d]. 2B [Ank05]. 2nd [Cha03b].

3rd [Lun03].

71 [BH08].

829 [BH08].

95 [Iha06]. 98 [NVP06].

Aalen [WS02]. Abecedarian [And08]. Aberrant [RS08]. Aberrations [New02]. Above [ZRM^{+}08]. Above-Ground [ZRM^{+}08]. Absence [Ros01]. Absent [IPC08]. Absolute [KW01c, KCO08]. Absorbers [LSQ03]. Abundance [Few04, Sch07e]. Abuse [LZHR01]. Accelerate [Liu03, LP05]. Accelerated [CW00, Hua00, KL08, PF06, SNQ09, ZL07a]. Accelerator [LSZH08]. Accelerometer [MAC^{+}06b]. Accounting [BBCL02, DT00, SCG^{+}07]. Accuracy [AD08, CCL09]. Accurate [Ano03g, CH06b, Hil08, HW00, LLL08a, SCS08a, YK04, ZS01]. ACE [EL05, GSCR04]. ACE-Inhibitor [GSCR04]. Achieve [Len02f]. Acid [MPE08]. Acquired [WT01]. Across [MMZ02]. Activation [CBCS07]. Active [XMN06, DJSW09]. Activity [CCDP09]. Actuarial [Bei04]. Acute [MRID01]. Adam [Bin06b, Pre03]. Adaptation [SSV05]. Adaptations [RR07]. Adapted [DSV04]. Adaptive [Ass07, Číź08, CRY09, DFC09, HS02b, HR03, JS05, KSG05, LL06a, LL06b, Len06, LJ06, Lin03b, LPP02, LA08, SY02, SC07, Yan01, ZYZS03, ZS01, Zou06, Pal08]. Addition [WBS07b]. Additional [FRS06, RT00]. Additive [AC07c, Che07a, Chr07, DZ03, DM00, FJ05, HSS01, HL05, KL00, Lit07, MY08b, OBMK07a, OBMK07b, Rup07, Woo04a, YPXH06, YLZ08]. Additively [Pui03]. Address [Lac09]. Adduct [MWL^{+}01]. Adhesion [HZZ^{+}08]. Adjudication [CK04]. Adjusted [BYE^{+}05, BY05, Edw05, GS07, Hol05, Kul06, Sha05, Tam05, Wes05, WS02]. Adjusting [Els02c, KM02, Rab02a, Rab02b, Sch02a]. Adjustment [Nem04, WIG^{+}04]. Adjustments [RT00]. Administrative [YZ05]. Adonis [Tro06]. Adrian [Ann05b, Nol03a]. Advance [HY07]. Advanced [Ank05, Ann06d, Hal03a, HM06, Ker06, NHS00, Ron05, Ber04, Hal03b, Sch07e, Wat07]. Advantage [GH08, Gel08, Hod08, Kat08]. Adversarial [IRB09]. Affairs [CH02, HC02, LNR03]. Affect [SRA08]. Affine [CE09, Ran00]. Affine-Invariant [Ran00]. Africa [CDR08a, CDR08b, Dun08, Pac08, Tho08]. After [GH08, Gel08, GSCR04, Hod08, KL06a, Kat08, SHY04b, Ano06l, Dra06, SW05b]. Afterword [Cas00a]. Against [LZHR01]. Age [IC05, LHN08, NLLH06, PCK06]. Aggregated [NN09]. Aggressive [WBBR05]. Aging [ZWT^{+}05]. Agreement [Cho06, LHSY02, Sch04b]. Agricultural [BHO07]. Agriculture [FKS02, LFKS03]. Ahrens [Cav06b]. Aid [PWM01]. AIDS [Wu02]. Air [DDZS02, DDH04, ZYDS08]. Aitchison [Ros06]. Aitkin [Hal06]. Akaike [NST07]. Ákos [All06]. al [BL07]. Alain [Teu03]. Alan [Ann06d]. Alastair [Aqu06]. Alex [Dur06]. Alexander [AC06b, Cho06, Kra04, Wil03a]. Alfred [Kie05]. Algebra [Imh03a, Imh03b, Jam02g, Mor05]. Algorithm [DTT08, HO02a, HO02b, HRV02, IZT08, Lia05, MH04, MY02, Por02, Rup02b, SL06, Szk03, VS08, XLB04]. Algorithms [Buo06, FG00, HO02a, HO02b, HRV02, IT02, Las07b, LBS06, MY02, Por02, RJ08, Ros03, Rue04, Rup02b, Teu08, AC05d, Mil06, Yan05b, Lüt03, Mün08]. All-or-None [Bak00]. Allan [Fil06]. Allelic [HZW07]. Allen [Cle03a]. Allenby [Ann06c]. Allison [Neu06]. Allocation [TRH07]. Allocative [KT05]. Allowing [BF02]. Along [SWDS03]. Alternating [GL06, Liu03]. Alternative [HHCK09, Kar07b, TIY03]. Alternatives [AKM01, BG01, LL04, SHH00, WQ07, WL00]. Altman [Van05]. Aman [Ron05]. Ambient [CHB^{+}08]. Ambroise [Gol05b]. American [BH08]. Amino [MPE08]. Amit [Rue04]. Among [PT09, SWS^{+}03, Fol00, RAHT06]. Amplification [RS09a]. Analog [BP01]. Analyses [Bow07, Cau01, Dmi05, Gro01, Hor01a, Hor01b, JFY07, Mag01, RHC07, RK01, SGL07, AC08d]. Analysing [Con08]. Analysis [AA00, AAK06, ADW07, AC01a, AC03c, AC04b, AC05c, AC05b, And05, Ano01j, Ano03i, AC03f, BP01, Ban05, BB01, BL07, Ber00b, BDS01, BBN^{+}04, BD03, Bin06b, BL06, BM01, Boo06, Boo07, BC03b, Bor03, Bra06a, BF01b, BGM^{+}05, Bro07, BD05, CP02, Car03a, Car03b, Car05a, CN02, CW00, CDHL05, CB09, CFK07, CE09, CM01, CM09b, CA02, CCL09, Cle03a, Cle03b, CK04, Cra07, CCDP09, DMJ^{+}08, Dav03, Dav06c, De 06, Dix06, DM06, DT00, ETST01, Elb08, EL00, ERS02, ES02, EL05, Els02a, FL04, FPH^{+}05, FHL07, FW07, Far01, FV09, FWP05, Fol00, FJS08, FR02, GFK07, Gen05a, GBT09, GRF03, Gre00, Gre07b, GL06, Gri02b, GWB08, GDOvS03, GDH09, Hav07, HI06, HRH02, HT04, Hoo07, HW01a, HM00a, HM00b, HM00c, Hu04, HZAG01, HWZ05]. Analysis [HZW07, HSB09, HR00, HZ08, IYSK08, IRB09, Jam02f, Jen02b, Jer06, JDvS02, JL09a, JHCN06, JKN08, Ker06, Kie08c, KU01a, KU01b, Koe02a, KG01a, Kra04, LL05a, Lar06, LL03a, Laz07, LB01, Len04a, Len05, LR04, Li08b, LY01a, LC01a, LY01b, LTMS02, LO06, LSGS04, Loc04, LC09, Lun07a, Lun07b, LCLC08, LM00, Mar01, MB02, Moe03a, Moe03b, Mor04, NSP00, OZ08, OMS^{+}09, Oak00, ORvSM01, OvSG05, Par07, PAS^{+}03, Par06, Par02b, PDG02a, PDG02b, PH08, Per06, PGLW00, PC00, Rag00a, Rag00b, Ram01a, Rao03b, Ree08, Rhe07, Rob03a, Rob03b, RSW09, RS09a, RS09b, RL02, Rub08, SBVW06, SSVV09, ST04, Sch03d, Sch07d, Sea02, SWP00, Sey05, SCS08b, SRA08, SSY07, Sin04, SMG^{+}05, Sma07, SCG^{+}07, Ste06, SPSZ05, SSL07, Tim03, TW07b, VPW08, Ver02, WC01, WW01]. Analysis [WLH04, WA06, WLH08, Woo04b, Xie03a, Xie03b, XSC08, Yan04b, YF08, YMW05, ZFZ03, ZYDS08, ZRM09, ZZIP07a, ZZIP07b, dGKS01, Ano04e, Buo06, Cav04, Dav06a, Gem06, Las06, Lun06, Mol05, Poo06, Rao03a, Rus05, Wan06, Yan04a, Zhu06, Bet03, BT08, Boo06, Car07, Dur06, Fie06, Gho07, Gre07a, Har07, Las07a, Sai06, Yan05a]. Analyst [Las08]. Analytical [PS06c, Tre06]. Analytics [Nta07]. Analyzing [Bak00, KMH05, MTR^{+}01, SNQ09, WQC01, Cho06, Daw03, Gol05b]. Anand [AC04d]. Anatoly [Kar06a]. Anchoring [vSH08]. Ancient [PM09]. Ancillary [RL01]. ANCOVA [AV01]. Anders [Cor05]. Andrè [Ber04]. Andrea [Ban06, Per06]. Andreas [Eas03]. Andrew [Bog04]. Angel [Jer06, Wik04]. Angle [KVZ07, SYCB08]. Angles [LDVT09, RFP09]. Anh [Gol05b]. Animal [Few04, Gia00, WS01a]. Ankle [RBP08]. Ann [Ram05]. Annette [Lan03]. Annotated [Str03b]. Anomaly [Hig09, LS09a, LS09b, Men09, WM09b]. Anopheline [HHS06]. ANOVA [AM01, AP04, Ano01p, DMRM04, GWKK01a, GWKK01b, GDOvS03, SME03, Whi01, Woo01, YW01]. Answers [Hil08, LLL08a, SCS08a]. Anthony [Ank05, Ban06, Dav03]. Anthracycline [GSCR04]. Anthrax [BB03]. Antigen [LTMS02, MPE08]. Antonio [Pic05]. Aparna [Kle06]. Applebaum [Nov05]. Application [ACK01, AM01, Ano01p, Ano02n, BHO07, BB03, BFV01, BRS^{+}00, CN02, CLLCL09, ES02, FKS02, FB00, FD08, FBV^{+}04, GM09, GWKK01a, GWKK01b, GLPS03, GMK^{+}04, GL08, HL00, HHS06, HT04, HZZ^{+}08, IPBS00, JFY07, JDvS02, KSZL02, KW01c, KT05, LSC04, LLW04, Li07, LTMS02, LFY^{+}08, LM09b, LDM09, LFKS03, NN09, OMS^{+}09, PDG02a, PDG02b, PPR09, Pui03, RGK01, RFP09, SWSM09, Sch03a, Sea02, SF07, SJ06, TW07a, Ver02, WT01, WSLC06, Wei08, WL00, Whi01, Woo01, Wu02, YF08, YL01, YW01, ZSB01, ZWT^{+}05, ZSCW06]. Applications [Ano02o, BLW01, Ber04, Boo06, BIK00, CCL^{+}08, CCPY09, CH06a, CW08, Che08, CLR06, Dav03, DJSW09, Dui16, FZ03, FSL^{+}06, Fra08, FKW01, GS01, Gol05a, HM03a, IYSK08, Imh03a, Imh03b, Kar07b, KW01a, KW01b, Kie07b, LW01a, Lar06, LW01b, LZ01, Lin04a, Lun07b, Lüt03, McN02b, MPK09, MICR06, PW05, Poo06, PV06, Ram01b, Rao03a, Rao03b, Ric04, Rob03b, Roo02b, Ros06, SBJRH05, Sch07c, SCC^{+}06, TW07b, WBS07a, YZ09, Zha05b, ZCI^{+}09, AC06c, Kra04, Len05, McN06, Ozt05, Rob03a, Sho05, Ste06, Sur04, Teu08, Von06, Wik06, Wil06b, Yan06, Yin04, Yor06, deL05, Bra08, Gen07b, Str07]. Applied [AHA01, CTG^{+}06, FHY07, Gil08b, Jam02g, LC09, Lun03, Sen08b, Tro06, Dob04, And05, Bro07, Cle06, Dig05, Eas03, Gho07, Hu04, Len06, Mol05, Mül05, Sch04a, Sun08, Ter04, Wes07]. Applying [Rog02a, Las08]. Approach [AC07b, Ano02j, Ano03g, AMW02, Bar08, BFHR03a, BFHR03b, BBCL02, BX03, BL06, BP08, CTG^{+}06, CM06b, CN08, CN05, EL00, EL05, FKC09, GA05, Gua06, HW00, Jam02e, Jam02f, Jam03b, JB01, Jes03b, JL06, KRS^{+}06, KW02, Kon06, Koo02, KE02, KZ03, Las08, Len02f, LL03b, LZ08, LC09, MB02, Mee00, MJB03, PR00, RGA01, SFPM04, SJ03, Tan04, Tan06a, TBJ08, WL05, WM09a, Wei08, WS01b, AC04b, Bai04, But06, Gol05a, Jam03a, Jes03a, Joe03, Sha07, Woo07, CH08, McN07, Yan08a]. Approaches [HRH02, MCG00, Ric04, Soo00, SjKH^{+}01, Kar07b, Cra07]. Approximate [Dui16, Fue07, RW05, Wu04, Zha05b]. Approximately [FW00b, SJ06].

Approximation [BP02, LLC07, Str00g, WL07a, Yin04, YZ09]. Approximations [Ano05j, AF01b, AF01a, Cai01, DKS06, Miy04, Mou01, Sol01, SJ06, Vid01]. Archeological [AHA01]. Arctic [SYCB08]. Area [Dat04, DP03, DP06, HT04, LH07, PT06, PS07, PS06c, RFS06, WF03]. Area-Level [HT04]. Areas [NC02, PS07]. ARFIMA [KOC07, Lin03b]. ARFIMA-GARCH [Lin03b]. Arising [GL06, Gus09, YC02]. ARMA [FRZ05]. ARMAX [PS05]. Armitage [Cav06a]. Arnaud [Dia03a]. Arnold [Lei05, Ann06d]. Array [GLN08]. Arrays [HWZ06, SL08, WIG^{+}04]. Arsenic [LSGS01, MICR06]. Artery [SSVV09]. Arthur [Cle03a, HK02]. Article [Ald08, DL08, Edw08, Fri02b, Sen08a, Sti08, Zab08b, Zab08a]. Artificial [Ram05]. Artificially [Yat05]. ASA [Lac09]. Asha [Wil06b]. Aspects [DM04]. Aspirin [BBN^{+}04]. Assessing [CM09a, GT09, HRC04, LHSY02, LA06, MPE08, Par06, Per06, SGH07, Sch03a, SS08, WD00]. Assessment [BG06b, FIC01, FB00, Kad07, LGS07, Mil07, RLL07a, RLL07b, ST02, SI02, SH06, Tre06, Wol07, YL01]. Asset [GM09]. Assignment [Bha09]. Assignments [Hil08, LLL08a, SCS08a]. Assisted [Chr07, JL06, LL07, Lit07, OBMK07a, OBMK07b, Rup07]. Associated [IC05, KMR09, KPD02, New02]. Association [BF02, BH08, CSCC06, CFK07, HJJ09, IYSK08, Li06, LZS^{+}06, LZ06, LDM09, NLLH06, Sab06, SAE06, TR06, WM09a, YF05, Mor05]. Assuming [FTW01]. Assumptions [KL08, PS06b]. Asthma [KRS^{+}06]. Astola [Gen05b]. Asymptotic [Ano05j, BB00a, Boo00, Cor02b, Dui16, Efr01, Eva00a, KL06b, LL05b, Mar00a, Miy04, Pic00, RvdVV00a, RvdVV00b, Ros03, She02a, Ste00c, Str00g, Vog02d, WL00, Zha05b]. Asymptotically [Zha04b]. Asymptotics [AA00, Rou06]. Athletics [EM08, Hil09]. Atkinson [Ban06]. Atlantic [Hig09, LS09a, LS09b, Men09, WM09b]. Atmosphere [Nyc00]. Atmospheric [Bri07]. Attack [HPC03]. Attacks [DM00]. Attention [BIK00]. Attitude [JR07]. Attitudes [GT09]. Attorneys [Las08]. Attributing [HB09, Ros02]. Attrition [ACK01]. Auction [LM09b]. Auctions [WJSS08]. Auditing [Mee03]. Augmented [Yat05]. Auto [GRF03]. Autocorrelated [XLCM03]. Automated [LR01]. Automatic [ORvSM01, Zha03, Cow07]. Autoregression [FF06, FRT09, HL06a, HW06, Kni06, KX04, KX06a, KXF^{+}06, KX06b, Rob06]. Autoregressive [BRS03, CR00b, KST01, NN09, PP05, RHC07, RFF04, SB05, WL01, XK09]. Auxiliary [Bak00, Efr05a, PW05, SW02c, WS01b]. Availability [BP08, CN08]. Aven [Ban05]. Average [BRS03, DHS08, GS07, HC03a, YL01]. Averaged [BHO07]. Averaging [CL03, CKN08, HC03b, HC06, Hof07, IR03b, Joh03, LM04, MICR06, RZ03, SD03, YY09]. Axel [Spe06a]. Axes [RBP08]. Ayyub [Loc04].

B [AC04a, And05, Ann05b, Bog04, Gen06, Kar03a, Len05, Ram05, Roh05, LZH09]. B-Splines [LZH09]. Back [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05b, Ano05c, Ano05d, Ano05e, Ano06b, Ano06c, Ano06d, Ano06e]. Background [WIG^{+}04]. Baddeley [Ann05b]. Bagging [IK08]. Bai [Ozt05]. Bailey [Mor05]. Balakrishnan [AC04a]. Balance [RRS07]. Balanced [Boo06, LPR01, PKC01, SO04]. Baldi [Emb03]. Ball [Sch02d]. Ballot [HI06]. Bands [LH07]. Bandwidth [GG08, GG07]. Banerjee [BH08]. Bani [Yan04c]. Bankruptcy [FS04b]. Barbara [Lin03a]. Bark [ZRM^{+}08]. Barker [Len04a]. Barnett [Wik06]. Barreto [Wie06]. Baseball [Koo02, Sch02d]. Based [ASFP09, AC04b, Ano01o, BF09, BT04, BZ00, CJS02, CCPY09, CP06, Cha03b, CSCC06, CV07, CCC09, CL08a, CG03b, CL08b, CF09, DHS08, DV03, EOU02, EL04, FR02, GGR07, GS06, HTX09, HH07, HI01, HHCK09, HQF08, IZ07, JM06, KJ07, KSN08, KS07a, KS07b, KVZ07, Kje00, KMR09, Kol03b, KF07, Lee00, Li06, LL07, LZS^{+}06, LZ06, MVY09, Mar07, MJvAC02, MVBC03b, MVBC03c, MAC^{+}06b, MCG00, OOK03, OS00, RD06, Sab06, SBVW06, SAE06, SSL03, SW02b, SYCB08, SPJ01, Swy02b, TSE02, TLW07b, TLW07a, TR06, VH02, WM09a, WH03b, WIG^{+}04, XWF05, YJ04, Zha04b, ZWS06, ZRM09, ZW08, Li08a, Nes07]. Basel [Hey08]. Baseline [FG04]. Basic [Fie00]. Basis [Ber00a, JM04, ZWL^{+}04]. Basketball [Har03]. Bausell [Len04a]. Bayes [BC03a, CL00, CD09, DX09, ETST01, Efr09, GR06a, HC01, JB01, KST01, KTZ00, Mar00b, SI02, YL05, Ann08]. Bayesian [Ank05, Bar08, Gen06, Mül05, Ott06, Sho05, ABPR06, Ano01j, Ano03i, AHA01, BG06a, BG06b, Ber00b, BG01, BDS01, BCR02, BBN^{+}04, BD09, BL06, BM01, BF01a, BFV01, BD05, CD07, CM06a, CTG^{+}06, CGR04, CNN07, CBL^{+}08, CS03b, Cow03a, CGRS01, CMRT09, De 06, DM06, DZ00, DC03, Dun05, DHSR08, DHE08, EL00, ERS02, EL05, EW08, FL09, FV09, GKM05, Gel04, Gen07a, Gen00a, Gen00b, GCK00, GMK^{+}04, GT07, GBT09, GC09a, Gil08b, GC09b, GG01, GL08, Gue08, GLN08, GDH09, Gus00, Hay00, HHS06, HBK^{+}09, Hig00, HC03b, HT04, HZW07, ICG02, IJS01, IR03a, JL08, JDvS02, KRS^{+}06, Kar07b, KSZL02, KL08, KG01b, KT05, LL03a, LB01, LDVT09, LW01b, LPM^{+}08, LPW05, LC09, MJW06, MG00, MD01, Mee03, MPE08, MW09, MICR06, MRID01, MSD04, NSP00, NC02]. Bayesian [NK00, OR01, PM09, PS06a, PC08, PAS^{+}03, PR00, QA09, RZ03, Raz00, Ree08, RSW09, RGA01, RMS09, SL08, STPN09a, STPN09b, Sco02, SR01, Sha07, SMG^{+}05, Sin07a, SPJ01, SF07, STNM09, TSV05, TI08, WM01, WRM07, WBS07a, WMNB01, Woo07, Wor00, XLB04, YL09, YY09, ZSCW06, Ann06b, Ann06c, Cow03b, Epi04, Gho06, Ram05, Ste04, Woo04b, Yan04a, Bar08]. Bayesian-Optimal [ABPR06]. Bayesians [Efr05b]. Bayo [Ste06]. Be [Cau01, FHY07, Fol00, Gro01, Hor01a, Hor01b, Mag01, RK01]. Beaver [DK01]. Beetle [ZRM^{+}08]. Before [Cle02c, GH08, Gel08, Hod08, Kat08]. Behavior [BCM02, Gol02c, WBBR05, WL00]. Behavioral [AC03a, AC03d, Sta08]. Behaviors [GT09]. Behind [GR06b]. Below [ZRM^{+}08]. Below- [ZRM^{+}08]. Benchmark [MICR06, PT06].

Benchmarking [RM09]. Bening [Lin04a]. Benjamin [Ano04e, Cav04]. Bennett [Cle03a]. Bent [CLR06]. Bent-Cable [CLR06]. Berc [Lüt03]. Bernard [Gem06]. Bernhard [Wil03a]. Bernoulli [AM01, Ano01p, CH06a, GWKK01a, GWKK01b, Whi01, Woo01, YW01]. Berry [Jam03a]. Best [FS06b, JL06, LPL06]. Between [CS04, FS04a, HS07a, LPL06, MWL^{+}01, Ric04, Ric08, Ros07, SWSM09, Ste01]. Beyond [Rhe03b, Wil03c, Rhe03a, Wil03a]. Bhaskara [Imh03a]. Bhat [Lun03]. Bhattacharyya [BH08]. Bias [BAW08, CR00b, DT00, GFK07, GP07, Gre03, JKN08, JKKW03, RM09, RFS06, SL06]. Biased [Ano02j, AMW02, BAW08, LJ00, SNQ09]. Biasing [LJ00]. Bikas [Not03a]. Bilal [Loc04]. Bilinear [Hof05]. Billboard [BF01a]. Bimal [Ozt05]. Bin [Yan04c]. Binary [AALT00, Bak00, CJ06b, Číź08, DKS06, FD08, GA01, HH06, HZZ^{+}08, KM01, LDM09, NC02, PPR09, PGLW00, SJ04, TRH07, YZ05, ZC07]. Binary-Choice [Číź08]. Binding [JL08]. Binomiai [Chi06]. Binomial [CIK08, CDR08a, CDR08b, Dun08, HCS06, Len03a, Pac08, Tho08, Len03b]. Bioassay [DNP05]. Bioavailability [Xie03a, Xie03b]. Bioconductor [Dav07a]. Bioequivalence [Xie03a, Xie03b, Bra08]. Bioinformatics [Mil06, Sch03b, Sch03c, Zhe06, Dav07a]. Biological [AC03a, AC03d, DR05, Kar06a, Len04a, Sch07e, YKL^{+}06]. Biologically [TH09]. Biology [Dav07a, Won00b, Gen04]. Biomarker [LTMS02]. Biomarkers [CP02, ZCPL08]. Biophysics [Kou08]. Biostatistical [Duv03a, Duv03b]. Biostatistics [Lyl06, Cav06a, Sin07a]. Bioterrorism [BB03]. Birth [NN09]. Bivariate [AAK06, CFK07, CDR08a, CDR08b, DW04, Dun08, HL06b, HZAG01, KST01, KR02, LDVT09, ORvSM01, Pac08, Pen00a, Pen00b, QvdLR06, Tho08, VH02, WW00a, WW00b, WW00c, Wei08]. Black [BBCL02]. Black-White [BBCL02]. Block [GL07, Mor07, Bai04]. Blocked [Rob04]. Blockers [XTSY01]. Blockstructures [NS01]. Blocs [GM08]. Bobrowski [Bin06b]. Böhning [CN08]. Bone [CLLCL09]. Bonferroni [PF08]. Bonferroni-Type [PF08]. Bono [Sch07b]. Book [Ala07, Alb00, Alb02, All06, AC01b, AC02a, AC02b, AC03a, AC03b, AC03c, AC04a, AC04b, AC04e, AC04c, AC04d, ACLJ^{+}04, AC05a, AC05c, AC05b, AC05d, ACE05, AC06c, AC06a, AC06b, AC07b, AC07c, AC07a, AC08a, AC08c, AC08d, AC08b, And05, Ank01, Ank05, Ann05a, Ann05b, Ann06d, Ann06a, Ann06b, Ann06c, Ann08, Ano00e, Ano01e, Ano01f, Ano01g, Ano01h, Ano02e, Ano04f, Ano04e, Ano05f, Ano05g, Ano06f, Ano06g, Ano06h, Ano06i, Ano07b, Ano07c, Ano07d, Ano07e, Ano08b, Ano08c, Ano09b, Ano09c, Ano09d, Ano09e, Aqu06, AC03f, Ass07, Bai04, Bai05, Bal05, Bal03, Ban06, Ban07, Ban05, Bar08, Bar01a, Bar01b, Bas03, Bei04, Ber04, Ber01, Ber00c, Bet03, BT08, Bin02]. Book [Bin06a, Bin06b, Bog04, Boi01, Bol02, Boo04, Boo06, Boo07, Bor03, Bor02, Bou07, Bra06a, Bra06b, Bra08, Bre05, Bre06, Bri07, Bro01, Bro07, Bue02, Buh00, Buo06, Bur08, But00, But06, Caf07, Cal01, Can00, Can07, Car01, Car03a, Car05a, Car05b, Car07, Car06, Cav04, Cav06b, Cav06a, Cav07, Cha03a, Cha03b, CH08, Che08, Chi06, Cho06, Chr01a, Chr01b, Cla03, Cle01, Cle02a, Cle02b, Cle03a, Cle06, Cli07, Coa00, Cog02, Cor02a, Cor02c, Cor05, Cow00a, Cow00b, Cow03b, Cow06, Cow07, Cra07, Dan03, Das02a, Dat04, Dav03, Dav06a, Dav01, Dav06c, Dav06b, DKC^{+}06, Dav07a, Dav07b, Daw03, Dea08, Del02a, Del06, Dev06, Dha01, Dia00a, Dia00b, Dia00c, Dia03a, Dia07, Dig05, Dix06, Dmi05, Dob02a, Dob03a, Dob03b, Dob04, Dob00, Dou06]. Book [Dow00, Dra06, Dua01, Dub01, Dur06, Duv03b, Dyk00, Eas03, Elb08, Ell01, Els02b, Emb01a, Emb01b, Emb01c, Emb02a, Emb02b, Emb02c, Emb03, Ens02, Epi04, Ern00, EP06, Eub07, Fel02a, Fel02b, Few04, Fie06, Fil06, Fra08, Fre04, Fri07, Fri02a, FS00, Fue03, Ful02, Gad00, Gal05, Gal07, Gan03a, Gan02, Gar00a, Gar00b, Gas03a, Gem06, Gen02, Gen07a, Gen06, Gen01, Gen07b, Gen04, Gen05b, Gen05a, Gey01, Ghe01, Gho01, Gho06, Gho07, Gil08a, Gil08b, Gol02a, Gol05a, Gol07, Gol02b, Gol05b, Gor08, Gra02, Gre07a, Gre07b, Gri02a, Gue08, Gun02, Hab03, Had05a, Had05b, Hal03b, Hal06, HSSM06, Hal07, Hal08, Har00b, Har07, Har08, Hav07, Haz00, Hel00, Hen02, Hes00, Hey08, Hin07, Hol00b, Hol00a, Hoo07, Hor02, Hu04, Hur00, Iha06, Imh03a]. Book [Ing00, Jam00a, Jam00b, Jam00c, Jam00d, Jam00e, Jam01a, Jam01b, Jam01c, Jam01d, Jam02a, Jam02b, Jam02c, Jam02d, Jam03a, Jen02a, Jen06, Jer06, Jes03a, Joe03, Jon00, Jon06, Jur06, Kac04, Kaf03, Kar02a, Kar02b, Kar02c, Kar03a, Kar06a, Kar06b, Kar06c, KHACM06, Kar07a, Kar07b, Ken03a, Ker06, Kie05, Kie06, Kie07a, Kie07b, Kie08c, Kie08b, Kie08a, Kim01a, Kim01b, Kim06, Kle06, Koe02b, Kok01, Kol02a, Kor04, Kor05, KL05, Kra08, Kra04, Kri05, Kue00, Lan00, Lan03, Lar02, Lar06, Las06, Las07b, Las07a, Las08, Lee03, Lee06, Lee08, Leh05, Lei05, Len02a, Len02b, Len02c, Len02d, Len02e, Len03b, Len04b, Len04a, Len05, Len06, Lev00, Li08a, Lin04a, Lin04b, Lin03a, Lin04d, Lin07, Lit00, Loc02a, Loc04, Lue05, Lun00a, Lun00b, Lun01, Lun02]. Book [LACPE03, Lun03, Lun05, Lun06, Lun07a, Lun07b, Lüt03, Lyl06, MTW00a, MTW00b, MTW00c, MTW00d, MTW00e, MTW00f, MTW00g, MTW00h, MTW00i, MTW00j, MTW00k, MTW00l, MTW00m, MTW00n, MTW00o, MTW00p, MTW00q, Mac01, Mac06a, Mad01, MWE^{+}04, MPL^{+}04, MDACL05, MCL05, Mal05, Man07, Mar03a, Mar05, Mat05, McC02b, McC02a, McC05, McG02, McN01, McN02a, McN04, McN06, McN07, Mey01, Mic06, Mil06, Mir05, Moe03a, Mol05, Mor03, Mor05, Mor04, Mül05, Mun04, Mün08, Mye07, Nea06, Nem04, Nes05, Nes07, Neu06, Nol03a, Not03a, Nov05, Nta07, Nus06, Ogd00a, Ogd00b, Ogd02, Ols02, Ott06, Oun08, Ozt05, PE00a, PE00b, PE00c, PE00d, PE00e, Pad00, Pal08, Par07, Par06, Par00, Par02a, Pea02, Ped01, Peñ06, Per02a, Per02b, PESJ03, PACP^{+}03, PEACL04, PLL^{+}05, Per06, PC06, Pic05, Poo04]. Book [Poo06, Pou04, Pra00, PS06c, Pre00, Pre02, Pre03, Qiu05, RLS00a, RLS00b, RLS00c, RVL00, Rac01, Ram05, Rao00, Rao03a, Rat00, Ree08, Rei00b, Rei02, Rei08, Rhe02, Rhe03a, Rhe07, Rhe08, Rid07, Rie06, Ril02, Rob00, Rob03a, Rob04, Rob08, Roc00, Rog02b, Roh05, Rol01, Rol00, Rol02, Ron00, Ron05, Roo02a, Ros04, Ros06, Rou06, Rue04, Ruk03a, Rup01, Rup02a, Rup04, Rus05, Rut00, Rya05, SRJ00a, SRJ00b, SRJ00c, SRJ00d, Sac03a, Sai06, SC00, San03a, Sch01, Sch02c, Sch07a, Sch07c, Sch03b, Sch06, Sch00a, Sch05, Sch04a, Sch00b, Sch02e, Sch04b, Sch07e, Sen01, Sey05, Sha07, She02b, Shi01, Sho05, Sig00, Sin06c, Sin07a, Sin07b, Sor00, Spa02a, Spe06a, Spe06b, Spi06, Sri02, Sta08, Sta01, Ste05a, Ste00b, Ste06, Ste04, Sto00, Sto08]. Book [Sto04, Str01a, Str00h, Str00i, Str00a, Str00b, Str00c, Str00j, Str00d, Str00e, Str01b, Str01c, Str01d, Str01e, Str01f, Str01g, Str02b, Str02a, Str03b, Str07, Str00l, Str02c, Stu00, Stu06, Sun02, Sun01, Sun08, Sun00, Sur04, Swa00, Swy02a, TJV00a, TJV00b, TJV00c, TJV00d, TJV00e, Tak00, Tay00, Tay03a, Ter04, Teu03, Teu08, The00, Tho02, Tim03, Tiw00, Tra05, Tri02a, Tro06, Tyl07, Tyl08, Van05, Vie02a, Vla08, Vog01a, Vog01b, Vog01c, Vog01d, Vog02a, Vog02b, Vog02e, Vog02c, Von06, Wal07, Wan01, Wan02, Wan06, Was01, Wat07, Weg00, Wel00, Wel01d, Wel01e, Wel01f, Wel01g, Wel01h, Wel01a, Wel01b, Wel01c, Wel01i, Wel01j, Wel01k, Wel01l, Wel01m, Wel01n, Wel01o, Wel01p, Wel01q, Wel01r, Wel01s, Wel01t, Wes07, Wie06, Wik04]. Book [Wik06, Wil03a, Wil04a, Wil04b, Wil06b, Wil06a, Win03, Wol08, Wol02a, Woo04b, Woo07, Xie03a, Yan04b, Yan04a, Yan05a, Yan06, Yan08a, Yan04c, Yan05b, Yan08b, Yin04, Ylv01, Yor06, Zal02, Zha06, ZS02a, Zhe06, Zhu06, Zim03, Ziv06, Zwi00, deL05, vdV01]. Boosting [BY03]. Bootstrap [Aba02, ADW07, Ann05a, Bhl02, Büh02, BP02, DKS06, Efr00, FHY07, GW05, GL07, Hal02, HH02, Lah02, Pol02, SBVW06, She02d, YW03, Can07]. Bootstrapping [LP06, PP05]. Borchers [Few04]. Borgelt [Car03a]. Boris [But06]. Borooah [Win03]. Boston [HCS06]. Bouchaud [Bre06]. Boundary [HS02a]. Bounded [Gan09, MRT05, MVF06, WQ07]. Bounded-Influence [MRT05, MVF06]. Bounding [BBS00]. Bounds [Ano00g, BJM06, Mao06]. Bouvier [AC04e]. Bowhead [Alt01, GP01, QM01, RCGT01a, RCGT01b, Zeh01]. Box [Spe06b, FTW01]. Boxplot [SGC05]. Boyd [Lue05]. Brain [SCG^{+}07, TW07a]. Branching [VS08, YSY08]. Brand [BCM02]. Braun [AC04b]. Break [DLRY06]. Breakdown [CVF06, HO02a, HO02b, HRV02, MY02, Por02, Rup02b]. Breaking [DXC08, IJ01, Lin04b]. Breakpoints [SW02a]. Breast [CAC^{+}08, IPBS00, SQC07]. Breastfeeding [MPK09]. Breeding [Gia00]. Brett [Len04a]. Brian [Daw03, Hal06, Lun06, Mor04]. Bridges [Sch02b]. Bridging [Sch03a]. Brief [Tsa00]. Brock [GGR07]. Broken [BFHR03a, BFHR03b, BX03, KZ03, MJB03]. Brownian [LM09a]. Brownian-Like [LM09a]. Brunsdon [Zim03]. Bryk [Dav03]. Buckland [Few04]. Bucklew [Pic05]. Building [FS04b, ZWL^{+}04]. Burlew [Bas03]. Burr [AC06a]. Business [Cle06]. Byrne [AC03b].

C [All06, AC04c, Ban06, Dev06, Duv03b, Fie06, Fre04, Gol05a, Imh03a, Kaf03, KL05, Lyl06, Mil06, Roh05, Sch06, Ste04, Yan04c, Yan05b, Mat05]. Cable [CLR06]. Calculation [SPJ01]. Calculations [BIK00]. Calculus [Ber04, KSG05, Nov05]. Calibrated [Bri04, GRG^{+}04, GRGB04, GLW^{+}06, GR06a, IC05, Rou04, TN04].

Calibrating [Zha03]. Calibration [Ano02m, BFV01, FHY07, HBK^{+}09, HGWR08, Hua02, MR05, PAS^{+}03, RMS09, WS01b]. California [HI06]. Calínski [Bai04]. Call [BGM^{+}05, WBS07a]. Callbacks [ELL00b]. Camille [Teu03]. Campaign [HB09, LZHR01]. Campbell [AC05a]. Campolongo [Per06]. Can [FHY07, Hil08, LLL08a, SCS08a]. Canadian [HS07b]. Cancer [Bak00, BP01, CAC^{+}08, CLLCL09, GMK^{+}04, GT07, GBT09, IPBS00, IC05, LTMS02, MFCF05, New02, RXS^{+}07, RRS07, SQC07, SRA08, SPJ01, SCC^{+}06, YTS08]. Capacity [Li07]. Capture [BF06, BP08, CN08, CBL^{+}08, Pol00]. Capture-Recapture [Pol00]. Carbon [BHO07, LSQ03]. Carcinogen [MWL^{+}01]. Carcinogen-Induced [MWL^{+}01]. Carcinogenesis [CG03b, Kol03b, MVBC03b, MVBC03c, WH03b]. Carcinogenicity [WS01a]. Care [Ano02k, Ano02i, GBM^{+}01, Gus09, LNR03, RZF06, RS09b, TEJ^{+}04, ZSCW06]. Carl [EP06]. Carl-Erik [EP06]. Carlo [All06, BL07, Dia03a, Fra08, Wie06, Ano00f, BZ00, CR00a, CM06b, CDHL05, Dia03b, FS01, Gan09, GDW04, GM06, HC01, HH07, IZ07, IT02, JHCN06, KJ07, KF07, LPD00, LS01, LW01b, Lia02, Lia05, LLC07, LLW01, Liu03, MPS00, Ros03, Tan04, TLW07b, TLW07a, dV04, Caf07]. Carol [Dig05, Zhu06]. Carotid [SSVV09]. Carriage [AALT00]. Carroll [Lee06]. Carryover [KS02, KS08]. Case [BFHR03a, BFHR03b, BX03, CCDH08, Che01, CAC^{+}08, CCC09, Eas03, EM05, Gre03, HR06, IK08, KZ03, KL04b, Lüt03, MJW06, MK06, MAC^{+}06b, MPRR04, MJB03, PM05, Ric04, Ric08, SSVV09, SI02, SYCB08, SMG^{+}05]. Case-Cohort [Che01, KL04b]. Case-Control [CCC09, MJW06, Ric04, Ric08, SMG^{+}05]. Castillo [Nem04]. Catch [KB04]. Catch-Effort [KB04]. Categorical [Ann06b, Bhl02, BT08, Büh02, DX09, Hal02, Lah02, MGG^{+}07, PAS^{+}03, Pol02, She02d, ZWS06, Ste06, Ste06]. CATS [SW05a, SW05b]. Causal [AC05a, CS00a, Cox00, Daw00a, Daw00b, EW08, Gre00, HBR01, HR06, HH06, HH08, IV04, JR08, LLL08b, Pea00, RG00, Rub00, Rub05, Sha00, SRA08, Sob00, Sob06, Tan06a, Tan06b, TEJ^{+}04, Was00, YL01, ZRM09]. Causality [QA09, Had05b]. Cause [MTR^{+}01, PCK06]. Cause-Age [PCK06]. cDNA [GRYB06, HWZ05]. CDs [McC02a]. Cell [FG00, Net09, RFP09, YSY08]. Cells [RSH^{+}02]. Censored [Ano02m, BB01, CDK05, CH02, Dix06, DFWJ06, Hel07, HC02, Hua02, KL08, LS09c, NVP06, Par07, Peñ06, Por03, QvdLR06, SC03, Str00f, SZ09, UCTW07, VH02, WW09, Yan04b, YZL08]. Censoring [Ano02j, AMW02, CLLCL09, LC09, OLT09, PF06, SSL07, WQC01, Wu02]. Censorship [Efr01, Tre06]. Census [BST03, EL00, EL05, HT04, Sch03a]. Center [BGM^{+}05, WBS07a, GLW^{+}06]. Central [Joe03, ZZ06]. Centuries [Dob02b]. Century [SRRW00, Nol03b, Sco02, Nol03a]. Cerebral [OMS^{+}09]. Cerioli [Ban06]. Cervical [SCC^{+}06]. Cessation [LDM09, LCLC08]. CGH [GLN08]. Chain [BL07, CR00a, CRY09, FS01, Gol05a, GM06, HC01, HH02, HH07, IZ07, JHCN06, KJ07, KF07, LPD00, Liu03, MSD04, Ros03, TLW07b, TLW07a]. Chains [Bhl02, Büh02, Emb03, Fry05, Hal02, Lah02, Pol02, She02d, Sis05, Teu08]. Challenges [Nyc00, Wei00]. Chalmond [Gem06]. Chambers [Fie06]. Chan [Wil06b]. Chance [Sch02d]. Change [And05, BGS03, Dra06, HT06, LL03a, MRID01]. Changepoint [DHS08, HS02a, SPJ01, SWDS03]. Changepoints [Ano02o]. Changes [DHSR08, HID09, RL02]. Channel [RTD01, dGKS01]. Chaotic [Gol02c]. Characteristic [BP01, CP02, CCPY09, DP03, DP06, LZ08, Pep00, RFS06]. Characterization [IC05, LSGS01]. Characterizations [PV06]. Characterize [MAC^{+}06b]. Characterizing [FD03, Pui03, SCC^{+}06]. Charges [RZF06]. Charles [Cow06, Lyl06, Moe03a, Ste06]. Charlton [Zim03]. Chart [HT06, MRW07]. Charting [HT06]. Charts [Wei08]. Checking [FRZ05, Zha03]. Checks [DKS06, Gal05]. Chemometrics [Sun08]. Chen [Ozt05, Yan04a, Yin04]. Chi [BC01, Dui16, Zha05b]. Chi-Squared [BC01]. Chi-Squared-Type [Dui16, Zha05b]. Chiang [Rus05]. Child [BM01]. Childhood [Gre03]. Chin [Rus05]. Chinook [NR02]. Choice [Ala07, Bal05, BFHR03a, BFHR03b, BX03, BCM02, Bro07, Číź08, Dea08, DH08, Efr04a, Frö08, KZ03, LLL08b, MJB03, OR01]. Cholesky [HRC04]. Cholesterol [ZWT^{+}05]. Chow [Xie03a]. Chris [Zim03]. Christensen [Hal03b]. Christian [Bei04, Car03a, Yor06]. Christophe [Gol05b]. Christopher [Yan04c]. Christy [Dur06]. Chronic [NSP00]. Chuang [Dur06]. Chuang-Stein [Dur06]. Chun [Ann06a]. Chung [Xie03a]. Circle [JP05, RFP09]. City [GFK07, BFHR03a, BFHR03b, BX03, KZ03, MJB03]. Claims [FV08, GFK07]. Class [BF06, BNV09, BP02, BFB00, CN02, CM06b, ES02, LTMS02, LTE08, Liu00, Mao04, OLT09, PDG02a, PDG02b, Sea02, SZ09, Ver02, YW03, ZYI05]. Classes [AKM01, Ano01m, LL05b, NL07, RHC07]. Classical [Ann06d, Ano01n, CT04, FS01, KB04]. Classification [BJM06, BY03, Cha04, CV07, CMRT09, DFS02, FJS08, FO09, HOS04, KS07a, KS07b, KL01, LLW04, Mar07, Qiu05, WRM07, WSP09, ZIP05, Yan04c]. Classifiers [HTX09, Wil03b, Wil03a]. Cleaning [CvA02, MH02, Sim02, SD02, WR02a, WR02b, Kar06b]. Cleary [Dav06b]. Climate [LL03a, STNM09]. Clinical [Ass07, CH08, Dmi05, Dur06, Har00a, LA08, Pal08, Ros06, SC03, Sin06c, Sto04, TSE02, YY09, Lin03a]. Clipped [KCO08]. Closed [Few04, Pui03]. Cloud [SYCB08]. Cluster [Bog04, FJS08, HB09, KMR09, PP01]. Cluster-Randomized [HB09]. Clustered [CCW02, CJ06a, DS05, GA01, HFZ05, KRS^{+}06, LL03b, LZ08, LC00, LC01b, SS04, WCL05]. Clustering [CL08a, DN07, DHE08, FR02, HHS06, JS03, JL08, LHNM08, MZ08, RD06, RSH^{+}02, SI02, SW02a, SW05b, TSV05, ZWS06, But06]. Clusters [DHSR08, SJ03, YC02, ZSB01]. Co [Har07, LSGS04]. Co-Integrated [Har07]. Co-Occurrence [LSGS04]. Coaching [Han04]. Codes [DM06, Gre03]. Coding [RY00]. Cody [Cla03]. Coefficient [CFL00, CFY00, CX08, CX09, CRW01, GKSB03, KS00, WLH08, WKL09, WX09, XZ07]. Coefficient-Dependent [KS00]. Coefficients [CW08, FZ09, OOK03, SW02b, TZW05, YPXH06, YLZ08]. Cognitive [Len04b]. Coherence [RW00]. Cohort [Che01, CAC^{+}08, CM09b, KL04b, MK06]. Cointegration [CH03]. Colin [McC02a]. Collaborators [Ano05a, Ano06a, Ano07a, Ano08a, Ano09a]. Collanist [Len03b]. Collapsed [vDP08]. Collection [MD01]. College [Har03]. Colon [CCDH08, CG03b, Kol03b, MWL^{+}01, MVBC03b, MVBC03c, WH03b]. Colonization [ZRM^{+}08]. Colton [Cav06a]. Combination [BBN^{+}04, BPB02, IPC08, LP05, SH06]. Combinations [LSC04, New02]. Combinatorics [Mor05]. Combined [LBS06]. Combining [EL00, FIC01, GY02, JM04, KR08, Mer04, MMZ02, RXS^{+}07, RT00, YY05]. Cometary [JKKW03]. Comment [AM01, Ald08, Alt01, Ano02n, BLW01, BX03, BR00, Boo00, Bri04, Bur04, CN02, CS00a, Cau01, CSCC06, CV07, CM01, Chr07, CG03a, CG03b, Cox00, Cra05, CvA02, Dem04, DLM04, DL08, Dun08, Edw05, Edw08, ES02, Els02c, Epp04, Eva00a, Eva00b, FW00a, FF06, FL09, FF09, Far01, FKW01, GS01, GC09a, GP01, Gro01, HL06a, Hal02, HW06, Hay00, HP04, HH07, Hig00, Hig09, Hil08, Hod08, Hol05, HW01a, HZ05, HRV02, HRV04, IZ07, Jia05, Kad07, Kal00, KJ07, KM02, Kat08, Kni06, Kol03a, Kol03b, KM05, KF07, KZ03, Lah02, Lav04, Laz07, Li00, Li06, LQ05, LZ01, LC01a, LZS^{+}06, Lin04c, Lit07, LLL08a, Lou05, Mag01, MG00, Mar00a, Mar07, MY02, Mar01, McC04, MH02, Men09, Mil07, MJB03, Pac08, Pea00, Pen00a, Pen00b, Pic00, Pol02]. Comment [Por02, QM01, Rag00a, Rag00b, Ram01a, Ram01b, Raz00, RK01, RG00, RRvdL00, Rob06, Rou04, Rub00, Rub08, Rup02b, Rup05, Rup07, Sab05, Sab06, SAE06, Sch02a, Sch09, Sch07d, Sea02, Sen08a, Ser04, SCS08b, Sha00, Sha05, She00b, SHY04a, She02d, Sim02, SD02, Sta09, Ste00c, Sti08, SW09, Tam05, TN04, Tho08, Tib05, TR06, Ver02, WC01, WW01, Was00, Wes05, Whi01, WM09b, Wol00, Wol07, Won00a, Woo01, Wor00, WH03a, WH03b, YW01, Zad04, Zeh01, Zha04a, Zha05a]. Comments [BYE^{+}05, EBD^{+}04, FPH^{+}05, GRG^{+}04, JL09b, KXF^{+}06, MMH^{+}04, Ram09, SBL^{+}04, SFK^{+}05b]. Communicating [Lac09]. Communications [McN02b]. Communities [DR05]. Community [BM01, LSGS01, LSGS04]. Companion [Lun06]. Comparative [AC05c, AC05b, HC01, ICLH05]. Comparing [Ano01n, BP01, Bra06a, FD08, Hil08, Koo02, LLL08a, PDK07, SCS08a, SGL07]. Comparison [ÁEdBCAM08, DFS02, HT06, Lee00, LZ08, LJZ04, SSL03, Kie05]. Comparisons [FZ09, HR03, JC07, LA06, SJS05, vdW02]. Competing [CS04, CFK07, Lit04, PF09]. Competing-Risks [CFK07]. Competition [Har03]. Complete [Peñ06, WS01b]. Complex [CN02, CGRS01, ES02, GR06a, NS02, PDG02a, PDG02b, Sea02, Ver02]. Complexity [Gol02c, Lee08, WS06]. Compliance [Bak00, GSCR04, JR08, LTE08]. Complications [CLLCL09]. Complier [YL01]. Complier-Average [YL01]. Compliers [Fol00]. Component [CM01, EHI08, Far01, Gri02b, KU01a, KU01b, LM00, Mar01, Ram01a, RO07, SSV05, Tim03, TIY03]. Components [BHPT06, EHI08, JL09a, LSC04, LS02, Ram00, SBVW06, SW02d, Kac04]. Composite [BB00a, BB00b, Boo00, Eva00a, Gua06, Mar00a, Pic00, RvdVV00a, RvdVV00b, Ste00c]. Composition [BGF01, DR05, PM09]. Compound [SC07]. Comprehensive [Rog02a]. Compression [RY00]. Computation [Lia02, Lia05, LLC07, LLW01, Kor05]. Computational [Bar08, CHR00a, Dav07a, Gen04, Hob00, IJ04, Lee03, Sch04a, Won00b, Bai05, Hav07, Jon06, Wik04]. Computationally [Ned09]. Computed [CCDH08]. Computer [BBK^{+}09, CGRS01, DM06, GL08, HBK^{+}09, HGWR08, Mor04, Swy02b]. Computer-Based [Swy02b]. Computing [Buo06, HC01, Li07, Sco02, Van05]. Concentration [CHB^{+}08]. Concept [LPR09]. Concepts [AC06b, RW00, Ros06, Wil04a]. Concise [AC03d, AC03a]. Concomitants [WSLC06]. Conditional [Ano00f, BG01, CD09, DZ03, FS04a, HM03a, HRL04, KST01, Kol02b, Mao04, MB05, MPS00, PSW05, PY00, Ric04, Ric08, RAHT06, SS00, VWNM02, Wei08, WL01, XK09, YJ04]. Conditionally [LL04, MRT05, RHC07]. Conditionals [SJ06]. Conditioned [MJvAC02]. Conditioning [RW00]. Conditions [YKL^{+}06]. Conducting [Cle03b, Cle03a]. Confidence [Ano00g, BYE^{+}05, BY05, CC09a, CFN01, Edw05, HIP06, Hol05, IWM04, KL06a, LSC04, Len06, LH07, SS09, Sha05, Tam05, Wes05, XSZ09, YK04]. Configural [Kra04]. Conformation [LDVT09]. Confounding [Gre03]. Congdon [Ann06b, Cow03b, Mül05]. Conjecture [Cle02c]. Conjoint [MB02]. Conserved [GL03]. Considerations [CL03, IR03b, Joh03, SD03, Swy02b]. Consistency [FPH^{+}05, JL09a, LPW05]. Consistent [FO09, Gua08, HW01b, KL06b, TPO03]. Consolidation [Li07]. Constant [NLLH06]. Constrained [NT05, STJ07, Mic06]. Constraint [AKS02, EM05]. Constraints [FG00, Liu00, PT06]. Constructed [WF03]. Constructing [LBS06, PW05, SS09]. Construction [AC05c, AC05b, ACS06, Dea08, LPD00]. Constructive [FS06a]. Consumer [IK08]. Consumption [vSH08]. Cont [Bin06a]. Contact [FW05]. Contaminant [LSGS04]. Contamination [KTZ00]. Contemporary [Dan03, Ott06]. Context [FS04a, GFK07]. Contingency [CM09a, Fie00, Jam02e, Lan05, MOJ05]. Continual [YY09]. Continuous [BF06, DV03, GDS03, GB00, GA01, LDM09, WZ05]. Contoured [GEG05]. Contours [Wei08]. Contrast [SWSM09]. Contrasts [Ano00g, WT07]. Contributions [Tho00]. Control [CAC^{+}08, CCC09, GW06, GG01, GRYB06, LPL06, LFGR04, MJW06, PGVW04, PDK07, Ric04, Ric08, SMG^{+}05, SRRW00, SC07, ZQ09, AC06a, Fre04, Gen05b, Nem04]. Controlled [FBV^{+}04, SQC07]. Controlling [WBS07b, Yek08].

Convergence [Ros03, SL06]. Convex [LM04, Lue05]. Convexity [BJM06]. Convolution [LSZH08, SSL03]. Cook [AC05a]. Coordinate [MJvAC02]. Copeland [Bra06a]. Copula [CFT06]. Core [Bar08]. Corner [MRŞK06]. Corporate [JRY04]. Corrected [NS02, Tsa03]. Correction [Ano00g, Ano00f, Ano01j, Ano01i, Ano02k, Ano02i, Ano02j, Ano02f, Ano02g, Ano02h, Ano03g, Ano05j, Ano05k, Ano05h, Ano05i, Ano06l, Ano06j, Ano06k, Ano08d, Ano09f, Ano09g, Dui16, FZ09, HW00, NVP06, NDKW01, SL06]. Corrections [Ano01n, Ano01m, Ano01o, Ano01p, Ano01k, Ano01l, Ano02m, Ano02n, Ano02o, Ano02l, Ano03g, Ano03j, Ano03h, Ano03i, Ano03e, Ano03f, Ano08e, Ano09h, Ano09i, DP06, MFCF05]. Correlated [Ano03i, BDS01, CL08b, FSL^{+}06, GA01, HT04, HRC04, KMR09, KK07, LL06c, PGH01, PT09, dGKS01]. Correlation [CHL06, CL08a, DM05, Efr07, Gil03, PWZZ09, SW02b, WC04, War08]. Correlation-Based [CL08a]. Correspondence [LM00]. Cortical [MMZ02]. Cost [Ano02m, Elb08, HY07, Hua02]. Cost-Effectiveness [Elb08]. Count [Dun05, IJ04, LZH09, PV06, TEB00, ZSG06]. Counterfactuals [CS00a, Cox00, Daw00a, Daw00b, Pea00, RG00, Rub00, Sha00, Was00]. Counting [SjKH^{+}01]. Counts [GT07, HCS06, LL06b, MS05, Sch03a]. County [DK01, RXS^{+}07]. County-Level [RXS^{+}07]. Couper [Nus06]. Course [Caf07, Gen00a, Gen00b, Hay00, Hig00, MG00, Mor03, Raz00, Vla08, Wor00, YKL^{+}06, Fil06]. Courtroom [Las08]. Covariables [SXZ07]. Covariance [Ano01m, Bur04, CvA02, DLM04, EBD^{+}04, Efr04b, FHL07, FW08, Gne02, KC01, KSN08, Koe02a, KL06b, LGS07, MH02, OZ08, OOK03, RLZ09, SHY04a, Sim02, SK02, SD02, Ste05b, WR02a, WR02b, WA06, War08, Zha04a]. Covariate [BAW08, CW01, EHT06, HM00a, HM00b, HM00c, KL00, LR04, QSZ08, Rag00a, Rag00b, SWSM09, SRL00, WA06, Wei08, Zuc05]. Covariate-Dependent [QSZ08, Wei08]. Covariates [Ano01o, Ano03g, BF06, Che02, Che04, CIS04, CZI07, CC09b, DZ03, Efr05a, GT09, HI01, HCS06, HL04, HW00, HW01b, LWRC04, MY08a, MZ08, QWP05, RL01, RL02, RAHT06, RT00, Wu02, Wu04, XPLT09, ZMP06]. Coverage [CH06b, CFN01, EL00, KL06a]. Cox [Ano03g, Kac04, Ano01o, Che02, CC09b, FTW01, HI01, HC06, HW00, TZW05, XPLT09]. Crash [ERS02]. Crashworthiness [BBK^{+}09]. Crisis [Mas04]. Cristianini [Las06]. Criteria [HC06, IZT08, Iri01a, KL04a, KE02, Tsa03]. Criterion [CHS^{+}03, Kim00, NST07]. Critical [Woo04b]. Cross [Bur04, DLM04, EBD^{+}04, Efr04b, FL09, GC09a, HRL04, Lee00, NLLH06, PAS^{+}03, STPN09a, STPN09b, SHY04a, Zha04a, Sto04]. Cross-Calibration [PAS^{+}03]. Cross-Over [Sto04]. Cross-Ratio [NLLH06]. Cross-Study [FL09, GC09a, STPN09a, STPN09b]. Cross-Validation [Bur04, DLM04, EBD^{+}04, Efr04b, HRL04, Lee00, SHY04a, Zha04a]. Crossed [CHRH01]. Crossover [HY04, HSY06, KS02, KS08, PC00]. Crystal [BNV09]. CSFE [HZ08]. CSFE-Labeling [HZ08]. Cull [TEB00]. Cumulative [MRW07]. Cure [CBCS07, LS09c, MY08a, OLT09, TIY03, YTS08, ZYI06]. Cured [LCLC08]. Current [BST03, DW04, Hob00, She00a, SMIO08, XLL04, Gen04]. Curve [BM08, CE09, CFN01, DP03, DP06, McN07, RFS06, Sch02d, TI08, WRM07, Zha03]. Curves [BFV01, CCDH08, HHS06, LZ08, LM04, ZW08]. Curvilinear [BG06b]. Cuscore [HT06]. CUSUM [Dra06]. Cycle [RFP09]. Cyclostationary [CV07, KS07a, KS07b, Mar07]. Cyganowski [Gan03a]. Cystic [YF08].

D [AC05a, And05, Bor03, Gen05a, Kac04, Mal05, Nes05, Neu06, Peñ06, Rie06, Rue04, Ruk03a, Yan04c]. D’Agostini [Woo04b]. Daily [KOC07]. Dani [Jes03a]. Daniel [Gen06, Mar05]. Dara [Ano03i]. Dark [WWP^{+}08]. Darkness [GR06b]. Dasu [Kar06b]. Data [AM01, AAK06, AKM01, AC01a, AC03e, And05, Ano01p, AMvW^{+}09, AALT00, AHA01, Bal03, Ban06, BF06, BF09, BBK^{+}09, BB01, BDS01, BIW00, Bha09, BT08, BD03, Boo04, Boo06, Bor03, Bow07, BHO07, BIK00, But06, BFB00, CCW02, CFJZ07, CDVY00, Car03a, Car03b, Car07, CN02, Cau01, Cha04, Che01, CJ06a, CAC^{+}08, CW01, CFK07, CJ06b, CM09b, CA02, Cla03, Cle06, CS00b, Con08, CK04, Cra05, Cra07, DR00, DN03, DS05, Dav03, Dav06a, Daw03, DKS06, DW04, Dix06, DM05, DFS02, DFWJ06, Dun03, Dun05, DX09, EOU02, EL04, ES02, EL05, EW08, FL04, FPHR05, FPH^{+}05, FSL^{+}06, FHL07, FW07, FW08, FW05, FG00, FM06, FS04b, Fue07, GWKK01a, GWKK01b, Gen00a, Gen00b, GDS03, Gne02, GY02, GB00, Gre07b, GL06, Gro01, GLN08, GRT04, HTX09]. Data [Hay00, HFZ05, Hel07, HSG02, Hig00, HCB06, HTO03, Hof05, HT04, HR06, HW01a, HM00a, HM00b, HM00c, Hor01a, Hor01b, Hu04, HW04, HZ05, HWZ05, HZW07, HSB09, HTZ^{+}09, ICG02, ICLH05, IZT08, IJ04, IR05, JS03, JvD03, JR07, Jer06, JDvS02, KRS^{+}06, Kar06b, KSN08, KM07, KMH05, KM01, Koe02a, KL08, KM05, LL05a, Lar06, Las07b, Laz07, LLW04, Len05, LASA03, LL03b, LL06c, LZ08, LC00, LY01a, LC01a, LY01b, LC01b, LTMS02, LHNM08, LS09c, LJ00, LPR09, LZH09, MZ08, Mag01, MG00, MGG^{+}07, Mar05, MK06, Men00, MFCF05, MCG00, NC02, Neu06, OZ08, OMS^{+}09, OS01, PGH01, PAS^{+}03, PDG02a, PDG02b, PDK07, Pen00a, Pen00b, PSH01, PPR09, Poo06, PS06c, PV06, QLS02, QZL09, QvdLR06, Rag00a, Rag00b, Raz00, RHC07, RK01, Rob08, RAHT06, Rup04, Sab05]. Data [SSVV09, San03a, San03b, SRC^{+}06, Sch07d, Sea02, SR01, Sey05, SW02b, She00a, SC03, SNQ09, SYCB08, SMIO08, SK02, SCG^{+}07, Ste06, SXZ07, SPSZ05, SSL07, SZ09, TSV05, TPO03, Tib05, TW07b, TEB00, Tre06, TIY03, VH02, Ver02, WW00a, WW00b, WW00c, WQC01, WT01, WC01, WW01, WCL05, WM09a, WBS07a, Whi01, Woo01, Woo04b, Wor00, WS01b, WZ02, XSC08, XLL04, XZ07, YF08, YMW05, YL01, YW01, YC02, YZL08, YKL^{+}06, YZ05, ZYI06, ZC07, Zha03, Zha05a, ZMAS05, ZWS06, ZSG06, ZWW08, Zhu06, ZZIP07a, ZZIP07b, dGKS01, vSH08, Yan04b, AC03b, Ann06b, Bet03, Car05b, Dig05, Dix06, Fie06, Gol05b, Kle06, Lin04b, Peñ06, Sai06, Wil06a, Yan05a, Zha06, AC04b, Boo07, Elb08, Hal07, Par07, Rei08, Sin07b].

Data-Driven [AKM01]. Data. [Boo06]. Databases [DM00]. Dataset [FWP05, SJ03]. Datasets [THC05]. Datta [Rou06]. David [AC03b, Bog04, Few04, Iha06, Kar06c, Lee06, Lyl06, McC05, Nov05, Sac03a, San03a, Sch06, Str03b, Tay03a, Yan04c, Yan05b, Yor06, Ziv06]. Davis [Ste06, Moe03a]. Davisson [Rie06]. Day [DC03]. Day-Specific [DC03]. Daytime [SYCB08]. DEA [Cau01, Gro01, Hor01a, Hor01b, Mag01, RK01]. Dead [HYFW05]. Death [TSE02]. Debajyoti [Yan04a]. Debt [JRY04]. Debugging [Bas03]. Decade [Sis05]. Deceased [FWP05]. Dechert [GGR07]. Decision [Ban05, Bro00a, LH09, Mee00, MPK09, Sac03a, Sac03b, SC07, SCC^{+}06, Woo07]. Decision-Oriented [Ban05]. Decisions [MD01, Rub05]. Declustering [ZOVJ02]. Decomposable [SR01]. Decomposition [CJS02, Hof07, HWZ06, HHCK09]. Decomposition-Based [HHCK09]. Decompositions [HSB09, IJS01, LO06]. Deconvolution [LS01]. Definite [ZCI^{+}09]. Degrees [Zha03]. Dekking [Dev06]. Del [deL05, Nem04]. Delay [ERS02, TBJ08]. Delays [MFCF05]. Deletion [XLB04]. Deletions [HBK^{+}09]. Demand [Dav06b, Li07]. Demidenko [Von06]. Demographic [EL00, EL05]. Demography [Xie00]. Demonstrate [Sob06]. Dempster [Rhe03a]. Denison [Yan04c, Bog04]. Dennis [Dix06]. Densities [DHSR08, HRL04, NN06]. Density [CAC^{+}08, CM01, CGR04, CHL03, Efr01, EHT06, Far01, FH00, FR02, Gen07b, ISZ07, IJS01, Kar03a, Kar03b, KU01a, KU01b, LDVT09, LPW05, LJ00, Mar01, Ram01a, SL08, SRB08, YL09, dV04]. Department [RAHT06, GFK07, LNR03]. Departure [TBJ08]. Dependence [BFD01, FKC09, JKN08, PAS^{+}03, PT09, SCG^{+}07]. Dependent [Bal03, Car05b, CRW01, CLLCL09, DMRM04, GG07, GS06, GDOvS03, HCS06, HZ08, KS00, Lob01, OLT09, PF06, QSZ08, RSM^{+}07, SWSM09, SMIO08, SMP03, WZ05, Wei08, ZCPL08, dGKS01]. Depression [STR08, TEJ^{+}04]. Deprivation [HT04]. Depth [Epp04, HP04, HRV04, Joe03, Kad07, LPR09, McC04, Mil07, MMH^{+}04, MM04, RLL07a, RLL07b, Ser04, Wol07]. Derivative [Bre06, Hin07]. Derivatives [CD05, LM09b, Mat05]. Derrick [Mal05]. Description [HY01a, Lee00]. Design [ABPR06, AC06c, BP01, BDZ06, BIK00, Che01, Dav06a, Dav06c, DFC09, DSV04, DJSW09, DMW05, DB09, HRS09, KSG05, Len02g, Lüt03, Mat05, Mor04, MSD04, Not03b, Pal08, PKC01, Rob04, Rub05, Rub08, SCS08b, Spe06b, WL00, Xie03b, Dou06, Ker06, Not03a, AC07b, AC07a, AC08a, Gen05a, Ker06, Kie08c, Xie03a]. Design-Adapted [DSV04]. Design-Adaptive [DFC09]. Designed [AC03c, Mor05]. Designing [IP02]. Designs [AAK06, AC05a, AC08b, Bai04, BDZ06, CAC^{+}08, DB03, DBPP08, DS08, FW00b, GA05, HY04, HSY06, KS02, KS08, LPP02, LP05, LFGR04, Mor07, PC00, SRL00]. Desmond [Kor05]. Desrosieres [Teu03]. Desu [Peñ06]. Detecting [DHS08, IR03a, MRID01, RM09, TW07a, Wal02, WH07]. Detection [DN03, DJSW09, GPT06, HT06, HS07b, HHCK09, Qiu07, Rue04, SYCB08, TBD^{+}03]. Deterioration [FIC01]. Determination [De 06, GM09, MJW06]. Determining [HL07, Hös03, Hss03, ZMAS05].

Deterministic [CW08, NDKW01, PR00, Sch07c]. Deterring [DM00]. Developing [Lac09, MB09, Iha06]. Development [Bra08, GH00, LP05, Ste01]. Developments [Fie00]. Deviation [KCO08]. Deviations [KW01c]. Dey [Mor03]. Diagnostic [AD08, BP08, CN08, CCL09, FRZ05, Gal05]. Diagnostics [WBBR05]. Dial [Men00]. Dichotomization [FTR04]. Dickson [Sch06]. Dictionary [GL03]. Dietary [CN02, ES02, PDG02a, PDG02b, Sea02, Ver02]. Dietrich [Kie05]. Difference [DM06, LPL06, RS09b]. Differences [And08, FS04a, SHH00]. Different [Fry05]. Differential [FL09, Gan03a, Gan03b, GC09a, LW08, STPN09a, STPN09b, WH07, WJSS08, vWvdWvdV08]. Differentially [IR03a, RWR03]. Difficulties [CHR00a]. Diffusion [FZ03, Kie06, Laz07, Sch07d, ZZIP07a, ZZIP07b, ZCI^{+}09]. Diffusion-Weighted [Laz07, Sch07d, ZZIP07a, ZZIP07b]. Diffusions [ASFP09]. Digital [BP01, LS01]. Dijksterhuis [Roh05]. Dimension [CB09, CN05, CF09, Hof07, IYSK08, LS02, LASA03, LW07, LWZ08, LL08, LHNM08, OR01, STJ07, WX08, Xia08, YW03]. Dimensional [CCL^{+}08, Cra05, DZ03, FPHR05, FPH^{+}05, HTX09, HGWR08, HZ05, KM05, MMB09, Sab05, SSV05, TSV05, Tib05, Wan09, Zha05a, ZMP06]. Dimensionality [LO06]. Dimensions [BNV09, JL09a, KCO08]. Dipak [Mor03]. Direct [MMZ02]. Directional [ADW07, BGS03, LW07, RBP08]. Directions [LLC00, ZYZS03]. Dirichlet [BC03a, GKM05, GS06, RDG08, TJBB06]. Dirty [ST04]. Disability [IS07, KP07, PAS^{+}03]. Disability-Free [IS07]. Disc [MW09]. Discipline [Bec00a]. Disclosure [DM00, Rei05]. Discovering [Mao04, New02, Mar05]. Discovery [BYE^{+}05, BY05, BH07, Edw05, GW06, GL03, GI07, Hol05, PGVW04, PZB09, RM09, Sha05, SC07, Tam05, TBD^{+}03, Wes05, Yek08, Spe06b]. Discrepancy [CN05]. Discrete [Boo07, CHL06, GB00, Liu00, MTR^{+}01, Pui03, Wil06a, Bal05]. Discriminant [FRS06, FR02, Gho07]. Discrimination [CP06, DFS02, HOS04, MTA07]. Discussion [Cai01, CL03, IR03b, Joh03, LH06, Mou01, Nad09, Net06, RZ03, SD03, Sol01, Tsa03, Vid01, WHT09, YK06]. Discussions [YKL^{+}06]. Disease [BP08, CP02, Cha04, CN08, GR02, IEMM08, KW02, Len05, NSP00, WM01, ZIP05]. Diseases [FW05]. Disequilibrium [LCT^{+}05]. Disparate [BIW00]. Dispersion [Joe03]. Disposal [ST02]. Dissimilarities [HCB06]. Distal [MWL^{+}01]. Distance [CS04, Jam03a, Jam03b, MTA07, RRS07, Sch07e, ZWS06]. Distance-Weighted [MTA07]. Distant [XLB04]. Distinct [LC09]. Distributed [IT02, KM01]. Distribution [AMP06, ABDL01, Ann06d, Ano03j, AKS02, BB00a, BBK03, Boo00, CS04, CIS04, CH06a, CD09, DM04, DFWJ06, Eva00a, FTW01, Gew04, Gil03, HM03a, HM06, HL06b, JKKW03, Ken03b, Kon06, KTZ00, LL04, LDVT09, Len03a, Len03b, Mar00a, McN02b, Pic00, Pre03, Ran00, RvdVV00a, RvdVV00b, Ste00c, Ken03a, Gol05a].

Distribution-Free [LL04, Pre03, Ran00]. Distributional [Aba02, KL08, Tan06a]. Distributions [ÁEdBCAM08, Ann06a, ACS06, Bei04, BNV09, CHR00a, CHJ08, Dui16, Fel02c, FS06a, GEG05, GKO05, Gre03, Gus09, Hös03, Hss03, IC05, JP05, LL05b, Liu00, MGT05, McN06, PWM01, Pui03, PV06, Ruk03b, ST04, She02a, TBJ08, VCR09, WWP^{+}08, YSY08, Zha05b, vdW02, AC04a, Ruk03a, Kar07a]. District [DK01]. Diverging [GR05]. Diversity [GDH09]. Dmitrienko [Dur06]. DNA [Gen05a, Ros04, AC01a, HBK^{+}09, ICG02, MWL^{+}01, TBD^{+}03, Dav07b]. Do [Gol05b, Sob06, ZSG06]. Dobson [Lan03]. Does [LL03a, Mas04, SRA08]. Dolphins [KB04]. Domain [FFJ07, JL06, OZ08]. Domain-Level [OZ08]. Donald [AC05a]. Donor [FWP05]. Dose [BT04, BDZ06, Che07b, DNP05, DBPP08, DDZS02, MICR06, TDGW01]. Dose-Finding [BT04, DBPP08]. Dose-Response [DDZS02]. Doses [LZHR01, Ros05, TL02]. Double [Che02]. Double-Semiparametric [Che02]. Doubly [CM06b, KL08, QSZ08, RM09, RFS06, Szk03]. Doucet [Dia03a]. Douglas [Poo04]. Downside [Mal05]. Downweighting [ZSB01]. Drager [Len03b]. Drawing [CB09]. Drinking [MICR06]. Driven [AKM01]. Driving [Ste01]. Dropouts [GT09, GRT04, RL02]. Drug [Bra08, LP05, LZHR01, TL02]. Drugs [Cau01, Gro01, Hor01a, Hor01b, Mag01, RK01]. Dual [LR00]. Due [Sch03a]. Duley [Lin03a]. Duncan [KL05]. Dupuis [Dob04]. Duration [ERS02]. Durbin [AC03f]. During [CCDP09]. Durrett [Ros04]. Dyadic [Hof05]. Dyk [MH04]. Dynamic [BF09, CX08, CX09, CW08, CJ06b, Dun03, FFJ07, FHLR05, FS01, GCK00, HL07, HT06, KM01, KP05, LPS01, LLW01, MRT05, MvdLRC01, MB09, OMS^{+}09, QA09, Sin07b]. Dynamical [DM05, XWF05]. Dynamically [Lia02]. Dynamics [BM08, CW08, LM09b, PMHB09, TEB00, WJSS08].

Early [And08]. EARTH [DTT08]. Earthquake [Sch03d, ZOVJ02]. eBay [WJSS08]. Ecological [Con08, LL03a, Rob08]. Econometrician [Tro06]. Econometricians [Vog02d]. Econometrics [Bal03, Imh03b, Lei05, Ott06, Ron05, Spa02b, Wie06, Wol08, Bai05, Imh03a, Ron05]. Economic [ACK01, BM08, Cau01, Gro01, Hor01a, Hor01b, IK08, Mag01, RK01]. Economics [Bei04, Jam02g, McN02b, Sch04a, Loc04]. Economy [Ber00a, GLPS03]. ed [Cha03b, Lun03, Vog02d]. Edge [DN03, Qiu07]. Edit [GDS03]. Edit/Multiple [GDS03]. edition [AC03c, Ber04, Boo04, Bor03, Car05a, Dav03, Hal03a, Lan03, Len04a, Rog02a, Rya05, Sto04, Tim03, Wol02b, Xie03b, AC04e, Leh05, Yan04b]. Editor [Ano03h, Ano03o, Bak07]. Editorial [Ano05a, Ano06a, Ano07a, Ano08a, Ano09a]. Education [BST03, KRS^{+}06, LHN08, Ste01]. Educational [SI02]. Edward [Gen05a, Lar06, Len04b]. Edwards [Str03b]. Effect [Ano02k, Ano02i, CC09b, CHRH01, ERS02, FTR04, Fol00, GBM^{+}01, GB00, HBR01, HQF08, McB02, MPE08, SWSM09, YL01]. Effective [DNP05, MSM05, OZ00, TL02, Zha03]. Effectiveness [Elb08, VPW08]. Effects [Aba02, And08, Ano02n, BLW01, BM01, BD05, CCW02, CSCC06, CH06b, CCL09, CLLCL09, DN07, FG04, FKW01, Gal07, GS01, GT09, Gus00, HB09, HSY06, HI06, Hof05, HH06, JC07, KW01a, KW01b, KT05, KS02, KS08, Li06, LZ01, LZS^{+}06, LZ06, LCLC08, Neu06, OS01, Ram01b, Ric08, RZF06, Ros02, RS08, Sab06, SAE06, SI02, SQC07, SS04, TR06, VWNM02, WA06, Wu02, WZ02, Wu04, ZLY05, ZL07b, ZYDS08, ZRM09, Cra07]. Efficacy [HPC03, KF02, LH09, TSE02]. Efficiency [ELL00b, FTR04, FZ09, KC01, KL04b, WLC02]. Efficient [Ano00f, CFL00, CFT06, CW08, CCC09, Číź08, DB03, GCK00, HSY06, MGT05, MC06, MMC07, MPS00, Ned09, PPR09, QSZ08, QvdLR06, SW02c, WCL05, Woo04a, XLCM03, YL05, YZ09, ZL07a, vdLHR02]. Effort [KB04]. Eggermont [Kar03a]. Elart [Len03b]. Eleanor [Nus06]. Election [DK01, PM05, PLM08, Han03, HI06]. Electorate [GM08]. Electrical [Car06, MMZ02]. Electricity [CS03b, KOC07]. Electroencephalographic [CCDP09]. Elementary [Gan03a, Gan03b, Kor04, Wat07, AC06a]. Elements [Emb03, Rup04, Lee03, Lun03, Yan08b]. Eli [Buo06]. Eliciting [GKO05]. Ellen [AC05d]. Elliott [Chi06]. Elliptical [MGT05]. Elliptically [GEG05]. Embrechts [AC06b]. Emergency [RAHT06]. Emission [CCDH08]. Emmanuel [Las07a]. Empirical [CL00, CCPY09, DM06, ETST01, Efr09, Eis00, HQF08, JB01, JL06, JYZ09, MJvAC02, NC08, QZL09, SI02, SXZ07, XZ07, YL05, ZS02b]. Employment [KP07]. Emulation [HBK^{+}09]. Encouragement [TEJ^{+}04]. Encyclopedia [Cav06a, Cav07, Sta08]. End [Sol00]. Endogenous [RAHT06]. Endpoints [CLQ03, LH09]. Energy [GLW^{+}06]. Engineering [AC06c, Cli07, Gol02c, Loc04, McN02b, Car06, Che08]. Engineers [AC08c, Ann06a, Len03a, Wes07, Len03b]. Enn [Ken03a]. Enrique [Nem04]. Ensembles [LS02, STNM09]. Enterococcus [HCS06]. Entropy [XLB04]. Enumeration [ZSB01]. Environment [McB02]. Environmental [BRS^{+}00, Dix06, Gut00, Hoo07, Poo04, Poo06, Sai06, SBJRH05, Tho00, Wik06]. Environments [KT09]. Enzyme [DMW05]. Epidemiological [Cha04, Kri05]. Epidemiology [Duv03a, Hu04, JHISR07, SBJRH05, Tho00, Cav06b, Duv03b, KL05, Spi06, Dav06a]. Episodic [HS07b, YF08]. Equal [Zha04b]. Equality [CL08b]. Equating [Bou07]. Equation [Cor05, LW08, McN07, SBJRH05, Ste01, WA00]. Equations [Gan03a, Gan03b, LC01b, VWNM02, WC04, WJSS08, ZWW08, Wil04b]. Equipment [PS06c]. Equivalence [FS06b, Ric04, Ric08, RS09b, Mun04]. Ergodic [Kie06]. Ergodicity [MH04]. Eric [Lyl06, McN04, Stu06]. Erik [EP06]. Erkki [Not03a]. Ernst [Dav06c]. Error [Ald08, BCR02, BBS00, Bur04, CDH09a, CDH09b, CW01, CT04, CC09b, CCDP09, DLM04, DL08, Edw08, EBD^{+}04, Efr04b, FF09, FTW01, GW05, HS02a, Hal08, HJ02, JHISR07, KL00, LL05b, LL03b, LR04, LW08, LC00, MC06, RFF04, Sch09, Sen08a, SHY04a, SRL00, Spi06, SS04, SB05, SRB08, Sta09, Ste00a, Sti08, SW09, WF03, WC09, Wu02, Yat05, YZ05, Zab08b, Zab08a, Zha04a]. Errors [DM07, DH08, DFC09, FRZ05, HRC04, HW01b, KMR09, KK07, LL09, RL01, Ros05, SXZ07, XLCM03, Zuc05]. Errors-in-Covariables [SXZ07]. Errors-in-Covariates [RL01]. Errors-in-Variables [DM07, DH08, DFC09]. Essay [Bro00a]. Essentials [She02c, Har08]. Estimate [EL04, HBR01, RXS^{+}07, ZSG06]. Estimated [WF03, WZ05]. Estimates [BBS00, CR00b, CAC^{+}08, CFN01, Efr09, GW05, KST01, OOK03, WWP^{+}08, WS02]. Estimating [AD08, XMN06, DR00, DDZS02, DR05, DFWJ06, FIC01, FG04, GH08, Gel08, GB00, HS02a, HPC03, HH07, HL00, Hod08, IZ07, IC05, JRY04, JC07, JLZ08, KJ07, Kat08, KF07, KT09, KP05, LPS01, LC01b, LO06, LM09b, LSQ03, MSZ02, MGG^{+}07, MPK09, Rei05, SWSM09, Sch07e, SSY07, Str00f, TLW07b, TLW07a, TIY03, TBJ08, VWNM02, WC04, Wil04b, ZWW08, ZZ06, Few04]. Estimation [AMP06, ACK01, Alt01, AKS02, BM09, BCM02, Bur04, CFL00, CS01, CH03, CDK05, CFT06, CZI07, CW08, CZC09, CRW01, CGR04, Chr07, Číź08, CHL03, CLLCL09, CBL^{+}08, CvA02, DLRY06, DM07, DLM04, DNP05, DP05, Efr01, EBD^{+}04, Efr04b, EHT06, FL04, FHL07, FFJ07, FW08, FW00b, FW05, FH00, FHLR05, FTW01, FR02, Fre04, FS01, Fry05, Fry07, Gil08b, GC09b, GP01, GR07, GR05, GRYB06, GL07, Gua08, Gua09, GM06, HM03a, HRL04, HYFW05, HFZ05, Hey08, HY01b, HCB06, HTO03, HL05, HWZ06, HW04, HQF08, HLS05, ISZ07, IS07, Iri01a, Iri01b, JS05, JM06, JL06, JM04, KF02, Kar03a, Kar03b, KC01, KSN08, KNF06, KMR09, Kri05, KL06b, KL04b, KR02, KTZ00, LP06, LDVT09, LWRC04, LW08, LPW05, LC00, Lit07, Liu00, LJ00, LHN08, LR06, LZH09, LM00, MMC07]. Estimation [MOJ05, MH02, MJvAC02, MR05, MICR06, MVF06, NLLH06, NC08, NS02, NS01, OBMK07a, OBMK07b, PSH01, PF06, PWZZ09, PPR09, PT06, PS07, PS06c, QLS02, Qiu07, QvdLR06, QM01, RCGT01a, RCGT01b, RBP08, RFS06, RFP09, Rup07, SSV05, SHY04a, Sim02, SW02c, SK02, SL06, SD02, SRL00, SRB08, Str07, TW07b, VS08, VR00, WA00, WR02a, WR02b, WCL05, WL05, WL07a, WX09, WS06, Woo04a, XLCM03, XK09, YSY08, Yan04c, YPXH06, YR02, YJ04, YL05, YR00, Zeh01, ZLY05, ZL07a, Zha04a, Zha04b, dV04, vdLHR02, Dat04, Kar03a, EP06, Sto08]. Estimations [CX08, CX09]. Estimator [CCB03, DFC09, GP07, KW01c, SWP00, SHH00, XLL04]. Estimators [CCC09, FZ03, HO02a, HO02b, HC03a, HRV02, KW01c, Lin03b, LBS06, MGT05, MC06, MK06, MY02, Mer04, Por02, Pui03, QWP05, Rup02b, SBVW06, SJ04, VH02, XPLT09, YR00]. Eugene [Von06]. Eun [Cho06]. Eva [Ann05b]. Evaluate [CP02]. Evaluating [FBV^{+}04, HR06, HR03, PSW05, UCTW07]. Evaluation [AD08, BZ00, EL05, IPC08, KRS^{+}06, Qiu05, SjKH^{+}01]. Evariste [Yor06]. Event [And05, BFB00, CK04, CLLCL09, ERS02, GB00, HS07b, HZAG01, HW04, Kle06, LTMS02, MGG^{+}07, NLLH06, PSH01, Pic05, RTD01, SGL07, WQC01, WT01]. Event-Dependent [CLLCL09]. Event-Related [RTD01]. Events [BB01, CLLCL09, ELL09, HW03, Len05, MSZ02, SMIO08, ZL07b, ZYDS08].

Everitt [Daw03, Lun06]. Evidence [Cle02c, Eva00b, Han03, Kal00, KM01, LA06, Roy00a, Roy00b, Wol00, Won00a]. Evolution [Las07a, Kim06, Ros04, Kar06a]. Evolutionary [LW01b, SWS^{+}03]. Ewens [Sch03b, Zhe06]. Exacerbations [YF08]. Exact [Ano00g, CJS02, RW05, Ros05, Wu04, vdW02, Boo07]. Examining [WBBR05]. Example [AC04b, GH08, Gel08, Hod08, Kat08]. Example-Based [AC04b]. Examples [AC04e, Lun07b]. Exceedance [GW06]. Excel [Wie06]. Exceptional [Dav07b]. Excess [CP06]. Exchange [ABDL01, Ano03j, FBV^{+}04, HS07b]. Exercise [IBPL04]. Exercises [Emb03, Jam02g]. Exhibiting [Gil03]. Existence [CIS04]. Existing [SGL07]. expanded [Xie03b]. Expectancy [Che07a, IS07]. Expected [EL04, MRŞK06, MSZ02, SS09]. Experiment [ETST01, Hil08, LLL08a, SCS08a]. Experimental [AC03a, AC03d, AC05a, AC07b, AC08a, AC08b, Bha09, DS08, EW08, Ker06, Len02f, Len04a, LFGR04, vSH08, AC05a]. Experimenters [Spe06b]. Experiments [AC03c, AC06c, BFHR03a, BFHR03b, BX03, Dea08, DM06, DB09, Hil08, HI06, HM00a, HM00b, HM00c, HZZ^{+}08, HZ08, JDvS02, Ker06, KZ03, Len02g, LLL08a, Mor05, MB09, MJB03, Rag00a, Rag00b, Ros07, Rub08, SCS08a, SCS08b, SSL03, TH09, TRH07, AC05c, AC05b, Mor04, Rob04]. Expert [Las08]. Explained [Dob04]. Exploratory [SSVV09, Jer06, Kar06b]. Exploring [GM08, Ban06]. Exponential [Ann06a, Ano05j, BS06, DM04, DV03, FTR04, KE02, Miy04, OS00]. Exponentially [DHS08, GS07]. Exposure [BIW00, BIK00, CHB^{+}08, KSZL02, SMG^{+}05]. Expressed [IR03a, RWR03]. Expression [CCL^{+}08, DFS02, FL09, GC09a, Gol05b, HWZ06, ICG02, MPRR04, STPN09a, STPN09b, WIG^{+}04, vWvdWvdV08]. Expressions [WH07]. Extended [BF02, JR08, KE02]. Extending [BC01, KS00, NST07, Gal07]. Extension [BF09, LSC04]. Extensions [GGR07, Had05a, LA08]. Extraction [CCDP09]. Extramarital [CH02, HC02]. Extrapolation [FW00b, NS02]. Extreme [CNN07, EM08, ELL09, GP07, Kie08b]. Extreme-Value [EM08]. Extremes [BCS07]. Eye [Cho06, VPW08, Kra04]. Eye-Movement [VPW08].

F [Ann06d, Boo04, Bra06a, Dev06, Mar03a, Str03b]. Fabio [Cha03a]. Face [HY07, Sob06]. Facing [Wei00]. Fackler [Sch04a]. Factor [CCL^{+}08, Fol00, FHLR05, FKC09, HL07, HT04, JL08, PMHB09]. Factorial [AAK06, AC07a, GA05]. Factors [BC03a, HL07, HC01, Mar00b, RXS^{+}07, ST02]. Fahrmeir [Cha03b]. Failure [BAW08, Bet03, Bor03, CCW02, CW01, GL06, HL04, HW03, HW04, KL08, MB05, Par07, Pen00a, Pen00b, SNQ09, WW00a, WW00b, WW00c, ZL07a, ZCPL08]. Failure-Time [Pen00a, Pen00b, WW00a, WW00b, WW00c]. False [BYE^{+}05, BY05, BH07, Edw05, GW06, Hol05, PGVW04, PZB09, RM09, Sha05, SC07, Tam05, TBD^{+}03, Wes05, Yek08]. Families [ACS06, AALT00, CM01, DV03, Far01, KU01a, KU01b, Mar01, Ram01a]. Family [BM01, FTR04, JP05]. Fan [All06, Bre05]. Fang [Len04a]. Far [LL03a]. Farebrother [Wil04a]. Farrell [Lin03a]. Fast [Ano01j, CHL06, SBVW06, Ste01, THC05]. FDG [OMS^{+}09]. Features [EG09, SSVV09]. Fecundability [DZ00]. Felsovályi [All06]. Fertility [DC03, WD00]. Fetal [Alt01, GP01, QM01, RCGT01a, RCGT01b, Zeh01]. Few [BL07, CN08]. Feynman [deL05]. Feynman-Kac [deL05]. Fibrosis [YF08]. Fiducial [EHI08, HIP06]. Field [Bri04, GRG^{+}04, GRGB04, OC06, Rou04, ST04, Ste04, TN04]. Fields [Gre03, Hig09, JNKD03, KSZL02, LS09a, LS09b, Men09, PGVW04, TW07a, WM09b, Yan06]. Fill [DT00]. Filter [LS01, Mye07]. Filtering [CM06b, FEG07, XLB04, dGKS01]. Filters [LZCC05]. Finance [Del02b, Fra08, Kor04, Lo00, McN02b, Rhe07, Vie02b, Ziv06, Chi06, Lin04a, Sch04a, Cli07]. Financial [Bre06, Dia07, FZ03, FW07, Hin07, Kor05, McN04, Bin06a]. Finasteride [SRA08]. Finding [BT04, CS09, DBPP08, SJ03]. Fine [RRS07, Car06]. Finite [CK08, CK09, DM06, Gol05a, Han06, IJS01, JL06, KC07, LL07, Lit04, Mee00, SW02c, SS04, Ron05]. Finite-Difference [DM06]. Finite-Sample [Ron05]. Firing [FGM^{+}08]. First [Caf07, Mor03]. Fisher [VWM01]. Fission [Gre07a]. Fit [AP04, DMW05, Esc06, FH01, HZ03, HGH08, Kim00, MOJ05, PR02, Pou04, SY06]. Fitting [Che01, Gua06, KR08, PWM01, TEB00]. Fitzmaurice [Mol05]. Fixed [BL07, JHCN06, Neu06]. Fixed-Width [BL07, JHCN06]. Flexible [CBCS07, FGM^{+}08, KL08]. Flight [TBJ08]. Flowgraph [Kle06]. Flows [QA09]. Floyd [Nus06]. Focused [CHS^{+}03, HC06, Tsa03]. Fokianos [Ano04e, Cav04]. Follow [BIK00, CTG^{+}06, IP02, RSM^{+}07]. Follow-Up [BIK00, IP02, CTG^{+}06, RSM^{+}07]. Food [Tre06]. Football [Har03]. Forecast [GLPS03, PT07]. Forecasting [BM08, Bri04, BS06, CD05, CM09b, CS03b, CGRS01, FHLR05, GRG^{+}04, GRGB04, GLW^{+}06, HY07, IK08, Li05, PCK06, QA09, Rou04, SW02d, TN04, Tri02b, Tsa00, WBS07a, Lei05, Dav06b]. Forecasts [PM05, PLM08]. Forest [AMvW^{+}09, Chr07, Lit07, OBMK07a, OBMK07b, PM09, Rup07, ZRM^{+}08]. Forests [IBPL04, LJ06, PM09]. Foreword [Cas00b]. Form [Gua09]. Formulae [deL05]. Forster [Ank05]. Fortran [Iha06]. Forward [Ban06, Wan09]. Fossil [PM09]. Fotheringham [Zim03]. Foundational [RW00]. Foundations [Jen02b, Par02b, Ban05, Hin07]. Fourier [CHL06, ZZ06]. Fowler [Nus06]. Fractal [Nes07]. Fractal-Based [Nes07]. Fraction [ZYI06]. Fractional [CH03]. Frailty [FWP05, LL03b]. Frame [LR00, LR06]. Framework [HT06, JFY07, Kul06, Li05, Li07, LW08, LCT^{+}05, MOJ05, TH09]. Francesca [Per06]. Francis [AC05c, AC05b, Dan03, Hal06]. Francois [Dob04]. Francois-A [Dob04]. Frank [Rao03a, Wie06]. Franses [Lun05]. Frantisek [Fue03]. Free [CS03a, HT06, HM03b, IS07, LL04, Pre03, Ran00]. Free-Knot [HM03b]. Freedom [Zha03]. Frees [Lar06]. Freitas [Dia03a]. French [Sac03a]. Frequencies [Els02c, KM02, Rab02a, Rab02b, Sch02a, SL06]. Frequency [FW07, HTO03, HW03, HZW07, HZZ^{+}08, Kra04, ZMAS05]. Frequentist [BG01, CL03, HC03a, IR03b, Joh03, Kar07b, SD03]. Frequentists [Efr05b]. Frey [AC06b]. Friedman [Rup04]. Fries [HK02]. Frisk [GFK07]. Front [Ano00h, Ano00i, Ano00j, Ano00k, Ano01q, Ano01r, Ano01s, Ano01t, Ano02p, Ano02q, Ano02r, Ano02s, Ano03k, Ano03l, Ano03m, Ano03n, Ano04g, Ano04h, Ano04i, Ano04j, Ano05l, Ano05m, Ano05n, Ano05o, Ano06m, Ano06n, Ano06o, Ano06p]. Frontier [HS02a]. Fu [Gol05a, Yin04]. Full [Han04, MOJ05, SJ06]. Full-Information [MOJ05]. Fully [Ano05j, Miy04, MCG00, WA06]. Function [DMW05, FHL07, Fry07, HM03a, HH07, HY01b, IZ07, Jam03a, Jam03b, KJ07, Kon06, KF07, KR02, LC00, LJ00, OS00, TLW07b, TLW07a, YJ04, ZWT^{+}05, vdLHR02]. Function-Based [HH07, IZ07, KJ07, KF07, TLW07b, TLW07a]. Functional [BBK^{+}09, BD09, Bow07, CFY00, CP06, CC09b, CM01, CL08a, CM09b, CSD09, Far01, GK07, Gen00a, Gen00b, Gre07b, Hay00, Hig00, HCB06, HW01b, HSB09, JS03, JS05, KU01a, KU01b, KR08, LL03b, LM04, LPR09, MZ08, MG00, Mar01, MAC^{+}06b, MSY06, MY08b, Ram00, Ram01a, Raz00, RO07, SSVV09, SF07, SME03, TPO03, TH09, Wor00, YF05, YMW05, ZW08, AC07b, Bin06b]. Functional-Coefficient [CFY00]. Functionally [DHE08]. Functioning [HID09]. Functions [Ano01m, BM08, CCPY09, CG03b, CLLCL09, Dem04, Gne02, GWB08, HL00, JLZ08, Kol03b, Lav04, LGS07, Lin04c, LM09b, MVBC03b, MVBC03c, SBL^{+}04, SB04, SW02c, Ste05b, WH03b, Zad04]. Fundamentals [Dmi05, Sin06c]. Future [CL00, Hil09, LL03a, O’F01, Sis05, Tsa00, Xie00]. Fuzzy [Dem04, Lav04, Lin04c, SBL^{+}04, SB04, Zad04].

G [AC05c, AC05b, Bog04, Poo06, Spe06b, Ste06, Yan04a, LPM^{+}08]. Gait [CV07, KS07a, KS07b, Mar07]. Gait-Based [CV07, KS07a, KS07b, Mar07]. Galaxy [Ken03a, Ken03b]. Game [Del02b, Sch02d]. Gamerman [Jes03a]. Gamma [IJ04]. Gap [BBCL02]. Gaps [ZRM^{+}08]. GARCH [FZ09, KOC07, Lin03b]. Garmt [Roh05]. Garrett [Mol05]. Gary [Ste06]. Gauri [Rou06]. Gaussian [BH08, LMP05, Pou04, BB76, DP05, GL08, KMH05, PS06a, SL08, Yan06, Oun08]. Geert [Dur06, Wil06a]. Gender [And08]. Gene [CCL^{+}08, DFS02, FL09, GC09a, Gol05b, GI07, HHS06, HJJ09, ICG02, IR05, IC05, KPD02, LFY^{+}08, MPRR04, STPN09a, STPN09b, TH09]. Genealogical [deL05]. GeneChip [WH07]. General [GM09, HL07, IV04, Kul06, LL05b, Len04a, LCT^{+}05, MMC07, WL00, WT07]. Generalizations [McN02b, Sur04]. Generalized [AC05a, BRS03, BS06, CR01, Cha03b, CHJ08, CZC09, Chr07, CKN08, CSD09, DS08, FHLR05, Gal07, GEG05, GM06, HIP06, HFZ05, HM03b, ICLH05, KE02, Lan03, Lia05, LC01b, Lit07, MGT05, McC00, MPE08, ML02, MVF06, NK00, NL07, OBMK07a, OBMK07b, Pui03, PC00, RLZ09, Rup07, Sin04, SJ06, VWM01, VWNM02, Woo04a, XK09, XSC08, YZ09, AC07c, Aqu06, Cor05, Cow06, Del06, Lin04a, Wil04b]. Generalizing [IV04]. Generated [OC06, ZRM^{+}08]. Generating [Kon06]. Generation [Str03a, Cow07]. Genes [DHE08, IR03a, IC05, RWR03, RFP09, SWS^{+}03, ZIP05]. Genetic [AC05d, Car05a, CSCC06, IPBS00, KL05, Li06, LZS^{+}06, LZ06, Sab06, SAE06, TR06, WM09a]. Genetics [Els02c, KM02, Rab02a, Rab02b, Sch02a, Wei00, Gue08]. Genome [Hav07, TBD^{+}03]. Genomic [New02]. Genomics [CCL^{+}08]. Genotyping [SL08]. Gentle [Lee03]. Geoadditive [HBF06]. Geof [Jon06]. Geoffrey [Gol05b]. Geographically [Zim03]. Geometric [FHW00, KE02, Mao06, MH04]. Geometrical [SSVV09]. George [Gen05a, Spe06b]. Geostatistical [Bri04, GRG^{+}04, GRGB04, HC07, KW02, PS06a, Rou04, TN04]. Geostatistics [Zha04b, Li08a]. Gerhard [Cha03b]. Gestation [Alt01, GP01, QM01, RCGT01a, RCGT01b, Zeh01]. Get [HB09]. Get-Out-the-Vote [HB09]. Getting [Gew04]. Geweke [Ott06]. Ghosh [Epi04]. Gibbs [Gel00, IJ01, vDP08]. Giesbrecht [AC05c, AC05b]. Gill [Van05]. Giné [Yor06]. Ginevan [Poo04]. Giudici [Cle06]. Giulio [Woo04b]. Givens [Jon06]. Glidden [Lyl06]. Glim4 [Hal06]. Global [PS06b, SFPM04, YZ09]. Gold [AD08, CCL09, IPC08, Rub08, SCS08b]. Goldstein [Leh05]. Good [Ann05a]. Goodness [DMW05, Esc06, FH01, FSL^{+}06, HGH08, Kim00, MOJ05, Pou04]. Goodness-of-Fit [DMW05, Esc06, FH01, MOJ05, Pou04]. Goos [Rob04]. Gordon [Dia03a]. Gotway [Dig05, Zhu06]. Gould [AC03a]. Gower [Roh05]. Gradient [BG06b]. Graduate [Fil06]. Grammar [Bra06b]. Grant [Sch03b, Zhe06]. Graph [CB09]. Graphical [Bra06a, Car03b, ZWT^{+}05, Car03a]. Graphically [CC00]. Graphics [AC04b, Bra06b]. Graphs [Kar06c]. Gray [Rie06]. Green [Mir05]. Greg [Ann06c]. Gregory [Lun03, Sch03b, Zhe06]. Ground [ZRM^{+}08]. Group [BF01b, LA08, PKC01, STR08, XTSY01]. Group-Randomized [STR08]. Grouped [CS09]. Groups [Gil03, IYSK08]. Groves [Nus06]. Growth [Alt01, GP01, MW09, QM01, RCGT01a, RCGT01b, WBBR05, Wei08, Zeh01]. Guide [All06, AC04e, Gil08a, Len04a, Par06, Per06, Rog02a, Tay03a, Tay03b, Cla03, Dur06, Hu04]. Guided [FM09, LS01]. Gulati [Lin04b]. Gumpertz [AC05c, AC05b]. Gupta [Ruk03a]. Gustafson [Spi06]. Gut [Fil06]. Gyrfi [Pre03].

H [Dev06, Gho06, Jon06, Mol05, Rhe03a, Str03b, Yan04c]. Haar [Fry07]. Haining [Sey05]. Hall [Aqu06]. Han [Yin04]. Han-Fu [Yin04]. Hand [San03a, McC05]. Handbook [AC03a, AC03d, Bas03, Len03a, Len03b, Lin04d, Wik04, AC04c, Ann06a, Cav06b]. Handling [Poo06]. Handwriting [Ram00]. Hans [Dav06b, Lun05]. Hansen [Yan04c]. Haplotype [CSCC06, CCC09, Els02c, KM02, Li06, LZS^{+}06, LZ06, Rab02a, Rab02b, Sab06, SAE06, Sch02a, TR06, WM09a]. Haplotype-Based [CCC09, WM09a]. Happy [RY00]. Harbor [HCS06]. Hardeo [Boo06]. Hardin [Wil04b]. Härdle [Spe06a]. Harmonic [Iri01b]. Harnessing [Boc08]. Harrell [Rao03a]. Harro [Pre03]. Harvard [Yan06]. Harvey [Leh05]. Hastie [Rup04]. Hastings [CJ01]. Haupt [AC05d, AC05d]. Having [RHC07]. Haystacks [CS09]. Hazard [BMDD07, CFJZ07, CAC^{+}08, CM09b, HC06, PSW05, Sin06a, Sin06b]. Hazards [Ano01o, BBN^{+}04, CW00, HI01, IJ04, KL00, QWP05, YLZ08, ZYI05]. Head [HY01b]. Health [AMvW^{+}09, BIW00, Dig05, FD08, Gre00, Had05a, HID09, RZF06, SRC^{+}06, ZSCW06, Lin04d]. Heart [IBPL04]. Hector [KB04]. Heikki [San03a]. Held [Yan06]. Helge [AC03c]. Helio [Jes03a]. Helsel [Dix06]. Henry [Kaf03]. Herbert [Gho06]. Herbrich [Wil03a]. Hermite [Pui03]. Hesketh [Cor05, Daw03]. Hessian [LLC00]. Heterogeneity [BF09, BL06, CTG^{+}06, CBL^{+}08, Els02c, KM02, KPD02, MAC^{+}06b, Rab02a, Rab02b, RGA01, Sch02a]. Heterogeneous [HTZ^{+}09, LR04, XSC08]. Heteroscedastic [AV01, AP04, DM07, FW00b, SRB08, WL01]. Heteroscedasticity [Gil03, XK09]. Heuristic [GG01]. Hidden [Alt07, BME08, GR02, GLN08, Sco02, SJS05, YKL^{+}06, dGKS01]. Hierarchical [BBK^{+}09, CG03b, FV08, GMK^{+}04, GRF03, GDH09, HHS06, Hob00, Kol03b, LL03a, Li05, LMP05, MW09, MVBC03b, MVBC03c, NC02, RZF06, RGA01, SI02, SPJ01, TJBB06, TI08, WSP09, WMNB01, WH03b, Yek08, ZWT^{+}05, ZSCW06, Dav03]. High [CCL^{+}08, CVF06, Cra05, DZ03, FPHR05, FPH^{+}05, FW07, HTX09, HO02a, HO02b, HGWR08, HZ05, HRV02, JL09a, KCO08, KM05, LHNM08, LCT^{+}05, MY02, MMB09, Por02, Rup02b, Sab05, SL08, SGH07, SSV05, STJ07, TSV05, Tib05, Wan09, Zha05a, ZMAS05, ZMP06]. High-Breakdown [CVF06, HO02a, HO02b, HRV02, MY02, Por02, Rup02b]. High-Density [SL08]. High-Dimension [LHNM08, STJ07]. High-Dimensional [CCL^{+}08, Cra05, FPHR05, FPH^{+}05, HTX09, HGWR08, HZ05, KM05, MMB09, Sab05, SSV05, TSV05, Tib05, Zha05a, ZMP06]. High-Frequency [FW07, ZMAS05]. High-Resolution [LCT^{+}05, SGH07]. Higham [Kor05]. Higher [BST03, Rou06, Str00g]. Higher-Order [Str00g]. Highest [YL09]. Highly [Mir05]. Hilbe [Wil04b]. Hilbert [LLZ07]. Hinde [Hal06]. Hinge [WL07b]. Hinkelmann [Ker06]. Hisashi [Bai05]. Historical [PDK07]. History [Str03b, Teu03, Tsa00]. HIV [Ano02k, Ano02i, CW08, GBM^{+}01, HR00, RAHT06]. HIV-1 [HR00, CW08]. HIV-Infected [RAHT06]. HIV-Positive [Ano02i, GBM^{+}01]. Hjort [Mir05]. Hoek [Chi06]. Hoeting [Jon06]. Holmes [Yan04c]. Homogeneity [CS04]. Homogeneous [Lan05]. Hong [KT09]. Horowitz [Fri02b]. Hot [BF01a]. Houdré [Yor06]. Hougaard [Yan05a]. Housing [Sob06]. Howe [Lüt03]. Howland [Wie06]. Hox [Rob03a]. Hrishikesh [Mal05]. Huet [AC04e]. Huffcutt [Cle03a]. Hui [Yan04a]. Human [CV07, KS07a, KS07b, Mar07, SME03, WD00]. Humberto [Wie06]. Hunter [Spe06b, Spe06b]. Hunting [PZB09]. Huzurbazar [Kle06]. Hybrid [LLL08b]. Hyperbolic [CHJ08]. Hypotheses [BKV01, CS09, Els02c, KM02, Mun04, Rab02a, Rab02b, Sch02a, Ann05a, Nea06]. Hypothesis [Ano05k, BZ00, DN07, Efr04a, FHY07, Len04a, Mar00b, RW05, SHC02].

Ian [Ros06]. Ibrahim [Yan04a]. Identically [IT02, KM01]. Identifiability [LM02]. Identification [CDK05, Kar07a, LS02, PWM01, PP01, Rei05, SSV05, SS08, ZYZS03]. Identifying [Che07b, CC00, TDGW01, TL02, ZLI^{+}09]. Identity [Mas04]. Ignorability [NC02]. II [Hey08]. III [Koe02a]. Ill [MJvAC02]. Ill-Conditioned [MJvAC02]. Illegal [Cau01, Gro01, Hor01a, Hor01b, Mag01, RK01]. Ilya [Gen05b]. Image [Ano01j, DN03, Gem06, JB01, KJG05, Lee00, Qiu07, QS07, Wan06]. Images [MMZ02, OC06, QS07, ZLI^{+}09]. Imaging [CCDH08, Gen00a, Gen00b, Hay00, Hig00, Laz07, MG00, Raz00, Sch07d, SF07, SCG^{+}07, Wor00, ZZIP07a, ZZIP07b, ZCI^{+}09]. Imbedding [Gol05a]. Immune [HHS06]. Immunodeficiency [WT01]. Impact [FTR04, Gre03, WBBR05]. Impacts [LL03a]. Imperfect [GSCR04]. Implementation [CIK08, Che07b, Gan09, HH07, IZ07, KJ07, KF07, TLW07b, TLW07a]. Implementing [MK06, TRH07]. Importance [GM06, Lia02, Ned09, OZ00, ST02, SSY07, TEB00]. Improve [ELL00b, JB01, Len02f, WS01a]. Improved [DDH04, HCB06, Tsa03]. Improvement [PF08]. Improving [KST01, KL04b]. Imputation [CS01, GDS03, JvD03, Mee00, QSZ08, RR07, SRC^{+}06, SW02b, YZ05, YR00]. Imputed [SS00]. Imputing [KSS08]. Incidence [BB76, BH08, DHS08, PDK07]. Inclusion [RJ08]. Income [SRC^{+}06]. Incompatible [GY02]. Incomplete [CK04, FIC01, JvD03, Mor07, PPR09, PGLW00, RAHT06, SR01, WM09a, YC02]. Incomplete-Data [WM09a]. Inconsistency [HO02a, HO02b, HRV02, LA06, MY02, Por02, Rup02b]. Inconsistent [Zha04b]. Incorporating [BL06, FRS06, Mal05, NT05, PDK07]. Increasing [Str00f]. Increment [BHO07]. Increment-Averaged [BHO07]. Incumbency [GH08, Gel08, Hod08, Kat08]. Independence [CM09a, DW04, GK07, MB05, TOR05]. Independent [FIC01, Fel02c, Gri02b, IT02, KM01, LZCC05, LFGR04, Mer04, WWP^{+}08]. Index [Ano08f, Ano09j, PSW05, WT09, Xia08, YR02]. Indexes [HWZ06]. Indirect [GR03]. Individual [YTS08]. Individuals [WM09a]. Indoor [McB02]. Induced [BAW08, MWL^{+}01]. Industry [Eas03, AC04c, Cle06]. Inefficiencies [KT05]. Inequalities [Yor06]. Inequality [Mic06]. Infected [RSH^{+}02, RAHT06]. Infectious [FW05]. Infer [TH09]. Inference [And08, Ann06d, Ano06l, BNV09, BCS07, BP02, CR01, CS00a, CSCC06, CCC09, CJ06b, CM01, CL03, CVF06, Cor02b, Cox00, Daw00a, Daw00b, DV03, DHSR08, Far01, GR03, GCK00, GSCR04, Han06, HH07, HI06, HR06, HW03, HH08, IZ07, IV04, IR03b, Jes03a, Jes03b, Jia05, JR08, Joh03, KJ07, Kie06, KU01a, KU01b, KX04, KF07, KT05, LPL06, LSZH08, LL03b, Li06, LQ05, LZS^{+}06, LZ06, Lin04b, Lit04, LR00, LLL08b, Lou05, MRT05, Mao06, Mar01, Mic06, NL07, OS00, PM09, Pea00, Pen00a, Pen00b, PR00, Ram01a, RTD01, RG00, Ros05, RAHT06, Rub00, Rub05, Rup04, Rup05, Sab06, SAE06, SS00, STJ07, Sha00, SD03, SHY04b, SQC07, STR08, Sob00, Sob06, SFK^{+}05b, SFK05a, SRL00, SXZ07, Tan06a, Tan06b, TLW07b, TLW07a, TR06, VH02, VVTR08, WW00a, WW00b, WW00c, WKL09]. Inference [Was00, Yan05b, YL01, ZYI05, ZSCW06, ZWW08, Sch05, Ank05, AC05a, Dav06c, Had05b, Ban07, Dra06, Eub07, Har08]. Inferences [CFL00, EM05, FJ05, KMR09, Wu04]. Inferential [CHR00a, DM04]. Inferring [Bha09, SWDS03]. Infilling [JNKD03]. Inflation [IK08]. Influence [MRT05, MVF06, WQ07]. Influential [ZSB01]. Information [Ano00l, Ano01w, Ano02t, Ano03q, Ano04o, Ano05s, Ano06u, Boc08, CLQ03, CHS^{+}03, EL00, EHT06, FIC01, FRS06, HC06, Iri01a, MOJ05, NT05, NST07, RXS^{+}07, RGK01, SW02c, Soo00, SJ03, Tsa03, WS01b, Yan08b, YR00, Lee08, Yan05b]. Information-Theoretic [SJ03]. Informational [IYSK08]. Informative [PS07, QLS02, SPSZ05, SSL07, WQC01]. Ingster [Pou04]. Inheritance [Hös03, Hss03]. Inhibitor [GSCR04]. Inhomogeneity [Kul06]. Inhomogeneous [GL07, Gua08, GWB08, Gua09, MSD04, WBS07a]. Injury [SjKH^{+}01]. Innovation [Spe06b]. Input [ST02]. Insertion [XLB04]. Insertion/Deletion [XLB04]. Instrumental [Aba02, CRC^{+}04, Sma07, Tan06b]. Insua [Sac03a]. Insurance [Ano02k, Ano02i, FV08, GBM^{+}01, Lin04a, Rhe07, Sch06]. Integer [FW00b]. Integer-Valued [FW00b]. Integrals [Lin07]. Integrated [Har07, Jes03a, Jes03b, OS00, Yan08a, ZMAS05]. Integration [FFJ07, Tan04, YPXH06, Stu06, Rhe08]. Intelligence [JDvS02, Ram05]. Intensity [Gua08, GWB08, Gua09, HYFW05, IJ04, PSW05]. Intensive [Hal07, RS09b]. Interacting [ABPR06, deL05]. Interaction [FTR04, FG04]. Interactions [Gus00, Sir06, VVTR08]. Interactive [LLC00]. Interdomain [BME08]. Interest [Bow07, JRY04]. Interference [HH08, MMZ02, Ros07, Sob06]. Interlaboratory [IWM04]. Intermediate [BB01]. Internal [RM09, SSVV09, SWSM09]. Internet [CS00b]. Interobserver [Sch04b]. Interpolations [Zha04b]. Interpretability [CL03, IR03b, Joh03, SD03]. Interpretation [BGF01]. Interpreting [AC03e, Rog02a, AC03b]. Interpurchase [BB76, BH08]. Interrelated [TH09]. Interval [DFWJ06, KL08, LS09c, LCT^{+}05, NC08, Par07]. Interval-Censored [DFWJ06, KL08, LS09c, Par07]. Intervals [BYE^{+}05, BY05, CFN01, EHI08, Edw05, HIP06, Hol05, IWM04, KL06a, LSC04, Sha05, Tam05, Wes05, XSZ09, Len06]. Intervention [And08, Fol00, LLL08b, QA09, WBBR05]. Interview [SRC^{+}06]. Intraday [CS03b]. Intrinsic [CM09a, CL08b, JKKW03, ZCI^{+}09]. Intrinsically [NN09]. Introduction [AC07c, Bau02, Bec00b, BT08, Dev06, Kie08a, Kor04, Kor05, Lam00, Lan03, Mar05, Mil06, RTW00a, Ree08, Rie06, Sch03c, Sin06b, Bin06b, Sch03b, Sin07a, Woo04b, Zhe06, Ziv06, Eub07, Fre04, Gil08b, Hab03, Hav07, Jen06, Kie08b, Mac06a, Pic05, Rhe08, Rob08, Sto08, Wal07]. Introductory [Gol02c, Ban07, Rhe07, Wie06]. Invariance [CT04, Gol07, LM02]. Invariant [Jam02f, Ran00]. Inverse [BB76, BH08, BC01, CN05, Gem06, LS01, LMP05, MJvAC02, NT05, SCC^{+}06, Szk03, ZMP06]. Inverse-Gaussian [LMP05]. Investigating [BM01, CTG^{+}06]. Investigations [LA08, Gen05a]. Investigators [Dmi05, Sin06c]. Investments [Mal05]. Involved [HHS06]. Involving [ACS06, DM06, LLL08b]. Ion [dGKS01]. IOS [PB04]. Irina [Pou04]. Irish [GM08]. Irregular [ISZ07]. Irregularly [Fue07, WC04]. ISBN [McC02a]. Isochores [FV09]. Isotonic [CD07, Dun05]. Isotropy [GSC04]. Issues [LHSY02, Man07, Rya00, Van05, WD00]. Item [Hab03, LFY^{+}08, SI02]. Item-Clustering [SI02]. Iterative [LR01]. IV [LSQ03]. Ivo [Hab03].

J [AC06b, Ann06d, AC03f, Bet03, Chi06, Dan03, Dat04, Dou06, Epi04, Fie06, Gas03a, Gol05b, Jam03a, Kac04, Kar06c, Ken03a, Kor05, Lan03, Lee06, Lin04b, McC05, Mic06, Mir05, Mor04, Nus06, Ros06, Rya05, Sch03b, Sch04a, Spe06b, Yan05b, Zhe06]. Jaakko [Gen05b]. Jackknife [CS01, JYZ09, YR00]. Jae [Bal03]. James [Dav06b, Fre04, Gol05a, Lee03, Mol05, Nus06, Pic05, KW01c, Sho05]. Jay [AC03a]. Jean [Bre06]. Jean-Philippe [Bre06]. Jeff [Van05]. Jeffreys [CIK08]. Jen [Xie03a]. Jen-Pei [Xie03a]. Jennifer [Jon06, Lin04d]. Jensen [Ann05b]. Jerald [Boo04]. Jerome [Rup04]. Jerzy [Gan03a]. Jesper [Sch05]. Jiahui [McN04]. Jiang [Gen04, Nes05]. Jianhong [Kar06a]. Jianqing [Bre05]. Jim [Ros06]. Jin [Ann06a, Dui16]. Jin-Ting [Dui16]. Job [YL01, ZRM09]. Job-Training [ZRM09]. Joel [Fri02b]. Joglekar [AC04d]. John [AC04b, AC06a, And05, Bor03, Chi06, Dav06c, Hal06, Las06, Mar03a, Ott06, Roh05, Ros06, Yan04b]. Johnson [Kar06b]. Joinpoint [GBT09]. Joint [BD09, BBK03, Gew04, GT09, GA01, HBR01, HZAG01, HW04, HID09, KT05, LTMS02, LDM09, PWM01, PWZZ09, RBP08, SL06, SJ06, Wu02, YTS08, ZYDS08]. Jointly [WT01]. Jolivet [AC04e]. Jolliffe [Tim03]. Jonathan [Ank05, ORvSM01]. Jones [BL07, Mil06]. Jong [Kri05]. Joop [Rob03a]. Jörg [Ste04]. Jos [Hu04]. Joseph [Iha06, Nea06, Wil04b, Yan04a]. Joshi [Mat05]. Journal [BH08]. Journey [PH00]. Jr [Cle03a, Jam03a, Nus06]. Jr. [Kaf03, Rao03a]. Judea [Had05b]. Judgment [FOD07, WSLC06]. Judith [And05]. Jump [ASFP09, Bin06a, FW07, Qiu07, Wan06]. Justice [DK01].

Kac [deL05]. Kageyama [Bai04]. Kalbfleisch [Bor03]. Kalman [Mye07]. Kamel [AC06c]. Kanevski [Sai06]. Kapadia [Wil06b]. Kaplan [MKCX04]. Karen [Bra06a]. Karl [Joe03]. Kay [Ros06]. Kedem [Ano04e, Cav04]. Keenan [All06]. Keith [Ann06d]. Kempthorne [Ker06]. Kendall [Ank05, Ann06d, MB05, Ank05, Ann06d]. Kenneth [Bal05, Car05a, Jam03a, Ste05a, Ter04]. Kent [PWM01]. Kernel [CHL03, GG08, LLZ07, SJ04, WLC02, Wil03a, Wil03b, XWF05, dV04, Las06].

Kernel-Based [XWF05]. Kernels [Wil03a, Wil03c]. Keung [Gal05]. Keven [Ram05]. Kevin [Len04a]. Khattree [AC04c]. Khuri [Ber04]. Kidney [FWP05]. Kim [Gol05b]. Kim-Anh [Gol05b]. Kindergarten [HR06]. Kinetic [DMW05]. Kirti [Not03a]. Klaas [Hab03]. Klaus [Ker06, Len03b]. Kleiber [Bei04]. Klein [Yan04b]. Kloeden [Gan03a]. Knot [HM03b, ZS01]. Knottnerus [Tra05]. Know [ZSG06]. Knowledge [Ala07, Ban05, BD03, Mar05]. Known [Gua09]. Koch [Ste06]. Koenker [Jur06]. Kohler [Pre03]. Kong [KT09]. Konstantinos [Ano04e, Cav04]. Koopman [AC03f]. Korb [Ram05]. Korn [Gen05a]. Korolev [Lin04a]. Kotz [McN06, Sur04, Bei04]. Kraaikamp [Dev06]. Krause [Eas03]. Kruse [Car03a]. Krzyzak [Pre03]. Kullback [WDL04]. Kung [Kri05]. Kung-Jong [Kri05]. Kunio [Jen06]. Kutoyants [Kie06].

L [AC05c, AC05b, AC05d, Bor03, Car06, Dev06, Fie06, Fri02b, Gen05a, Iha06, Jer06, Nea06, Sch04a, Yan04b]. Labeling [HZ08, SGC05, YL09]. Lack [AP04, HZ03, PR02, SY06]. Lack-of-Fit [AP04, HZ03, SY06]. Lahiri [Car05b]. Laird [Mol05]. Lance [Dig05]. Landau [Lia05]. Lange [Car05a, Ste05a, Ter04]. Laplace [Str00g, Ano05j, Li08b, McN02b, Miy04]. Large [AA00, CHL06, Efr04a, Efr07, Efr09, Fue07, HDW07, JC07, KL06a, KSN08, KW01c, LSC04, LS02, LSQ03, MZ08, RLZ09, SW02d, SJ06, Teu03, WSP09]. Large-Sample [KL06a, LSC04, SJ06]. Large-Scale [Efr04a, Efr07, Efr09, JC07, LSQ03, MZ08]. LaRiccia [Kar03a]. Larose [Mar05]. Lasso [PC08, Zou06, WL07a, ZQ09]. Laszlo [Pre03]. Latent [AALT00, BF06, BF09, CN02, CA02, CBCS07, Cor05, Dun03, ES02, HRH02, HZAG01, JR07, LTMS02, LCLC08, LM00, MVF06, PDG02a, PDG02b, SI02, Sea02, Ver02, ZFZ03, McN07]. Latini [Poo06]. Lauder [Ros06]. Laundering [DJSW09]. Laurencelle [Dob04]. Laurent [Emb03]. Lavine [Mac06a]. Law [Eis00]. Lawal [Ste06]. Lawless [Boo04]. Lawson [Bog04]. Layer [Kad07, Mil07, RLL07a, RLL07b, Wol07]. Lead [BIK00]. Learn [CRY09]. Learning [DJSW09, EW08, LS06, Rup04, STZW03, Wil03b, Wil03c, Yan05b, Wil03a, Oun08, Bur08, Wil03a]. Least [KVZ07, KW01c, RO07, SHH00, VWM01, WL07a]. Lee [Rie06, Bal03, Gho06]. Left [Tre06]. Lehmann [Nea06]. Leibler [WDL04]. Leland [Bra06b]. Lelia [Lin03a]. Lemm [Ste04]. Lemmon [Iha06]. Lemuel [Dmi05, Sin06c, Wil06b]. Length [Ano02j, AMW02, BAW08, Bhl02, Büh02, CH06a, Hal02, HY01a, Lah02, Lee00, Pol02, SNQ09, She02d]. Length-Biased [Ano02j, AMW02, BAW08, SNQ09]. Leonhard [Yan06]. Lepkowski [Nus06]. Letter [Ano03h, Ano03o, Bak07, HK02]. Leukemia [Gre03, HSG02]. Level [HT04, Kje00, MMC07, MPE08, OZ08, PKC01, RXS^{+}07, RGK01]. Level-Based [Kje00]. Levels [AA00, CHB^{+}08, CNN07, KTZ00, Mei07, MWL^{+}01, SWSM09, ZWT^{+}05]. Levenbach [Dav06b]. Lévy [CCPY09, Nov05]. Li [Gal05]. Licensure [ERS02]. Lid [Len04a, Mir05]. Lieven [Lue05]. Life [AC08c, Bre00, IS07, KR02, SZ09]. Lifetables [CM09b]. Lifetime [Ano02m, Boo04, BF01a, Hua02, PF06]. Lifetimes [Cha03a]. Lift [Joe03, OOK03]. Like [CHL03, LM09a, MJvAC02]. Likelihood [Ano01o, BC03a, BB01, BR00, BM09, CCPY09, CSCC06, CR00b, Che01, CZI07, CP07, CZC09, CJ01, CF09, DV03, FHW00, FW00a, FL01, Fue07, Gua06, HI01, HQF08, Iri01a, Jia05, JYZ09, Kad07, Kar03a, Kar03b, KSN08, KE02, Li00, Li06, LQ05, LZS^{+}06, LZ06, Lou05, MRW07, Mil07, MJvAC02, MV00, MvdV00, NC08, NN06, OS00, Pui03, QZL09, RLL07a, RLL07b, Rei00a, RW00, RRvdL00, Rup05, Sab06, SAE06, She00b, SFK^{+}05b, SFK05a, SXZ07, Tan04, TR06, VR00, Wal02, WL05, WM09a, War08, WZ05, Wol07, XLL04, XZ07, YJ04, ZLY05, ZWL^{+}04, ZRM09, ZS02b, Zuc05, Gue08]. Likelihood-Based [Ano01o, CSCC06, CF09, DV03, HI01, HQF08, KSN08, Li06, LZS^{+}06, LZ06, Sab06, SAE06, TR06, YJ04, ZRM09]. Likelihood-Like [MJvAC02]. Likelihoods [Ric08, VCR09, dV04]. Likert [JR07]. Lim [Ann06a]. Limit [Fel02c, Kar06a]. Limitation [DM00]. Limited [Bal03, MOJ05]. Limited- [MOJ05]. Limits [CH06b, Gus09]. Linden [Dou06]. Lindsey [Gas03a]. Linear [Ann06d, Ann08, CFJZ07, CR01, Cha03b, CJ06a, CW01, CT04, Chr00b, CVF06, CKN08, Cow06, Dav03, DS08, EHI08, Esc06, Eub07, FW00b, FRS06, Fie00, FG00, Gal07, GR06a, GW05, Hal03a, Hal03b, HFZ05, HRC04, HJJ09, ICLH05, IT02, JM06, KVZ07, KMR09, KPY03, Lan03, Lan05, LSC04, LWRC04, LL09, LJZ04, LH07, Lyl06, McC00, MPE08, ML02, Mor03, MVF06, MH07, MW05, NK00, NL07, PS06b, Rao03a, Rao03b, Sen08b, She00a, Sin04, SS08, SB05, SJ06, SY06, VWNM02, WL00, WS05, XSC08, YLZ08, YZL08, YR02, YJ04, YY05, YL05, ZYZS03, Dou06, Gil08a, Yan08a]. Linearity [HS02b]. Linearization [VWM01]. Lines [Ano00g]. Link [CDVY00]. Linkage [BL06, LR01, LCT^{+}05, Rei08, SSY07]. Linked [LL05a]. Linker [BME08]. Linking [Bou07]. Lisa [Gen05a]. Liski [Not03a]. List [SWSM09]. Litigation [Las08]. Liu [Xie03a]. Liver [SWSM09]. Lives [AC08c]. Loa [CDR08a, CDR08b, Dun08, Pac08, Tho08, CDR08a, CDR08b, Dun08, Pac08, Tho08]. Load [CS03b, KP05]. Load-Sharing [KP05]. Local [BB01, CW08, CB09, DFC09, EG09, GG07, Iri01a, Iri01b, LLW00, QS07, RSW09, SFPM04, WKL09, WZ02, XLCM03, YJ04]. Locality [WLC02]. Locally [CP06, MGT05, MC06, QvdLR06, RL01, WW09, vdLHR02]. Locating [Ano02o]. Location [CDK05, Epp04, GG07, HP04, HRV04, MRT05, McC04, MMH^{+}04, MM04, Ser04]. Location-Dependent [GG07]. Location-Scale [CDK05, Epp04, HP04, HRV04, McC04, MMH^{+}04, MM04, Ser04]. Loci [JFY07]. Log [BFB00, Chr00b, Fie00, HJJ09, SS08]. Log-Linear [Chr00b, Fie00, HJJ09, SS08]. Log-Rank [BFB00]. Logistic [Ano00f, HW01b, LM09a, Lyl06, MPS00, Rao03a, Rao03b, SRL00]. Logit [BF09, Win03]. Long [CHL06, HLS05, JKKW03, Lin03b, Rus05, TBJ08]. Long-Memory [CHL06, Lin03b]. Long-Period [JKKW03]. Long-Term [TBJ08]. Longest [CH06a, Kon06]. Longevity [KG01a]. Longford [PS06c]. Longitudinal [And05, AALT00, BF09, CJ06b, Cor05, Cra07, DM05, Dun03, FL04, FHL07, FW08, FBV^{+}04, GB00, GRT04, Hal07, HW01a, Hu04, IEMM08, KRS^{+}06, LZ08, LY01a, LC01a, LY01b, LTMS02, LTE08, LDM09, MGG^{+}07, MTR^{+}01, Mol05, Neu06, OS01, PAS^{+}03, PPR09, PGLW00, RL02, RSM^{+}07, SJS05, SPJ01, SK02, SPSZ05, SSL07, WT01, WC01, WW01, WCL05, WLC02, Wil06a, WZ02, XZ07, YMW05, YL01, YC02, YTS08, ZC07, Zha06, ZW08, vdLHR02, Car07, Cla03, Lar06]. Longitudinal/Clustered [WCL05]. Look [Ber00b]. Lopuhaä [Dev06]. Loss [XMN06, BY03, HZW07, JM04, PT07, WL07b]. Losses [SCC^{+}06]. Lou [Gol05a]. Louis [Dob04]. Low [LHNM08, STJ07]. Low-Sample [LHNM08, STJ07]. Lower [CH06b, Mao06, NR02]. Luck [CR08]. Ludwig [Cha03b]. Lui [Kri05]. Lumelskii [Sur04]. Lundström [EP06]. Lymphatic [RSH^{+}02].

M [AC04e, AC04d, Bas03, Dev06, Gen05a, Imh03a, Kor04, Lee06, Loc04, Mol05, Nus06, Peñ06, Rhe03a, Rie06, Sch06, Sch04b, Stu06, Wie06, Wil04b, LP06, Men00]. M.-A [AC04e]. Machine [Bur08, Oun08]. Machines [LLW04, MD01, WS07, Wil03a, Wil03c, WL07b]. MacKay [Yan05b]. Macroeconomic [GLPS03]. Magnetic [Gen00a, Gen00b, Gre03, Hay00, Hig00, Laz07, MG00, Raz00, Sch07d, SF07, Wor00, ZZIP07a, ZZIP07b, ZLI^{+}09]. Magnetometry [AHA01]. Magnetoresistance [HY01b]. Maignan [Sai06]. Main [SRL00]. Maindonald [AC04b]. Maintenance [MSM05]. Maize [LC09]. Making [Woo07]. Malaria [MPE08]. Malawi [BM01]. Mallick [Yan04c]. Mammographic [CAC^{+}08]. Mammography [BP01]. Managed [ZSCW06]. Management [AC06b, CHJ08, Man07, Rhe03a, Rhe03b, Bre06, Dav06b, Lüt03, Dia07]. Mandal [Not03a]. Manifest [Cho06]. Manifolds [CV07, KS07a, KS07b, Mar07]. Mannila [San03a]. MANOVA [Gho07]. Manufacturing [NHS00, AC04d]. Many [FHY07, ZSG06]. MAPLE [Gan03b, Gan03a]. Mapping [GR02, KSZL02, LCT^{+}05, NSP00, PM09, TW07a, ZW08]. Mappings [CHL03]. Marc [Bre06]. Marchette [Kar06c]. Marcia [AC05c, AC05b]. Marco [Ban06, Per06]. Margaret [Qiu05]. Margin [WSP09]. Marginal [BFD01, BF02, BF06, BC03a, CJ01, HH02, HBR01, IJS01, MTR^{+}01, MvdLRC01, WCL05, WLC02, YPXH06, YC02]. Margins [BFD01]. Mari [Had05a]. Marianna [Sur04]. Mario [Boo06, Sch04a]. Mark [Dav06a, Mat05, Yan04c]. Marked [CM06b, HL06b, IJ04, McB02, Sch07c]. Marker [NLLH06]. Markers [DC03, IEMM08, SPJ01]. Market [Mal05]. Marketing [RA00, Ann06c]. Markets [Cau01, Gro01, Hor01a, Hor01b, Mag01, RK01]. Markov [BL07, Emb03, Gol05a, Yan06, Alt07, AALT00, BME08, BF09, Bhl02, Büh02, CR00a, FS01, Fry05, GR02, GM06, GLN08, Hal02, HC01, HH02, HH07, IZ07, JHCN06, KJ07, KF07, Lah02, LPD00, Liu03, MSD04, Pol02, Ros03, ST04, Sco02, SJS05, She02d, Sis05, TLW07b, TLW07a, YKL^{+}06, dGKS01, Teu08]. Marlene [Spe06a]. Martin [Nol03a, Zim03]. Martinez [Jer06, Wik04, Ken03a, Jer06, Wik04]. Martingales [Emb03, Lin07]. Masonry [FIC01]. Mass [CP06]. Massive [THC05]. Matched [Ric04, Ric08, Ros02, RRS07, RS09a, SMG^{+}05]. Matching [Han04, Kar07b, LPR01, LZHR01, Rou06, RT00, Sch07a, TBD^{+}03, WS02]. Mate [LHN08]. Material [HT04]. Mathematica [McC02a]. Mathematical [Hin07, Kor04, Loc02b, ST02, Car05a, McC02a, Wil06b]. Mathematics [Kor05, Loc02b]. MATLAB [Jer06, Wik04]. Matrices [RLZ09, War08, ZCI^{+}09]. Matrix [CHL06, DXC08, Imh03b, Jam02g, KC01, KL06b, OOK03, Ruk03b, SK02, Imh03a, Ruk03a]. Matrixes [FW08]. Matter [Ano00a, Ano00b, Ano00c, Ano00d, Ano00h, Ano00i, Ano00j, Ano00k, Ano01a, Ano01b, Ano01c, Ano01d, Ano01q, Ano01r, Ano01s, Ano01t, Ano02a, Ano02b, Ano02c, Ano02d, Ano02p, Ano02q, Ano02r, Ano02s, Ano03a, Ano03b, Ano03c, Ano03d, Ano03k, Ano03l, Ano03m, Ano03n, Ano04a, Ano04b, Ano04c, Ano04d, Ano04g, Ano04h, Ano04i, Ano04j, Ano05b, Ano05c, Ano05d, Ano05e, Ano05l, Ano05m, Ano05n, Ano05o, Ano06b, Ano06c, Ano06d, Ano06e, Ano06m, Ano06n, Ano06o, Ano06p, CHB^{+}08, LL03a, WWP^{+}08]. Maura [Ste06]. Maximization [Jia05, LQ05, Lou05, Rup05, SFK^{+}05b, SFK05a]. Maximized [MPK09]. Maximizing [ST04]. Maximum [BM09, CR00b, CZI07, Che07b, Kar03b, KE02, Pui03, TDGW01, TL02, VR00, Wal02, WL05, XLL04, ZLY05, Kar03a]. Mazliak [Emb03]. McClure [Dav06c]. McCulloch [Cow06, Ann06c, Lyl06]. McDonald [Van05]. McLachlan [Gol05b]. MCMC [Gue08, CTG^{+}06, CRY09, PGH01, Fra08]. McNeil [AC06b]. McShane [Gen05a]. Mean [Ald08, CLLCL09, DL08, Edw08, FG04, HT06, Hig09, KR02, LPL06, LS09a, LS09b, LO06, Men09, MvdLRC01, Sen08a, Sti08, Str00f, SZ09, WF03, WM09b, Yat05, Zab08b, Zab08a, ZZ06]. Means [Bra06a, Dra06, EL04, FS04a, JL06, LNR03, SS00]. Measure [FGM^{+}08, IYSK08, OC06, Tay03a, Tay03b, Kie07a]. Measure-Theoretic [Tay03a, Tay03b]. Measured [BIW00, BNV09, CC09b, CRW01, LC00, Ros05, Wu02]. Measurement [BCR02, BBS00, BST03, CDH09a, CDH09b, CW01, CT04, CCDP09, EL00, FF09, JHISR07, KL00, KT05, LL03b, LR04, LW08, LL09, MC06, Sch09, SRL00, SB05, SRB08, Sta09, Ste00a, SW09, WC09, YZ05, Poo04, Hal08, McC05, Spi06]. Measurements [LC09, Moe03b, WLH08, Moe03a]. Measures [DMRM04, Dem04, ERS02, HZAG01, Lav04, Lin04c, Lyl06, PAS^{+}03, SBL^{+}04, SB04, Stu06, WC04, Zad04, ZYDS08, Lin07, Sch04b]. Medfly [KG01a]. Medhi [Rya05]. Media [Kar06a, LZHR01]. Median [FEG07, HTX09, HS02b, KG01b]. Median-Based [HTX09]. Medical [Ano02m, Bre00, Daw03, Gas03a, Gas03b, Hua02, Len04a, Qiu05, Woo07]. Medicine [Ros06]. Meester [Dev06]. Meetings [O’F01]. Meier [MKCX04].

Melding [PR00]. Melendres [Lüt03]. Melvin [Yan04b]. Membership [Dem04, Lav04, Lin04c, SBL^{+}04, SB04, Zad04]. Memory [CHL06, HLS05, Lin03b, ORvSM01]. Meng [MH04]. Menopause [NLLH06]. Menten [DB03, DMW05]. Mervyn [Mic06]. Mesoscale [Bri04, GRG^{+}04, GRGB04, Rou04, TN04]. Messy [Koe02a]. Meta [BD05, CCL09, Cle03a, Cle03b, DT00]. Meta-Analysis [BD05, CCL09, DT00]. Metanalysis [BBN^{+}04]. Metastatic [CLLCL09]. Meteorological [Hin07]. Meters [MD01]. Methadone [RL02]. Method [Aqu06, Bal03, Bra06a, Bri04, Che01, Che02, Del06, DT00, FGM^{+}08, GRG^{+}04, GRGB04, GLW^{+}06, KST01, LWZ08, LLW00, RWR03, Rou04, TN04, THC05, WZ05, YY09, Zuc05]. Method-of-Moments [Bal03]. Methodology [Bec00a, FBV^{+}04, Pep00, Spa02b, Yek08, Nus06]. Methods [AV01, ADW07, AC03a, AC03d, AC04d, Ano00f, Ano01o, AHA01, Bai05, BP01, Bal05, Boo04, Boo06, BC03b, Bou07, Bra08, Bri07, Car03a, Car03b, Car05a, Car05b, CCPY09, CDHL05, Cle06, Cra07, Dav03, Dia03a, Dia03b, DHS08, DFS02, Duv03a, Duv03b, Eis00, EL04, FFJ07, FM09, FH00, FM06, FD08, GKO05, Gue08, Had05a, HC01, HI01, HMS00, HW01b, ICLH05, IJ01, IJ04, Jam03a, Jam03b, Jen02b, KL04a, Kie05, Kim06, KS00, KL05, Kra04, Las06, Len06, LHSY02, LZH09, Lyl06, Man07, Mar03a, Mar03b, MPS00, Moe03a, Moe03b, MWL^{+}01, NK00, Nem04, Neu06, Pal08, Peñ06, RW05, RAHT06, Rus05, Sch03b, Sch03c, Sco02, SMIO08, Str00g, Tan06b, Tyl07, WLC02, Wik06, YW03, YC02, Zhe06, Zhu06, ZZ06, Ann08, AC03f, Bre05, Cho06, Lin04d]. Methods [Ree08, Dea08, Tyl08]. Methylation [LFY^{+}08]. Metropolis [CJ01, LLW00]. Micah [Van05]. Michael [Buo06, Gen04, Gen05a, Mac06a, Poo04, Pre03, Van05]. Michaelis [DB03, DMW05]. Michel [Sai06]. Michele [Bas03]. Mick [Nus06]. Microarray [Cra05, DB09, ETST01, FPHR05, FPH^{+}05, Gen05a, Gol05b, HZ05, HWZ05, ICG02, IR05, KM05, LLW04, LC09, QS07, Sab05, TH09, Tib05, TW07b, WH07, YKL^{+}06, Zha05a, Gen05b]. Microarrays [Dav06c, GRYB06, IR03a, MPRR04, RWR03]. Microchips [AC01a]. Microdata [Rei05, SS08]. Microscopy [HY01b]. Microsimulation [RMS09]. Microsoft [Wie06]. Mielke [Jam03a]. Migon [Jes03a]. Miguel [Boo06]. Mikhail [Sai06]. Millard [Eas03]. Millennium [PH00]. Miller [Lun03]. Millions [LPS01]. Min [Nes05]. Ming [Yan04a]. Ming-Hui [Yan04a]. Minimal [KL06a]. Minimax [FW00b]. Minimaxity [Str00k]. Minimum [CN05, HY01a, Lee00, LH07, PY00, RRS07, TL02]. Mining [But06, Car03b, Cle06, FS04b, Kar06b, Rup04, San03b, Car03a, Mar05, San03a, Las07b]. Miranda [Sch04a]. Mirkin [But06]. Misaligned [MCG00]. Misclassification [SRL00, Spi06]. Mises [LDVT09]. Misleading [Eva00b, Kal00, Roy00a, Roy00b, Wol00, Won00a]. Mismeasured [HL04, MY08a]. Missense [ZIP05]. Missing [Ano01o, Bak00, BIK00, Che02, Che04, CIS04, CZI07, CAC^{+}08, DR00, HI01, HM00a, HM00b, HM00c, ICLH05, IZT08, KRS^{+}06, LL08, LWRC04, MK06, Men00, Poo06, QWP05, QSZ08, QZL09, Rag00a, Rag00b, RL02, SRC^{+}06, SMG^{+}05, WLH04, Wu04, XPLT09, YL01, ZWW08, PS06c]. Missing-Data [ICLH05, IZT08]. Missingness [AAK06]. Misspecification [PB04]. Misspecified [Els02c, KM02, Rab02a, Rab02b, Sch02a]. Mitigate [LL03a]. Mitochondrial [HBK^{+}09]. Mitzenmacher [Buo06]. Mixed [Alt07, Ano02n, BLW01, BHO07, CCW02, CVF06, Cow06, Cra07, EHI08, FKW01, Gal07, GS01, Gil08a, Hof05, IT02, Kad07, KW01a, KW01b, KS02, KS08, LZ01, LA06, LCLC08, Mil07, MAC^{+}06b, NK00, Ram01b, RLL07a, RLL07b, Sin04, SME03, SJ06, VWNM02, Wol07, Wu02, WZ02, Wu04, Von06]. Mixed-Effects [Ano02n, BLW01, CCW02, FKW01, GS01, Hof05, KW01a, KW01b, LZ01, LCLC08, Ram01b, VWNM02, Wu02, WZ02, Wu04, Cra07]. Mixed-Layer [Kad07, Mil07, RLL07a, RLL07b, Wol07]. Mixed-Model [SME03]. Mixing [GKM05, Wal02, Yan01]. Mixture [BC03a, BM09, CHR00a, CS04, CK08, CK09, FS01, Fry05, GCK00, GRT04, GI07, HJ02, HL06b, HZ08, KC07, LR01, LW01b, LMP05, NST07, OLT09, Ric04, RSW09, SL08, SPJ01, TIY03, WBBR05, WL01, WS06, YL09]. Mixtures [AC08d, CC00, Dui16, Han06, IJS01, LPM^{+}08, LPW05, PWM01, Zha05b]. MLE [CIS04]. MM [SBVW06]. Mobility [Sob06]. Möbius [CHL03]. Möbius-Like [CHL03]. Model [Ano01o, Ano06l, AALT00, BME08, BB76, BF09, BH08, BG01, BBK03, BM09, BMDD07, BF01a, BD05, CJS02, Cha04, CR00b, Che01, CDK05, CH06b, CT04, Chr07, CL03, CS03a, CBL^{+}08, CKN08, Cra05, DMRM04, DB03, DKS06, DP05, DZ00, EHI08, FZ03, FL04, FPHR05, FPH^{+}05, FL09, FM09, FWP05, FB00, FHLR05, FTW01, FS04b, FR02, FKC09, Gal07, Gen00a, Gen00b, GT09, GC09a, GA01, GRT04, GDH09, GL03, HL07, HY01a, Hay00, HBK^{+}09, HI01, Hig00, HGWR08, Hig09, HC03a, HC03b, HC06, Hof07, HL05, HRC04, Hua00, HZAG01, HOS04, HZ05, HWZ05, HC07, HZ08, IZT08, Iri01a, IJS01, IR03a, IR03b, IPC08, JS05, JRY04, JR07, JL06, Joh03, JKN08, KL06a, KL04a, KVZ07, KST01, KM01, KL08, KM05, KS02, LS09a, LS09b, Len04a, LL07, LFY^{+}08, LTE08, Lit04]. Model [Lit07, LS09c, LCLC08, MY08a, MG00, MH04, McB02, McK05, Men09, MW09, MTR^{+}01, MR05, MICR06, Mor03, MCG00, MRID01, MW05, NL07, NDKW01, OS01, OBMK07a, OBMK07b, Pen00a, Pen00b, PS06b, PSW05, PB04, RZ03, RD06, Raz00, RTD01, Ric04, RBP08, RZF06, Rup07, Sab05, SG00, STPN09a, STPN09b, SI02, SY02, SD03, SHY04b, SH06, SME03, SWDS03, Sur04, TZW05, Tib05, WW00a, WW00b, WW00c, WA06, WWP^{+}08, WX09, WM09b, WS05, WL01, Wor00, WS01b, Wu02, WIG^{+}04, Xia08, XWF05, XZ07, YY09, YTS08, ZLY05, ZL07a, ZFZ03, Zha04b, ZWL^{+}04, Zha05a, ZSCW06, Ann06d, Li08a]. Model-Assisted [Chr07, JL06, LL07, Lit07, OBMK07a, OBMK07b, Rup07]. Model-Based [CJS02, FR02, LL07, MCG00, RD06, WIG^{+}04, Zha04b, Li08a]. Model-Calibration [WS01b]. Model-Free [CS03a]. Model-Guided [FM09]. Model-Independent [WWP^{+}08]. Modeling [And05, AMvW^{+}09, BFD01, BF02, BNV09, BM08, CCL^{+}08, CM09b, CD09, CG03b, CNN07, CBCS07, Cor05, CS03b, Cra07, CDR08a, CDR08b, DMJ^{+}08, DR05, DC03, Dun08, DX09, Fan00, FL04, FV08, GLPS03, GKSB03, GKM05, Gel04, GL08, GA01, GLN08, GI07, Gus00, Hal03a, HJ02, HL06b, HSG02, Hil09, HM03b, HW03, HW04, HID09, HZZ^{+}08, IEMM08, KW02, Kol03b, KG01b, Li05, LMP05, MFCF05, MVBC03b, MVBC03c, NR02, NL07, Nta07, Pac08, PS06a, PGH01, Rao03a, Rao03b, Rub05, SGH07, STNM09, Ste01, SFPM04, Tho08, Tre06, WM01, WT01, WJSS08, WMNB01, WH03b, ZYDS08, ZRM^{+}08, Bog04, Hal03b, Gem06, Lee08, McN04, Wal07, Woo07, Zha06]. Modelling [Bin06a, Cha03b, Cow03a, Hal06, PMHB09, Sai06, Cow03b, Mül05]. Models [Aba02, ACK01, AV01, Alt07, AC07c, Ann06b, Ano02n, Ano04e, BGS03, BG06b, BF02, BF06, BC03a, BLW01, BB00a, BB00b, BBK^{+}09, BRS03, BDZ06, BD09, Boo04, Boo00, Boo06, Bow07, BHO07, BB03, BRS^{+}00, BKV01, CFL00, CFY00, CCW02, CX08, CX09, CR01, Car03a, Car03b, CDH09a, CDH09b, Cav04, Cha03a, Cha03b, CS04, CW00, Che02, Che04, CIS04, CFT06, CJ06a, CZI07, CW08, CK08, CIK08, CK09, CW01, CP07, CC09b, CZC09, CRW01, CJ06b, CE09, Chi06, Chr00b, Chr07, CCL09, Číź08, CVF06, CKN08, CHRH01, Dan03, Dav03, DLRY06, Dav07b, DDH04, Dou06, DS08, Dun03, Esc06, Eub07, Eva00a, FH01, FJ05, FF09, FW00b, FTR04, FW05, Fie00, FS06b, FG00, FRZ05, FKW01, FS01, GM09, GS01, Gas03a, Gas03b, Gen06, GCK00, GBT09, Gil08a, GC09b]. Models [GL08, GRF03, GR02, Gua06, Gus09, Had05b, Hal08, Han06, HL06b, HFZ05, HL00, HBF06, HBR01, Hob00, Hof05, HS02b, HJJ09, HGH08, ICG02, ICLH05, IT02, IJ04, IWM04, JM06, JLZ08, KW01a, KW01b, KW04, KC07, Kie05, KB04, Kle06, KMR09, KR08, Kol02b, KOC07, Kra04, KT05, Lan03, Lan05, LR01, LB01, LSZH08, LL03b, LL06c, LW01b, LWRC04, LW08, LL09, LZ01, LWY01, LTMS02, LHSY02, Lin04a, Lin03b, Lit07, LJZ04, LDM09, LM00, MC06, MMC07, MRT05, Mar00a, MB02, McC00, McN07, MPE08, ML02, MAC^{+}06b, MVF06, MY08b, MvdLRC01, NST07, NC02, NK00, NC08, NN09, OBMK07a, OBMK07b, OS00, OLT09, PLM08, PS05, PF06, PH08, PWZZ09, PT06, Pic00, PW05, Pol00, PR00, Pui03, Ram01b, Rao03a, Rao03b, RL01, RvdVV00a, RvdVV00b, Ros04, Rue04, Rup07, RMS09, Rya05]. Models [SL08, ST02, SSV05, SBJRH05, Sch09, Sch03d, Sco02, SJS05, SR01, Sen08b, SNQ09, Sho05, Sin04, Sin07b, SPJ01, SS08, STNM09, SB05, Sta09, Ste00a, Ste00c, SMP03, SXZ07, SJ06, SY06, SW09, SZ09, TEJ^{+}04, Teu08, TEB00, TIY03, UCTW07, VS08, VWNM02, Von06, WA00, WC04, WBBR05, WQ07, WLH08, WKL09, WZ05, Wil04a, Woo04a, Wu02, WZ02, Wu04, XK09, XSC08, XLL04, YSY08, YF05, Yan08a, YPXH06, YR02, YY05, YL05, YKL^{+}06, ZYI05, ZYI06, ZL07b, ZC07, ZYZS03, ZWT^{+}05, ZCI^{+}09, ZLI^{+}09, dGKS01, Leh05, Cor05, Cow06, Fue03, Lun05, Lyl06, Nes05, Per06, Pou04, Spe06a, Win03, Fri07, Gal07, Hal07, Par06, Rob08, Vla08, Wil06a]. Modern [AC07a, Dev06, Efr00, Hal08, Len04a, PS06c, AC08a]. Modes [Ano05j, Lit04, Miy04]. Modified [LSC04]. Moeschberger [Yan04b]. Mohamed [Sch04b]. Molecular [Gen04, Kim06, Won00b]. Molenaar [Hab03]. Molenberghs [Dur06, Wil06a]. Møller [Sch05]. Moment [ACK01]. Moments [Bal03, GR05, Aqu06, Del06]. Monahan [Mar03a]. Money [DJSW09]. Monitoring [AMvW^{+}09, CH08, ELL09, LL06b, LNR03, Sin06c, Gor08]. Monotone [LZH09]. Monte [All06, BL07, Caf07, Dia03a, Fra08, Wie06, Ano00f, BZ00, CR00a, CM06b, CDHL05, Dia03b, FS01, Gan09, GDW04, GM06, HC01, HH07, IZ07, IT02, JHCN06, KJ07, KF07, LPD00, LS01, LW01b, Lia02, Lia05, LLC07, LLW01, Liu03, MPS00, Ros03, Tan04, TLW07b, TLW07a, dV04]. Monthly [PKC01]. Moral [deL05]. Morphology [Ano01j]. Morphometric [IYSK08]. Morphometrics [LM02]. Mortality [Ano02k, Ano02i, BM01, CHB^{+}08, CM09b, DDZS02, DDH04, GT07, GBM^{+}01, IBPL04, KST01, NSP00, PCK06, SWSM09]. Mosler [Joe03]. Mosquitoes [HHS06]. Motif [XLB04]. Motifs [JL08]. Movement [VPW08]. Moving [BRS03, DHS08, Fry05, GS07]. Moyé [Dmi05, Sin06c, Wil06b]. Much [LL03a]. Mukerjee [Rou06]. Müller [Kie05, Spe06a]. Multi [FW07, SYCB08]. Multi-Angle [SYCB08]. Multi-Scale [FW07]. Multicategory [LLW04, LS06, PT09]. Multicenter [BMDD07]. Multiclass [WS07]. Multicourse [TSE02]. Multidimensional [CB09, Dun03, GB00, LFY^{+}08, OR01, Sch03d]. Multifactor [CE09]. Multifold [PS05]. Multigranular [KJG05]. Multigroup [IR05]. Multilevel [Cor05, CSD09, HR06, OZ08, Rob03b, Leh05, Rob03a]. Multilocus [GDH09]. Multinomial [CC09a, Net09, Win03]. Multiparameter [CP07, CZC09]. Multiperiod [GG01]. Multiple [AD08, And08, BT04, BYE^{+}05, BY05, CS09, CHRH01, DN07, Edw05, ELL09, FKS02, Fol00, FKC09, GDS03, Hol05, HZAG01, IYSK08, IPC08, JvD03, JC07, Koo02, LFY^{+}08, LLW00, LJZ04, LR06, LFKS03, MZ08, Mer04, MMZ02, MTR^{+}01, MPRR04, RR07, RW05, SRC^{+}06, Sha05, SL06, SWDS03, TDGW01, TL02, Tam05, WBBR05, WSP09, WS01a, Wes05, WT07, Woo04a, Xia08, YKL^{+}06, vdW02, Dmi05]. Multiple-Cause [MTR^{+}01]. Multiple-Frame [LR06]. Multiple-Index [Xia08]. Multiple-Output [FKS02, Koo02, LFKS03]. Multiple-Try [LLW00]. Multiplicative [IJ04]. Multiply [VVTR08]. Multipopulation [GDH09]. Multiresolution [BMDD07]. Multisample [PC00]. Multiscale [KJG05, Wal02]. Multiset [LCL09]. Multisite [SG00]. Multistate [Hua00, Kle06]. Multivariate [AM01, Ano01p, ACS06, Ban06, BF09, BM09, CD07, CFJZ07, CJS02, Cha04, CH03, CFT06, CA02, CS03a, DN03, DM05, DX09, GPT06, GWKK01a, GWKK01b, GDS03, HS07a, HM03b, KL08, LNR03, LL04, LASA03, LWZ08, Lun06, MH04, OvSG05, Pai09, PGH01, Ran00, RZF06, RL02, SBVW06, TOR05, VH02, VWM01, WSLC06, Wei08, Whi01, WS05, Woo01, Yan05a, YF05, YK04, YW01, ZC07, ZYZS03, ZQ09, vdLHR02, Cha03b, Gen06, Joe03, McN06]. Multiway [KL01]. Mun [Cho06]. Murphy [Len04a]. Murray [Hal06, McC02a]. Murthy [Nes05]. Musical [Iri01b]. Mutations [FD03, ZIP05]. Muzaffar [AC06c]. Myors [Len04a]. Myoung [Bal03]. Myoung-Jae [Bal03].

N [AC04a, Car05b, Dat04, Kar03a, Nes05, LP06]. Nabendu [Ann06a]. Nadarajah [McN06]. Nagar [Ruk03a]. Naish [Teu03]. Nalini [Mor03]. Nan [Mol05]. Nando [Dia03a]. Nanoscale [Kou08]. Nanostructures [DMJ^{+}08]. Narayan [Lun03]. National [DDZS02, FWP05, SRC^{+}06, SI02]. Natural [Gor08, HI06, SO04]. Near [XSZ09]. Nearest [CS01, CvA02, CMRT09, LJ06, MH02, Sim02, SD02, WR02a, WR02b]. Nearest-Neighbor [CS01, CvA02, CMRT09, MH02, Sim02, SD02, WR02a, WR02b]. Nearly [Ros03]. Nearly-Periodic [Ros03]. Needle [FBV^{+}04]. Needles [CS09]. Negative [HCS06, PAS^{+}03]. Neighbor [CS01, CRY09, CvA02, CMRT09, MH02, RHC07, Sim02, SD02, WR02a, WR02b]. Neighbors [LJ06]. Neil [Dia03a, Mil06]. Nello [Las06]. Nelson [Bra06a, Len05, WS02]. Nelson-Aalen [WS02]. Neonatal [RS09b]. Nested [LTE08, MK06, RDG08]. Net [CH06a]. Network [XMN06, CDVY00, EW08, HRH02, HGH08, LL06b, QA09]. Networks [Bau02, GMK^{+}04, GI07, Kou08, LW01a, SFPM04, ZSG06, Gho06, Rue04]. Neural [GMK^{+}04, Gho06, LW01a, SFPM04]. Neuroimaging [Bow07]. Neuronal [FGM^{+}08]. Neurons [HBK^{+}09]. Nevzorov [AC04a]. Newman [Duv03b]. Next [Str03a]. Nicholas [PS06c]. Nicholson [Ram05]. Nielsen [Kim06]. Night [PM05]. Nigra [HBK^{+}09]. Nils [Mir05]. Nitrogen [FKS02, LFKS03]. NNVE [CvA02, MH02, Sim02, SD02, WR02a, WR02b]. Noise [DM00, JB01, ZLI^{+}09, dGKS01]. Noisy [ZMAS05]. Non [PS06a]. Non-Gaussian [PS06a]. Nonclassical [BBS00]. Noncompliance [FG04, LTE08, YL01]. Nonconcave [FL01]. Noncorrelation [HS07a]. Nondetects [Dix06]. None [Bak00]. Nonidentified [Gus09]. Nonignorable [KRS^{+}06, QLS02, RL02]. Nonlinear [AC04e, Ano02n, BLW01, BKV01, CFY00, CRC^{+}04, CV07, CB09, Esc06, FKW01, GS01, Gas03b, GT09, GDW04, Hal08, HS07b, HM03b, KS07a, KS07b, KW01a, KW01b, KW04, KT05, LW01a, LZ01, Mar07, MRID01, Ram01b, SSV05, SMP03, SXZ07, SFPM04, VWM01, VWNM02, Wu02, Wu04, YZ09, Bre05, Gas03a, Yan04c]. Nonnull [JC07]. Nonoverlapping [MW09]. Nonparametric [ASFP09, AV01, AAK06, Ano03g, Ano05k, AKS02, BBCL02, BG01, Bre05, CX08, CX09, CRC^{+}04, CDH09a, CDH09b, Che04, CDK05, CHJ08, CFK07, CD09, CG03b, CCDP09, CFN01, DM07, DSV04, DNP05, DTT08, DX09, DT00, Efr05a, EHT06, Els02c, Fan00, FJ05, FF09, FM09, FM06, FOD07, Fry07, Gal07, GA05, GG08, GKM05, GC09b, GSC04, Gua08, Gua09, Hab03, HM03a, HMS00, HW01a, HM00a, HM00b, HM00c, HL05, HW00, HLS05, Jam02e, Jen06, KM02, KW04, Kol03b, KPD02, KR02, KTZ00, LGS07, LZ08, LPW05, LC00, LY01a, LC01a, LY01b, Lin04b, LJ00, MK06, MR05, MWL^{+}01, MVBC03b, MVBC03c, Ned09, OMS^{+}09, PSH01, Pre03, Rab02a, Rab02b, Rag00a, Rag00b, Rid07, Ros05, Sch02a, Sch09, She02a, SHC02, Sta09, SW09, TOR05, Tre06, WC01, WW01, WL05]. Nonparametric [WA06, WLH08, WLC02, WH03b, XLCM03, ZW08, Che08, Cra07, Gre07b, Peñ06, Pou04, Spe06a, Wol08]. Nonparametric-Correction [Ano03g, HW00]. Nonparametrics [Gho06, LDVT09, Epi04]. Nonproportional [BBN^{+}04]. Nonrandom [Hil08, LLL08a, PM05, SCS08a]. Nonrandomized [HBR01, Hil08, LLL08a, SCS08a]. Nonresponse [MTR^{+}01, NC02, OZ08, QLS02, EP06]. Nonseparable [Ano01m, Gne02]. Nonsmooth [CK08, CK09]. Nonstandard [LP06]. Nonstationary [BME08, DLRY06, EL04, FO09, HCS06, HOS04, KMH05, Lin03b, ORvSM01, OvSG05, SG00, VR00]. Nonuniform [Cow07]. Norm [WS07]. Normal [BM09, CH06b, EHI08, FHY07, FRS06, HRC04, HTZ^{+}09, LL06c, LPW05, SJ06, War08]. Normality [She02a, Kaf03]. Normalization [Cra05, FPHR05, FPH^{+}05, HZ05, HWZ05, KM05, Sab05, Tib05, Zha05a]. Normalized [LMP05]. Normalizing [RWR03]. Normals [OMS^{+}09]. Normative [ZWT^{+}05]. North [Hig09, LS09a, LS09b, Men09, WM09b]. Northwest [MRŞK06]. Note [DNP05, Hil09, KC01, KK07, SY06]. Notz [Mor04]. Nowcasting [XWF05]. Nripes [Not03a]. Nualart [Yor06]. Nuclear [ST02]. Nucleotide [GDH09, SWDS03]. Nuisance [LP06, VCR09]. Null [BB00a, BB00b, Boo00, Efr04a, Els02c, Eva00a, JC07, KM02, Mar00a, Pic00, Rab02a, Rab02b, RvdVV00a, RvdVV00b, Sch02a, Ste00c]. Nullity [FZ09]. Number [AA00, HL07, MSZ02, Mao06, SW02d, SJ03]. Numbers [Teu03]. Numerical [JKN08, Mar03b, NDKW01, Mar03a, Van05]. Nutritional [JHISR07].

Object [LSQ03, Rue04]. Objective [Ano03i, BDS01, CM06a, Sho05]. Observation [HS02a, SPSZ05, SSL07]. Observational [Han04, HRS09, HR06, LZHR01, Ros02, RRS07, RS09a, RS09b, Wol02b]. Observations [AM01, Ano01p, CZ07, EOU02, GK07, GWKK01a, GWKK01b, HZ08, ISZ07, JM06, KTZ00, Whi01, WS05, Woo01, YW01]. Observed [DN03, LM09b]. Observing [Eva00b, Kal00, Roy00a, Roy00b, Wol00, Won00a]. Obstructive [NSP00]. Occurrence [DR05, LSGS01, LSGS04, And05]. Occurrences [Sch03d, ZOVJ02]. Ocean [Kad07, Mil07, RLL07a, RLL07b, WMNB01, Wol07]. Odds [Che01, ZLY05]. Offen [Dur06]. Offspring [YSY08]. O’Gorman [Len06]. O’Hagan [Ank05]. Ojeda [Boo06]. Oligonucleotide [HWZ06, WIG^{+}04]. Oliver [Dan03, Zhu06]. Ombach [Gan03a]. Omnibus [WQ07]. On-Demand [Li07]. Oncogenic [LFY^{+}08]. One [AP04, FHLR05, LL04, PKC01, YW03]. One-Level [PKC01]. One-Sided [FHLR05, LL04]. One-Way [AP04]. Online [LM09b]. Only [BP08, CN08, Del02b, SGL07]. Onset [IC05]. Operating [BP01, CP02, DP03, DP06, LZ08, Pep00, RFS06]. Operational [Nta07]. Ophthalmology [AM01, Ano01p, GWKK01a, GWKK01b, Whi01, Woo01, YW01]. Optimal [ABPR06, AC07b, Bha09, BDZ06, BF01b, Dea08, DMW05, DBPP08, Dou06, DM00, HS07a, HSY06, HS02b, HC07, KS02, KS08, Li07, LFGR04, MD01, MRT05, MVY09, Mor07, MSD04, MPRR04, Not03a, Not03b, SS09, SH06, TW07b, TRH07, THC05, Rob04]. Optimality [HY04, HR03, ICS03, LH07, PT07]. Optimization [FD08, Fre04, LLW00, Wil03a, Wil03c, Lue05, Ste05a]. Optimizing [FG00, MRŞK06]. Optimum [AC08b]. Option [FM09, Kor05]. Options [Kor04]. Oracle [FL01, SC07, Zou06]. Orbits [JKKW03]. Ord [Ann06d]. Order [BF02, CK08, CK09, Efr01, GS06, GWB08, Gua09, HM03a, Kim00, Mic06, OS00, Rou06, RFP09, SSL03, Str00g, VWNM02, WSLC06]. Order-Based [GS06]. Order-Preserving [HM03a]. Ordered [BFD01, Win03]. Ordering [EM05]. Ordinal [GC09b, KRS^{+}06, PAS^{+}03, ZFZ03]. Ordinary [Had05a]. Orientations [BNV09]. Oriented [Ban05]. Oscar [Ker06]. Oscillations [Mei07]. Other [Kor04, Len05]. Outbreak [BB03]. Outcome [Bak00, FD08, HM00a, HM00b, HM00c, Rag00a, Rag00b, RSM^{+}07, WZ05, ZCPL08, KRS^{+}06]. Outcome-Dependent [RSM^{+}07, WZ05]. Outcomes [CHRH01, HRC04, HH06, MTR^{+}01, NN09, RL02, Rub05, SGL07]. Outlier [Ano01n, GPT06, MW05, SGC05, TBD^{+}03]. Outliers [CC00]. Outmigrating [NR02]. Output [BL07, Bri04, CJ01, FKS02, GRG^{+}04, GRGB04, HGWR08, JHCN06, Koo02, LFKS03, Rou04, TN04]. Outputs [FKS02, LFKS03]. Ovarian [RRS07]. Overcoming [RGA01]. Overdispersed [CKN08]. Overdispersion [ZSG06]. Overidentifying [Sma07]. Overlap [MRŞK06]. Overview [LL05b, Sin06b]. Ozone [Mei07, SGH07].

P [Dav06b, Dev06, Eas03, Kac04, Kar03a, Lee06, Mal05, Nea06, Not03a, Nus06, Sch05, Spe06b, Van05, Yan04b]. PA [DK01]. Paap [Lun05]. Packet [RTD01]. Padgett [Lin04b]. Padhraic [San03a]. Pair [Ric08]. Pair-Matched [Ric08]. Pal [Ann06a]. Palta [Had05a]. Panel [IJ04, KM01, Lar06, LZH09, Bal03]. Paolo [Cle06, Emb03]. Paradis [Las07a]. Parallel [CRY09]. Parallel-Chain [CRY09]. Parameter [AMP06, BBS00, CC09a, DH08, GR05, LW01b, LW08, Woo04a]. Parameterization [Gel04]. Parameters [BYE^{+}05, BY05, DV03, Edw05, Hey08, HTO03, Hol05, LP06, RFP09, Sha05, Tam05, VCR09, Wes05]. Parametric [Ann05a, Ano02n, BG01, Bre05, Che04, DP06, EG09, FH01, FTW01, GBT09, Gua09, MWL^{+}01, Lin04b]. Pareto [AMP06]. Parking [MD01]. Parsimonious [SK02]. Part [OS01, RZF06, ZSCW06]. Partial [Ala07, AD08, CJ06a, FPH^{+}05, HFZ05, JR08, PWZZ09, RO07, Zuc05, vWvdWvdV08, Kar07a]. Partially [CFJZ07, DR00, DN03, FBV^{+}04, LWRC04, LL09, YLZ08, YR02, vDP08]. Particle [ABPR06, LZCC05, deL05]. Particulate [CHB^{+}08]. Partitioning [LH09, TPO03]. Parts [Jia05, LQ05, Lou05, Rup05, SFK^{+}05b, SFK05a]. Pascal [Cle02c]. Passerini [Poo06]. Past [CL00, O’F01, Xie00]. Paths [LH09, WSP09, YZ09]. Pathways [LFY^{+}08]. Patients [YF08]. Patilea [CN08]. Pattern [GRT04, Kar06c, Las06, TBJ08, Bur08, Las07b]. Pattern-Mixture [GRT04]. Patterns [Gol05a, GL07, GL03, LPS01, LCLC08, Mat05]. Paul [AC06b, Jam03a, Neu06, Sch04a, Spi06, Tra05]. Pavel [Mil06]. Pay [MD01]. Pearl [Had05b]. Peat [Lin04d]. Peer [Bha09]. Pei [Xie03a]. Peihua [Wan06]. Penalized [FL01, JRY04, JLZ08, Kar03a, Kar03b, KK07, MZ08, WL05, War08, YR02]. Penalties [Bur04, DLM04, EBD^{+}04, Efr04b, SHY04a, Zha04a]. Penalty [CK08, CK09]. Penetrances [Hös03, Hss03]. Pensky [Sur04]. People [ZSG06]. Pepe [Qiu05]. Perception [SME03]. Perfect [FOD07]. Performance [HT06, HC03b, Koo02, RZ03]. Performances [Hil09]. Period [JKKW03]. Periodic [KOC07, Ros03, Lun05]. Periodogram [Li08b]. Periodontal [RHC07]. Permutation [BF01b, Jam03b, WL00, Ann05a, Jam03a]. Permutations [GRF03]. Perry [And08]. Perspective [BGM^{+}05, Hob00, Ban05, Gen07a, Hal08]. Perspectives [STJ07, Sis05, Tra05]. Perturbation [Bri04, GRG^{+}04, GRGB04, Rou04, TN04]. PET [OMS^{+}09]. Peter [Ann06b, Ann06c, Bet03, Bin06a, Bra06a, Cav06a, Cow03b, Gan03a, Mir05, Rob04]. Pevzner [Mil06]. PGA [CR08]. Pharmaceutical [Eas03, GH00]. Pharmacologic [CCDH08]. Pharmacology [Kra08]. Pharmacometrics [Kra08]. Phase [CCB03, KNF06, LP05, RFP09, YY09]. Phenomena [Gol02c]. Phenotype [KPD02]. Phenotypes [JFY07]. Philip [Lun05]. Philippe [Bre06]. Phillip [Ann05a, Yan05a]. Phones [MD01]. Photon [CCDH08]. Phylogenetic [LPD00, SWDS03]. Phylogenetics [Las07a]. Physical [HID09]. Physician [TEJ^{+}04]. Physics [Bre06, Vie02b]. Piecewise [KMH05, NLLH06, Sch07c]. Pierce [Dan03]. Pierre [Emb03, deL05]. Pigeot [Cav06b]. Pivotal [ADW07]. Placebo [TPO03]. Plan [Kje00]. Planning [Li07, AC05c, AC05b]. Plant [Dan03]. Play [CR08]. Players [Koo02]. Plot [Rob04]. PLUS [AC04e, Daw03, Eas03, McN04]. Plus(R) [Lun06]. PM [KSZL02]. Pneumococcal [AALT00, CTG^{+}06]. Point [Dra06, Gua06, GL07, Gua08, GWB08, Gua09, HL06b, IJ04, LWY01, McB02, Nes07, NC08, Sch07c, Sch03d, Sch05, WM01, Wat07, Sch07c]. Points [KM01]. Pointwise [CFN01]. Poisson [Lin04a, BIW00, CM06b, FWP05, FRT09, HYFW05, RAHT06, WBS07a]. Police [GFK07, RM09]. Policy [FD08, GFK07, HR06, KM07]. Political [Bec00a, RGK01]. Politics [Teu03]. Pollard [Tay03a]. Pollen [PM09]. Polling [RGK01]. Pollution [DDZS02, DDH04]. Polya [HJ02, Han06]. Polymorphisms [DHE08, GDH09]. Polynomial [DFC09, EG09, GG07, WA00, WZ02, XLCM03]. Population [Ano02k, Ano02i, BST03, CBL^{+}08, Els02c, GBM^{+}01, Had05a, HTZ^{+}09, IC05, JL06, KM02, LL07, Lit04, MMC07, Mee00, MAC^{+}06b, Rab02a, Rab02b, RGK01, Sch02a, SW02c, SS04, TEB00, XSZ09]. Population-Calibrated [IC05]. Population-Level [MMC07]. Populations [YSY08, Few04, Sch07e]. Portmanteau [GK07, PR02]. Positive [Ano02i, BFD01, GBM^{+}01, ZCI^{+}09]. Positive-Definite [ZCI^{+}09]. Positives [BP08, CN08]. Possibly [JM06]. Post [Dra06, HDS06, ZSB01]. Post-Processing [HDS06]. Posterior [CHR00a, CIS04, Gew04, Gus09, HDS06, Iri01a, ST04, She02a, SJ06, YL09, dV04]. Postinfection [HH06]. Postrandomization [SGL07]. Postseason [Har03]. Poststratification [GM06, RGK01, WSLC06, YR00]. Poststratifying [RGK01]. Posttest [HQF08]. Posttransplant [SWSM09]. Potential [Rub05, Sin06a, Sin06b]. Potentials [RTD01]. Potters [Bre06]. Poursat [AC04e]. Power [Boc08, BIK00, DM04, HR03, ICS03, Len04a, MH07, YZL08, Len04a]. Power-Transformed [YZL08]. Powerful [PR02]. PowerPoint [Len04b]. pp [McC02a]. Prabhakar [Nes05]. Practical [AC04e, Bar08, Dur06, Gil08a, Hu04, Kar07b, Len04a, Swy02b, AC05d]. Practice [Dav06b, Dia03b, Gri02b, McC05, Par06, Per06, RY00, Dia03a, Sey05, Gre07b, Sch07a, Wol08]. Practices [MSM05, RL02, Bou07]. Pranab [Mic06]. Pravastatin [BBN^{+}04]. Pravastatin-Aspirin [BBN^{+}04]. Precedence [AKS02]. Precipitation [CNN07]. Preclinical [GH00]. Predicting [BBK^{+}09, Mao04, SS04]. Prediction [AC08c, BME08, BHPT06, Bur04, CDH09a, CDH09b, DLM04, EBD^{+}04, Efr04b, Efr09, EOU02, FF09, FD08, GT07, GR07, GR06a, JL06, KSZL02, MVY09, NDKW01, NS01, Par02b, Sch09, SHY04a, Sta09, SW09, THC05, UCTW07, YTS08, Zha04a, Qiu05, Rup04]. Predictions [Fue03]. Predictive [FS04b, HDS06, ML02, PS05, PY00, ZCPL08]. Predictor [DHSR08, IBPL04, Lan05, LC00]. Predictors [BD09, BS06, LL07, LL08, LM09a, SW02d, WF03]. Preferences [LHN08]. Preliminary [Str07]. Prentice [Bor03]. Preparing [Mal05]. Preschool [And08]. Presence [Efr05a, HS02a, HYFW05, HLS05, KS08, LTE08, RFS06, RFF04, SW02c, SRB08, SSL07, VCR09, Wal02, XSZ09]. Present [CL00, Xie00]. Preserving [HM03a]. Presidential [Han03, Lac09]. Presmoothing [HCB06]. Pretest [HQF08]. Prevalence [CDR08a, CDR08b, Dun08, Pac08, RXS^{+}07, Tho08]. Prevent [HR00]. Prevention [Bak00, SQC07, SjKH^{+}01]. Preventive [WBBR05]. Price [HY07, IK08, MKCX04, WJSS08]. Prices [Ano02o, KOC07]. Pricing [Bre06, FM09, GM09, Mat05]. Primary [LH09, TEJ^{+}04]. Primate [JDvS02]. Primer [AC04a, Bro07, Mye07]. Principal [BHPT06, BFHR03a, BFHR03b, BX03, CCDH08, CM01, Far01, FBV^{+}04, JR08, JL09a, KU01a, KU01b, KZ03, LLC00, LM00, Mar01, MJB03, Ram01a, RO07, SBVW06, Soo00, SW02d, ZRM09, Tim03]. Principle [HY01a]. Principles [Gri02b, San03b, Sho05, Gol07, San03a]. Prior [CIK08, Gre03, ICS03, NT05, VCR09]. Priors [CL08b, GC09b, IJ01, LPM^{+}08, LMP05, RHC07, Rou06, WDL04, WS01a]. Priouret [Emb03]. Prison [ZSG06]. Pro [Sch07b]. Probabilistic [Bri04, Buo06, Car06, GRG^{+}04, GRGB04, GLW^{+}06, Jen02b, Rou04, TN04, Gol07, Kie07b]. Probabilities [CH06b, Das02b]. Probability [Bin06b, Cha03a, Cle02c, Del02b, Dem04, Dev06, Eva00b, Gan03a, Gan03b, GKO05, HRL04, Iri01a, KL06a, Kal00, Kar07a, Kie07a, Lav04, Lin04c, Mao04, Net09, PS07, Roy00a, Roy00b, She02c, SBL^{+}04, SB04, Spa02b, Tay03b, Wol00, Won00a, Zad04, Tay03a, Ter04, Buo06, Car06, Fil06, Ros04, Rou06].

Probable [Ald08, DL08, Edw08, Sen08a, Sti08, Zab08b, Zab08a]. Probit [FB00, GA01, Win03]. Problem [BFV01, DM07, DFC09, FH00, Geo00, Mee03]. Problems [ADW07, BCR02, CCB03, DH08, Efr09, Gem06, IZT08, JM04, Man07, MJvAC02, QZL09, Szk03, Roh05]. Procedure [GL07, Lee00]. Procedures [Che07b, FL04, HH07, HR03, IZ07, KJ07, KF07, KSG05, SGC05, TDGW01, TL02, TLW07b, TLW07a]. Proceed [SRRW00]. Process [AALT00, BGS03, BG06b, BC03a, Dav06b, DXC08, GKM05, GL08, Gua06, LTMS02, Lob01, MSZ02, McB02, Nem04, RDG08, Sch07c, Sch03d, SRRW00, Str00f, SjKH^{+}01, VS08, WBS07a, ZQ09]. Processes [Bin06b, CM06b, CCPY09, CV07, Cli07, Fri07, GEG05, GKSB03, GS06, Gua08, GWB08, Gua09, HL06b, HW04, IJ04, KS07a, KS07b, KMH05, LWY01, LDM09, Lun03, Lun07a, Mar07, MSY06, Nov05, Oun08, PCK06, PY00, TJBB06, Vie02b, YSY08, YF05, Bin06a, Kie06, Sch05, Hoo07, Nes07, Sch07c]. Processing [HDS06, Wan06, Rie06, Can07]. Procrustes [Roh05]. Product [Len05]. Production [FKS02, LFKS03]. Profession [Mas04]. Professionals [Lac09]. Profile [BR00, FW00a, LKF05, Li00, MV00, MvdV00, RRvdL00, She00b]. Profiled [Tsa03]. Profiles [MAC^{+}06b, MRID01]. Profiling [GR06b, RZF06, TPO03, ZSCW06]. Progenitor [YSY08]. Prognosis [FSL^{+}06, FG04]. Prognostic [RT00, ZCPL08]. Program [FBV^{+}04, GBT09, KRS^{+}06]. Programmer [Cla03]. Programs [Bas03, SjKH^{+}01, ZRM09]. Progress [Sis05, SI02]. Progression [FD03, IEMM08]. Projection [CL08a, GPT06]. Projections [PP01]. Projective [LWZ08]. Projects [And08]. Propensity [DR00, IV04, RT00, Tan06a]. Proper [GR07]. Properties [CIK08, FL01, ICS03, KP05, LL05b, LGS07, SWP00, Zou06]. Property [Pui03, WDL04]. Proportion [GW06, JC07]. Proportional [Ano01o, Che01, GWB08, HI01, IJ04, QWP05, ZLY05]. Proportions [CL08b, Bra06a]. Proposed [BP01]. Propriety [CIS04]. Prospects [Fan00]. Prostate [GMK^{+}04, LTMS02, SRA08, YTS08]. Prostate-Specific [LTMS02]. Protein [BME08, LDVT09, XLB04]. Proton [LSZH08]. Provider [RZF06]. Providers [ZSCW06]. Proximal [MWL^{+}01]. Proximity [CB09, McB02]. Proxy [PM09]. Pseudo [VR00, VCR09, Zuc05]. Pseudo-Likelihoods [VCR09]. Pseudo-Maximum [VR00]. Pseudoscore [CCB03]. Pseudovariables [WBS07b]. Psychiatric [LNR03]. Psychometrics [Bro00b]. Public [Dig05, LL03a]. Publication [DT00]. Pulmonary [NSP00, YF08, ZWT^{+}05]. Purchase [BB76, BH08, BBK03]. Pursuit [GPT06, ZWL^{+}04]. Pythagorean [Tra05].

Q [Gen04]. Qiu [Wan06]. Qiwei [Bre05]. Quadrant [BFD01]. Quadratic [JM04, SW02c]. Quality [AC06a, Gen05b, GRYB06, Len02f, Man07, Nem04, Poo04, Rei08, ZYDS08]. Quantal [DNP05]. Quantification [McC05]. Quantify [JKN08]. Quantile [CX08, CX09, CH02, EG09, FF06, GP07, HL06a, HM03a, HW06, HZ03, HC02, HL05, Kni06, KG01a, KX04, KX06a, KXF^{+}06, KX06b, LLZ07, MH07, PH08, PF09, Rob06, WW09, Wei08, WC09, XK09, YZL08, Jur06]. Quantiles [DZ03, MS05, Por03, NVP06]. Quantitative [All06, AC03b, AC03e, Gue08, HHS06, JFY07, Kra08, Lin04d, LCT^{+}05, AC06b, Had05a]. Quantity [BBK03]. Quasi [AC05a, LSQ03, MB05, RL01, WM09a]. Quasi-Experimental [AC05a]. Quasi-Independence [MB05]. Quasi-Likelihood [WM09a]. Quasi-Score [RL01]. Quasi-Stellar [LSQ03]. Queueing [BGM^{+}05, Rya05, Vla08]. Queueing-Science [BGM^{+}05].

R [AC04b, AC04e, AC04c, AC05a, Aqu06, Bra06a, Cow06, Dav07a, Dix06, Epi04, Fie06, Gal07, Hu04, Iha06, Jer06, Kac04, Las07a, Lee06, Len04b, Len04a, Lun06, Lun07b, Not03a, Sch03b, Wik04, Wil04a, AC07c, Har07, Tyl07]. R&D [AC04d]. Rabe [Cor05, Daw03]. Rabe-Hesketh [Cor05, Daw03]. Race [Sch03a]. Racial [GFK07, GR06b, RM09]. Racing [GRF03]. Radar [XWF05]. Radhakrishna [Imh03a]. Radiance [LLW04]. Radius [GEG05]. Radmacher [Gen05a]. Raftery [Nol03a]. Raghavarao [Peñ06]. Rahul [Rou06]. Railroad [MSM05]. Rainfall [SG00]. Ralf [Wil03a]. Rama [Bin06a]. Ramamoorthi [Epi04]. Random [AAK06, BM01, Boo06, BD05, CIS04, CH06b, CZI07, CCL09, CHRH01, Cow07, DMRM04, Efr01, Fel02c, Gol02c, GRT04, HSY06, Hil08, Kar06a, KSZL02, Kie08a, KT05, LJ06, LLL08a, LM04, MW09, OS01, PGVW04, PCK06, RTD01, Ric08, Ros05, ST04, SCS08a, SS04, TW07a, WLH04, Yan06, ZLY05, ZL07b, Kar06c]. Random-Effects [BD05, KT05, OS01, Ric08]. Randomization [Ass07, Bai04, GSCR04, HI06, HR03, STR08]. Randomized [Bak00, BFHR03a, BFHR03b, BX03, BF01b, Buo06, GSCR04, HB09, Har00a, Hil08, HM00a, HM00b, HM00c, HR00, KZ03, LLL08a, MJB03, Rag00a, Rag00b, Ros07, Rub08, SCS08a, SCS08b, SQC07, STR08, Sob06, TEJ^{+}04]. Randomness [Kul06, Pai09]. Randy [AC05d]. Rank [AKM01, BFB00, CS03a, DS05, GGR07, Hel07, JDvS02, KMR09, OOK03, Ozt02, PF06, WKL09, vdW02]. Rank-Based [GGR07, KMR09]. Rank-Sum [DS05]. Ranked [FM06, Ozt02, Ozt05]. Ranked-Set [FM06, Ozt02]. Rankings [FOD07, Gil03]. Ranks [XSZ09]. Rao [Imh03a, AC04c, Dat04, Imh03a]. Rare [BFB00, Pic05]. Rasch [Ric04]. Rasmus [Kim06, Sch05]. Rate [ABDL01, Ano03j, BYE^{+}05, BY05, CBCS07, DHS08, Edw05, HS07b, Hol05, HS02b, IBPL04, JRY04, KT09, MY08a, PZB09, Sha05, SC07, Tam05, Wes05, Yek08]. Rate-Adjusted [BY05, Edw05, Hol05, Sha05, Tam05, Wes05]. Rate-Optimal [HS02b]. Rated [HID09]. Rater [Cho06]. Raters [AD08]. Rates [BGS03, BH07, BL06, XMN06, Bra06a, CM09b, GBT09, HPC03, KST01, NSP00, PDK07, RXS^{+}07, RM09, Ros03, Sch03a, TIY03, TBD^{+}03]. Ratio [FHW00, NLLH06]. Ratios [BP02]. Ratto [Per06]. Raudenbush [Dav03]. Ravishanker [Mor03]. Raz [ORvSM01]. RC [BF02]. REACT [Ber00a].

Readings [LTMS02, Str03b]. Reagle [Mal05]. Real [LPS01, LW01b]. Real-Parameter [LW01b]. Realized [ABDL01, Ano03j]. Reappraising [KG01a]. Reasoning [Car06, Had05b, Sha07, Woo04b, Teu03]. Reassessment [CMRT09, YY09]. Recall [HI06]. Recapture [BF06, BP08, CN08, CBL^{+}08, Pol00]. Receiver [BP01, CP02, DP03, DP06, LZ08, Pep00, RFS06]. Receiver-Operating [RFS06]. Receptor [PGH01]. Recidivism [Sir06]. Recognition [Bur08, CV07, KS07a, KS07b, Las07b, Mar07, Rue04, Kar06c]. Reconstruction [DN03]. Record [Hil09, LR01, Lin04b, PM09, Rei08]. Record-Breaking [Lin04b]. Records [EM08, JNKD03]. Recounts [Han03]. Recovery [But06, CLQ03, IBPL04]. Recurrences [Len05]. Recurrent [CLLCL09, HW03, HW04, PSH01, SMIO08, WQC01, YF08, ZL07b, ZYDS08, Len05]. Recursive [BPB02, Sco02]. Reduced [CS03a, GP07]. Reduced-Bias [GP07]. Reduction [CR00b, CB09, CP07, CN05, CF09, IYSK08, JB01, LS02, LASA03, LW07, LWZ08, LL08, WX08, Xia08, Yat05, YW03]. Reestimation [SC03]. Reevaluation [And08]. Reexamination [FZ03]. Reference [HT06, NK00]. Reference-Free [HT06]. Reflectivities [XWF05]. Reg [KOC07]. Reg-ARFIMA [KOC07]. Regime [GLW^{+}06]. Regime-Switching [GLW^{+}06]. Regimes [IV04, MvdLRC01, MB09]. Region [Bow07, KPD02]. Regional [CRY09, DDZS02]. Regions [BME08, CC09a, Joe03, MWL^{+}01, PY00, SS09, YK04]. Registration [KR08, TI08]. Registry [MFCF05]. Regression [AV01, AC04e, Ano00g, Ano00f, Ano02m, Ano03g, BCR02, BIW00, BFV01, BRS^{+}00, BY03, BC01, CFY00, CD07, CFJZ07, CRC^{+}04, CCB03, Cha04, CR00b, Che01, Che02, Che04, CIS04, CDK05, CJ06a, Che07a, CZI07, CIK08, CH02, CLR06, Číź08, CC00, CS03a, CN05, Cra07, CSD09, DM07, DSV04, DP03, DP06, Dun05, Efr05a, EG09, FH01, Fue03, Gal07, GBT09, GG07, GW05, GI07, Gus00, HM03a, HDW07, HJ02, HSS01, HO02a, HO02b, HZ03, HL00, Hel07, HC06, HM03b, HC02, HW01a, HS02b, HL05, HL04, HW00, HW01b, Hua02, HRV02, JLZ08, KW04, KC07, KVZ07, KG01a, KG01b, KS00, KPY03, KL00, LL05a, LL05b, LLC00, LW07, LLZ07, LW08, LY01a, LC01a, LY01b, LC01b, LFY^{+}08, LM09a, LJZ04, LH07, LS09c, MMC07, MY02, MPS00, MMB09, Mer04, MH07, MW05]. Regression [NT05, NST07, NVP06, Neu06, OOK03, Ozt02, PH08, PF09, PWZZ09, Por02, Por03, QWP05, Rao03a, Rao03b, RO07, Rup02b, SSV05, SW02b, She00a, SJ04, Sin07b, Sma07, SRL00, SHH00, SPSZ05, SSL07, Tan06b, TZW05, Tro06, UCTW07, VWM01, Wal07, WC01, WW01, WLH04, Wan06, WQ07, WX08, Wan09, WW09, WT09, WZ05, WC09, WLC02, XLCM03, XPLT09, XLL04, Yan01, YLZ08, YZL08, YY05, ZS01, ZZ06, ZMP06, ZCI^{+}09, ZLI^{+}09, Zuc05, Had05a, Jen06, Jur06, Lee06, Pre03, Zim03, Ano04e, Cav04, Lyl06, Rao03a]. Regularity [OC06]. Regularization [AF01b, AF01a, Cai01, Mou01, Sol01, Vid01, War08, Wil03a, Wil03c]. Regularized [HSB09, YZ09]. Regulation [HHS06]. Regulatory [GI07].

Rejoinder [AF01a, BFHR03b, BB00a, BYE^{+}05, BY05, Bhl02, CDH09b, CDR08b, Daw00b, EBD^{+}04, Efr04b, FPHR05, FPH^{+}05, GWKK01b, GRG^{+}04, GRGB04, Gel08, Gen00b, HO02b, HC03b, HM00a, HM00c, Hor01b, JL09b, KS07b, KW01b, KU01b, KXF^{+}06, KX06b, LS09a, LY01b, LZ06, MM04, MVBC03a, MVBC03c, MvdV00, OBMK07b, PDG02b, Rab02b, RLL07a, RCGT01b, RvdVV00b, Roy00b, STPN09b, SCS08b, SBL^{+}04, SB04, SFK^{+}05b, SFK05a, TLW07a, WW00b, WW00c, WR02b, YKL^{+}06, YK06, Zab08b, ZZIP07a, MMH^{+}04]. Rekab [AC06c]. Related [Ann06a, DHE08, Els02a, LNR03, RTD01, Str00g, WM09a]. Relating [CHB^{+}08]. Relation [Gre03, WM01]. Relations [RHC07]. Relationship [MWL^{+}01, Ste01]. Relationships [DDZS02, TH09]. Relative [CAC^{+}08, FZ09, IBPL04, KL04b, ST02]. Relative-Risk [KL04b]. Reliability [KP05, Law00, SSL03, AC08c, Gen07a]. Religion [LHN08]. Remarks [BL07, CN08]. Renyan [Nes05]. Repairs [Len05]. Repeated [ERS02, FEG07, HZAG01, LC09, Lyl06, Moe03a, Moe03b, WC04, WLH08]. Repeatedly [CRW01]. Replicates [GRYB06, HL04]. Replication [KNF06, SY06]. Reporting [MFCF05]. Representation [FS06a]. Representations [FO09, Gen07b]. Reproducing [LLZ07]. Resampling [BZ00, Gan09, HO02a, HO02b, HRV02, LWZ08, Liu03, MY02, Por02, Rup02b, Car05b]. Resampling-Based [BZ00]. Research [GH00, Len04a, Lin04d, Sto04, Tsa00]. Researchers [All06]. Reservoir [EOU02]. Residual [KR02, KTZ00, Sch03d, SZ09, WBBR05]. Residuals [HRC04]. Residue [OMS^{+}09]. Resistant [Ano01n]. Resolution [LCT^{+}05, SGH07]. Resolutions [BIW00]. Resonance [Gen00a, Gen00b, Hay00, Hig00, Laz07, MG00, Raz00, Sch07d, SF07, Wor00, ZZIP07a, ZZIP07b, ZLI^{+}09]. Resource [Gor08]. Resources [Chr07, Lit07, OBMK07a, OBMK07b, Rup07, SO04]. Respiratory [ZYDS08]. Responders [TPO03]. Response [Ass07, BDZ06, DDZS02, DHSR08, Fri02b, GB00, Gre03, Hab03, HHS06, HK02, HR03, LASA03, LFY^{+}08, LFKS03, MVY09, MPK09, SS04, TRH07, WLH04, Ala07, AC08d]. Response-Adaptive [Ass07, HR03]. Response-Maximized [MPK09]. Responses [GA01, KSS08, LWZ08, PGLW00, QSZ08]. Restoration [Qiu07]. Restricted [BDZ06, CR00b]. Restrictions [BF02, Efr01, RFP09, Sma07, Mic06]. Results [Fie00, GRF03]. Retention [HR06]. Retrospective [WS02]. Return [CNN07]. rev [Vog02d]. Reverse [Ste01]. Review [Alb00, Alb02, All06, AC01b, AC02a, AC02b, AC03a, AC03b, AC03c, AC04a, AC04b, AC04e, AC04c, AC04d, ACLJ^{+}04, AC05a, AC05b, AC05d, ACE05, AC06c, AC06a, AC06b, And05, Ank01, Ank05, Ann05a, Ann05b, Ann06a, Ann06b, Ann06c, Aqu06, AC03f, Bai04, Bai05, Bal05, Bal03, Ban06, Ban05, Bar01a, Bar01b, Bas03, Bei04, Ber01, Ber00c, Bet03, Bin02, Bin06a, Bin06b, Bog04, Boi01, Bol02, Bor03, Bor02, Bra06a, Bra06b, Bre05, Bre06, Bro01, Bue02, Buh00, Buo06, But00, But06, Cal01, Can00, Car01, Car03a, Car05a, Car05b, Car06, Cav04, Cav06b, Cav06a, Cha03a, Cha03b, Chi06, Cho06, Chr01a, Chr01b, Cla03, Cle01, Cle02a, Cle02b, Cle03a, Cle06, Coa00, Cog02, Cor02a, Cor02c, Cor05]. Review [Cow00a, Cow00b, Cow03b, Cow06, Dan03, Das02a, Dat04, Dav03, Dav06a, Dav01, Dav06c, Dav06b, DKC^{+}06, Daw03, Del02a, Dev06, Dha01, Dia00a, Dia00b, Dia00c, Dia03a, Dig05, Dix06, Dmi05, Dob02a, Dob03a, Dob03b, Dob04, Dob00, Dou06, Dow00, Dra06, Dua01, Dub01, Dur06, Duv03b, Dyk00, Eas03, Ell01, Els02b, Emb01a, Emb01b, Emb01c, Emb02a, Emb02b, Emb02c, Emb03, Ens02, Epi04, Ern00, EP06, Fel02a, Fel02b, Few04, Fie06, Fil06, Fre04, Fri02a, FS00, Fue03, Ful02, Gad00, Gal05, Gan03a, Gan02, Gar00a, Gar00b, Gas03a, Gem06, Gen02, Gen06, Gen01, Gen04, Gen05b, Gen05a, Gey01, Ghe01, Gho01, Gho06, Gol02a, Gol05a, Gol02b, Gol05b, Gra02, Gri02a, Gun02, Hab03, Had05a, Had05b, Hal03b, Hal06, HSSM06, HC01, Har00b, Haz00, Hel00]. Review [Hen02, Hes00, Hol00b, Hol00a, Hor02, Hu04, Hur00, ICLH05, Iha06, Imh03a, Ing00, Jam00a, Jam00b, Jam00c, Jam00d, Jam00e, Jam01a, Jam01b, Jam01c, Jam01d, Jam02a, Jam02b, Jam02c, Jam02d, Jam03a, Jen02a, Jen06, Jer06, Jes03a, Joe03, Jon00, Jon06, Jur06, Kac04, Kaf03, Kar02a, Kar02b, Kar02c, Kar03a, Kar06a, Kar06b, Kar06c, KHACM06, Ken03a, Ker06, Kie05, Kie06, Kim01a, Kim01b, Kim06, Kle06, Koe02b, Kok01, Kol02a, Kor04, Kor05, KL05, Kra04, Kri05, Kue00, Lan00, Lan03, Lar02, Lar06, Las06, Lee03, Lee06, Leh05, Lei05, Len02a, Len02b, Len02c, Len02d, Len02e, Len03b, Len04b, Len04a, Len05, Len06, Lev00, Lin04a, Lin04b, Lin03a, Lin04d, Lit00, Loc02a, Loc04, Lue05, Lun00a, Lun00b, Lun01, Lun02, LACPE03, Lun03, Lun05, Lun06]. Review [Lüt03, Lyl06, MTW00a, MTW00b, MTW00c, MTW00d, MTW00e, MTW00f, MTW00g, MTW00h, MTW00i, MTW00j, MTW00k, MTW00l, MTW00m, MTW00n, MTW00o, MTW00p, MTW00q, Mac01, Mac06a, Mad01, MWE^{+}04, MPL^{+}04, MDACL05, MCL05, Mal05, Mar03a, Mar05, Mat05, McC02b, McC02a, McC05, McG02, McN01, McN02a, McN04, McN06, Mey01, Mic06, Mil06, Mir05, Moe03a, Mol05, Mor03, Mor05, Mor04, Mül05, Mun04, Nea06, Nem04, Nes05, Neu06, Nol03a, Not03a, Nov05, Nus06, Ogd00a, Ogd00b, Ogd02, Ols02, Ott06, Ozt05, PE00a, PE00b, PE00c, PE00d, PE00e, Pad00, Par00, Par02a, Pea02, Ped01, Peñ06, Per02a, Per02b, PESJ03, PACP^{+}03, PEACL04, PLL^{+}05, Per06, PC06, Pic05, Poo04, Poo06, Pou04, Pra00, PS06c, Pre00, Pre02, Pre03, Qiu05, RLS00a, RLS00b, RLS00c, RVL00, Rac01, Ram05, Rao00, Rao03a, Rat00]. Review [Rei00b, Rei02, Rhe02, Rhe03a, Rie06, Ril02, Rob00, RW00, Rob03a, Rob04, Roc00, Rog02b, Roh05, Rol01, Rol00, Rol02, Ron00, Ron05, Roo02a, Ros04, Ros06, Rou06, Rue04, Ruk03a, Rup01, Rup02a, Rup04, Rus05, Rut00, SRJ00a, SRJ00b, SRJ00c, SRJ00d, Sac03a, Sai06, SC00, SBJRH05, San03a, Sch01, Sch02c, Sch03b, Sch06, Sch00a, Sch05, Sch00b, Sch02e, Sch04b, Sen01, Sey05, She02b, Shi01, Sho05, Sig00, Sin06c, Sor00, Spa02a, Spe06a, Spe06b, Spi06, Sri02, Sta01, Ste05a, Ste00b, Ste06, Ste04, Sto00, Sto04, Str01a, Str00h, Str00i, Str00a, Str00b, Str00c, Str00j, Str00d, Str00e, Str01b, Str01c, Str01d, Str01e, Str01f, Str01g, Str02b, Str02a, Str03b, Str00l, Str02c, Stu00, Stu06, Sun02, Sun01, Sun00]. Review [Sur04, Swa00, Swy02a, TJV00a, TJV00b, TJV00c, TJV00d, TJV00e, Tak00, Tay00, Tay03a, Ter04, The00, Tho02, Tim03, Tiw00, Tra05, Tri02a, Tro06, Van05, Vie02a, Vog01a, Vog01b, Vog01c, Vog01d, Vog02a, Vog02b, Vog02e, Vog02c, Von06, Wan01, Wan02, Wan06, Was01, Weg00, Wel00, Wel01d, Wel01e, Wel01f, Wel01g, Wel01h, Wel01a, Wel01b, Wel01c, Wel01i, Wel01j, Wel01k, Wel01l, Wel01m, Wel01n, Wel01o, Wel01p, Wel01q, Wel01r, Wel01s, Wel01t, Wie06, Wik04, Wik06, Wil03a, Wil04a, Wil04b, Wil06b, Wil06a, Win03, Wol02a, Woo04b, Xie03a, Yan04b, Yan04a, Yan05a, Yan06, Yan04c, Yan05b, Yin04, Ylv01, Yor06, Zal02, Zha06, ZS02a, Zhe06, Zhu06, Zim03, Ziv06, Zwi00, deL05, vdV01, Ala07, AC05c, AC07b, AC07c, AC07a, AC08a, AC08c, AC08d, AC08b]. review [Ann06d, Ann08, Ano04e, Ass07, Ban07, Bar08, Ber04, BT08, Boo04, Boo06, Boo07, Bou07, Bra08, Bri07, Bro07, Bur08, Caf07, Can07, Car07, Cav07, CH08, Che08, Cli07, Cow07, Cra07, Dav07a, Dav07b, Dea08, Del06, Dia07, Elb08, Eub07, Fra08, Fri07, Gal07, Gen07a, Gen07b, Gho07, Gil08a, Gil08b, Gol07, Gor08, Gre07a, Gre07b, Gue08, Hal07, Hal08, Har07, Har08, Hav07, Hey08, Hin07, Hoo07, Kar07a, Kar07b, Kie07a, Kie07b, Kie08c, Kie08b, Kie08a, Kra08, Las07b, Las07a, Las08, Lee08, Li08a, Lin07, Lun07a, Lun07b, Man07, McN07, Mün08, Mye07, Nes07, Nta07, Oun08, Pal08, Par07, Par06, Ree08, Rei08, Rhe07, Rhe08, Rid07, Rob08, Rya05, Sch07a, Sch07c, Sch04a, Sch07e, Sha07, Sin07a, Sin07b, Sta08, Sto08, Str07, Sun08, Teu03]. review [Teu08, Tyl07, Tyl08, Vla08, Wal07, Wat07, Wes07, Wol08, Woo07, Yan08a, Yan08b]. Reviews [Ano00e, Ano01e, Ano01f, Ano01g, Ano01h, Ano01u, Ano02e, Ano03p, Ano04f, Ano04k, Ano04l, Ano04m, Ano04n, Ano05f, Ano05g, Ano05p, Ano05q, Ano05r, Ano06f, Ano06g, Ano06h, Ano06i, Ano06q, Ano06r, Ano06s, Ano06t, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano08b, Ano08c, Ano08g, Ano08h, Ano08i, Ano09b, Ano09c, Ano09d, Ano09e, ACLJ^{+}04, ACE05, DKC^{+}06, Dob03a, Dob03b, HSSM06, KHACM06, LACPE03, MWE^{+}04, MPL^{+}04, MDACL05, MCL05, PESJ03, PACP^{+}03, PEACL04, PLL^{+}05, PC06]. revised [Xie03b]. Revisit [McN02b]. Reweighting [XPLT09]. Riani [Ban06]. Richard [Gen05a, Lun05]. Richardson [Mir05]. Richness [WL05]. Rick [Ros04]. Ridge [AC08d, War08]. Right [Ano02j, AMW02, Gew04, QvdLR06]. Right-Censored [QvdLR06]. Rios [Sac03a]. Risk [AC06b, Ban05, BJM06, Bre06, CHJ08, Dia07, ERS02, FB00, Fol00, Gan09, Gen07a, GS07, Hey08, IRB09, IBPL04, Jen02b, KW02, KSS08, KL04b, LPR01, Lüt03, Mal05, RXS^{+}07, Rhe03a, Rhe03b, Sch06, SPJ01, SS08, Tre06, WM01, Kri05, Loc04, Rhe03a]. Risk-Adjusted [GS07]. Risks [CFK07, ELL09, Nta07, PF09, Rei05, Kie05]. River [NR02]. Rob [Ann06c]. Robert [Chi06, Nus06, Rie06, Rup04, Sey05, Zha06]. Robust [BKV01, CR01, CCC09, Číź08, CLLCL09, CvA02, DMJ^{+}08, De 06, DB03, FW00b, FB00, GR03, GRYB06, HFZ05, HMS00, KVZ07, KL06b, LL07, MRW07, MVY09, MH02, MVF06, MW05, NL07, PH00, PC00, QSZ08, RM09, RFS06, SBVW06, Sim02, Sin04, SD02, VVTR08, WR02a, WR02b, WQ07, WS06, WL07b, Tyl07, Tyl08]. Robust-Likelihood [MRW07]. Robustness [CM09a]. Roger [Jur06, Nus06]. Role [Sch02d]. Romano [Nea06]. Ron [Cla03]. Ronald [Hal03b]. Root [KX04, PP05]. Rose [McC02a]. Rosemary [Mor05]. Rosenblatt [ACS06]. Ross [Bor03, Kor04]. Rossi [Ann06c]. Rotation [PKC01, RBP08]. Rotations [BNV09]. Router [CDVY00]. Rowe [Gen06]. Rudiger [AC06b]. Rudolf [Car03a]. Rue [Yan06]. Ruin [Das02b, Sch06]. Rule [MRŞK06, SCC^{+}06]. Rules [Ano01n, FRS06, FJS08, GR07, MPK09, SC07, UCTW07]. Run [CH06a]. Running [CR00a]. Runs [Gol05a, Kon06, Pai09, Roo02b]. Ruppert [Lee06, Ziv06]. Rustem [Lüt03].

S [AC04e, AC03f, Bra06a, Car05b, Dav03, Jes03a, Lun06, Moe03a, NVP06, Sho05, Ste06, Daw03, Eas03, McN04]. S-PLUS [AC04e, Daw03, Eas03, McN04]. S-Plus(R) [Lun06]. Saar [Ken03a]. Sacramento [NR02]. Saddlepoint [BP02, Str00g]. Safe [OZ00, TDGW01, TL02]. Safety [ST02]. Sahai [Boo06]. Salmon [NR02]. Saltelli [Per06]. Sample [ADW07, CN02, CC09a, De 06, EL04, EM05, ES02, FS06b, FM06, Hua00, KL06a, LSC04, LHNM08, MJW06, MPRR04, PKC01, PDG02a, PDG02b, Sea02, STJ07, SW02b, SC03, SJ06, Ver02, Ron05, Tra05]. Sampled [JS03]. Sampler [LKF05, LCL09]. Samplers [HH07, IZ07, KJ07, KF07, TLW07b, TLW07a, vDP08]. Samples [Efr01, GDH09, HRS09, KW01c, LM09b, MRŞK06, Ozt02, PM05, SS04, Yat05]. Sampling [Ano02j, AMW02, BAW08, CCB03, Gel00, GM06, IJ01, KNF06, Kje00, Lia02, Lit04, LLW00, Mee00, MR05, Ned09, OZ00, Ozt05, PS07, QLS02, RRS07, Sch07e, SSY07, SO04, TEB00, WZ05, Gor08, Mün08]. Samuel [Bei04, McN06, Sur04]. Sandwich [KC01]. Sankar [Rou06]. Sanpei [Bai04]. Santner [Mor04]. Saralees [McN06]. Sarcoma [BT04]. Särndal [EP06]. SARS [KT09]. SAS [Bas03, Cla03, Ste06, Cle03b, Cle03a, Neu06, All06, AC08b]. Sasha [Gan03a]. SAS(R) [Dur06]. SAT [Han04]. Satellite [LLW04, SYCB08]. Satterthwaite [BP02]. Saying [vSH08]. Scale [CDK05, Efr04a, Efr07, Efr09, Epp04, FW07, HP04, HRV04, JC07, LSQ03, MZ08, MRT05, McC04, MMH^{+}04, MM04, RGA01, Ser04, SWP00, XSC08]. Scales [LC09, ZMAS05]. Scaling [Bou07, CB09, JM06, OR01]. Scan [CZ07, HTZ^{+}09, Jam02h]. Scanning [HY01b]. Scans [Roo02b]. Scatterplot [Ber00a]. Schabenberger [Dan03, Zhu06]. Schafer [Iha06]. Schedules [GL06]. Scheinkman [GGR07]. Schemes [CBCS07, Mor05, ZS01]. Schölkopf [Wil03a]. School [BFHR03a, BFHR03b, BX03, KZ03, MJB03]. Schools [CTG^{+}06]. Science [BGM^{+}05, Che08, Cle02c, Cli07, GC09b, Gol02c, Kra08, Lin04d, Sta08]. Sciences [AC03d, Bre00, Gre00, Sob00, AC03a, Bei04, Dan03, Lar06, Len04a, Bri07, Cav07]. Scientific [GM09, Par06, Per06, Sha07]. Scientist [Van05]. Scientists [Efr05b, Len03a, Len03b, Ann06a, Gil08b, AC08c, Sun08, Wes07]. Scope [KS00]. Score [CT04, IV04, NS02, RL01, RT00]. Scores [DR00, Tan06a]. Scoring [GR07, VWM01]. Screening [BP01, BP08, CN08, MB09, RXS^{+}07, SPJ01, SCC^{+}06, Wan09]. Sea [Hig09, LS09a, LS09b, Men09, WM09b]. Sean [All06]. Search [Fre04, HDW07, VPW08, Ban06]. Searching [XLB04]. Searle [Cow06]. Seasonal [KOC07]. Second [GWB08, Gua09, VWNM02, AC03c, AC04e, Ber04, Boo04, Bor03, Car05a, Dav03, Hal03a, Lan03, Len04a, Rog02a, Rya05, Sto04, Tim03, Wol02b, Xie03b, Yan04b]. Second-Order [GWB08, Gua09, VWNM02]. Secondary [HPC03, LH09]. Seeding [Har03]. SEER [GBT09]. Segmentation [KJG05, Lee00, QS07]. Segregation [ZFZ03]. Seismic [EOU02]. Seismology [VS08, VJ00]. Select [YW03]. Selected [BYE^{+}05, BY05, Edw05, HY04, Hol05, PS07, Sha05, SGL07, Tam05, Wes05]. Selecting [TSE02]. Selection [Ano06l, CM06a, CK08, CK09, CD09, CL03, CKN08, DTT08, DHE08, FL01, FL04, FG04, FS04b, FD08, FJS08, GG08, Geo00, GG07, GI07, HY01a, Har03, Hof07, HC07, IZT08, Iri01a, IJS01, IR03a, IR03b, IR05, Joh03, JLZ08, JKKW03, KL06a, KL04a, KC07, KVZ07, Kim00, LNR03, LB01, LPM^{+}08, LL09, MPE08, ML02, MW05, Pen00a, Pen00b, RD06, SR01, SY02, SD03, SHY04b, SH06, SF07, TSV05, WW00a, WW00b, WW00c, WLH08, WL00, WBS07b, YL05, ZWL^{+}04, ZS01]. Selective [LFY^{+}08, Lo00]. Self [HID09, KS02, KS08, TPO03]. Self-Carryover [KS08]. Self-Consistent [TPO03]. Self-Rated [HID09]. Semi [Ano02n, DP06]. Semi-Parametric [Ano02n, DP06]. Semicontinuous [JvD03, OS01]. Semilinear [Cra05, FPHR05, FPH^{+}05, HZ05, HWZ05, KM05, Sab05, Tib05, Zha05a]. Semiparametric [BLW01, BD09, BHO07, BCM02, BD05, CCW02, CP02, Che01, Che02, CH03, Che04, CFT06, CJ06b, CKN08, DP03, DDH04, Dun05, FL04, FHL07, FW08, FKW01, GS01, GT07, GBT09, HW01a, HL04, ISZ07, IT02, JRY04, JFY07, JLZ08, JM04, KW01a, KW01b, KG01b, LB01, LL06c, LZ01, LY01a, LC01a, LY01b, LC01b, LWY01, LS09c, LZH09, MGT05, MC06, MMC07, MK06, OLT09, PMHB09, Pen00a, Pen00b, PPR09, Ram01b, She02a, SNQ09, SMG^{+}05, Spe06a, SPSZ05, WW00a, WW00b, WW00c, WC01, WW01, WLH04, WCL05, XSC08, XLL04, ZYI06, ZL07b, Zuc05, Lee06, Tro06]. Sen [Mic06]. Senn [Sto04]. Sensitivity [HRS09, HY01b, RS09a, RS09b, SGL07, SRA08, Sma07, Par06, Per06]. Separation [Gen06]. Sequence [GL03, Ros04]. Sequences [MPE08, SWDS03]. Sequential [CDHL05, Che07b, DJSW09, Dia03b, DS08, DB09, Gan09, LA08, PF08, TEB00, TRH07, Dia03a]. Sequestration [BHO07]. Serially [PT09]. Series [Ano04e, AC03f, Bhl02, BC03b, Bre05, BRS^{+}00, Büh02, BS06, CFY00, Cav04, CP06, CR00b, CHL06, CGR04, Cor02b, DLRY06, DDH04, Els02a, Esc06, FS04a, FO09, Fue03, GPT06, GDW04, Hal02, HS07a, Har07, HCS06, HOS04, HZZ^{+}08, IK08, Lah02, LW01a, LS02, Li08b, Lun05, Lun07b, McN04, ORvSM01, OvSG05, PP05, PMHB09, Par02b, PR02, PS05, PW05, Pol02, SWP00, She02d, Sol00, SFPM04, Tri02b, Tsa00, VR00, VWM01, ZYDS08, Gal05]. Serologic [FW05]. Server [Li05]. Service [AC08c]. Set [FM06, LPR01, LH07, Ozt02, Ozt05, PWM01]. Seth [Wil06b]. Sets [Dem04, IYSK08, KSN08, Kie08a, Lav04, Lin04c, SBL^{+}04, SB04, Zad04, Kie08a]. Several [LJZ04]. Severity [SRA08]. Sexual [GT09]. Shadish [AC05a]. Shah [Not03a]. Shaikh [AC06c]. Shape [ADW07, CV07, KS07a, KS07b, Mar07, Mic06]. Shapes [Jam02f]. Shared [LS02]. Sharing [KP05]. Shawe [Las06]. Shawe-Taylor [Las06]. Shayle [Cow06]. Shein [Xie03a]. Shein-Chung [Xie03a]. Sheldon [Kor04]. Shiboski [Lyl06]. Shmulevich [Gen05b]. Short [Lin03b, TBJ08]. Short- [Lin03b]. Short-Term [TBJ08]. Shotgun [HDW07]. Shoukri [Sch04b]. Should [Cau01, Gro01, Gus09, Hor01a, Hor01b, Mag01, RK01]. Shrinkage [CCC09, RJ08, TW07b, WX09, Yat05]. Sided [FHLR05, LL04, LHN08]. Sieve [Bhl02, Büh02, CZI07, Hal02, Lah02, Pol02, She02d, XLL04]. Sigma [AC04d]. Sign [LL04, Ran00]. Signal [Ano05k, Can07, CCDP09, Gen06, Rie06, SHC02]. Signals [BH07, Iri01b, LS01, TW07a]. Signature [McK05]. Significance [CH06a, Efr07, Ell00a, Len06, LHNM08, PZB09]. Sijtsma [Hab03]. Silvapulle [Mic06]. SIMEX [DH08, LL03b]. Similarities [LHN08]. Similarity [SWS^{+}03]. Simon [Gen05a, Sac03a]. Simple [BKV01, GS07, KPY03, Len04a, LH07]. Simpler [Ran00]. Simplex [Liu00]. Simulation [Fre04, Liu03, MSD04, NS02, PR00, Sch05, Wie06, Pic05, Dia07, Fra08]. Simulations [Bal05]. Simulator [CHB^{+}08]. Simulators [CGRS01, Gew04]. Simultaneous [CS09, CBL^{+}08, Efr04a, Efr07, ELL09, FHY07]. Singapore [KT09]. Singer [And05, Nus06]. Single [CCDH08, GDH09, RTD01, YR02]. Single-Channel [RTD01]. Single-Index [YR02]. Single-Nucleotide [GDH09]. Single-Photon [CCDH08]. Singular [Hof07, HWZ06, HHCK09, HSB09]. Sinha [Ozt05, Yan04a]. Siniha [Not03a]. Sinusoidal [SL06]. SIV [RSH^{+}02]. Sivo [All06]. Six [AC04d]. Sixten [EP06]. Size [Bei04, CC09a, De 06, DR05, LHNM08, MJW06, MPRR04, SS09, STJ07, SC03]. SiZer [HM06]. Skeletal [CLLCL09]. Skew [DM04, MGT05]. Skew-Elliptical [MGT05]. Skewed [FS06a]. Skill [CR08]. Skinner [Fie06]. Skrondal [Cor05]. Slab [IR05]. Sleep [CCDP09]. SLEX [HOS04, OvSG05]. Slice [FH00]. Sliced [BC01, WX08, ZMP06]. Small [CC09a, Efr01, NC02, PT06, PS07, PS06c, WF03, Dat04]. Small-Area [PT06, PS07, PS06c]. Smith [McC02a, Bet03]. Smoking [LDM09, LCLC08, PPR09, PGLW00]. Smola [Wil03a]. Smolts [NR02]. Smooth [CX08, CX09, CC09b, DSV04, GG07, Gus00, HLS05, YK04]. Smoothed [Hel07, Szk03]. Smoothers [Ber00a, KE02]. Smoothing [AM01, Ano01p, BCR02, CRW01, DN03, DH08, FEG07, GWKK01a, GWKK01b, GDW04, GDOvS03, KW04, KK07, QS07, RSW09, ST04, SW05b, Whi01, Woo01, Woo04a, YW01, Zha03]. Smoothing-Parameter [DH08]. Smoothly [KCO08]. Smyth [San03a]. Sneh [Lin04b]. Social [Gil08b, GC09b, HRH02, HGH08, Lar06, Len04a, Sir06, Sob00, Van05, ZSG06, Ala07].

Sociology [Raf00]. Soft [BT04]. Software [Iha06, Ste01, Gil08a]. Soil [Dan03]. Soils [BHO07]. Sojourn [Hua00]. Solomon [Kac04]. Solution [CHL06, Mee03, YZ09]. Solutions [Dav07a, Jam02g, Man07]. Solved [Emb03]. Some [CT04, DV03, LFGR04, PV06, RW00, Tho00, Tra05, WD00]. Songs [BF01a]. Sophia [Cor05, Daw03]. Sound [Iri01b]. Source [Gen06, LSGS01, McB02, WM01]. Sources [ZLI^{+}09]. Space [AC03f, BC03b, Gne02, GLW^{+}06, HRH02, Hoo07, KSZL02, LGS07, PSW05, PT06, SGH07, Ste05b, SMP03, VS08, ZOVJ02, dV04]. Spaced [Fue07]. Spaces [BDZ06, LLZ07]. Spall [Fre04]. Spanning [Liu03]. Sparse [CCL^{+}08, JNKD03, PWZZ09, TW07a, YMW05]. Sparsely [JS03, LM09b]. Sparsity [JL09a]. Spatial [Ano05k, BGS03, BG06b, BH07, BIW00, CA02, CNN07, CDR08a, CDR08b, DN03, Dig05, Dun08, EL04, Fue07, GKSB03, GKM05, GY02, GSC04, Gua06, GL07, Gua08, GWB08, Gua09, HSG02, HTZ^{+}09, JNKD03, JKN08, KSN08, KW02, KSZL02, KMH05, Kul06, NC08, Pac08, RHC07, Sai06, Sch05, SHC02, SF07, SCG^{+}07, SWDS03, Tho08, THC05, Zhu06, ZRM^{+}08, Bog04, Sey05]. Spatial-Prediction [THC05]. Spatial-Temporal [ZRM^{+}08]. Spatially [Ano03i, BDS01, GKSB03, HT04, LL06c, MCG00, SO04, ZS01]. Spatio [Ano01m, Hig09, LS09a, LS09b, Men09, PM09, WM09b, XWF05]. Spatio-Temporal [Ano01m, Hig09, LS09a, LS09b, Men09, PM09, WM09b, XWF05]. Spatiotemporal [AMvW^{+}09, Bow07, NN09, PLM08, WMNB01]. Special [BP01]. Species [BGF01, DR05, Mao06, WL05, ZRM^{+}08]. Specific [DC03, LTMS02, PCK06]. Specification [GC09b, Kol02b]. Specifying [MK06]. SPECT [SCG^{+}07]. Spectral [CGR04, GDOvS03, ISZ07, RSW09]. Spectroscopic [BFV01]. Speeds [Fry05]. Sperlich [Spe06a]. Spike [IR05, JFY07]. Spizzichino [Cha03a]. Splicing [HHCK09]. Spline [AM01, Ano01p, CRW01, GWKK01a, GWKK01b, GDOvS03, JRY04, KW04, KK07, WLC02, Whi01, Woo01, YW01, YR02, ZYZS03]. Splines [BCR02, CD07, HM03b, LZH09, RSW09, ZS01]. Split [FSL^{+}06, HRS09, Rob04]. Split-Plot [Rob04]. Splits [KL01]. Splitstone [Poo04]. Sports [Loc02b]. Spot [KOC07]. Spotted [QS07]. Springer [McC02a]. Springer-Verlag [McC02a]. SPSS [Ste06]. Squared [BC01, Dui16, WF03, Yat05, Zha05b]. Squares [RO07, VWM01, WL07a]. SSA [Els02a]. Stability [Ros01, Kar06a]. Stable [Woo04a]. Stage [Cha04, HTO03, KSG05, LPP02, MK06]. Stage-Frequency [HTO03]. Standard [AD08, CCL09, FZ09, GW05, IPC08, Rub08, SCS08b]. State [AC03f, BC03b, FFJ07, LCLC08, PT06, SRRW00, SMP03, dGKS01, dV04]. State-Dependent [SMP03, dGKS01]. State-Domain [FFJ07]. State-Space [AC03f, BC03b, SMP03, dV04]. Stated [Dea08]. Stateline [GLW^{+}06]. Stationary [Ano01m, Bhl02, Büh02, CP06, Gne02, Hal02, Lah02, Lin03b, PW05, Pol02, She02d]. Statistic [HTZ^{+}09, OS00]. Statistical [AC04a, AC05c, AC05b, AC06a, Ban07, BH08, BGF01, Bor03, Bre06, BB03, Bro00a, BGM^{+}05, Car05a, Cav07, Cha03b, CDHL05, CZC09, Cle06, Cor02b, Cow03b, Cow03a, Dan03, DMJ^{+}08, DM00, Els02c, Eub07, Eva00b, Frö08, GKO05, Gil08a, Har08, Hin07, Iha06, Jam02f, Jes03b, Kal00, KM02, Kar06c, KJG05, Kol02b, Las08, Laz07, LPL06, Lee08, Leh05, Len06, Li07, LHSY02, LHNM08, LCT^{+}05, Mac06a, Man07, McK05, Mee03, Mic06, Moe03b, Mun04, Nea06, NDKW01, O’F01, OC06, ORvSM01, Par07, Peñ06, Qiu05, Rab02a, Rab02b, Rie06, Roy00a, Roy00b, Rup04, Rya00, Sac03b, ST04, Sch02a, Sch02b, Sch03c, Sch07d, STJ07, Sen08b, SWP00, SRRW00, Szk03, TH09, Teu03, TBJ08, Tyl07, Van05, VVTR08, VJ00, Wei00, Wil04a, Wol00, Won00a, ZZIP07a, ZZIP07b, ZQ09]. Statistical [dGKS01, Len04a, AC03c, AC04e, AC04d, AC06c, Bei04, Boo04, Bri07, CH08, Dob04, Elb08, Hal06, Hoo07, Jes03a, Kar07b, Kie06, Kim06, KL05, Kri05, Lun07a, Moe03a, Nem04, Poo04, Ros06, Rus05, Sac03a, Sch07a, Sch03b, Sch05, Sin06c, Zhe06, Zhu06]. Statistician [Ell00a, Sch07b, PS06c]. Statisticians [Dob02b, Ann05b]. Statistics [ADW07, AC04c, Ank05, Ann06d, Ann06c, Ann08, Bai05, Ber04, BD03, Boc08, Bre00, CZ07, Che08, Dav07b, Dev06, Dig05, Dix06, Eas03, Efr00, FHW00, Gas03b, Gen06, Gia00, Gil08b, Gol02c, Gua09, GH00, Gut00, Imh03a, Imh03b, Jam02h, Ken03b, Lac09, Las08, Law00, Mar03b, Mas04, McC02a, NHS00, Nol03b, PC00, Raf00, RTW00b, Rog02a, RA00, Sch02d, Sho05, Spi06, Sta08, Str03a, Tho00, Tyl08, WSLC06, Wat07, Wes07, Wik04, Wik06, Wil06b, vdW02, Bar08, Gas03a, Jon06, Lee03, Mar03a, Ott06, Str03b, Dav06c, Gre07a, Ken03a, Lei05, Nol03a, Rid07, Spe06b, Ziv06]. Status [BP08, CN08, DW04, She00a, XLL04]. Stefan [Mun04, Spe06a]. Stefano [Per06]. Stein [Dur06, Str07, KW01c]. Stein-Type [Str07, KW01c]. Stellar [LSQ03]. Step [CH02, HC02]. Stepdown [RW05]. Stephen [All06, Dav03, Duv03b, Few04, Lue05, Lyl06, Sto04]. Stepwise [Che07b]. Stereological [BCS07]. Stereology [Ann05b]. Sterility [DZ00]. Steven [Ann06d, Eas03]. Stewart [Zim03]. Stick [DXC08, IJ01]. Stick-Breaking [DXC08]. Still [CR00a]. Stochastic [AKS02, Bau02, Bin06b, CM06b, DSV04, EM05, FD08, Fre04, FO09, Gan03a, Gan03b, GL03, HDW07, HBK^{+}09, Hös03, Hss03, Kie05, Kou08, LW01a, LLC07, Lun03, Lun07a, LCLC08, Mir05, Nov05, NS01, PY00, Rhe07, Rhe08, Str00f, Vie02b, ZOVJ02, Cli07, Fri07, Kie08c, Rya05, Yin04, Yor06]. Stochastics [Kor05]. Stock [Ano02o, Mal05]. Stokes [Ste06]. Stop [GFK07]. Stop-and-Frisk [GFK07]. Stopping [Str00f]. Stops [GR06b, RM09]. Stoyan [Kie05]. Strata [LBS06]. Strategic [Len02f]. Strategies [DHS08, Rao03a, Rao03b, TSE02]. Strategy [DM00]. Stratification [BFHR03a, BFHR03b, BX03, FBV^{+}04, JR08, KZ03, MJB03, ZRM09]. Stratified [KNF06, Kje00]. Stratospheric [Mei07]. Streaky [CR08]. Streams [LL06b]. Strength [Sur04]. Stress [Hey08, Sur04]. Stress-Strength [Sur04]. Strictly [GR07]. STRIDE [Cau01, Gro01, Hor01a, Hor01b, Mag01, RK01]. Stroke [PAS^{+}03]. Structural [CJS02, Cor05, DLRY06, FG04, HSS01, HBR01, McN07, SSV05, SBJRH05, Ste01, WA00, WS05]. Structure [BF09, Els02a, JRY04, LSQ03, ZSG06]. Structured [JHISR07, LLC00, Mir05, OZ08]. Structures [EW08, FWP05]. Stuart [Ann06d, Spe06b]. Student [Ald08, DL08, Edw08, FHY07, MH04, Sen08a, Sti08, Zab08b, Zab08a]. Studies [AD08, All06, BMDD07, Bra08, Cha04, CSCC06, CCC09, DBPP08, Eas03, GL06, HRS09, KL04b, Li06, LZS^{+}06, LZ06, MJW06, MK06, OMS^{+}09, Ric04, Ric08, Ros02, RS09a, RSM^{+}07, Sab06, SAE06, SYCB08, SMG^{+}05, Sob06, TR06, WH07, WM09a, Wol02b, Wu02, Xie03b, YTS08, vdLHR02, Xie03a]. Study [BFHR03a, BFHR03b, BX03, BIK00, CCDH08, CLQ03, CHRH01, Dav06a, FL09, GH08, Gel08, GMK^{+}04, GC09a, Gre03, Han04, HHS06, Hod08, HR06, HQF08, IK08, Kat08, KB04, KZ03, LZHR01, MMZ02, MAC^{+}06b, MJB03, OC06, RRS07, RS09b, SSVV09, STPN09a, STPN09b, SI02, SRL00, WL00, ZWT^{+}05]. Study/Validation [SRL00]. Studying [LFY^{+}08, SME03]. Stulajter [Fue03]. Sturdy [GP07]. Style [Len04b]. Subgradient [NN06]. Subject [AKS02, HSY06, LFGR04, MTR^{+}01, PT06, RFP09, YL01]. Subjective [Kad07, Mil07, RLL07a, RLL07b, Wol07, Cha03a, Sho05]. Subjects [MMZ02]. Submodels [DP05]. Subsampled [GM06]. Subsampling [EL04, ELL00b, GSC04, HL00]. Subset [SGL07]. Subspace [CL08a, Liu03, ZZ06]. Subspace-Spanning [Liu03]. Substantia [HBK^{+}09]. Sue [AC05d]. Sufficient [CN05, CF09, LL08]. Sullivan [Qiu05]. Sum [DS05, MRW07]. Summaries [MMC07]. Summarizing [HT04]. Sums [Fel02c]. Superefficiency [Ber00a]. Supervised [BHPT06]. Support [LLW04, WS07, Wil03a, Wil03c, WL07b]. Suppression [FG00]. Supremacy [Net09]. Surface [FW05, Hig09, LS09a, LS09b, Men09, Qiu07, WMNB01, WM09b]. Surfaces [AC08d]. Surplus [FKS02, LFKS03]. Surrogate [CLQ03]. Surveillance [DHS08, KPY03]. Survey [AHA01, BST03, CN02, EL00, ELL00b, ES02, EL05, FW05, Fie06, KSS08, Lo00, MRŞK06, MR05, OZ08, PDG02a, PDG02b, PS06c, QLS02, SRC^{+}06, Sea02, SW02b, SS08, Tra05, Ver02, WS01b, ZSB01, Nus06]. Surveys [EP06, FIC01, LR00, LR06, Mer04, MTR^{+}01, RXS^{+}07, ZSB01]. Survival [BB01, BBN^{+}04, Bet03, BMDD07, BFB00, CFJZ07, FSL^{+}06, FWP05, HSG02, HBF06, IPBS00, KG01a, LL03b, LR04, LL06c, Lyl06, MK06, NR02, Oak00, OLT09, PH08, Rao03a, Rao03b, SC03, Sin07b, SZ09, TIY03, Yan04a, Yan05a, YF05, YTS08, ZYI06, Zuc05, vdLHR02, Yan04b]. Survivors [UCTW07]. Susceptibility [IPBS00]. Suslina [Pou04]. SV40 [LFY^{+}08]. Swishchuk [Kar06a]. Switching [FS01, GLW^{+}06, Kie08c]. Sylvia [Mir05]. Symbolic [BD03]. Symmetric [BNV09, JP05]. Symmetries [Gol07]. Symptoms [ZYDS08]. Synchronization [LM04]. Synchrony [FGM^{+}08]. Syndrome [WT01]. Synergism [XTSY01]. Synthesis [DMJ^{+}08]. System [KP05, NR02, Ste06]. Systems [ABPR06, CHL06, CGRS01, GR06a, Kar06a, Kie08c, LSGS01, LSGS04, deL05, Mir05].

T [AC05a, AC06a, Bog04, Few04, Mar05, Nol03a, PS06c, Tim03]. Table [Jam02e]. Tables [BFD01, CM09a, CDHL05, Dob04, Fie00, Lan05, MOJ05]. Tabular [FG00]. Tactile [SME03]. Tadeusz [Bai04]. Tail [AKM01, WT09]. Takezawa [Jen06]. Tale [ZMAS05]. Tamoxifen [SQC07]. Tampering [DK01]. Tamraparni [Kar06b]. Tanizaki [Bai05]. Tankov [Bin06a]. Tanner [Nol03a]. Tao [Gen04]. Tapering [KSN08]. Tarantola [Per06]. Tau [MB05]. Tauberian [Kie07b]. Taylor [Las06, VWM01]. Teams [Har03]. Technical [KT05]. Technique [Fry07]. Techniques [AC06b, Can07, Dia07, Els02a, GT07, Rob03a, Rob03b, Yan04b, Bas03, Rei08]. Technological [HY07]. Telegraphic [Ano01u, Ano03p, Ano04k, Ano04l, Ano04m, Ano04n, Ano05p, Ano05q, Ano05r, Ano06q, Ano06r, Ano06s, Ano06t, Ano07f, Ano07g, Ano07h, Ano07i, Ano08g, Ano08h, Ano08i, ACLJ^{+}04, ACE05, DKC^{+}06, Dob03a, Dob03b, HSSM06, KHACM06, LACPE03, MWE^{+}04, MPL^{+}04, MDACL05, MCL05, PESJ03, PACP^{+}03, PEACL04, PLL^{+}05, PC06]. Telephone [BGM^{+}05]. Temperature [Hig09, LS09a, LS09b, Men09, WM09b]. Temporal [Ano01m, Hig09, LS09a, LS09b, Men09, PM09, WM09b, XWF05, ZRM^{+}08]. Temporally [PGH01]. Tennis [KM01]. Tensor [ZCI^{+}09]. Tensors [Laz07, Sch07d, ZZIP07a, ZZIP07b]. Teratogenesis [CHRH01]. Terje [Ban05]. Term [JRY04, TBJ08]. Termination [CLLCL09, SMIO08]. Terrence [Car06]. Test [BP08, BC01, CN08, CS03a, Dou06, GK07, GGR07, GSC04, HZ03, HS02b, KPD02, LL04, PR02, PB04, Ran00, Str07, TDGW01, TL02, WL00, vSH08, vWvdWvdV08, Dra06, Bou07]. Testing [AC08c, Ano02o, Ano05k, BG01, BKV01, CS09, CS04, DN07, DW04, Efr04a, Efr07, Els02c, FTR04, FS04a, FZ09, FKC09, GG08, GR06b, Jam02e, KM02, Kim00, LH09, Lob01, Mar00b, MB05, MOJ05, MPRR04, Net09, PT07, PT09, Pou04, Rab02a, Rab02b, RW05, RFF04, Sch02a, SHC02, Swy02b, WA06, WT07, XTSY01, YPXH06, Hey08, Kaf03, Mun04, Nea06]. Tests [Aba02, ASFP09, AP04, AKM01, Ann05a, BP08, BZ00, BPB02, BF01b, BFB00, CM09a, CT04, CN08, CCL09, CL08b, DS05, Esc06, FH01, FHY07, FZ09, FOD07, Gan09, Gew04, HS07a, HSS01, KL06b, Kul06, Len06, Lin03b, Pai09, PP05, PF08, Qiu05, SY06, TOR05, WQ07, WL00, WS01a, Len04a]. Texts [Ron05]. Their [Ano02n, BLW01, CHL03, FKW01, GS01, KW01a, KW01b, LZ01, Lin04a, McN06, Pui03, Ram01b]. Theodore [Cav06a, Kar06b]. Theorems [Kar06a, Kie07b]. Theoretic [Mee00, Soo00, SJ03, Tay03a, Tay03b]. Theoretical [FPH^{+}05]. Theory [AC07a, Ank05, Ann06d, Ann08, Ass07, Boo06, Bro00a, CLR06, Cor02b, Dea08, EM08, Gol05a, Gre07b, HM06, Kar07b, Kie07a, Kie08b, Kie08a, LSC04, LLW04, Len02g, LLW01, LPP02, McC05, Ozt05, Par02b, Pre03, Ree08, RY00, Rya05, Sac03b, Sch07a, Sch07c, Sey05, Ste01, Stu06, Sur04, SCC^{+}06, Tra05, Tyl08, Vog02d, Von06, Wol08, Yan06, Yan05b, Emb03, Hab03, Mor03, Sac03a, Ste04, Bre06, Kie07a, Str07, Yan08b]. Therapeutic [TSE02]. Thick [FH00]. Thick-Slice [FH00]. Thinned [GL07]. third [Leh05]. Thode [Kaf03]. Thomas [AC05a, Len06, Mor04, KL05]. Thoughts [Ber00b]. Three [Ano00g, BNV09, CH02, HC02]. Three-Step [CH02, HC02]. Threshold [DV03]. Thresholding [ELL09, JB01, KS00, RLZ09]. Thresholds [LL06a, LL06b]. Throught [Mac06a]. Thy [CRY09]. Tibshirani [Rup04]. Ties [XSZ09]. Time [Ano04e, Bhl02, BC03b, Bor03, Bre05, BRS^{+}00, Büh02, BS06, CFY00, CCW02, CDVY00, Cav04, CP06, CR00b, CHL06, CW08, CW01, CGR04, CK04, Cor02b, DLRY06, DDH04, ERS02, Els02a, Esc06, FFJ07, FS04a, FD03, FO09, Fue03, GPT06, Gal05, Gen00a, Gen00b, GT09, GLW^{+}06, GDW04, GB00, GL06, GDOvS03, Hal02, HS07a, Har07, Hay00, HYFW05, Hig00, HCS06, HL04, HW03, HOS04, HW04, HZZ^{+}08, IK08, Kle06, KL08, KT09, Lah02, LW01a, LPS01, LS02, LGS07, Li08b, LTE08, LM04, Lun05, MG00, MGG^{+}07, McN04, Mei07, ORvSM01, OvSG05, PP05, PKC01, Par07, PMHB09, Par02b, Pen00a, Pen00b, PR02, PS05, PW05, Pol02, Raz00, RAHT06, SGH07, SWSM09, SWP00, SNQ09, SGL07, She02d, Sol00, Ste05b, Str00f, SFPM04, TZW05, Tri02b, Tsa00, VS08, VR00, WW00a, WW00b, WW00c]. Time [WT01, Wor00, YKL^{+}06, ZL07a, ZMAS05, ZYDS08, ZCPL08, Gne02, Hoo07, KSZL02, PSW05, ZOVJ02, AC03f, Lun07a, Lun07b]. Time- [FFJ07]. Time-Course [Gen00a, Gen00b, Hay00, Hig00, MG00, Raz00, Wor00]. Time-Dependent [GDOvS03, HCS06, SWSM09, ZCPL08]. Time-Series [Tri02b]. Time-To-Event [ERS02, CK04, GB00, Kle06, SGL07]. Time-Varying [CDVY00, CW08, KT09, LTE08, RAHT06, TZW05]. Time-Warped [LM04]. Timed [WC04]. Times [BB76, BH08, BAW08, Hua00, HZAG01, IP02, MD01, MB05, SPSZ05, SSL07]. Timing [BBK03, LPS01]. Ting [Dui16]. Tissue [BT04, RSH^{+}02]. Today [Ber00b]. Toeplitz [CHL06]. Tolerance [CH06b, KSS08]. Tolerated [Che07b]. Tomography [XMN06, CCDH08, CDVY00]. Tomorrow [Ber00b]. Tools [AC04e, Bas03, LHSY02, Poo04, AC06b]. Tooth [FSL^{+}06]. Topics [Gen04, Not03b, Kor04, Not03a]. Total [LL07, MSZ02]. Tour [CR08]. Tourangeau [Nus06]. Toutenburg [AC03c]. Toxicities [BT04]. Toxicology [FB00, Rya00]. Track [Gre07a]. Tracker [DM00]. Tracks [MSM05]. Tractable [BNV09]. Traditional [Len04a]. Traffic [CS00b, GR06b, QA09, SjKH^{+}01]. Train [Bal05]. Training [And08, YL01, ZRM09]. Trait [Dun03, JFY07]. Traits [LCT^{+}05, ZFZ03]. Trajectories [WBBR05]. Transcription [JL08]. Transdimensional [Sis05]. Transform [CHL06, HTO03]. Transformation [CW01, FTW01, LL06c, LWY01, MY08a, MH07, SW05b, SNQ09, ZYI06, ZL07b]. Transformations [MVY09]. Transformed [SZ09, YZL08, ZYI05]. Transition [ASFP09, HL06b, ZC07]. Transition-Based [ASFP09]. Transitional [BRS^{+}00]. translated [Teu03]. Translocation [SW02a]. Transmission [AALT00, CTG^{+}06, HR00, KT09]. Transplant [FWP05]. Treating [TEJ^{+}04]. Treatment [Aba02, Ala07, CLLCL09, DN07, Fol00, FBV^{+}04, Frö08, GB00, GL06, HR03, HQF08, IV04, LPL06, LLL08b, LA06, MB09, Ros01, Ros02, RRS07, RS08, RL02, SWSM09, YZ05]. Treatments [HBR01, KS08, STR08]. Tree [Han06, LLC00, LPD00]. Tree-Structured [LLC00]. Treed [GL08]. Trees [FSL^{+}06, HJ02, KL01]. Trend [Hig09, HLS05, LS09a, LS09b, Men09, RFF04, TBJ08, WM09b]. Trends [Kad07, Mei07, Mil07, PM09, PPR09, PGLW00, RLL07a, RLL07b, SGH07, Wol07]. Trevor [Rup04]. Trial [Bak00, BT04, Fol00, GSCR04, Har00a, HR00, LDM09, SQC07, STR08]. Trials [BF01b, CH08, Dmi05, Dur06, IWM04, LA08, LLL08b, SC03, Sin06c, Sto04, TEJ^{+}04, TSE02, YY09, Lin03a, Ass07, Pal08]. Trim [DT00]. Trimmed [ÁEdBCAM08, SHH00]. Tropical [CDR08a, CDR08b, Dun08, Pac08, Tho08, WMNB01]. True [IWM04]. Truncated [AMP06, WL07b, Yan04b]. Truncation [MB05]. Truth [Spa02b]. Try [LLW00]. Tufte [Len04b]. Tumor [PDK07]. Tumors [DFS02]. Tutz [Cha03b]. Twisk [Hu04]. Two [BFD01, BP08, CM09a, CCB03, Cha04, CN08, CCL09, CL08b, EHI08, FM06, Hua00, HWZ05, HSB09, KNF06, KSG05, KT09, KS08, LPP02, LHN08, MK06, OS01, PCK06, RXS^{+}07, RHC07, RZF06, TIY03, ZMAS05, ZSCW06]. Two-Component [EHI08, TIY03]. Two-Part [OS01, RZF06, ZSCW06]. Two-Phase [CCB03, KNF06]. Two-Sample [FM06, Hua00]. Two-Sided [LHN08]. Two-Stage [Cha04, KSG05, LPP02, MK06]. Two-Way [BFD01, CM09a, HWZ05, HSB09]. Twofold [KST01]. Type [Dui16, KW01c, PF08, Str07, VS08, Zha05b].

U [Lun03]. U.K. [GLPS03]. U.S. [BB03, GT07, IK08, LSGS01]. Ullah [Ron05]. Ultra [Wan09]. Ultra-High [Wan09]. Unascertainable [Ano03g, HW00]. Unbalanced [Boo06, EHI08, Fry07, GA05]. Unbiased [KL01, MRT05, WC04]. Uncertainties [GLPS03]. Uncertainty [Gol02c, NC02, STNM09]. Unconditional [Ano02j, AMW02]. Uncontrolled [Gre03]. Uncorrelated [FRZ05, Lob01]. Understanding [Dev06, FHW00, MWL^{+}01, Nyc00]. Undervotes [Han03]. Undesirable [FKS02, LFKS03]. Unemployed [YL01]. Unequal [JM06]. Unfolding [JR07]. Unified [CH08, GA05, HT06, JFY07, LZ08, LPP02, MB02, MOJ05, WL07a]. Uniform [CFN01, PF08]. Uniformly [Gan09]. Unifying [LCT^{+}05]. Unimodal [DFWJ06]. Union [RY00]. Unit [KX04, PP05]. Units [Ros07, RS09b]. Univariate [Efr05a, FS06a, SJ04]. Universal [HY04]. Universe [LSQ03]. Unknown [DV03, JM06, LJ00, PT07]. Unmixing [WS05, Gen06]. Unsupervised [Ano01j]. untitled [Alb00, Alb02, AC01b, AC02a, AC02b, Ank01, Ano01v, Bar01a, Bar01b, Ber01, Ber00c, Bin02, Boi01, Bol02, Bor02, Bro01, Bue02, Buh00, But00, Cal01, Can00, Car01, Chr01a, Chr01b, Cle01, Cle02a, Cle02b, Coa00, Cog02, Cor02a, Cor02c, Cow00a, Cow00b, Das02a, Dav01, Del02a, Dha01, Dia00a, Dia00b, Dia00c, Dob02a, Dob00, Dow00, Dua01, Dub01, Dyk00, Ell01, Els02b, Emb01a, Emb01b, Emb01c, Emb02a, Emb02b, Emb02c, Ens02, Ern00, Fel02a, Fel02b, Fri02a, FS00, Ful02, Gad00, Gan02, Gar00a, Gar00b, Gen02, Gen01, Gey01, Ghe01, Gho01, Gol02a, Gol02b, Gra02, Gri02a, Gun02, Har00b, Haz00, Hel00, Hen02, Hes00, Hol00b, Hol00a, Hor02, Hur00, Ing00, Jam00a, Jam00b, Jam00c, Jam00d, Jam00e, Jam01a, Jam01b]. untitled [Jam01c, Jam01d, Jam02a, Jam02b, Jam02c, Jam02d, Jen02a, Jon00, Kar02a, Kar02b, Kar02c, Kim01a, Kim01b, Koe02b, Kok01, Kol02a, Kue00, Lan00, Lar02, Len02a, Len02b, Len02c, Len02d, Len02e, Lev00, Lit00, Loc02a, Lun00a, Lun00b, Lun01, Lun02, MTW00a, MTW00b, MTW00c, MTW00d, MTW00e, MTW00f, MTW00g, MTW00h, MTW00i, MTW00j, MTW00k, MTW00l, MTW00m, MTW00n, MTW00o, MTW00p, MTW00q, Mac01, Mad01, McC02b, McG02, McN01, McN02a, Mey01, Ogd00a, Ogd00b, Ogd02, Ols02, PE00a, PE00b, PE00c, PE00d, PE00e, Pad00, Par00, Par02a, Pea02, Ped01, Per02a, Per02b, Pra00, Pre00, Pre02, RLS00a, RLS00b, RLS00c, RVL00, Rac01, Rao00, Rat00, Rei00b, Rei02, Rhe02, Ril02, Rob00, Roc00, Rog02b, Rol01, Rol00, Rol02, Ron00, Roo02a, Rup01, Rup02a, Rut00, SRJ00a, SRJ00b, SRJ00c, SRJ00d, SC00]. untitled [Sch01, Sch02c, Sch00a, Sch00b, Sch02e, Sen01, She02b, Shi01, Sig00, Sor00, Spa02a, Sri02, Sta01, Ste00b, Sto00, Str01a, Str00h, Str00i, Str00a, Str00b, Str00c, Str00j, Str00d, Str00e, Str01b, Str01c, Str01d, Str01e, Str01f, Str01g, Str02b, Str02a, Str00l, Str02c, Stu00, Sun02, Sun01, Sun00, Swa00, Swy02a, TJV00a, TJV00b, TJV00c, TJV00d, TJV00e, Tak00, Tay00, The00, Tho02, Tiw00, Tri02a, Vie02a, Vog01a, Vog01b, Vog01c, Vog01d, Vog02a, Vog02b, Vog02e, Vog02c, Wan01, Wan02, Was01, Weg00, Wel00, Wel01d, Wel01e, Wel01f, Wel01g, Wel01h, Wel01a, Wel01b, Wel01c, Wel01i, Wel01j, Wel01k, Wel01l, Wel01m, Wel01n, Wel01o, Wel01p, Wel01q, Wel01r, Wel01s, Wel01t, Wol02a, Ylv01, Zal02, ZS02a, Zwi00, vdV01]. Upfal [Buo06]. Usage [RGA01]. Use [CHL03, RAHT06]. Used [Cau01, Gro01, Hor01a, Hor01b, Mag01, RK01]. Useful [IK08]. User [Tay03a, Tay03b]. Uses [SWP00]. Using [AC04b, Ano05j, BP08, BM01, XMN06, CCDH08, CHB^{+}08, CM01, CN08, Cle03a, Cle03b, CGRS01, DR00, Dav07a, Daw03, DH08, DFS02, Dur06, Eas03, EL05, Far01, FWP05, FBV^{+}04, FO09, Fue03, GT07, GBT09, Gil08a, GSC04, GL03, HGWR08, HWZ06, HOS04, HSB09, IZT08, IR03a, IJ04, KMH05, KU01a, KU01b, KTZ00, LR01, LSZH08, LDVT09, LPD00, LS02, LL03b, LW08, LC01b, LSQ03, LZH09, LCLC08, LM00, Mar01, Miy04, MAC^{+}06b, Neu06, NC08, PM09, PGH01, PCK06, PAS^{+}03, PLM08, PP01, PW05, PC00, QA09, Ram01a, RHC07, RWR03, Rub05, SS08, Ste06, SW02d, SjKH^{+}01, Tan06a, Tan06b, TEB00, VS08, WRM07, WJSS08, WS01a, Wie06, WS01b, XSC08, YW03, YTS08, YR00, ZRM09, ZSG06, ZQ09, dGKS01, Man07].

V [AC04a, Epi04, Kar03a, Kle06, Lyl06]. Vaccine [HPC03, HH06, KF02]. Validation [BBK^{+}09, Bur04, DLM04, EBD^{+}04, Efr04b, FZ03, HRL04, Hey08, Lee00, PS05, PS06b, SHY04a, SRL00, SXZ07, Zha04a]. Valuation [Hin07, Kor05]. Value [EM08, Hof07, HWZ06, HHCK09, HSB09, Kie08b, Rhe03a, Rhe03b]. Valued [FW00b]. Values [BB00a, BB00b, Boo00, EL04, Eva00a, HDS06, IWM04, Mar00a, Mar00b, MMB09, Pic00, RvdVV00a, RvdVV00b, SSY07, Ste00c, ZCPL08]. Vandenberghe [Lue05]. Vani [Win03]. VaR [GP07]. Variability [HR03, Sch03a]. Variable [Aba02, Bak00, Bhl02, BL06, Büh02, CM06a, Cho06, CA02, CD09, Cor05, CKN08, DTT08, FL01, FS04b, FD08, Geo00, GI07, Hal02, HZAG01, JR07, JLZ08, KC07, Lah02, LPM^{+}08, LL09, LM00, ML02, MVF06, NS02, Pol02, RJ08, RD06, RGK01, SI02, She02d, SF07, TSV05, WLH08, Wan09, WBS07b, YL05, ZFZ03, ZWL^{+}04]. Variable-Length [Bhl02, Büh02, Hal02, Lah02, Pol02, She02d]. Variables [CRC^{+}04, CRW01, DM07, DH08, DFC09, EOU02, FJS08, GB00, JvD03, LFGR04, PT09, PW05, Sma07, Tan06b, YZ05, Bal03]. Variance [AA00, Ano02o, Boo06, CS01, CP07, CvA02, DP05, EHI08, EL04, FWP05, GL07, Gua09, KNF06, LSC04, LO06, LBS06, MH02, MSY06, Ros03, SWP00, Sim02, SD02, Sto08, WR02a, WR02b, YJ04, YR00, Kac04]. Variances [SMP03, TW07b, WF03]. Variate [Cow07, Ruk03a, Ruk03b]. Variation [GH08, Gel08, HSG02, Hod08, Kat08, KW02, Ram00, SWDS03]. Variogram [NC08]. Varying [CFL00, CDVY00, CW08, CRW01, EL04, GKSB03, GT09, KT09, LTE08, RAHT06, TZW05, WLH08, WKL09, WX09, XSC08, XZ07, YPXH06, YLZ08]. Varying-Coefficient [CFL00, WLH08]. Varying-Scale [XSC08]. Vector [LLW04, WS07, Wil03a, Wil03c, WL07b]. Vectors [Fel02c, ZWS06]. Vedel [Ann05b]. Vehicle [BBK^{+}09]. Veil [GR06b]. Vending [MD01]. Verbeke [Wil06a]. Verification [McK05, RFS06]. Verlag [McC02a]. Versions [FH00]. Versus [BG01]. Vertical [HR00, Gen07b]. Vestrup [Stu06]. Veterans [LNR03]. via [ABPR06, CS04, CL08a, CN05, CvA02, FL01, FB00, Gho06, HL06b, Hös03, Hss03, LLC00, MRŞK06, MB05, MH02, NS02, RSW09, Sim02, SD02, WR02a, WR02b, XLB04, ZWL^{+}04]. Vic [Wik06]. Vicent [Ken03a]. Victor [Lin04a]. View [O’F01, Wat07]. Vignette [VJ00]. Vignettes [RTW00b]. Vinod [Mal05]. Viral [AC01a, FD03]. Vision [O’F01]. Visualizing [Wil04a]. Vital [Sch03a]. Vittinghoff [Lyl06]. Vladimir [Lin04a]. Vol [Ank05, Ann06d, Boo06, Kar03a, Kar03b, Ker06, Koe02a, Loc02b]. Volatility [ABDL01, Ano03j, FFJ07, FW07, ZMAS05]. Volume [Ano00l, Ano01w, Ano02t, Ano03q, Ano04o, Ano05s, Ano06u, Ano08f, Ano09j, PY00]. Vote [DK01, HB09]. Voting [GM08]. Vouchers [BFHR03a, BFHR03b, BX03, KZ03, MJB03].

W [Cav06b, Dav03, Gen05a, Gol05a, Hab03, Hu04, Jam03a, Lar06, Len06, Ros06, Str03b, Wil04a, Wil04b]. Waagepetersen [Sch05]. Wages [SR08]. Wai [Gal05]. Wait [SWSM09]. Wait-List [SWSM09]. Walk [Pre03]. Waller [Dig05]. Walter [Dur06, Few04]. Wand [Lee06]. Wang [Lia05, McN04]. Waning [KF02]. War [SR08]. Ware [Mol05]. Warped [LM04]. Warren [Sch03b, Zhe06]. Waste [ST02]. Water [LSGS01, LSGS04, Man07, MICR06]. Waters [LSGS01]. Wave [OC06]. Wavelet [AF01b, AF01a, BFV01, Cai01, CG03b, FO09, JB01, Kol03b, KS00, MVBC03b, MVBC03c, MAC^{+}06b, Mou01, RTD01, SWP00, Sol01, Vid01, WH03b, ZW08]. Wavelet-Based [CG03b, Kol03b, MVBC03b, MVBC03c, MAC^{+}06b, WH03b, ZW08]. Wavelets [DSV04, WRM07]. Way [AP04, BFD01, CM09a, HWZ05, HSB09]. Wayne [Len05]. Wealth [BBCL02]. Weather [Bri04, CD05, GRG^{+}04, GRGB04, NDKW01, Rou04, TN04, XWF05, Hin07]. Web [Li05]. Wei [Gen05b]. Weibull [Nes05]. Weighted [BC01, BFB00, CZ07, Che01, DHS08, FEG07, GS07, HTZ^{+}09, IJ04, Laz07, Lia02, MTA07, QWP05, Sch07d, WW09, ZZIP07a, ZZIP07b, Zim03, dV04]. Weighting [LLW01, Tan06b]. Weiss [Zha06]. Welcoming [Bec00a]. Well [EOU02]. Wellek [Mun04]. Wells [Nol03a]. Wendy [Jer06, Wik04]. Wenyaw [Wil06b]. Werwatz [Spe06a]. Whales [Alt01, GP01, QM01, RCGT01a, RCGT01b, Zeh01]. White [BBCL02]. Whittle [VR00]. Whole [TBD^{+}03]. Whole-Genome [TBD^{+}03]. Wicksell [FH00]. Width [BL07, JHCN06]. Wildfire [PSW05]. Wilkinson [Bra06b]. Willett [And05]. William [AC05a, Lin04b, Mor04, Spe06b]. Williams [Mor04]. Willingness [LL03a]. Wim [Dou06]. Wind [GLW^{+}06, OC06]. Wind-Generated [OC06]. Window [HL00]. Winds [WMNB01]. Winfred [Cle03a]. Winston [Cle03a]. Wire [Gre03]. Wit [Dav06c]. Within [GM08, PS07]. Without [CS00a, CCL09, Cox00, Daw00a, Daw00b, FTW01, KL06b, LC00, Pea00, RGK01, RG00, Rub00, Sha00, SY06, Was00]. Without/With [LC00]. Witnesses [Las08]. Wludyka [Bra06a]. Wolfgang [Spe06a]. Wombling [BG06a, BG06b]. Women [GT09, RAHT06]. Woodward [Dav06a]. Wooi [Ann06a]. Words [Dav07b, Dav07b]. Working [WC04]. Workload [Li05, Li07]. World [Loc02b, McC05]. Worst [Lüt03, Mal05]. Worst-Case [Lüt03]. Would [Fol00]. Would-Be [Fol00]. Wright [Gen05a]. Wu [Dra06, Kar06a].

Xie [Nes05]. xiii [McC02a]. Xitao [All06]. Xu [Gen04].

Yali [Rue04]. Yan [Sur04]. Yanhong [Dra06]. Yao [Bre05]. Yatchew [Tro06]. Yea [vSH08]. Yea-Saying [vSH08]. Year [RTW00b, Sch03a, UCTW07]. Years [CR00a]. Yield [BM08, CE09, Hil08, LLL08a, SCS08a]. Ying [Gen04]. Yingdong [Gen05a]. York [BFHR03a, BFHR03b, BX03, GFK07, KZ03, McC02a, MJB03]. Young [Cho06, GT09]. Yu [Len04a, Lin04a, Pou04, Yan04c]. Yu-Fang [Len04a]. Yury [Kie06].

Zehua [Ozt05]. Zellner [Lei05]. Zhang [Gen05b, Dui16, Gen04]. Zhao [Gen05a]. Zhidong [Ozt05]. Zivot [McN04]. Zonoid [Joe03]. Zucchini [Few04].

References

Akritas:2000:AAV

[AA00] Michael Akritas and Steven Arnold. Asymptotics for analysis of variance when the number of levels is large. Journal of the American Statistical Association, 95(449):212–226, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/Akritas1.htm;   http://www.jstor.
org/stable/2669539.

Akritas:2006:NAF

[AAK06] Michael G. Akritas, Efi S. Antoniou, and Jouni Kuha. Nonparametric analysis of factorial designs with random missingness: Bivariate data. Journal of the American Statistical Association, 101(476):1513–1526, December 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27639769.

Auranen:2000:TPC

[AALT00] Kari Auranen, Elja Arjas, Tuija Leino, and Aino K. Takala. Transmission of pneumococcal carriage in families: a latent Markov process model for binary longitudinal data. Journal of the American Statistical Association, 95(452):1044– 1053, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669741.

Abadie:2002:BTD

[Aba02] Alberto Abadie. Bootstrap tests for distributional treatment effects in instrumental variable models. Journal of the American Statistical Association, 97(457):284–292, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085782;   reqidx=/cw/asa/01621459/v97n457/
s25/p284.

Andersen:2001:DRE

[ABDL01] Torben G. Andersen, Tim Bollerslev, Francis X. Diebold, and Paul Labys. The distribution of realized exchange rate volatility. Journal of the American Statistical Association, 96(453):42–55, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670339;   reqidx=/cw/
asa/01621459/v96n453/s5/p42.

Amzal:2006:BOD

[ABPR06] Billy Amzal, Frédéric Y. Bois, Eric Parent, and Christian P. Robert. Bayesian-optimal design via interacting particle systems. Journal of the American Statistical Association, 101 (474):773–785, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590734.

Amaratunga:2001:ADV

[AC01a] Dhammika Amaratunga and Javier Cabrera. Analysis of data from viral DNA microchips. Journal of the American Statistical Association, 96(456):1161–1170, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085879;   reqidx=/cw/asa/01621459/v96n456/
s3/p1161.

Anderson-Cook:2001:BRB

[AC01b] C. M. Anderson-Cook. Book review: [untitled]. Journal of the American Statistical Association, 96(454):776–777, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670312.

Anderson-Cook:2002:BRBa

[AC02a] C. M. Anderson-Cook. Book review: [untitled]. Journal of the American Statistical Association, 97(459):928, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085747.

Anderson-Cook:2002:BRBb

[AC02b] Christine M. Anderson-Cook. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1216, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085864.

Anderson-Cook:2003:BRBa

[AC03a] Christine M. Anderson-Cook. Book review: Concise Handbook of Experimental Methods for the Behavioral and Biological Sciences, by Jay E. Gould. Journal of the American Statistical Association, 98(461):249, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045213.

Anderson-Cook:2003:BRBb

[AC03b] Christine M. Anderson-Cook. Book review: Interpreting Quantitative Data, by David Byrne. Journal of the American Statistical Association, 98(462):492, June 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30045259.

Anderson-Cook:2003:BRBc

[AC03c] Christine M. Anderson-Cook. Book review: Statistical Analysis of Designed Experiments (second edition), by Helge Toutenburg. Journal of the American Statistical Association, 98 (463):770–771, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045312.

Anderson-Cook:2003:CHE

[AC03d] Christine M. Anderson-Cook. Concise handbook of experimental methods for the behavioral and biological sciences. Journal of the American Statistical Association, 98(461):249–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s26/p249.

Anderson-Cook:2003:IQD

[AC03e] Christine M. Anderson-Cook. Interpreting quantitative data. Journal of the American Statistical Association, 98(462):492–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s23/p492.

Aruoba:2003:BRB

[AC03f] S. Borağan Aruoba and Sean D. Campbell. Book review: Time Series Analysis by State-Space Methods, by J. Durbin; S. J. Koopman. Journal of the American Statistical Association, 98(461):255–256, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045223.

Anderson-Cook:2004:BRBa

[AC04a] Christine M. Anderson-Cook. Book review: A Primer on Statistical Distributions, by N. Balakrishnan; V. B. Nevzorov. Journal of the American Statistical Association, 99(466):568, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s39/p568;   http:
//www.jstor.org/stable/27590425.

Anderson-Cook:2004:BRBb

[AC04b] Christine M. Anderson-Cook. Book review: Data Analysis and Graphics Using R: An Example-Based Approach, by John Maindonald; John Braun. Journal of the American Statistical Association, 99(467):901–902, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s47/p901;   http://www.jstor.org/stable/27590463.

Anderson-Cook:2004:BRBd

[AC04c] Christine M. Anderson-Cook. Book review: Handbook of Statistics 22: Statistics in Industry, by R. Khattree; C. R. Rao. Journal of the American Statistical Association, 99(467):904, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s48/p902;   http:
//www.jstor.org/stable/27590468.

Anderson-Cook:2004:BRBe

[AC04d] Christine M. Anderson-Cook. Book review: Statistical Methods for Six Sigma in R&D and Manufacturing, by Anand M. Joglekar. Journal of the American Statistical Association, 99(468):1205–1206, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/

vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s36/p1205;   http://www.jstor.org/stable/27590503.

Anderson-Cook:2004:BRBc

[AC04e] Christine M. Anderson-Cook. Book review: Statistical Tools for Nonlinear Regression: A Practical Guide with S-PLUS and R Examples, second edition, by S. Huet; A. Bouvier; M.-A. Poursat; E. Jolivet. Journal of the American Statistical Association, 99(467):902–903, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s52/p904;   http://www.jstor.org/stable/27590464.

Anderson-Cook:2005:BRBa

[AC05a] Christine M. Anderson-Cook. Book review: Experimental and Quasi-Experimental Designs for Generalized Causal Inference, by William R. Shadish; Thomas D. Cook; Donald T. Campbell. Journal of the American Statistical Association, 100(470):708, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s36/p708;   http://miranda.
asa.catchword.org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/
01621459/v100n470/s37/p708;   http://www.jstor.org/
stable/27590595.

Anderson-Cook:2005:BRBc

[AC05b] Christine M. Anderson-Cook. Book review: Planning, Construction, and Statistical Analysis of Comparative Experiments, by Francis G. Giesbrecht; Marcia L. Gumpertz. Journal of the American Statistical Association, 100(470):708–709, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590596.

Anderson-Cook:2005:BRBb

[AC05c] Christine M. Anderson-Cook. Book review: Planning, Construction, and Statistical Analysis of Comparative Experiments. Francis G. Giesbrecht and Marcia L. Gumpertz. Journal of the American Statistical Association, 100(470):708–709, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.

org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/
s36/p708;   http://miranda.asa.catchword.org/vl=1515603/
cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s37/p708.

Anderson-Cook:2005:BRBd

[AC05d] Christine M. Anderson-Cook. Book review: Practical Genetic Algorithms, by Randy L. Haupt; Sue Ellen Haupt. Journal of the American Statistical Association, 100(471):1099, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590650.

Anderson-Cook:2006:BRBb

[AC06a] Christine M. Anderson-Cook. Book review: Elementary Statistical Quality Control, by John T. Burr. Journal of the American Statistical Association, 101(473):402, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047488.

Anderson-Cook:2006:BRBc

[AC06b] Christine M. Anderson-Cook. Book review: Quantitative Risk Management: Concepts, Techniques, and Tools, by Alexander J. McNeil; Rudiger Frey; Paul Embrechts. Journal of the American Statistical Association, 101(476):1731–1732, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27639804.

Anderson-Cook:2006:BRBa

[AC06c] Christine M. Anderson-Cook. Book review: Statistical Design of Experiments with Engineering Applications, by Kamel Rekab; Muzaffar Shaikh. Journal of the American Statistical Association, 101(473):396–397, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047477.

Anderson-Cook:2007:BRBc

[AC07a] Christine M. Anderson-Cook. Book review: A Modern Theory of Factorial Design. Journal of the American Statistical Association, 102(478):765, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2007:BRBa

[AC07b] Christine M. Anderson-Cook. Book review: Functional Approach to Optimal Experimental Design. Journal of the American Statistical Association, 102(477):386, March 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2007:BRBb

[AC07c] Christine M. Anderson-Cook. Book review: Generalized Additive Models: An Introduction With R. Journal of the American Statistical Association, 102(478):760–761, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2008:BRBa

[AC08a] Christine M. Anderson-Cook. Book review: Modern Experimental Design. Journal of the American Statistical Association, 103(481):426–427, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2008:BRBd

[AC08b] Christine M. Anderson-Cook. Book review: Optimum Experimental Designs, With SAS. Journal of the American Statistical Association, 103(483):1327–1328, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2008:BRBb

[AC08c] Christine M. Anderson-Cook. Book review: Reliability, Life Testing, and the Prediction of Service Lives for Engineers and Scientists. Journal of the American Statistical Association, 103 (482):887–888, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2008:BRBc

[AC08d] Christine M. Anderson-Cook. Book review: Response Surfaces, Mixtures, and Ridge Analyses. Journal of the American Statistical Association, 103(482):888, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2005:BRBe

[ACE05] Christine M. Anderson-Cook and Paul Embrechts. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 100(471):1099–1100, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590651.

Abowd:2001:MEA

[ACK01] John M. Abowd, Bruno Crépon, and Francis Kramarz. Moment estimation with attrition: An application to economic models. Journal of the American Statistical Association, 96 (456):1223–1231, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085885;   reqidx=/cw/
asa/01621459/v96n456/s9/p1223.

Anderson-Cook:2004:BRBf

[ACLJ^{+}04] Christine M. Anderson-Cook, Russell V. Lenth, Murray Jorgensen, Robert B. Lund, Jane C. Lindsey, and RVL. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 99(468):1209–1210, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590509.

Arnold:2006:FMD

[ACS06] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Families of multivariate distributions involving the Rosenblatt construction. Journal of the American Statistical Association, 101(476):1652–1662, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639779.

Albert:2008:EDA

[AD08] Paul S. Albert and Lori E. Dodd. On estimating diagnostic accuracy from studies with multiple raters and partial gold standard evaluation. Journal of the American Statistical Association, 103(481):61–73, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Amaral:2007:PBM

[ADW07] G. J. A. Amaral, I. L. Dryden, and Andrew T. A. Wood. Pivotal bootstrap methods for k-sample problems in directional statistics and shape analysis. Journal of the American Statistical Association, 102(478):695–707, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alvarez-Esteban:2008:TCD

[ÁEdBCAM08] Pedro César Álvarez-Esteban, Eustasio del Barrio, Juan Antonio Cuesta-Albertos, and Carlos Matrán. Trimmed comparison of distributions. Journal of the American Statistical Association, 103(482):697–704, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Antoniadis:2001:RWAb

[AF01a] A. Antoniadis and J. Fan. Regularization of wavelet approximations: Rejoinder. Journal of the American Statistical Association, 96(455):964–967, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2670242.

Antoniadis:2001:RWAa

[AF01b] Anestis Antoniadis and Jianqing Fan. Regularization of wavelet approximations. Journal of the American Statistical Association, 96(455):939–955, September 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670237;   reqidx=/cw/asa/01621459/v96n455/s14/p939.

Aykroyd:2001:BMA

[AHA01] Robert G. Aykroyd, John G. B. Haigh, and Gayle T. Allum. Bayesian methods applied to survey data from archeological magnetometry. Journal of the American Statistical Association, 96(453):64–76, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670341;   reqidx=/cw/
asa/01621459/v96n453/s7/p64.

Albers:2001:DDR

[AKM01] Willem Albers, Wilbert C. M. Kallenberg, and Felix Martini. Data-driven rank tests for classes of tail alternatives. Journal of the American Statistical Association, 96(454):685–696, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670306;   reqidx=/cw/asa/01621459/v96n454/
s27/p685.

Arcones:2002:NED

[AKS02] Miguel A. Arcones, Paul H. Kvam, and Francisco J. Samaniego. Nonparametric estimation of a distribution subject to a stochastic precedence constraint. Journal of the American Statistical Association, 97(457):170–182, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085772;   reqidx=/cw/asa/01621459/v97n457/
s15/p170.

Alamar:2007:BRB

[Ala07] Ben Alamar. Book review: Social Choice With Partial Knowledge of Treatment Response. Journal of the American Statistical Association, 102(480):1481, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:2000:BRB

[Alb00] James H. Albert. Book review: [untitled]. Journal of the American Statistical Association, 95(450):679–680, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669420.

Albert:2002:BRB

[Alb02] Jim Albert. Book review: [untitled]. Journal of the American Statistical Association, 97(457):354–355, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085790.

Aldrich:2008:CSA

[Ald08] John Aldrich. Comment: On Student’s 1908 article “The Probable Error of a Mean”. Journal of the American Statistical Association, 103(481):8–11, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Allen:2006:BRB

[All06] Jeff M. Allen. Book review: SAS for Monte Carlo Studies: A Guide for Quantitative Researchers, by Xitao Fan; Ákos Felsővályi; Stephen A. Sivo; Sean C. Keenan. Journal of the American Statistical Association, 101(476):1728– 1729, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639799.

Altman:2001:EFG

[Alt01] Naomi S. Altman. Estimation of fetal growth and gestation in bowhead whales: Comment. Journal of the American Statistical Association, 96(455):923–925, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670232.

Altman:2007:MHM

[Alt07] Rachel MacKay Altman. Mixed hidden Markov models. Journal of the American Statistical Association, 102(477):201–210, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aerts:2001:SSA

[AM01] Marc Aerts and Geert Molenberghs. Smoothing spline ANOVA for multivariate Bernoulli observations, with application to ophthalmology data: Comment. Journal of the American Statistical Association, 96(453):155–157, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670353.

Aban:2006:PET

[AMP06] Inmaculada B. Aban, Mark M. Meerschaert, and Anna K. Panorska. Parameter estimation for the truncated Pareto distribution. Journal of the American Statistical Association, 101(473):270–277, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047455.

Augustin:2009:MSF

[AMvW^{+}09] Nicole H. Augustin, Monica Musio, Klaus von Wilpert, Edgar Kublin, Simon N. Wood, and Martin Schumacher. Modeling spatiotemporal forest health monitoring data. Journal of the American Statistical Association, 104(487):899–911, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Asgharian:2002:LBS

[AMW02] Masoud Asgharian, Cyr Emile M’Lan, and David B. Wolfson. Length-biased sampling with right censoring: An unconditional approach. Journal of the American Statistical Association, 97(457):201–209, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085775;   reqidx=/cw/
asa/01621459/v97n457/s18/p201.

Anderson:2005:BRB

[And05] Carolyn J. Anderson. Book review: Applied Longitudinal Data Analysis: Modeling Change and Event Occurrence, by Judith D. Singer; John B. Willett. Journal of the American Statistical Association, 100(469):352–353, March 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n469/s43/p353;   http:
//www.jstor.org/stable/27590549.

Anderson:2008:MIG

[And08] Michael L. Anderson. Multiple inference and gender differences in the effects of early intervention: a reevaluation of the Abecedarian, Perry Preschool, and early training projects. Journal of the American Statistical Association, 103(484): 1481–1495, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ankenman:2001:BRB

[Ank01] Bruce Ankenman. Book review: [untitled]. Journal of the American Statistical Association, 96(453):348, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670381.

Ankerst:2005:BRB

[Ank05] D. Pauler Ankerst. Book review: Kendall’s Advanced Theory of Statistics, Vol. 2B: Bayesian Inference, by Kendall; Anthony O’Hagan; Jonathan Forster. Journal of the American Statistical Association, 100(472):1465–1466, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590700.

Annis:2005:BRBa

[Ann05a] David H. Annis. Book review: Permutation, Parametric, and Bootstrap Tests of Hypotheses, by Phillip Good. Journal of the American Statistical Association, 100(472):1457– 1458, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590686.

Annis:2005:BRBb

[Ann05b] David H. Annis. Book review: Stereology for Statisticians, by Adrian Baddeley; Eva B. Vedel Jensen. Journal of the American Statistical Association, 100(472):1462, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590693.

Annis:2006:BRBb

[Ann06a] Charles Annis. Book review: Handbook of Exponential and Related Distributions for Engineers and Scientists, by Nabendu Pal; Chun Jin; Wooi K. Lim. Journal of the American Statistical Association, 101(476):1734–1735, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639809.

Annis:2006:BRBc

[Ann06b] David H. Annis. Book review: Bayesian Models for Categorical Data, by Peter Congdon. Journal of the American Statistical Association, 101(474):844–845, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590741.

Annis:2006:BRBd

[Ann06c] David H. Annis. Book review: Bayesian Statistics and Marketing, by Peter Rossi; Greg Allenby; Rob McCulloch. Journal of the American Statistical Association, 101(476):1732– 1733, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639806.

Annis:2006:BRBa

[Ann06d] David H. Annis. Book review: Kendall’s Advanced Theory of Statistics, Vol. 1: Distribution Theory, Kendall’s Advanced Theory of Statistics, Vol. 2A: Classical Inference and the Linear Model, by Alan Stuart; J. Keith Ord; Steven F. Arnold. Journal of the American Statistical Association, 101(476): 1721, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639788.

Annis:2008:BRB

[Ann08] David H. Annis. Book review: Bayes Linear Statistics: Theory and Methods. Journal of the American Statistical Association, 103(483):1319, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2000:BMa

[Ano00a] Anonymous. Back matter. Journal of the American Statistical Association, 95(449):??, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669595.

Anonymous:2000:BMb

[Ano00b] Anonymous. Back matter. Journal of the American Statistical Association, 95(450):??, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669443.

Anonymous:2000:BMc

[Ano00c] Anonymous. Back matter. Journal of the American Statistical Association, 95(451):??, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669515.

Anonymous:2000:BMd

[Ano00d] Anonymous. Back matter. Journal of the American Statistical Association, 95(452):??, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669816.

Anonymous:2000:BR

[Ano00e] Anonymous. Book reviews. Journal of the American Statistical Association, 95(449):327–??, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2000:CEM

[Ano00f] Anonymous. Correction: Efficient Monte Carlo methods for conditional logistic regression. Journal of the American Statistical Association, 95(451):1020, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669514.

Anonymous:2000:CEC

[Ano00g] Anonymous. Correction: Exact confidence bounds for all contrasts of three or more regression lines. Journal of the American Statistical Association, 95(451):1020, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669513.

Anonymous:2000:FMa

[Ano00h] Anonymous. Front matter. Journal of the American Statistical Association, 95(449):327, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669516.

Anonymous:2000:FMb

[Ano00i] Anonymous. Front matter. Journal of the American Statistical Association, 95(450):674, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669371.

Anonymous:2000:FMc

[Ano00j] Anonymous. Front matter. Journal of the American Statistical Association, 95(451):1006, September 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669444.

Anonymous:2000:FMd

[Ano00k] Anonymous. Front matter. Journal of the American Statistical Association, 95(452):1369, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669738.

Anonymous:2000:VI

[Ano00l] Anonymous. Volume information. Journal of the American Statistical Association, 95(452):1383–1395, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669737.

Anonymous:2001:BMa

[Ano01a] Anonymous. Back matter. Journal of the American Statistical Association, 96(453):??, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670396.

Anonymous:2001:BMb

[Ano01b] Anonymous. Back matter. Journal of the American Statistical Association, 96(454):??, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670333.

Anonymous:2001:BMc

[Ano01c] Anonymous. Back matter. Journal of the American Statistical Association, 96(455):??, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2670273.

Anonymous:2001:BMd

[Ano01d] Anonymous. Back matter. Journal of the American Statistical Association, 96(456):??, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085944.

Anonymous:2001:BRa

[Ano01e] Anonymous. Book reviews. Journal of the American Statistical Association, 96(453):339–??, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
reqidx=/cw/asa/01621459/v96n453/s27/p339.

Anonymous:2001:BRb

[Ano01f] Anonymous. Book reviews. Journal of the American Statistical Association, 96(454):775–??, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
reqidx=/cw/asa/01621459/v96n454/s33/p775.

Anonymous:2001:BRc

[Ano01g] Anonymous. Book reviews. Journal of the American Statistical Association, 96(455):1133–??, September 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://rosina.catchword.com/cgi-bin/linker?ini=asa&
amp;   reqidx=/cw/asa/01621459/v96n455/s31/p1133.

Anonymous:2001:BRd

[Ano01h] Anonymous. Book reviews. Journal of the American Statistical Association, 96(456):1522–??, December 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://rosina.catchword.com/cgi-bin/linker?ini=asa&
amp;   reqidx=/cw/asa/01621459/v96n456/s31/p1522.

Anonymous:2001:Ca

[Ano01i] Anonymous. Correction. Journal of the American Statistical Association, 96(453):356–??, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
reqidx=/cw/asa/01621459/v96n453/s28/p356.

Anonymous:2001:CBM

[Ano01j] Anonymous. Correction: Bayesian morphology: Fast unsupervised Bayesian image analysis. Journal of the American Statistical Association, 96(453):356, March 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670395.

Anonymous:2001:Cb

[Ano01k] Anonymous. Corrections. Journal of the American Statistical Association, 96(454):784–??, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
reqidx=/cw/asa/01621459/v96n454/s34/p784.

Anonymous:2001:Cc

[Ano01l] Anonymous. Corrections. Journal of the American Statistical Association, 96(456):1534–??, December 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://rosina.catchword.com/cgi-bin/linker?ini=asa&
amp;   reqidx=/cw/asa/01621459/v96n456/s32/p1534.

Anonymous:2001:CCN

[Ano01m] Anonymous. Corrections: Classes of nonseparable, spatiotemporal stationary covariance functions. Journal of the American Statistical Association, 96(454):784, June 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670331.

Anonymous:2001:CCC

[Ano01n] Anonymous. Corrections: Comparing classical and resistant outlier rules. Journal of the American Statistical Association, 96(456):1534, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085942.

Anonymous:2001:CLB

[Ano01o] Anonymous. Corrections: Likelihood-based methods for missing covariates in the Cox proportional hazards model. Journal of the American Statistical Association, 96(456):1534, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085943.

Anonymous:2001:CSS

[Ano01p] Anonymous. Corrections: Smoothing spline ANOVA for multivariate Bernoulli observations, with application to ophthalmology data. Journal of the American Statistical Association, 96(454):784, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670332.

Anonymous:2001:FMa

[Ano01q] Anonymous. Front matter. Journal of the American Statistical Association, 96(453):339, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2670334.

Anonymous:2001:FMb

[Ano01r] Anonymous. Front matter. Journal of the American Statistical Association, 96(454):775, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670274.

Anonymous:2001:FMc

[Ano01s] Anonymous. Front matter. Journal of the American Statistical Association, 96(455):1133, September 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670218.

Anonymous:2001:FMd

[Ano01t] Anonymous. Front matter. Journal of the American Statistical Association, 96(456):1522, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085876.

Anonymous:2001:TR

[Ano01u] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 96(455):1141–??, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-
bin/linker?ini=asa&amp;   reqidx=/cw/asa/01621459/
v96n455/s32/p1141.

Anonymous:2001:U

[Ano01v] Anonymous. [untitled]. Journal of the American Statistical Association, 96(456):1526, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085922.

Anonymous:2001:VI

[Ano01w] Anonymous. Volume information. Journal of the American Statistical Association, 96(456):1535–1543, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085875.

Anonymous:2002:BMa

[Ano02a] Anonymous. Back matter. Journal of the American Statistical Association, 97(457):??, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085812.

Anonymous:2002:BMb

[Ano02b] Anonymous. Back matter. Journal of the American Statistical Association, 97(458):??, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085700.

Anonymous:2002:BMc

[Ano02c] Anonymous. Back matter. Journal of the American Statistical Association, 97(459):??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085751.

Anonymous:2002:BMd

[Ano02d] Anonymous. Back matter. Journal of the American Statistical Association, 97(460):??, December 2002. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085874.

Anonymous:2002:BR

[Ano02e] Anonymous. Book reviews. Journal of the American Statistical Association, 97(457):352–??, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
reqidx=/cw/asa/01621459/v97n457/s31/p352.

Anonymous:2002:Ca

[Ano02f] Anonymous. Correction. Journal of the American Statistical Association, 97(457):367–??, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
reqidx=/cw/asa/01621459/v97n457/s32/p367.

Anonymous:2002:Cc

[Ano02g] Anonymous. Correction. Journal of the American Statistical Association, 97(459):929–??, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s37/p928;   http://giorgio.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/s38/
p929.

Anonymous:2002:Cd

[Ano02h] Anonymous. Correction. Journal of the American Statistical Association, 97(460):1218–??, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s43/p1215;   http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s44/p1218.

Anonymous:2002:CEIb

[Ano02i] Anonymous. Correction: Effect of insurance on mortality in an HIV-positive population in care. Journal of the American Statistical Association, 97(460):1218, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085873.

Anonymous:2002:CLB

[Ano02j] Anonymous. Correction: Length-biased sampling with right censoring: An unconditional approach. Journal of the American Statistical Association, 97(459):929, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085750.

Anonymous:2002:CEIa

[Ano02k] Anonymous. Correction: The effect of insurance on mortality in an HIV + population in care. Journal of the American Statistical Association, 97(457):367, March 2002. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/3085811.

Anonymous:2002:Cb

[Ano02l] Anonymous. Corrections. Journal of the American Statistical Association, 97(458):661–??, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s24/p649;   http:/
/giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s25/p660;   http:/
/giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s26/p661.

Anonymous:2002:CCR

[Ano02m] Anonymous. Corrections: Calibration regression of censored lifetime medical cost. Journal of the American Statistical Association, 97(458):661, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085698.

Anonymous:2002:CCS

[Ano02n] Anonymous. Corrections: Comment on “semi-parametric nonlinear mixed-effects models and their application”. Journal of the American Statistical Association, 97(458):661, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085697.

Anonymous:2002:CTL

[Ano02o] Anonymous. Corrections: Testing and locating variance changepoints with applications to stock prices. Journal of the American Statistical Association, 97(458):661, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085699.

Anonymous:2002:FMa

[Ano02p] Anonymous. Front matter. Journal of the American Statistical Association, 97(457):352, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085752.

Anonymous:2002:FMb

[Ano02q] Anonymous. Front matter. Journal of the American Statistical Association, 97(458):649, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085649.

Anonymous:2002:FMc

[Ano02r] Anonymous. Front matter. Journal of the American Statistical Association, 97(459):918, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085701.

Anonymous:2002:FMd

[Ano02s] Anonymous. Front matter. Journal of the American Statistical Association, 97(460):1200, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085814.

Anonymous:2002:VI

[Ano02t] Anonymous. Volume information. Journal of the American Statistical Association, 97(460):1219–1228, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085813.

Anonymous:2003:BMa

[Ano03a] Anonymous. Back matter. Journal of the American Statistical Association, 98(461):??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045231.

Anonymous:2003:BMb

[Ano03b] Anonymous. Back matter. Journal of the American Statistical Association, 98(462):??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045273.

Anonymous:2003:BMc

[Ano03c] Anonymous. Back matter. Journal of the American Statistical Association, 98(463):??, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045325.

Anonymous:2003:BMd

[Ano03d] Anonymous. Back matter. Journal of the American Statistical Association, 98(464):??, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045364.

Anonymous:2003:Ca

[Ano03e] Anonymous. Corrections. Journal of the American Statistical Association, 98(462):501–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s34/p499;   http:/
/fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s35/p501.

Anonymous:2003:Cb

[Ano03f] Anonymous. Corrections. Journal of the American Statistical Association, 98(463):779–??, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s52/p779.

Anonymous:2003:CCR

[Ano03g] Anonymous. Corrections: Cox regression with accurate covariates unascertainable: a nonparametric-correction approach. Journal of the American Statistical Association, 98(463): 779, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045323.

Anonymous:2003:CLE

[Ano03h] Anonymous. Corrections: Letter to the editor. Journal of the American Statistical Association, 98(462):501, June 2003.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045272.

Anonymous:2003:COB

[Ano03i] Anonymous. Corrections: Objective Bayesian analysis of spatially correlated dara. Journal of the American Statistical Association, 98(463):779, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045324.

Anonymous:2003:CDR

[Ano03j] Anonymous. Corrections: The distribution of realized exchange rate volatility. Journal of the American Statistical Association, 98(462):501, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045271.

Anonymous:2003:FMa

[Ano03k] Anonymous. Front matter. Journal of the American Statistical Association, 98(461):??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045187.

Anonymous:2003:FMb

[Ano03l] Anonymous. Front matter. Journal of the American Statistical Association, 98(462):??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045233.

Anonymous:2003:FMc

[Ano03m] Anonymous. Front matter. Journal of the American Statistical Association, 98(463):??, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045275.

Anonymous:2003:FMd

[Ano03n] Anonymous. Front matter. Journal of the American Statistical Association, 98(464):??, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045328.

Anonymous:2003:LE

[Ano03o] Anonymous. Letter to the editor. Journal of the American Statistical Association, 98(461):261–??, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s42/p259;   http://fidelio.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/s43/
p261.

Anonymous:2003:TR

[Ano03p] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 98(464):1087–??, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s41/p1087.

Anonymous:2003:VI

[Ano03q] Anonymous. Volume information. Journal of the American Statistical Association, 98(464):??, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30045327.

Anonymous:2004:BMa

[Ano04a] Anonymous. Back matter. Journal of the American Statistical Association, 99(465):??, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590384.

Anonymous:2004:BMb

[Ano04b] Anonymous. Back matter. Journal of the American Statistical Association, 99(466):??, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590429.

Anonymous:2004:BMc

[Ano04c] Anonymous. Back matter. Journal of the American Statistical Association, 99(467):??, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590470.

Anonymous:2004:BMd

[Ano04d] Anonymous. Back matter. Journal of the American Statistical Association, 99(468):??, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590510.

Anonymous:2004:BRB

[Ano04e] Anonymous. Book review: Regression Models for Time Series Analysis. Benjamin Kedem and Konstantinos Fokianos. Journal of the American Statistical Association, 99(465):299–??, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s36/p299.

Anonymous:2004:BR

[Ano04f] Anonymous. Book reviews. Journal of the American Statistical Association, 99(467):897–??, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s41/p897.

Anonymous:2004:FMa

[Ano04g] Anonymous. Front matter. Journal of the American Statistical Association, 99(465):??, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590346.

Anonymous:2004:FMb

[Ano04h] Anonymous. Front matter. Journal of the American Statistical Association, 99(466):??, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590386.

Anonymous:2004:FMc

[Ano04i] Anonymous. Front matter. Journal of the American Statistical Association, 99(467):??, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590431.

Anonymous:2004:FMd

[Ano04j] Anonymous. Front matter. Journal of the American Statistical Association, 99(468):??, December 2004. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590473.

Anonymous:2004:TRa

[Ano04k] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 99(465):300–301, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s37/p300.

Anonymous:2004:TRb

[Ano04l] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 99(466):570–573, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s42/p570.

Anonymous:2004:TRc

[Ano04m] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 99(467):905–908, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s53/p905.

Anonymous:2004:TRd

[Ano04n] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 99(468):1209–1210, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s42/p1209.

Anonymous:2004:VI

[Ano04o] Anonymous. Volume information. Journal of the American Statistical Association, 99(468):??, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590472.

Anonymous:2005:EC

[Ano05a] Anonymous. 2005 editorial collaborators. Journal of the American Statistical Association, 100(472):1470–1481, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2005:BMa

[Ano05b] Anonymous. Back matter. Journal of the American Statistical Association, 100(469):??, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590557.

Anonymous:2005:BMb

[Ano05c] Anonymous. Back matter. Journal of the American Statistical Association, 100(470):??, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590606.

Anonymous:2005:BMc

[Ano05d] Anonymous. Back matter. Journal of the American Statistical Association, 100(471):??, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590653.

Anonymous:2005:BMd

[Ano05e] Anonymous. Back matter. Journal of the American Statistical Association, 100(472):??, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590702.

Anonymous:2005:BRa

[Ano05f] Anonymous. Book reviews. Journal of the American Statistical Association, 100(471):1090, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2005:BRb

[Ano05g] Anonymous. Book reviews. Journal of the American Statistical Association, 100(472):1456, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2005:Ca

[Ano05h] Anonymous. Correction. Journal of the American Statistical Association, 100(470):716–718, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s46/p716.

Anonymous:2005:Cb

[Ano05i] Anonymous. Correction. Journal of the American Statistical Association, 100(471):1101, September 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2005:CFE

[Ano05j] Anonymous. Correction: Fully exponential Laplace approximations using asymptotic modes. Journal of the American Statistical Association, 100(471):1101, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590652.

Anonymous:2005:CNH

[Ano05k] Anonymous. Correction: Nonparametric hypothesis testing for a spatial signal. Journal of the American Statistical Association, 100(470):716–718, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590605.

Anonymous:2005:FMa

[Ano05l] Anonymous. Front matter. Journal of the American Statistical Association, 100(469):??, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590512.

Anonymous:2005:FMb

[Ano05m] Anonymous. Front matter. Journal of the American Statistical Association, 100(470):??, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590559.

Anonymous:2005:FMc

[Ano05n] Anonymous. Front matter. Journal of the American Statistical Association, 100(471):??, September 2005. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590608.

Anonymous:2005:FMd

[Ano05o] Anonymous. Front matter. Journal of the American Statistical Association, 100(472):??, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590656.

Anonymous:2005:TRa

[Ano05p] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 100(469):357–??, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s50/
p357.

Anonymous:2005:TRb

[Ano05q] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 100(470):713–715, June 2005. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s45/
p713.

Anonymous:2005:TRd

[Ano05r] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 100(472):1466–1469, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2005:VI

[Ano05s] Anonymous. Volume information. Journal of the American Statistical Association, 100(472):??, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590655.

Anonymous:2006:EC

[Ano06a] Anonymous. 2006 editorial collaborators. Journal of the American Statistical Association, 101(476):1737–1741, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:BMa

[Ano06b] Anonymous. Back matter. Journal of the American Statistical Association, 101(473):??, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047496.

Anonymous:2006:BMb

[Ano06c] Anonymous. Back matter. Journal of the American Statistical Association, 101(474):??, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590764.

Anonymous:2006:BMc

[Ano06d] Anonymous. Back matter. Journal of the American Statistical Association, 101(475):??, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590823.

Anonymous:2006:BMd

[Ano06e] Anonymous. Back matter. Journal of the American Statistical Association, 101(476):??, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27639811.

Anonymous:2006:BRa

[Ano06f] Anonymous. Book reviews. Journal of the American Statistical Association, 101(473):387, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:BRb

[Ano06g] Anonymous. Book reviews. Journal of the American Statistical Association, 101(474):843, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:BRc

[Ano06h] Anonymous. Book reviews. Journal of the American Statistical Association, 101(475):1306, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:BRd

[Ano06i] Anonymous. Book reviews. Journal of the American Statistical Association, 101(476):1718, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:Ca

[Ano06j] Anonymous. Correction. Journal of the American Statistical Association, 101(473):407, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:Cb

[Ano06k] Anonymous. Correction. Journal of the American Statistical Association, 101(475):1321, September 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:CIA

[Ano06l] Anonymous. Correction: Inference after model selection. Journal of the American Statistical Association, 101(475): 1321, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590822.

Anonymous:2006:FMa

[Ano06m] Anonymous. Front matter. Journal of the American Statistical Association, 101(473):??, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047434.

Anonymous:2006:FMb

[Ano06n] Anonymous. Front matter. Journal of the American Statistical Association, 101(474):??, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590704.

Anonymous:2006:FMc

[Ano06o] Anonymous. Front matter. Journal of the American Statistical Association, 101(475):??, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590766.

Anonymous:2006:FMd

[Ano06p] Anonymous. Front matter. Journal of the American Statistical Association, 101(476):??, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27639753.

Anonymous:2006:TRa

[Ano06q] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 101(473):405, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:TRb

[Ano06r] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 101(474):859, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:TRc

[Ano06s] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 101(475):1318–1320, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:TRd

[Ano06t] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 101(476):1735–1736, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2006:VI

[Ano06u] Anonymous. Volume information. Journal of the American Statistical Association, 101(476):??, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639752.

Anonymous:2007:EC

[Ano07a] Anonymous. 2007 editorial collaborators. Journal of the American Statistical Association, 102(480):1487–1491, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:BRa

[Ano07b] Anonymous. Book reviews. Journal of the American Statistical Association, 102(477):379, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:BRb

[Ano07c] Anonymous. Book reviews. Journal of the American Statistical Association, 102(478):757, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:BRc

[Ano07d] Anonymous. Book reviews. Journal of the American Statistical Association, 102(479):1073, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:BRd

[Ano07e] Anonymous. Book reviews. Journal of the American Statistical Association, 102(480):1472, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:TRa

[Ano07f] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 102(477):391–393, March 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:TRb

[Ano07g] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 102(478):768–769, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:TRc

[Ano07h] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 102(479):1080–1083, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2007:TRd

[Ano07i] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 102(480):1481–1484, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:EC

[Ano08a] Anonymous. 2008 editorial collaborators. Journal of the American Statistical Association, 103(484):1730–1734, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:BRa

[Ano08b] Anonymous. Book reviews. Journal of the American Statistical Association, 103(481):424, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:BRb

[Ano08c] Anonymous. Book reviews. Journal of the American Statistical Association, 103(484):1705–1728, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:Ca

[Ano08d] Anonymous. Correction. Journal of the American Statistical Association, 103(482):896, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:Cb

[Ano08e] Anonymous. Corrections. Journal of the American Statistical Association, 103(484):1729, December 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:IV

[Ano08f] Anonymous. Index to volume 103 (2008). Journal of the American Statistical Association, 103(484):1735–1743, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:TRa

[Ano08g] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 103(481):433–436, March 2008. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:TRb

[Ano08h] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 103(482):891–893, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2008:TRc

[Ano08i] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 103(483):1328–1330, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:EC

[Ano09a] Anonymous. 2009 editorial collaborators. Journal of the American Statistical Association, 104(488):1729–1734, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:BRa

[Ano09b] Anonymous. Book reviews. Journal of the American Statistical Association, 104(485):409–426, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:BRb

[Ano09c] Anonymous. Book reviews. Journal of the American Statistical Association, 104(486):855–871, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:BRc

[Ano09d] Anonymous. Book reviews. Journal of the American Statistical Association, 104(487):1272–1291, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:BRd

[Ano09e] Anonymous. Book reviews. Journal of the American Statistical Association, 104(488):1713–1727, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:Ca

[Ano09f] Anonymous. Correction. Journal of the American Statistical Association, 104(485):427, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:Cd

[Ano09g] Anonymous. Correction. Journal of the American Statistical Association, 104(488):1728, December 2009. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:Cb

[Ano09h] Anonymous. Corrections. Journal of the American Statistical Association, 104(486):872, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:Cc

[Ano09i] Anonymous. Corrections. Journal of the American Statistical Association, 104(487):1292, September 2009. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2009:IV

[Ano09j] Anonymous. Index to volume 104 (2009). Journal of the American Statistical Association, 104(488):1735–1744, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:2004:HOW

[AP04] M. G. Akritas and N. Papadatos. Heteroscedastic oneway ANOVA and lack-of-fit tests. Journal of the American Statistical Association, 99(466):368–382, April 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/s7/
p368;   http://www.jstor.org/stable/27590393.

Aquila:2006:BRB

[Aqu06] Rosario Dell’ Aquila. Book review: Generalized Method of Moments, by Alastair R. Hall. Journal of the American Statistical Association, 101(475):1309–1310, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590806.

Ait-Sahalia:2009:NTB

[ASFP09] Yacine Aït-Sahalia, Jianqing Fan, and Heng Peng. Nonparametric transition-based tests for jump diffusions. Journal of the American Statistical Association, 104(487):1102–1116, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Assaid:2007:BRB

[Ass07] Christopher Assaid. Book review: The Theory of Response-Adaptive Randomization in Clinical Trials. Journal of the American Statistical Association, 102(480):1479, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:2001:AMH

[AV01] Michael G. Akritas and Ingrid Van Keilegom. ANCOVA methods for heteroscedastic nonparametric regression models. Journal of the American Statistical Association, 96(453):220–232, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670361;   reqidx=/cw/asa/01621459/v96n453/
s17/p220.

Bailey:2004:BRB

[Bai04] R. A. Bailey. Book review: Block Designs: A Randomization Approach, by Tadeusz Calínski; Sanpei Kageyama. Journal of the American Statistical Association, 99(466):558–560, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s25/p558;   http://www.jstor.org/stable/27590411.

Baiocchi:2005:BRB

[Bai05] Giovanni Baiocchi. Book review: Computational Methods in Statistics and Econometrics, by Hisashi Tanizaki. Journal of the American Statistical Association, 100(470):705–707, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s34/
p705;   http://www.jstor.org/stable/27590593.

Baker:2000:ARC

[Bak00] Stuart G. Baker. Analyzing a randomized cancer prevention trial with a missing binary outcome, an auxiliary variable, and all-or-none compliance. Journal of the American Statistical Association, 95(449):43–50, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
Baker.htm;   http://www.jstor.org/stable/2669521.

Baker:2007:LE

[Bak07] Stuart G. Baker. Letter to the Editor. Journal of the American Statistical Association, 102(477):394, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baltagi:2003:BRB

[Bal03] Badi H. Baltagi. Book review: Panel Data Econometrics: Method-of-Moments and Limited Dependent Variables, by Myoung-Jae Lee. Journal of the American Statistical Association, 98(463):769–770, September 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30045310.

Ballman:2005:BRB

[Bal05] Karla V. Ballman. Book review: Discrete Choice Methods with Simulations, by Kenneth E. Train. Journal of the American Statistical Association, 100(469):351–352, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s42/
p352;   http://www.jstor.org/stable/27590548.

Banks:2005:BRB

[Ban05] David L. Banks. Book review: Foundations of Risk Analysis: A Knowledge and Decision-Oriented Perspective, by Terje Aven. Journal of the American Statistical Association, 100(470):703–704, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s30/p703;   http://www.
jstor.org/stable/27590589.

Banerjee:2006:BRB

[Ban06] Tathagata Banerjee. Book review: Exploring Multivariate Data with the Forward Search, by Anthony C. Atkinson; Marco Riani; Andrea Cerioli. Journal of the American Statistical Association, 101(473):398, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047480.

Banerjee:2007:BRB

[Ban07] Tathagata Banerjee. Book review: Introductory Statistical Inference. Journal of the American Statistical Association, 102 (480):1474, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barbour:2001:BRB

[Bar01a] A. D. Barbour. Book review: [untitled]. Journal of the American Statistical Association, 96(454):780, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670321.

Bartlett:2001:BRB

[Bar01b] Roy F. Bartlett. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1134–1135, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670259.

Barber:2008:BRB

[Bar08] Jarrett J. Barber. Book review: Bayesian Core: A Practical Approach to Computational Bayesian Statistics. Journal of the American Statistical Association, 103(481):432–433, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bassett:2003:BRB

[Bas03] Roland Bassett. Book review: Debugging SAS Programs: A Handbook of Tools and Techniques, by Michele M. Burlew. Journal of the American Statistical Association, 98(464):1087, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s40/p1087;   http://www.
jstor.org/stable/30045361.

Bauerle:2002:ISN

[Bau02] Nicole Bauerle. Introduction to stochastic networks. Journal of the American Statistical Association, 97(460):1212–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n460/s39/p1212.

Bergeron:2008:CBI

[BAW08] Pierre-Jérôme Bergeron, Masoud Asgharian, and David B. Wolfson. Covariate bias induced by length-biased sampling of failure times. Journal of the American Statistical Association, 103(482):737–742, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banerjee:1976:PIM

[BB76] A. K. Banerjee and G. K. Bhattacharyya. A purchase incidence model with inverse Gaussian interpurchase times. Journal of the American Statistical Association, 71(356):823– 829, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286845. See remarks [BH08].

Bayarri:2000:ADV

[BB00a] M. J. Bayarri and J. O. Berger. Asymptotic distribution of P values in composite null models: Rejoinder. Journal of the American Statistical Association, 95(452):1168–1170, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669756.

Bayarri:2000:VCN

[BB00b] M. J. Bayarri and James O. Berger. P values for composite null models. Journal of the American Statistical Association, 95(452):1127–1142, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669749.

Bebchuk:2001:LLA

[BB01] Judith D. Bebchuk and Rebecca A. Betensky. Local likelihood analysis of survival data with censored intermediate events. Journal of the American Statistical Association, 96(454):449–457, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670283;   reqidx=/cw/
asa/01621459/v96n454/s9/p449.

Brookmeyer:2003:SMB

[BB03] Ron Brookmeyer and Natalie Blades. Statistical models and bioterrorism: Application to the U.S. anthrax outbreak. Journal of the American Statistical Association, 98 (464):781–788, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s1/p781;   http://www.
jstor.org/stable/30045329.

Barsky:2002:ABW

[BBCL02] Robert Barsky, John Bound, Kerwin Kofi Charles, and Joseph P. Lupton. Accounting for the black-white wealth gap: a nonparametric approach. Journal of the American Statistical Association, 97(459):663–673, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s1/p663;   http://www.jstor.org/stable/3085702.

Boatwright:2003:MJD

[BBK03] Peter Boatwright, Sharad Borle, and Joseph B. Kadane. A model of the joint distribution of purchase quantity and timing. Journal of the American Statistical Association, 98(463):564–572, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s7/p564;   http://
www.jstor.org/stable/30045282.

Bayarri:2009:PVC

[BBK^{+}09] M. J. Bayarri, James O. Berger, Marc C. Kennedy, Athanasios Kottas, Rui Paulo, Jerry Sacks, John A. Cafeo, Chin-Hsu Lin, and Jian Tu. Predicting vehicle crashworthiness: Validation of computer models for functional and hierarchical data. Journal of the American Statistical Association, 104(487):929–943, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berry:2004:BSA

[BBN^{+}04] Scott M. Berry, Donald A. Berry, Kannan Natarajan, Chen-Sheng Lin, Charles H. Hennekens, and Rene Belder. Bayesian survival analysis with nonproportional hazards: Metanalysis of combination pravastatin-aspirin. Journal of the American Statistical Association, 99(465):36–44, March 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=

1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/s5/
p36;   http://www.jstor.org/stable/27590351.

Black:2000:BPE

[BBS00] Dan A. Black, Mark C. Berger, and Frank A. Scott. Bounding parameter estimates with nonclassical measurement error. Journal of the American Statistical Association, 95(451):739– 748, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669454.

Bura:2001:ESI

[BC01] Efstathia Bura and R. Dennis Cook. Extending sliced inverse regression: the weighted chi-squared test. Journal of the American Statistical Association, 96(455):996–1003, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670245;   reqidx=/cw/asa/01621459/v96n455/
s17/p996.

Basu:2003:MLB

[BC03a] Sanjib Basu and Siddhartha Chib. Marginal likelihood and Bayes factors for Dirichlet process mixture models. Journal of the American Statistical Association, 98 (461):224–235, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s22/p224;   http://
www.jstor.org/stable/30045209.

BoraanAruoba:2003:TSA

[BC03b] S. Bora an Aruoba and Sean D. Campbell. Time series analysis by state-space methods. Journal of the American Statistical Association, 98(461):255–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s36/p255.

Briesch:2002:SEB

[BCM02] Richard A. Briesch, Pradeep K. Chintagunta, and Rosa L. Matzkin. Semiparametric estimation of brand choice behavior. Journal of the American Statistical Association, 97 (460):973–982, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s5/p973;   http://
www.jstor.org/stable/3085819.

Berry:2002:BSR

[BCR02] Scott M. Berry, Raymond J. Carroll, and David Ruppert. Bayesian smoothing and regression splines for measurement error problems. Journal of the American Statistical Association, 97(457):160–169, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085771;   reqidx=/cw/
asa/01621459/v97n457/s14/p160.

Bortot:2007:ISE

[BCS07] P. Bortot, S. G. Coles, and S. A. Sisson. Inference for stereological extremes. Journal of the American Statistical Association, 102(477):84–92, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Billard:2003:SDS

[BD03] L. Billard and E. Diday. From the statistics of data to the statistics of knowledge: Symbolic data analysis. Journal of the American Statistical Association, 98(462):470–487, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s18/p470;   http://www.jstor.org/stable/30045255.

Burr:2005:BSM

[BD05] Deborah Burr and Hani Doss. A Bayesian semiparametric model for random-effects meta-analysis. Journal of the American Statistical Association, 100(469):242–251, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s28/
p242;   http://www.jstor.org/stable/27590534.

Bigelow:2009:BSJ

[BD09] Jamie L. Bigelow and David B. Dunson. Bayesian semiparametric joint models for functional predictors. Journal of the American Statistical Association, 104(485):26–36, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:2001:OBA

[BDS01] James O. Berger, Victor De Oliveira, and Bruno Sansó. Objective Bayesian analysis of spatially correlated data. Journal of the American Statistical Association, 96(456): 1361–1374, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085905;   reqidx=/cw/
asa/01621459/v96n456/s18/p1361.

Biedermann:2006:ODD

[BDZ06] Stefanie Biedermann, Holger Dette, and Wei Zhu. Optimal designs for dose–response models with restricted design spaces. Journal of the American Statistical Association, 101 (474):747–759, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590732.

Beck:2000:PMW

[Bec00a] Nathaniel L. Beck. Political methodology: a welcoming discipline. Journal of the American Statistical Association, 95 (450):651–654, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669411.

Becker:2000:I

[Bec00b] Mark P. Becker. [introduction]. Journal of the American Statistical Association, 95(450):628–629, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669405.

Beirlant:2004:BRB

[Bei04] Jan Beirlant. Book review: Statistical Size Distributions in Economics and Actuarial Sciences, by Christian Kleiber; Samuel Kotz. Journal of the American Statistical Association, 99(468):1206–1207, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/

vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s37/p1206;   http://www.jstor.org/stable/27590504.

Beran:2000:RSS

[Ber00a] Rudolf Beran. REACT scatterplot smoothers: Superefficiency through basis economy. Journal of the American Statistical Association, 95(449):155–171, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
Beran.htm;   http://www.jstor.org/stable/2669535.

Berger:2000:BAL

[Ber00b] James O. Berger. Bayesian analysis: a look at today and thoughts of tomorrow. Journal of the American Statistical Association, 95(452):1269–1276, December 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669768.

Berry:2000:BRB

[Ber00c] Donald A. Berry. Book review: [untitled]. Journal of the American Statistical Association, 95(449):339, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669571.

Berk:2001:BRB

[Ber01] Kenneth N. Berk. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1531, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085932.

Berenhaut:2004:BRB

[Ber04] Kenneth S. Berenhaut. Book review: Advanced Calculus with Applications in Statistics (second edition), by Andrè I. Khuri. Journal of the American Statistical Association, 99(467):903–904, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s51/p903;   http:
//www.jstor.org/stable/27590467.

Betensky:2003:BRB

[Bet03] Rebecca A. Betensky. Book review: Analysis of Failure and Survival Data, by Peter J. Smith. Journal of the American Statistical Association, 98(463):771–772, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045314.

Bradlow:2001:BLM

[BF01a] Eric T. Bradlow and Peter S. Fader. A Bayesian lifetime model for the “Hot 100” Billboard songs. Journal of the American Statistical Association, 96(454):368–381, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670276;   reqidx=/cw/asa/01621459/v96n454/s2/p368.

Braun:2001:OPT

[BF01b] Thomas M. Braun and Ziding Feng. Optimal permutation tests for the analysis of group randomized trials. Journal of the American Statistical Association, 96(456):1424–1432, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085910;   reqidx=/cw/asa/01621459/v96n456/
s23/p1424.

Bartolucci:2002:ERA

[BF02] Francesco Bartolucci and Antonio Forcina. Extended RC association models allowing for order restrictions and marginal modeling. Journal of the American Statistical Association, 97(460):1192–1199, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s22/p1192;   http:/
/www.jstor.org/stable/3085841.

Bartolucci:2006:CLM

[BF06] Francesco Bartolucci and Antonio Forcina. A class of latent marginal models for capture–recapture data with continuous covariates. Journal of the American Statistical Association, 101(474):786–794, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590735.

Bartolucci:2009:MED

[BF09] Francesco Bartolucci and Alessio Farcomeni. A multivariate extension of the dynamic logit model for longitudinal data based on a latent Markov heterogeneity structure. Journal of the American Statistical Association, 104(486):816–831, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buyske:2000:CWL

[BFB00] Steven Buyske, Richard Fagerstrom, and Zhiliang Ying Steven Buyske. A class of weighted log-rank tests for survival data when the event is rare. Journal of the American Statistical Association, 95(449):249–258, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
Buyske.htm;   http://www.jstor.org/stable/2669542.

Bartolucci:2001:PQD

[BFD01] Francesco Bartolucci, Antonio Forcina, and Valentino Dardanoni. Positive quadrant dependence and marginal modeling in two-way tables with ordered margins. Journal of the American Statistical Association, 96(456):1497–1505, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085916;   reqidx=/cw/asa/01621459/v96n456/
s29/p1497.

Barnard:2003:PSAa

[BFHR03a] John Barnard, Constantine E. Frangakis, Jennifer L. Hill, and Donald B. Rubin. Principal stratification approach to broken randomized experiments: a case study of school choice vouchers in New York City. Journal of the American Statistical Association, 98(462):299–311, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s5/p299;   http://www.jstor.org/stable/30045238.

Barnard:2003:PSAb

[BFHR03b] John Barnard, Constantine E. Frangakis, Jennifer L. Hill, and Donald B. Rubin. Principal stratification approach to broken randomized experiments: a case study of school choice vouchers in New York City [with rejoinder]. Journal of the American Statistical Association, 98(462):320–323, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045242.

Brown:2001:BWR

[BFV01] P. J. Brown, T. Fearn, and M. Vannucci. Bayesian wavelet regression on curves with application to a spectroscopic calibration problem. Journal of the American Statistical Association, 96(454):398–408, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670278;   reqidx=/cw/
asa/01621459/v96n454/s4/p398.

Berger:2001:BCF

[BG01] James O. Berger and Alessandra Guglielmi. Bayesian and conditional frequentist testing of a parametric model versus nonparametric alternatives. Journal of the American Statistical Association, 96(453):174–184, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670357;   reqidx=/cw/
asa/01621459/v96n453/s13/p174.

Banerjee:2006:BW

[BG06a] Sudipto Banerjee and Alan E. Gelfand. Bayesian wombling. Journal of the American Statistical Association, 101(476): 1487–1501, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banerjee:2006:BWC

[BG06b] Sudipto Banerjee and Alan E. Gelfand. Bayesian wombling: Curvilinear gradient assessment under spatial process models. Journal of the American Statistical Association, 101 (476):1487–1501, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639767.

Billheimer:2001:SIS

[BGF01] Dean Billheimer, Peter Guttorp, and William F. Fagan. Statistical interpretation of species composition. Journal of the American Statistical Association, 96(456):1205–1214, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085883;   reqidx=/cw/asa/01621459/v96n456/
s7/p1205.

Brown:2005:SAT

[BGM^{+}05] Lawrence Brown, Noah Gans, Avishai Mandelbaum, Anat Sakov, Haipeng Shen, Sergey Zeltyn, and Linda Zhao. Statistical analysis of a telephone call center: a queueingscience perspective. Journal of the American Statistical Association, 100(469):36–50, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s5/p36;   http://www.
jstor.org/stable/27590517.

Banerjee:2003:DRC

[BGS03] Sudipto Banerjee, Alan E. Gelfand, and C. F. Sirmans. Directional rates of change under spatial process models. Journal of the American Statistical Association, 98 (464):946–954, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s20/p946;   http://www.
jstor.org/stable/30045341.

Benjamini:2007:FDR

[BH07] Yoav Benjamini and Ruth Heller. False discovery rates for spatial signals. Journal of the American Statistical Association, 102(480):1272–1281, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bemmaor:2008:BBP

[BH08] Albert C. Bemmaor and Daniel Hoppe. On Banerjee and Bhattacharyya (1976), “A Purchase Incidence Model With Inverse Gaussian Interpurchase Times,” Journal of the American Statistical Association, 71, 823–829. Journal of the American Statistical Association, 103(482):894–895, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BB76].

Bhattacharya:2009:IOP

[Bha09] Debopam Bhattacharya. Inferring optimal peer assignment from experimental data. Journal of the American Statistical Association, 104(486):486–500, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhlmann:2002:SBV

[Bhl02] Peter Bhlmann. [sieve bootstrap with variable-length Markov chains for stationary categorical time series]: Rejoinder. Journal of the American Statistical Association, 97(458):466–471, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085662.

Breidt:2007:SMM

[BHO07] F. Jay Breidt, Nan-Jung Hsu, and Stephen Ogle. Semiparametric mixed models for increment-averaged data with application to carbon sequestration in agricultural soils. Journal of the American Statistical Association, 102(479):803–812, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bair:2006:PSP

[BHPT06] Eric Bair, Trevor Hastie, Debashis Paul, and Robert Tibshirani. Prediction by supervised principal components. Journal of the American Statistical Association, 101(473):119–137, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047444.

Brown:2000:PCD

[BIK00] C. Hendricks Brown, Alka Indurkhya, and Sheppard G. Kellam. Power calculations for data missing by design: Applications to a follow-up study of lead exposure and attention. Journal of the American Statistical Association, 95(450): 383–395, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/2000/June/Brown.htm;   http:/
/www.jstor.org/stable/2669375.

Bingham:2002:BRB

[Bin02] Nick Bingham. Book review: [untitled]. Journal of the American Statistical Association, 97(457):365, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085806.

Bingham:2006:BRBa

[Bin06a] N. H. Bingham. Book review: Financial Modelling with Jump Processes, by Rama Cont; Peter Tankov. Journal of the American Statistical Association, 101(475):1315–1316, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590815.

Bingham:2006:BRBb

[Bin06b] N. H. Bingham. Book review: Functional Analysis for Probability and Stochastic Processes: An Introduction, by Adam Bobrowski. Journal of the American Statistical Association, 101(476):1728, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639798.

Best:2000:SPR

[BIW00] Nicola G. Best, Katja Ickstadt, and Robert L. Wolpert. Spatial Poisson regression for health and exposure data measured at disparate resolutions. Journal of the American Statistical Association, 95(452):1076–1088, December 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669744.

Bartlett:2006:CCR

[BJM06] Peter L. Bartlett, Michael I. Jordan, and Jon D. McAuliffe. Convexity, classification, and risk bounds. Journal of the American Statistical Association, 101(473):138–156, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047445.

Bunzel:2001:SRT

[BKV01] Helle Bunzel, Nicholas M. Kiefer, and Timothy J. Vogelsang. Simple robust testing of hypotheses in nonlinear models. Journal of the American Statistical Association, 96 (455):1088–1096, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670254;   reqidx=/cw/
asa/01621459/v96n455/s26/p1088.

Biswas:2006:BAI

[BL06] Swati Biswas and Shili Lin. A Bayesian approach for incorporating variable rates of heterogeneity in linkage analysis. Journal of the American Statistical Association, 101 (476):1341–1351, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639755.

Bednorz:2007:FRF

[BL07] Witold Bednorz and Krzysztof Latuszyński. A few remarks on “Fixed-Width Output Analysis for Markov Chain Monte Carlo” by Jones et al. Journal of the American Statistical Association, 102(480):1485–1486, December 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [JHCN06].

Bates:2001:SNM

[BLW01] Douglas M. Bates, Mary J. Lindstrom, and Grace Wahba. [semiparametric nonlinear mixed-effects models and their applications]: Comment. Journal of the American Statistical Association, 96(456):1292–1293, December 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/3085899.

Bolstad:2001:ICM

[BM01] William M. Bolstad and Samuel O. Manda. Investigating child mortality in Malawi using family and community random effects: a Bayesian analysis. Journal of the American Statistical Association, 96(453):12–19, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670336;   reqidx=/cw/
asa/01621459/v96n453/s2/p12.

Bowsher:2008:DEF

[BM08] Clive G. Bowsher and Roland Meeks. The dynamics of economic functions: Modeling and forecasting the yield curve. Journal of the American Statistical Association, 103(484): 1419–1437, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boldea:2009:MLE

[BM09] Otilia Boldea and Jan R. Magnus. Maximum likelihood estimation of the multivariate normal mixture model. Journal of the American Statistical Association, 104(488):1539–1549, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bouman:2007:MHM

[BMDD07] Peter Bouman, Xiao-Li Meng, James Dignam, and Vanja Dukić. A multiresolution hazard model for multicenter survival studies. Journal of the American Statistical Association, 102(480):1145–1157, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bae:2008:PPI

[BME08] Kyounghwa Bae, Bani K. Mallick, and Christine G. Elsik. Prediction of protein interdomain linker regions by a nonstationary hidden Markov model. Journal of the American Statistical Association, 103(483):1085–1099, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bingham:2009:MIM

[BNV09] Melissa A. Bingham, Daniel J. Nordman, and Stephen B. Vardeman. Modeling and inference for measured crystal orientations and a tractable class of symmetric distributions for rotations in three dimensions. Journal of the American Statistical Association, 104(488):1385–1397, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bock:2008:SHP

[Boc08] Mary Ellen Bock. Statistics: Harnessing the power of information. Journal of the American Statistical Association, 103 (484):1331–1333, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bognar:2004:BRB

[Bog04] Matthew A. Bognar. Book review: Spatial Cluster Modeling, by Andrew B. Lawson; David G. T. Denison. Journal of the American Statistical Association, 99(467):898–899, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/

vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s42/p898;   http://www.jstor.org/stable/27590458.

Boik:2001:BRB

[Boi01] Robert J. Boik. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1528–1529, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085927.

Bolthausen:2002:BRB

[Bol02] Erwin Bolthausen. Book review: [untitled]. Journal of the American Statistical Association, 97(457):360, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085798.

Boos:2000:ADV

[Boo00] Dennis D. Boos. Asymptotic distribution of P values in composite null models: Comment. Journal of the American Statistical Association, 95(452):1163–1164, December 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669753.

Boone:2004:BRB

[Boo04] E. Boone. Book review: Statistical Models and Methods for Lifetime Data (second edition), by Jerald F. Lawless. Journal of the American Statistical Association, 99(465):298–299, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s35/p298;   http:
//www.jstor.org/stable/27590381.

Booth:2006:BRB

[Boo06] James G. Booth. Book review: Analysis of Variance for Random Models: Theory, Methods, Applications, and Data Analysis, Vol. 1: Balanced Data, Analysis of Variance for Random Models: Theory, Methods, Applications, and Data Analysis, Vol. 2: Unbalanced Data., by Hardeo Sahai; Mario Miguel Ojeda. Journal of the American Statistical Association, 101(475):1310, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590807.

Booth:2007:BRB

[Boo07] James G. Booth. Book review: Exact Analysis of Discrete Data. Journal of the American Statistical Association, 102 (478):764–765, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Borkowf:2002:BRB

[Bor02] Craig B. Borkowf. Book review: [untitled]. Journal of the American Statistical Association, 97(458):651, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085680.

Borgan:2003:BRB

[Bor03] Ørnulf Borgan. Book review: The Statistical Analysis of Failure Time Data (second edition), by John D. Kalbfleisch; Ross L. Prentice. Journal of the American Statistical Association, 98(464):1084–1085, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s37/p1084;   http://www.jstor.org/stable/30045358.

Boucher:2007:BRB

[Bou07] Thomas R. Boucher. Book review: Test Equating, Scaling, and Linking: Methods and Practices. Journal of the American Statistical Association, 102(478):762–763, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bowman:2007:SMR

[Bow07] F. Dubois Bowman. Spatiotemporal models for region of interest analyses of functional neuroimaging data. Journal of the American Statistical Association, 102(478):442–453, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baker:2001:PDA

[BP01] Stuart G. Baker and Paul F. Pinsky. A proposed design and analysis for comparing digital and analog mammography: Special receiver operating characteristic methods for cancer screening. Journal of the American Statistical Association, 96(454):421–428, June 2001. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670280;   reqidx=/cw/
asa/01621459/v96n454/s6/p421.

Butler:2002:SAB

[BP02] Ronald W. Butler and Marc S. Paolella. Saddlepoint approximation and bootstrap inference for the Satterthwaite class of ratios. Journal of the American Statistical Association, 97(459):836–846, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s15/p836;   http://
www.jstor.org/stable/3085725.

Bohning:2008:CRA

[BP08] Dankmar Böhning and Valentin Patilea. A capture–recapture approach for screening using two diagnostic tests with availability of disease status for the test positives only. Journal of the American Statistical Association, 103(481):212–221, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See remarks [CN08].

Brannath:2002:RCT

[BPB02] Werner Brannath, Martin Posch, and Peter Bauer. Recursive combination tests. Journal of the American Statistical Association, 97(457):236–244, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085778;   reqidx=/cw/
asa/01621459/v97n457/s21/p236.

Bickel:2000:PLC

[BR00] Peter J. Bickel and Ya’acov Ritov. On profile likelihood: Comment. Journal of the American Statistical Association, 95 (450):466–468, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669387.

Bradstreet:2006:BRBa

[Bra06a] Thomas E. Bradstreet. Book review: The Analysis of Means: A Graphical Method for Comparing Means, Rates, and Proportions, by Peter R. Nelson; Peter S. Wludyka; Karen A.

F. Copeland. Journal of the American Statistical Association, 101(474):848–849, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590748.

Bradstreet:2006:BRBb

[Bra06b] Thomas E. Bradstreet. Book review: The Grammar of Graphics, by Leland Wilkinson. Journal of the American Statistical Association, 101(476):1719–1720, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639786.

Bradstreet:2008:BRB

[Bra08] Thomas E. Bradstreet. Book review: Bioequivalence Studies in Drug Development: Methods and Applications. Journal of the American Statistical Association, 103(483):1319–1320, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breslow:2000:SLM

[Bre00] Norman E. Breslow. Statistics in the life and medical sciences. Journal of the American Statistical Association, 95 (449):281–282, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669546.

Breidt:2005:BRB

[Bre05] Jay Breidt. Book review: Nonlinear Time Series: Nonparametric and Parametric Methods, by Jianqing Fan; Qiwei Yao. Journal of the American Statistical Association, 100(469):348–349, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s37/p348;   http://www.
jstor.org/stable/27590543.

Breymann:2006:BRB

[Bre06] Wolfgang Breymann. Book review: Theory of Financial Risk and Derivative Pricing: From Statistical Physics to Risk Management, by Jean-Philippe Bouchaud; Marc Potters. Journal of the American Statistical Association, 101(474):850–852, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590750.

Briggs:2004:CCP

[Bri04] William Briggs. Comment: Calibrated probabilistic mesoscale weather field forecasting: The geostatistical output perturbation method. Journal of the American Statistical Association, 99(467):586–587, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n467/s3/p586.

Briggs:2007:BRB

[Bri07] William (Matt) Briggs. Book review: Statistical Methods in the Atmospheric Sciences. Journal of the American Statistical Association, 102(477):380, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:2000:ESD

[Bro00a] Lawrence D. Brown. An essay on statistical decision theory. Journal of the American Statistical Association, 95(452):1277– 1281, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669769.

Browne:2000:P

[Bro00b] Michael W. Browne. Psychometrics. Journal of the American Statistical Association, 95(450):661–665, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669413.

Brown:2001:BRB

[Bro01] Lawrence Brown. Book review: [untitled]. Journal of the American Statistical Association, 96(454):778, June 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670316.

Brownstone:2007:BRB

[Bro07] David Brownstone. Book review: Applied Choice Analysis: A Primer. Journal of the American Statistical Association, 102 (477):390, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brumback:2000:TRM

[BRS^{+}00] Babette A. Brumback, Louise M. Ryan, Joel D. Schwartz, Lucas M. Neas, Paul C. Stark, and Harriet A. Burge. Transitional regression models, with application to environmental time series. Journal of the American Statistical Association, 95(449):16–27, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/Brumback.htm;   http:/
/www.jstor.org/stable/2669519.

Benjamin:2003:GAM

[BRS03] Michael A. Benjamin, Robert A. Rigby, and D. Mikis Stasinopoulos. Generalized autoregressive moving average models. Journal of the American Statistical Association, 98(461):214–223, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s21/p214;   http://
www.jstor.org/stable/30045208.

Burman:2006:GEP

[BS06] Prabir Burman and Robert H. Shumway. Generalized exponential predictors for time series forecasting. Journal of the American Statistical Association, 101(476):1598–1606, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27639775.

Black:2003:MHE

[BST03] Dan Black, Seth Sanders, and Lowell Taylor. Measurement of higher education in the census and current population survey. Journal of the American Statistical Association, 98(463):545–554, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s5/p545;   http://
www.jstor.org/stable/30045280.

Bekele:2004:DFB

[BT04] B. Nebiyou Bekele and Peter F. Thall. Dose-finding based on multiple toxicities in a soft tissue sarcoma trial. Journal of the American Statistical Association, 99(465):26–35, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/s4/
p26;   http://www.jstor.org/stable/27590350.

Bilder:2008:BRB

[BT08] Christopher R. Bilder and Joshua M. Tebbs. Book review: An Introduction to Categorical Data Analysis. Journal of the American Statistical Association, 103(483):1323, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buerle:2002:BRB

[Bue02] Nicole Buerle. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1212–1213, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085858.

Buhlmann:2000:BRB

[Buh00] Hans Buhlmann. Book review: [untitled]. Journal of the American Statistical Association, 95(450):686, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669427.

Buhlmann:2002:SBV

[Büh02] Peter Bühlmann. Sieve bootstrap with variable-length Markov chains for stationary categorical time series. Journal of the American Statistical Association, 97(458):443–456, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s8/p443;   http://www.jstor.org/stable/3085657.

Buot:2006:BRB

[Buo06] Max Buot. Book review: Probability and Computing: Randomized Algorithms and Probabilistic Analysis, by Michael Mitzenmacher; Eli Upfal. Journal of the American Statistical Association, 101(473):395–396, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047475.

Burman:2004:CEP

[Bur04] Prabir Burman. Comment: The estimation of prediction error: Covariance penalties and cross-validation. Journal of the American Statistical Association, 99(467):632–633, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s10/p632.

Burr:2008:BRB

[Bur08] Thomas Burr. Book review: Pattern Recognition and Machine Learning. Journal of the American Statistical Association, 103 (482):886–887, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Butler:2000:BRB

[But00] Ronald W. Butler. Book review: [untitled]. Journal of the American Statistical Association, 95(450):682–683, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669424.

Buttrey:2006:BRB

[But06] Samuel E. Buttrey. Book review: Clustering for Data Mining: A Data Recovery Approach, by Boris Mirkin. Journal of the American Statistical Association, 101(474):854–855, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590755.

Berk:2003:PSA

[BX03] Richard A. Berk and Hongquan Xu. Principal stratification approach to broken randomized experiments: a case study of school choice vouchers in New York City [with comment]. Journal of the American Statistical Association, 98 (462):318–320, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045241.

Buhlmann:2003:BLR

[BY03] Peter Bühlmann and Bin Yu. Boosting with the L_{2} loss: Regression and classification. Journal of the American Statistical Association, 98(462):324–339, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s6/p324;   http://www.jstor.org/stable/30045243.

Benjamini:2005:RFD

[BY05] Yoav Benjamini and Daniel Yekutieli. Rejoinder: False discovery rate-adjusted multiple confidence intervals for selected parameters. Journal of the American Statistical Association, 100(469):90–93, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s14/p90.

Benjamini:2005:FDR

[BYE^{+}05] Yoav Benjamini, Daniel Yekutieli, Don Edwards, Juliet Popper Shaffer, Ajit C. Tamhane, Peter H. Westfall, and Burt Holland. False discovery rate: Adjusted multiple confidence intervals for selected parameters [with comments, rejoinder]. Journal of the American Statistical Association, 100(469):71–81, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s8/p71;   http://www.
jstor.org/stable/27590520.

Boos:2000:MCE

[BZ00] Dennis D. Boos and Ji Zhang. Monte Carlo evaluation of resampling-based hypothesis tests. Journal of the American Statistical Association, 95(450):486–492, June 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
2000/June/Boos.htm;   http://www.jstor.org/stable/
2669393.

Christensen:2002:LVA

[CA02] William F. Christensen and Yasuo Amemiya. Latent variable analysis of multivariate spatial data. Journal of the American Statistical Association, 97(457):302–317, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/

stable/3085784;   reqidx=/cw/asa/01621459/v97n457/
s27/p302.

Chen:2008:BCR

[CAC^{+}08] Jinbo Chen, Rajeev Ayyagari, Nilanjan Chatterjee, David Y. Pee, Catherine Schairer, Celia Byrne, Jacques Benichou, and Mitchell H. Gail. Breast cancer relative hazard estimates from case–control and cohort designs with missing data on mammographic density. Journal of the American Statistical Association, 103(483):976–988, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Caffo:2007:BRB

[Caf07] Brian Caffo. Book review: A First Course in Monte Carlo. Journal of the American Statistical Association, 102(478):758, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2001:RWA

[Cai01] T. Tony Cai. Regularization of wavelet approximations: Discussion. Journal of the American Statistical Association, 96 (455):960–962, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670240.

Callas:2001:BRB

[Cal01] Peter Callas. Book review: [untitled]. Journal of the American Statistical Association, 96(453):349–351, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670383.

Candes:2000:BRB

[Can00] Emmanuel J. Candes. Book review: [untitled]. Journal of the American Statistical Association, 95(450):682, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669423.

Canty:2007:BRB

[Can07] Angelo J. Canty. Book review: Bootstrap Techniques for Signal Processing. Journal of the American Statistical Association, 102(480):1478–1479, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carey:2001:BRB

[Car01] V. J. Carey. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1135–1136, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670260.

Carey:2003:BRB

[Car03a] Vincent J. Carey. Book review: Graphical Models: Methods for Data Analysis and Mining, by Christian Borgelt; Rudolf Kruse. Journal of the American Statistical Association, 98(461):253–254, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045220.

Carey:2003:GMM

[Car03b] Vincent J. Carey. Graphical models: Methods for data analysis and mining. Journal of the American Statistical Association, 98(461):253–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s33/p253.

Carey:2005:BRBa

[Car05a] Vincent J. Carey. Book review: Mathematical and Statistical Methods for Genetic Analysis (second edition), by Kenneth Lange. Journal of the American Statistical Association, 100(470):712, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s42/p712;   http://miranda.
asa.catchword.org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/
01621459/v100n470/s43/p712;   http://www.jstor.org/
stable/27590601.

Carey:2005:BRBb

[Car05b] Vincent J. Carey. Book review: Resampling Methods for Dependent Data, by S. N. Lahiri. Journal of the American Statistical Association, 100(470):712–713, June 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s42/p712;   http:
//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/

rpsv/cw/asa/01621459/v100n470/s43/p712;   http://www.
jstor.org/stable/27590601;   http://www.jstor.org/
stable/27590602.

Carlton:2006:BRB

[Car06] Matthew Carlton. Book review: Probability and Probabilistic Reasoning for Electrical Engineering, by Terrence L. Fine. Journal of the American Statistical Association, 101(473):396, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047476.

Carey:2007:BRB

[Car07] Vincent J. Carey. Book review: Longitudinal Data Analysis. Journal of the American Statistical Association, 102(479): 1075, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Casella:2000:A

[Cas00a] George Casella. Afterword. Journal of the American Statistical Association, 95(452):1368, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669790.

Casella:2000:F

[Cas00b] George Casella. Foreword. Journal of the American Statistical Association, 95(452):1269, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669767.

Caulkins:2001:SDS

[Cau01] Jonathan P. Caulkins. [should the DEA’s STRIDE data be used for economic analyses of markets for illegal drugs?]: Comment. Journal of the American Statistical Association, 96 (456):1263–1264, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085889.

Cavanaugh:2004:BRB

[Cav04] Joseph Cavanaugh. Book review: Regression Models for Time Series Analysis, by Benjamin Kedem; Konstantinos Fokianos. Journal of the American Statistical Association, 99(465):299, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590382.

Cavanaugh:2006:BRBb

[Cav06a] Joseph Cavanaugh. Book review: Encyclopedia of Biostatistics, by Peter Armitage; Theodore Colton. Journal of the American Statistical Association, 101(474):844, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590740.

Cavanaugh:2006:BRBa

[Cav06b] Joseph Cavanaugh. Book review: Handbook of Epidemiology, by W. Ahrens; I. Pigeot. Journal of the American Statistical Association, 101(473):402–403, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047490.

Cavanaugh:2007:BRB

[Cav07] Joseph Cavanaugh. Book review: Encyclopedia of Statistical Sciences. Journal of the American Statistical Association, 102 (479):1074–1075, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2009:LMS

[CB09] Lisha Chen and Andreas Buja. Local multidimensional scaling for nonlinear dimension reduction, graph drawing, and proximity analysis. Journal of the American Statistical Association, 104(485):209–219, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cooner:2007:FCR

[CBCS07] Freda Cooner, Sudipto Banerjee, Bradley P. Carlin, and Debajyoti Sinha. Flexible cure rate modeling under latent activation schemes. Journal of the American Statistical Association, 102 (478):560–572, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Corkrey:2008:BCR

[CBL^{+}08] Ross Corkrey, Steve Brooks, David Lusseau, Kim Parsons, John W. Durban, Philip S. Hammond, and Paul M. Thompson. A Bayesian capture–recapture population model with simultaneous estimation of heterogeneity. Journal of the American Statistical Association, 103(483):948–960, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:2000:IRO

[CC00] R. Dennis Cook and Frank Critchley. Identifying regression outliers and mixtures graphically. Journal of the American Statistical Association, 95(451):781–794, September 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669462.

Chafai:2009:CRM

[CC09a] Djalil Chafaï and Didier Concordet. Confidence regions for the multinomial parameter with small sample size. Journal of the American Statistical Association, 104(487):1071–1079, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2009:CMS

[CC09b] Yu-Jen Cheng and Ciprian M. Crainiceanu. Cox models with smooth functional effect of covariates measured with error. Journal of the American Statistical Association, 104(487): 1144–1154, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2003:PER

[CCB03] Nilanjan Chatterjee, Yi-Hau Chen, and Norman E. Breslow. A pseudoscore estimator for regression problems with twophase sampling. Journal of the American Statistical Association, 98(461):158–168, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s16/p158;   http://
www.jstor.org/stable/30045203.

Chen:2009:SER

[CCC09] Yi-Hau Chen, Nilanjan Chatterjee, and Raymond J. Carroll. Shrinkage estimators for robust and efficient inference in haplotype-based case-control studies. Journal of the American Statistical Association, 104(485):220–233, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Caffo:2008:CSP

[CCDH08] Brian S. Caffo, Ciprian M. Crainiceanu, Lijuan Deng, and Craig W. Hendrix. A case study in pharmacologic colon imaging using principal curves in single-photon emission computed tomography. Journal of the American Statistical Association, 103(484):1470–1480, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crainiceanu:2009:NSE

[CCDP09] Ciprian M. Crainiceanu, Brian S. Caffo, Chong-Zhi Di, and Naresh M. Punjabi. Nonparametric signal extraction and measurement error in the analysis of electroencephalographic activity during sleep. Journal of the American Statistical Association, 104(486):541–555, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carvalho:2008:HDS

[CCL^{+}08] Carlos M. Carvalho, Jeffrey Chang, Joseph E. Lucas, Joseph R. Nevins, Quanli Wang, and Mike West. Highdimensional sparse factor modeling: Applications in gene expression genomics. Journal of the American Statistical Association, 103(484):1438–1456, December 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chu:2009:REM

[CCL09] Haitao Chu, Sining Chen, and Thomas A. Louis. Random effects models in a meta-analysis of the accuracy of two diagnostic tests without a gold standard. Journal of the American Statistical Association, 104(486):512–523, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2009:ELM

[CCPY09] Ngai Hang Chan, Song Xi Chen, Liang Peng, and Cindy L. Yu. Empirical likelihood methods based on characteristic functions with applications to Lévy processes. Journal of the American Statistical Association, 104(488):1621–1630, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2002:SME

[CCW02] T. Cai, S. C. Cheng, and L. J. Wei. Semiparametric mixedeffects models for clustered failure time data. Journal of the American Statistical Association, 97(458):514–522, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s13/p514;   http://www.jstor.org/stable/3085667.

Campbell:2005:WFW

[CD05] Sean D. Campbell and Francis X. Diebold. Weather forecasting for weather derivatives. Journal of the American Statistical Association, 100(469):6–16, March 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n469/s2/p6;   http://
www.jstor.org/stable/27590514.

Cai:2007:BMI

[CD07] Bo Cai and David B. Dunson. Bayesian multivariate isotonic regression splines. Journal of the American Statistical Association, 102(480):1158–1171, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chung:2009:NBC

[CD09] Yeonseung Chung and David B. Dunson. Nonparametric Bayes conditional distribution modeling with variable selection. Journal of the American Statistical Association, 104 (488):1646–1660, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:2009:NPM

[CDH09a] Raymond J. Carroll, Aurore Delaigle, and Peter Hall. Nonparametric prediction in measurement error models. Journal of the American Statistical Association, 104(487):993–1003, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:2009:RNP

[CDH09b] Raymond J. Carroll, Aurore Delaigle, and Peter Hall. Rejoinder: Nonparametric prediction in measurement error models. Journal of the American Statistical Association, 104(487): 1013–1014, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2005:SMC

[CDHL05] Yuguo Chen, Persi Diaconis, Susan P. Holmes, and Jun S. Liu. Sequential Monte Carlo methods for statistical analysis of tables. Journal of the American Statistical Association, 100(469):109–120, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s16/p109;   http://www.
jstor.org/stable/27590522.

Chen:2005:NIE

[CDK05] Songnian Chen, Gordon B. Dahl, and Shakeeb Khan. Nonparametric identification and estimation of a censored location-scale regression model. Journal of the American Statistical Association, 100(469):212–221, March 2005. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s25/
p212;   http://www.jstor.org/stable/27590531.

Crainiceanu:2008:BBS

[CDR08a] Ciprian M. Crainiceanu, Peter J. Diggle, and Barry Rowlingson. Bivariate binomial spatial modeling of Loa loa prevalence in tropical Africa. Journal of the American Statistical Association, 103(481):21–37, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Pac08, Dun08, Tho08, CDR08b].

Crainiceanu:2008:RBB

[CDR08b] Ciprian M. Crainiceanu, Peter J. Diggle, and Barry Rowlingson. Rejoinder: Bivariate binomial spatial modeling of Loa loa prevalence in tropical Africa. Journal of the American Statistical Association, 103(481):43, March 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CDR08a, Pac08, Dun08, Tho08].

Cao:2000:TVN

[CDVY00] Jin Cao, Drew Davis, Scott Vander Wiel, and Bin Yu. Timevarying network tomography: Router link data. Journal of the American Statistical Association, 95(452):1063–1075, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669743.

Chib:2009:AMA

[CE09] Siddhartha Chib and Bakhodir Ergashev. Analysis of multifactor affine yield curve models. Journal of the American Statistical Association, 104(488):1324–1337, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:2009:LBS

[CF09] R. Dennis Cook and Liliana Forzani. Likelihood-based sufficient dimension reduction. Journal of the American Statistical Association, 104(485):197–208, March 2009. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2007:PLH

[CFJZ07] Jianwen Cai, Jianqing Fan, Jiancheng Jiang, and Haibo Zhou. Partially linear hazard regression for multivariate survival data. Journal of the American Statistical Association, 102 (478):538–551, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2007:NAA

[CFK07] Yu Cheng, Jason P. Fine, and Michael R. Kosorok. Nonparametric association analysis of bivariate competing-risks data. Journal of the American Statistical Association, 102 (480):1407–1415, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2000:EEI

[CFL00] Zongwu Cai, Jianqing Fan, and Runze Li. Efficient estimation and inferences for varying-coefficient models. Journal of the American Statistical Association, 95(451):888–902, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669472.

Cummins:2001:CIN

[CFN01] David J. Cummins, Tom G. Filloon, and Douglas Nychka. Confidence intervals for nonparametric curve estimates: Toward more uniform pointwise coverage. Journal of the American Statistical Association, 96(453):233–246, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/

stable/2670362;   reqidx=/cw/asa/01621459/v96n453/
s18/p233.

Chen:2006:EES

[CFT06] Xiaohong Chen, Yanqin Fan, and Viktor Tsyrennikov. Efficient estimation of semiparametric multivariate copula models. Journal of the American Statistical Association, 101 (475):1228–1240, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590797.

Cai:2000:FCR

[CFY00] Zongwu Cai, Jianqing Fan, and Qiwei Yao. Functionalcoefficient regression models for nonlinear time series. Journal of the American Statistical Association, 95(451):941– 956, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669476.

Clyde:2003:C

[CG03a] Merlise Clyde and Edward I. George. Comment. Journal of the American Statistical Association, 98(463):584–??, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n463/s9/p584.

Clyde:2003:WBN

[CG03b] Merlise Clyde and Edward I. George. Wavelet-based nonparametric modeling of hierarchical functions in colon carcinogenesis [with comment]. Journal of the American Statistical Association, 98(463):584–585, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045284.

Choudhuri:2004:BES

[CGR04] Nidhan Choudhuri, Subhashis Ghosal, and Anindya Roy. Bayesian estimation of the spectral density of a time series. Journal of the American Statistical Association, 99 (468):1050–1059, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/

rpsv/cw/asa/01621459/v99n468/s17/p1050;   http://www.
jstor.org/stable/27590484.

Craig:2001:BFC

[CGRS01] Peter S. Craig, Michael Goldstein, Jonathan C. Rougier, and Allan H. Seheult. Bayesian forecasting for complex systems using computer simulators. Journal of the American Statistical Association, 96(454):717–729, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670309;   reqidx=/cw/
asa/01621459/v96n454/s30/p717.

Chernozhukov:2002:TSC

[CH02] Victor Chernozhukov and Han Hong. Three-step censored quantile regression and extramarital affairs. Journal of the American Statistical Association, 97(459):872–882, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085728.

Chen:2003:SEM

[CH03] Willa W. Chen and Clifford M. Hurvich. Semiparametric estimation of multivariate fractional cointegration. Journal of the American Statistical Association, 98(463): 629–642, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s15/p629;   http://
www.jstor.org/stable/30045290.

Chen:2006:DLL

[CH06a] Jihong Chen and Xiaoming Huo. Distribution of the length of the longest significance run on a Bernoulli net and its applications. Journal of the American Statistical Association, 101(473):321–331, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047460.

Chen:2006:LTL

[CH06b] Shun-Yi Chen and Bernard Harris. On lower tolerance limits with accurate coverage probabilities for the normal random effects model. Journal of the American Statistical Association, 101(475):1039–1049, September 2006. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590781.

Chattopadhyay:2008:BRB

[CH08] Somesh Chattopadhyay and Thomas Hammerstrom. Book review: Statistical Monitoring of Clinical Trials: A Unified Approach. Journal of the American Statistical Association, 103 (481):430–431, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chaganty:2003:BRB

[Cha03a] N. Rao Chaganty. Book review: Subjective Probability Models for Lifetimes, by Fabio Spizzichino. Journal of the American Statistical Association, 98(463):771, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30045313.

Chappell:2003:BRB

[Cha03b] Rick Chappell. Book review: Multivariate Statistical Modelling Based on Generalized Linear Models (2nd ed.), by Ludwig Fahrmeir; Gerhard Tutz. Journal of the American Statistical Association, 98(463):776–777, September 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30045321.

Chatterjee:2004:TSR

[Cha04] Nilanjan Chatterjee. A two-stage regression model for epidemiological studies with multivariate disease classification data. Journal of the American Statistical Association, 99(465):127–138, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s13/p127;   http:
//www.jstor.org/stable/27590359.

Calder:2008:RAP

[CHB^{+}08] Catherine A. Calder, Christopher H. Holloman, Steven M. Bortnick, Warren Strauss, and Michele Morara. Relating ambient particulate matter concentration levels to mortality using an exposure simulator. Journal of the American Statistical Association, 103(481):137–148, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2001:WSL

[Che01] Hua Yun Chen. Weighted semiparametric likelihood method for fitting a proportional odds regression model to data from the case-cohort design. Journal of the American Statistical Association, 96(456):1446–1457, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085912;   reqidx=/cw/asa/01621459/v96n456/
s25/p1446.

Chen:2002:DSM

[Che02] Hua Yun Chen. Double-semiparametric method for missing covariates in Cox regression models. Journal of the American Statistical Association, 97(458):565–576, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s18/p565;   http://www.jstor.org/stable/3085672.

Chen:2004:NSM

[Che04] Hua Yun Chen. Nonparametric and semiparametric models for missing covariates in parametric regression. Journal of the American Statistical Association, 99(468):1176–1189, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s28/p1176;   http://www.jstor.org/stable/27590495.

Chen:2007:AER

[Che07a] Ying Qing Chen. Additive expectancy regression. Journal of the American Statistical Association, 102(477):153–166, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheung:2007:SIS

[Che07b] Ying Kuen Cheung. Sequential implementation of stepwise procedures for identifying the maximum tolerated dose. Journal of the American Statistical Association, 102(480):1448– 1461, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chernick:2008:BRB

[Che08] Michael R. Chernick. Book review: Nonparametric Statistics, With Applications to Science and Engineering. Journal of the American Statistical Association, 103(483):1326–1327, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiu:2006:BRB

[Chi06] Wai F. Chiu. Book review: Binomiai Models in Finance, by John Van Der Hoek; Robert J. Elliott. Journal of the American Statistical Association, 101(476):1731, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639803.

Chen:2008:NRM

[CHJ08] Ying Chen, Wolfgang Härdle, and Seok-Oh Jeong. Nonparametric risk management with generalized hyperbolic distributions. Journal of the American Statistical Association, 103 (483):910–923, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clements:2003:MLM

[CHL03] Adam Clements, Stan Hurn, and Kenneth Lindsay. Möbiuslike mappings and their use in kernel density estimation. Journal of the American Statistical Association, 98(464):993–1000, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v98n464/s24/p993;   http:
//www.jstor.org/stable/30045345.

Chen:2006:CMD

[CHL06] Willa W. Chen, Clifford M. Hurvich, and Yi Lu. On the correlation matrix of the discrete Fourier transform and the fast solution of large Toeplitz systems for long-memory time series. Journal of the American Statistical Association, 101 (474):812–822, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590737.

Choudhary:2006:BRB

[Cho06] Pankaj K. Choudhary. Book review: Analyzing Rater Agreement: Manifest Variable Methods, by Alexander von Eye; Eun Young Mun. Journal of the American Statistical Association, 101(474):854, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590753.

Celeux:2000:CID

[CHR00a] Gilles Celeux, Merrilee Hurn, and Christian P. Robert. Computational and inferential difficulties with mixture posterior distributions. Journal of the American Statistical Association, 95(451):957–970, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669477.

Christensen:2000:LLL

[Chr00b] Ronald Christensen. Linear and log-linear models. Journal of the American Statistical Association, 95(452):1290– 1293, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669772.

Christensen:2001:BRB

[Chr01a] Ronald Christensen. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1524–1525, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085920.

Christoffersen:2001:BRB

[Chr01b] Peter F. Christoffersen. Book review: [untitled]. Journal of the American Statistical Association, 96(453):347–348, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670380.

Christman:2007:CMA

[Chr07] Mary C. Christman. Comment: Model-assisted estimation of forest resources with generalized additive models. Journal of the American Statistical Association, 102(478):411–412, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Coull:2001:CRE

[CHRH01] Brent A. Coull, James P. Hobert, Louise M. Ryan, and Lewis B. Holmes. Crossed random effect models for multiple outcomes in a study of teratogenesis. Journal of the American Statistical Association, 96(456):1194–1204, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085882;   reqidx=/cw/asa/01621459/v96n456/
s6/p1194.

Claeskens:2003:FIC

[CHS^{+}03] Gerda Claeskens, Nils Lid Hjort, Xiaotong Shen, Daniel P. Dougherty, Wesley O. Johnson, Hemant Ishwaran, J. Sunil Rao, R. Dennis Cook, Lexin Li, Chih-Ling Tsai, Adrian E. Raftery, Yingye Zheng, Nils Lid Hjort, and Gerda Claeskens. The focused information criterion. Journal of the American Statistical Association, 98(464):900–916, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s12/p900;   http://www.jstor.org/stable/30045340.

Chen:2008:PIJ

[CIK08] Ming-Hui Chen, Joseph G. Ibrahim, and Sungduk Kim. Properties and implementation of Jeffreys’s prior in binomial regression models. Journal of the American Statistical Association, 103(484):1659–1664, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2004:PPD

[CIS04] Ming-Hui Chen, Joseph G. Ibrahim, and Qi-Man Shao. Propriety of the posterior distribution and existence of the MLE for regression models with covariates missing at random. Journal of the American Statistical Association, 99(466):421–438, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s12/p421;   http:
//www.jstor.org/stable/27590398.

Cizek:2008:REA

[Číź08] Pavel Číźek. Robust and efficient adaptive estimation of binary-choice regression models. Journal of the American Statistical Association, 103(482):687–696, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chib:2001:MLM

[CJ01] Siddhartha Chib and Ivan Jeliazkov. Marginal likelihood from the Metropolis–Hastings output. Journal of the American Statistical Association, 96(453):270–281, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670365;   reqidx=/cw/asa/01621459/v96n453/
s21/p270.

Chen:2006:PLR

[CJ06a] Kani Chen and Zhezhen Jin. Partial linear regression models for clustered data. Journal of the American Statistical Association, 101(473):195–204, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047449.

Chib:2006:ISD

[CJ06b] Siddhartha Chib and Ivan Jeliazkov. Inference in semiparametric dynamic models for binary longitudinal data. Journal of the American Statistical Association, 101(474):685–700, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590727.

Casals:2002:EMM

[CJS02] José Casals, Miguel Jerez, and Sonia Sotoca. An exact multivariate model-based structural decomposition. Journal of the American Statistical Association, 97(458):553–564, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s17/p553;   http://www.jstor.org/stable/3085671.

Cook:2004:ATE

[CK04] Thomas D. Cook and Michael R. Kosorok. Analysis of time-to-event data with incomplete event adjudication. Journal of the American Statistical Association, 99(468): 1140–1152, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s25/p1140;   http://www.
jstor.org/stable/27590492.

Chen:2008:OSF

[CK08] Jiahua Chen and Abbas Khalili. Order selection in finite mixture models with a nonsmooth penalty. Journal of the American Statistical Association, 103(484):1674–1683, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2009:OSF

[CK09] Jiahua Chen and Abbas Khalili. Order selection in finite mixture models with a nonsmooth penalty. Journal of the American Statistical Association, 104(485):187–196, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cottet:2008:VSM

[CKN08] Remy Cottet, Robert J. Kohn, and David J. Nott. Variable selection and model averaging in semiparametric overdispersed generalized linear models. Journal of the American Statistical Association, 103(482):661–671, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carlin:2000:EBP

[CL00] Bradley P. Carlin and Thomas A. Louis. Empirical Bayes: Past, present and future. Journal of the American Statistical Association, 95(452):1286–1289, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669771.

Cook:2003:DII

[CL03] Dennis Cook and Lexin Li. Discussion: Inference and interpretability considerations in frequentist model averaging and selection. Journal of the American Statistical Association, 98(464):925–928, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s16/p925.

Chiou:2008:CBF

[CL08a] Jeng-Min Chiou and Pai-Ling Li. Correlation-based functional clustering via subspace projection. Journal of the American Statistical Association, 103(484):1684–1692, December 2008.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Consonni:2008:TBI

[CL08b] Guido Consonni and Luca La Rocca. Tests based on intrinsic priors for the equality of two correlated proportions. Journal of the American Statistical Association, 103(483):1260–1269, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clarke:2003:BRB

[Cla03] William R. Clarke. Book review: Longitudinal Data and SAS: A Programmer’s Guide, by Ron Cody. Journal of the American Statistical Association, 98(463):773, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045316.

Cleary:2001:BRB

[Cle01] Rick Cleary. Book review: [untitled]. Journal of the American Statistical Association, 96(454):779, June 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670318.

Cleary:2002:BRBa

[Cle02a] Richard J. Cleary. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1201–1202, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085843.

Cleary:2002:BRBb

[Cle02b] Richard J. Cleary. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1216, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085863.

Cleary:2002:SCE

[Cle02c] Richard J. Cleary. The science of conjecture: Evidence and probability before Pascal. Journal of the American Statistical Association, 97(460):1201–??, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-

bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s24/p1201.

Cleary:2003:BRB

[Cle03a] Richard J. Cleary. Book review: Conducting Meta Analysis Using SAS, by Winfred Arthur, Jr.; Winston Bennett, Jr.; Allen I. Huffcutt. Journal of the American Statistical Association, 98(461):254, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045221.

Cleary:2003:CMA

[Cle03b] Richard J. Cleary. Conducting meta analysis using SAS. Journal of the American Statistical Association, 98(461):254–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s34/p254.

Cleary:2006:BRB

[Cle06] Richard J. Cleary. Book review: Applied Data Mining: Statistical Methods for Business and Industry, by Paolo Giudici. Journal of the American Statistical Association, 101 (475):1317–1318, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590820.

Cline:2007:BRB

[Cli07] Daren B. H. Cline. Book review: Stochastic Processes in Science, Engineering, and Finance. Journal of the American Statistical Association, 102(480):1476–1477, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:2009:REM

[CLLCL09] Richard J. Cook, Jerald F. Lawless, Lajmi Lakhal-Chaieb, and Ker-Ai Lee. Robust estimation of mean functions and treatment effects for recurrent events under event-dependent censoring and termination: Application to skeletal complications in cancer metastatic to bone. Journal of the American Statistical Association, 104(485):60–75, March 2009. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2003:IRS

[CLQ03] Song Xi Chen, Denis H. Y. Leung, and Jing Qin. Information recovery in a study with surrogate endpoints. Journal of the American Statistical Association, 98(464): 1052–1062, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s29/p1052;   http://www.
jstor.org/stable/30045350.

Chiu:2006:BCR

[CLR06] Grace Chiu, Richard Lockhart, and Richard Routledge. Bentcable regression theory and applications. Journal of the American Statistical Association, 101(474):542–553, June 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590716.

Chiou:2001:IDF

[CM01] Jeng-Min Chiou and Hans-Georg Muller. Inference for density families using functional principal component analysis: Comment. Journal of the American Statistical Association, 96 (454):534–537, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670292.

Casella:2006:OBV

[CM06a] George Casella and Elías Moreno. Objective Bayesian variable selection. Journal of the American Statistical Association, 101(473):157–167, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047446.

Centanni:2006:MCA

[CM06b] Silvia Centanni and Marco Minozzo. A Monte Carlo approach to filtering for a class of marked doubly stochastic Poisson processes. Journal of the American Statistical Association, 101(476):1582–1597, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639774.

Casella:2009:ARI

[CM09a] George Casella and Elías Moreno. Assessing robustness of intrinsic tests of independence in two-way contingency tables. Journal of the American Statistical Association, 104 (487):1261–1271, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiou:2009:MHR

[CM09b] Jeng-Min Chiou and Hans-Georg Müller. Modeling hazard rates as functional data for the analysis of cohort lifetables and mortality forecasting. Journal of the American Statistical Association, 104(486):572–585, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cucala:2009:BRN

[CMRT09] Lionel Cucala, Jean-Michel Marin, Christian P. Robert, and D. M. Titterington. A Bayesian reassessment of nearestneighbor classification. Journal of the American Statistical Association, 104(485):263–273, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carriquiry:2002:LCA

[CN02] Alicia L. Carriquiry and Sarah M. Nusser. [latent class analysis of complex sample survey data: Application to dietary data]: Comment. Journal of the American Statistical Association, 97(459):729–731, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085709.

Cook:2005:SDR

[CN05] R. Dennis Cook and Liqiang Ni. Sufficient dimension reduction via inverse regression: a minimum discrepancy approach. Journal of the American Statistical Association, 100(470):410–428, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s6/p410;   http:
//www.jstor.org/stable/27590565.

Chu:2008:FRC

[CN08] Haitao Chu and Lei Nie. A few remarks on “A Capture– Recapture Approach for Screening Using Two Diagnostic Tests With Availability of Disease Status for the Test Positives Only” by Böhning and Patilea. Journal of the American Statistical Association, 103(484):1518–1519, December 2008.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BP08].

Cooley:2007:BSM

[CNN07] Daniel Cooley, Douglas Nychka, and Philippe Naveau. Bayesian spatial modeling of extreme precipitation return levels. Journal of the American Statistical Association, 102(479): 824–840, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Coakley:2000:BRB

[Coa00] Clint W. Coakley. Book review: [untitled]. Journal of the American Statistical Association, 95(449):332–333, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669565.

Cogley:2002:BRB

[Cog02] Timothy Cogley. Book review: [untitled]. Journal of the American Statistical Association, 97(457):364, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085804.

Conquest:2008:AED

[Con08] Loveday Conquest. Analysing ecological data. Journal of the American Statistical Association, 103(483):1317, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Corcoran:2002:BRB

[Cor02a] Jem N. Corcoran. Book review: [untitled]. Journal of the American Statistical Association, 97(457):360, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085797.

Corradi:2002:ATS

[Cor02b] Valentina Corradi. Asymptotic theory of statistical inference for time series. Journal of the American Statistical Association, 97(460):1208–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s35/p1208.

Corradi:2002:BRB

[Cor02c] Valentina Corradi. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1208–1209, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085854.

Corrado:2005:BRB

[Cor05] Luisa Corrado. Book review: Generalized Latent Variable Modeling: Multilevel, Longitudinal, and Structural Equation Models, by Anders Skrondal; Sophia Rabe-Hesketh. Journal of the American Statistical Association, 100(470):710–711, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s39/
p710;   http://www.jstor.org/stable/27590598.

Cowles:2000:BRBa

[Cow00a] Mary Kathryn Cowles. Book review: [untitled]. Journal of the American Statistical Association, 95(450):675–676, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669416.

Cowles:2000:BRBb

[Cow00b] Mary Kathryn Cowles. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1011–1012, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669495.

Cowles:2003:BSM

[Cow03a] Mary Kathryn Cowles. Bayesian statistical modelling. Journal of the American Statistical Association, 98(461):256–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s37/p256.

Cowles:2003:BRB

[Cow03b] Mary Kathryn Cowles. Book review: Bayesian Statistical Modelling, by Peter Congdon. Journal of the American Statistical Association, 98(461):256–257, March 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30045224.

Cowles:2006:BRB

[Cow06] Mary Kathryn Cowles. Book review: Generalized, Linear, and Mixed Models, by Charles E. McCulloch; Shayle R. Searle. Journal of the American Statistical Association, 101(476): 1724, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639793.

Cowles:2007:BRB

[Cow07] Mary Kathryn Cowles. Book review: Automatic Nonuniform Random Variate Generation. Journal of the American Statistical Association, 102(479):1079–1080, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cox:2000:CIC

[Cox00] D. R. Cox. Causal inference without counterfactuals: Comment. Journal of the American Statistical Association, 95 (450):424–425, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669378.

Cai:2002:SRO

[CP02] Tianxi Cai and Margaret Sullivan Pepe. Semiparametric receiver operating characteristic analysis to evaluate biomarkers for disease. Journal of the American Statistical Association, 97(460):1099–1107, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s15/p1099;   http:/
/www.jstor.org/stable/3085834.

Chandler:2006:DLS

[CP06] Gabriel Chandler and Wolfgang Polonik. Discrimination of locally stationary time series based on the excess mass functional. Journal of the American Statistical Association, 101 (473):240–253, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047453.

Cheng:2007:VRM

[CP07] Ming-Yen Cheng and Liang Peng. Variance reduction in multiparameter likelihood models. Journal of the American Statistical Association, 102(477):293–304, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cappe:2000:MCM

[CR00a] Olivier Cappe and Christian P. Robert. Markov chain Monte Carlo: 10 years and still running! Journal of the American Statistical Association, 95(452):1282–1286, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669770.

Cheang:2000:BRA

[CR00b] Wai-Kwong Cheang and Gregory C. Reinsel. Bias reduction of autoregressive estimates in time series regression model through restricted maximum likelihood. Journal of the American Statistical Association, 95(452):1173–1184, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669758.

Cantoni:2001:RIG

[CR01] Eva Cantoni and Elvezio Ronchetti. Robust inference for generalized linear models. Journal of the American Statistical Association, 96(455):1022–1030, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670248;   reqidx=/cw/asa/01621459/v96n455/
s20/p1022.

Connolly:2008:SLS

[CR08] Robert A. Connolly and Richard J. Rendleman, Jr. Skill, luck, and streaky play on the PGA tour. Journal of the American Statistical Association, 103(481):74–88, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Craig:2005:CSH

[Cra05] Bruce A. Craig. Comment: Semilinear high-dimensional model for normalization of microarray data. Journal of the American Statistical Association, 100(471):799–800, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crainiceanu:2007:BRB

[Cra07] Ciprian M. Crainiceanu. Book review: Nonparametric Regression Methods for Longitudinal Data Analysis. Mixed-effects Modeling Approaches. Journal of the American Statistical Association, 102(478):761, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:2004:NNR

[CRC^{+}04] Raymond J. Carroll, David Ruppert, Ciprian M. Crainiceanu, Tor D. Tosteson, and Margaret R. Karagas. Nonlinear and nonparametric regression and instrumental variables. Journal of the American Statistical Association, 99(467):736–750, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s22/p736;   http:
//www.jstor.org/stable/27590444.

Chiang:2001:SSE

[CRW01] Chin-Tsang Chiang, John A. Rice, and Colin O. Wu. Smoothing spline estimation for varying coefficient models with repeatedly measured dependent variables. Journal of the American Statistical Association, 96(454):605–619, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670300;   reqidx=/cw/asa/01621459/v96n454/
s21/p605.

Craiu:2009:LTN

[CRY09] Radu V. Craiu, Jeffrey Rosenthal, and Chao Yang. Learn from thy neighbor: Parallel-chain and regional adaptive MCMC. Journal of the American Statistical Association, 104(488): 1454–1466, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Casella:2000:CIC

[CS00a] George Casella and Stephen P. Schwartz. Causal inference without counterfactuals: Comment. Journal of the American Statistical Association, 95(450):425–427, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669379.

Cleveland:2000:ITD

[CS00b] William S. Cleveland and Don X. Sun. Internet traffic data. Journal of the American Statistical Association, 95(451):979– 985, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669481.

Chen:2001:JVE

[CS01] Jiahua Chen and Jun Shao. Jackknife variance estimation for nearest-neighbor imputation. Journal of the American Statistical Association, 96(453):260–269, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670364;   reqidx=/cw/asa/01621459/v96n453/
s20/p260.

Cook:2003:MFT

[CS03a] R. Dennis Cook and C. Messan Setodji. A model-free test for reduced rank in multivariate regression. Journal of the American Statistical Association, 98(462):340–351, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s7/p340;   http://www.jstor.org/stable/30045244.

Cottet:2003:BMF

[CS03b] Remy Cottet and Michael Smith. Bayesian modeling and forecasting of intraday electricity load. Journal of the American Statistical Association, 98(464):839–849, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/s7/
p839;   http://www.jstor.org/stable/30045335.

Charnigo:2004:THM

[CS04] Richard Charnigo and Jiayang Sun. Testing homogeneity in a mixture distribution via the L^{2} distance between competing models. Journal of the American Statistical Association, 99(466):488–498, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s18/p488;   http:
//www.jstor.org/stable/27590404.

Cai:2009:STG

[CS09] T. Tony Cai and Wenguang Sun. Simultaneous testing of grouped hypotheses: Finding needles in multiple haystacks. Journal of the American Statistical Association, 104(488): 1467–1481, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2006:CLB

[CSCC06] Nilanjan Chatterjee, Christine Spinka, Jinbo Chen, and Raymond J. Carroll. Comment: Likelihood-based inference on haplotype effects in genetic association studies. Journal of the American Statistical Association, 101(473):108–111, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crainiceanu:2009:GMF

[CSD09] Ciprian M. Crainiceanu, Ana-Maria Staicu, and Chong-Zhi Di. Generalized multilevel functional regression. Journal of the American Statistical Association, 104(488):1550–1561, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2004:ISS

[CT04] Chi-Lun Cheng and Chih-Ling Tsai. The invariance of some score tests in the linear model with classical measurement error. Journal of the American Statistical Association, 99(467):805–809, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s28/p805;   http:
//www.jstor.org/stable/27590450.

Cauchemez:2006:IHP

[CTG^{+}06] Simon Cauchemez, Laura Temime, Didier Guillemot, Emmanuelle Varon, Alain-Jacques Valleron, Guy Thomas, and Pierre-Yves Boëlle. Investigating heterogeneity in pneumococcal transmission: a Bayesian MCMC approach applied to a follow-up of schools. Journal of the American Statistical Association, 101(475):946–958, September 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590774.

Chellappa:2007:CGB

[CV07] Rama Chellappa and Ashok Veeraraghavan. Comment: Gaitbased human recognition by classification of cyclostationary processes on nonlinear shape manifolds. Journal of the American Statistical Association, 102(480):1126–1127, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Croux:2002:NNV

[CvA02] Christophe Croux and Stefan van Aelst. [nearest-neighbor variance estimation (NNVE): Robust covariance estimation via nearest-neighbor cleaning]: Comment. Journal of the American Statistical Association, 97(460):1006–1009, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085822.

Copt:2006:HBI

[CVF06] Samuel Copt and Maria-Pia Victoria-Feser. High-breakdown inference for mixed linear models. Journal of the American Statistical Association, 101(473):292–300, March 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047457.

Chen:2000:AAH

[CW00] Ying Qing Chen and Mei-Cheng Wang. Analysis of accelerated hazards models. Journal of the American Statistical Association, 95(450):608–618, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/2000/June/Chen.htm;
http://www.jstor.org/stable/2669403.

Cheng:2001:LTM

[CW01] Su-Chun Cheng and Naisyin Wang. Linear transformation models for failure time data with covariate measurement error. Journal of the American Statistical Association, 96(454):706–716, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670308;   reqidx=/cw/
asa/01621459/v96n454/s29/p706.

Chen:2008:ELE

[CW08] Jianwei Chen and Hulin Wu. Efficient local estimation for time-varying coefficients in deterministic dynamic models with applications to HIV-1 dynamics. Journal of the American Statistical Association, 103(481):369–384, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2008:NQE

[CX08] Zongwu Cai and Xiaoping Xu. Nonparametric quantile estimations for dynamic smooth coefficient models. Journal of the American Statistical Association, 103(484):1595–1608, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2009:NQE

[CX09] Zongwu Cai and Xiaoping Xu. Nonparametric quantile estimations for dynamic smooth coefficient models. Journal of the American Statistical Association, 104(485):371–383, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2007:SSW

[CZ07] Hock Peng Chan and Nancy Ruonan Zhang. Scan statistics with weighted observations. Journal of the American Statistical Association, 102(478):595–602, June 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2009:SEG

[CZC09] Ming-Yen Cheng, Wenyang Zhang, and Lu-Hung Chen. Statistical estimation in generalized multiparameter likelihood models. Journal of the American Statistical Association, 104(487): 1179–1191, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2007:SML

[CZI07] Qingxia Chen, Donglin Zeng, and Joseph G. Ibrahim. Sieve maximum likelihood estimation for regression models with covariates missing at random. Journal of the American Statistical Association, 102(480):1309–1317, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Daniels:2003:BRB

[Dan03] Michael Daniels. Book review: Contemporary Statistical Models for the Plant and Soil Sciences, by Oliver Schabenberger; Francis J. Pierce. Journal of the American Statistical Association, 98(464):1080–1082, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s33/p1080;   http://www.jstor.org/stable/30045354.

Dassios:2002:BRB

[Das02a] Angelos Dassios. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1211–1212, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085857.

Dassios:2002:RP

[Das02b] Angelos Dassios. Ruin probabilities. Journal of the American Statistical Association, 97(460):1211–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s38/p1211.

Datta:2004:BRB

[Dat04] Gauri Sankar Datta. Book review: Small Area Estimation, by J. N. K. Rao. Journal of the American Statistical Association, 99(465):294–295, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s30/p294;   http:
//www.jstor.org/stable/27590376.

Davis:2001:BRB

[Dav01] William W. Davis. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1529–1530, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085928.

Davidian:2003:BRB

[Dav03] Marie Davidian. Book review: Hierarchical Linear Models: Applications and Data Analysis Methods (second edition), by Stephen W. Raudenbush; Anthony S. Bryk. Journal of the American Statistical Association, 98(463):767–768, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045307.

Davies:2006:BRB

[Dav06a] Simon Davies. Book review: Epidemiology: Study Design and Data Analysis, by Mark Woodward. Journal of the American Statistical Association, 101(473):404–405, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047493.

Davis:2006:BRBb

[Dav06b] Ginger M. Davis. Book review: Forecasting: Practice and Process for Demand Management, by Hans Levenbach; James P. Cleary. Journal of the American Statistical Association, 101(476):1732, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639805.

Davis:2006:BRBa

[Dav06c] J. Wade Davis. Book review: Statistics for Microarrays: Design, Analysis and Inference, by Ernst Wit; John McClure. Journal of the American Statistical Association, 101(474): 855–856, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590757.

Davis:2007:BRBa

[Dav07a] J. Wade Davis. Book review: Bioinformatics and Computational Biology Solutions Using R and Bioconductor. Journal of the American Statistical Association, 102(477):388–389, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davis:2007:BRBb

[Dav07b] J. Wade Davis. Book review: DNA, Words and Models: Statistics of Exceptional Words. Journal of the American Statistical Association, 102(480):1475, December 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dawid:2000:CICa

[Daw00a] A. P. Dawid. Causal inference without counterfactuals. Journal of the American Statistical Association, 95(450): 407–424, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/2000/June/Dawid.htm;   http:/
/www.jstor.org/stable/2669377.

Dawid:2000:CICb

[Daw00b] A. P. Dawid. Causal inference without counterfactuals: Rejoinder. Journal of the American Statistical Association, 95 (450):444–448, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669385.

Dawson:2003:BRB

[Daw03] Jeffrey D. Dawson. Book review: Analyzing Medical Data Using S-PLUS, by Brian Everitt; Sophia Rabe-Hesketh. Journal of the American Statistical Association, 98(463):773– 774, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045317.

Dette:2003:RED

[DB03] Holger Dette and Stefanie Biedermann. Robust and efficient designs for the Michaelis–Menten model. Journal of the American Statistical Association, 98(463):679–686, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s20/p679;   http://www.jstor.org/stable/30045295.

Durrieu:2009:SDM

[DB09] Gilles Durrieu and Laurent Briollais. Sequential design for microarray experiments. Journal of the American Statistical Association, 104(486):650–660, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2008:ODD

[DBPP08] Holger Dette, Frank Bretz, Andrey Pepelyshev, and José Pinheiro. Optimal designs for dose-finding studies. Journal of the American Statistical Association, 103(483):1225–1237, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dunson:2003:BMM

[DC03] David B. Dunson and Bernardo Colombo. Bayesian modeling of markers of day-specific fertility. Journal of the American Statistical Association, 98(461):28–37, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s4/p28;   http://www.jstor.org/stable/30045191.

Dominici:2004:IST

[DDH04] Francesca Dominici, Aidan MC Dermott, and Trevor J. Hastie. Improved semiparametric time series models of air pollution and mortality. Journal of the American Statistical Association, 99(468):938–948, December 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/s4/
p938;   http://www.jstor.org/stable/27590477.

Dominici:2002:APM

[DDZS02] Francesca Dominici, Michael Daniels, Scott L. Zeger, and Jonathan M. Samet. Air pollution and mortality: Estimating regional and national dose-response relationships. Journal of the American Statistical Association, 97(457):100–111, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085762;   reqidx=/cw/asa/01621459/v97n457/
s10/p100.

DeSantis:2006:SSD

[De 06] Fulvio De Santis. Sample size determination for robust Bayesian analysis. Journal of the American Statistical Association, 101(473):278–291, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047456.

Dean:2008:BRB

[Dea08] Angela Dean. Book review: The Construction of Optimal Stated Choice Experiments: Theory and Methods. Journal of the American Statistical Association, 103(483):1320–1321, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delbaen:2002:BRB

[Del02a] F. Delbaen. Book review: [untitled]. Journal of the American Statistical Association, 97(459):923, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085737.

Delbaen:2002:PFO

[Del02b] F. Delbaen. Probability and finance: It’s only a game! Journal of the American Statistical Association, 97(459): 923–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s27/p923.

deLaubenfels:2005:BRB

[deL05] Ralph deLaubenfels. Book review: Feynman-Kac Formulae: Genealogical and Interacting Particle Systems with Applications, by Pierre Del Moral. Journal of the American Statistical Association, 100(472):1460, December 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590690.

DellAquila:2006:BRB

[Del06] Rosario Dell’Aquila. Book review: Generalized Method of Moments. Journal of the American Statistical Association, 101 (475):1309–1310, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dempster:2004:CMF

[Dem04] A. P. Dempster. Comment: Membership functions and probability measures of fuzzy sets. Journal of the American Statistical Association, 99(467):882–884, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s38/p882.

Devore:2006:BRB

[Dev06] Jay Devore. Book review: A Modern Introduction to Probability and Statistics: Understanding Why and How, by F. M. Dekking; C. Kraaikamp; H. P. Lopuhaä; L. E. Meester. Journal of the American Statistical Association, 101(473): 393–394, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047473.

Delaigle:2009:DAL

[DFC09] Aurore Delaigle, Jianqing Fan, and Raymond J. Carroll. A design-adaptive local polynomial estimator for the errors-in-variables problem. Journal of the American Statistical Association, 104(485):348–359, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dudoit:2002:CDM

[DFS02] Sandrine Dudoit, Jane Fridlyand, and Terence P. Speed. Comparison of discrimination methods for the classification of tumors using gene expression data. Journal of the American Statistical Association, 97(457):77–87, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
3085760;   reqidx=/cw/asa/01621459/v97n457/s8/p77.

Dumbgen:2006:EUD

[DFWJ06] Lutz Dümbgen, Sandra Freitag-Wolf, and Geurt Jongbloed. Estimating a unimodal distribution from interval-censored data. Journal of the American Statistical Association, 101 (475):1094–1106, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590787.

deGunst:2001:SAI

[dGKS01] M. C. M. de Gunst, H. R. Künsch, and J. G. Schouten. Statistical analysis of ion channel data using hidden Markov models with correlated state-dependent noise and filtering. Journal of the American Statistical Association, 96 (455):805–815, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;

http://www.jstor.org/stable/2670221;   reqidx=/cw/
asa/01621459/v96n455/s3/p805.

Delaigle:2008:USS

[DH08] Aurore Delaigle and Peter Hall. Using SIMEX for smoothingparameter choice in errors-in-variables problems. Journal of the American Statistical Association, 103(481):280–287, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dhaene:2001:BRB

[Dha01] Geert Dhaene. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1527–1528, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085925.

Dunson:2008:BSC

[DHE08] David B. Dunson, Amy H. Herring, and Stephanie M. Engel. Bayesian selection and clustering of polymorphisms in functionally related genes. Journal of the American Statistical Association, 103(482):534–546, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dong:2008:SSD

[DHS08] Yuping Dong, A. S. Hedayat, and B. K. Sinha. Surveillance strategies for detecting changepoint in incidence rate based on exponentially weighted moving average methods. Journal of the American Statistical Association, 103(482):843–853, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dunson:2008:BIC

[DHSR08] David B. Dunson, Amy H. Herring, and Anna Maria Siega-Riz. Bayesian inference on changes in response densities over predictor clusters. Journal of the American Statistical Association, 103(484):1508–1517, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:2000:BRBa

[Dia00a] Persi Diaconis. Book review: [untitled]. Journal of the American Statistical Association, 95(449):342–343, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669576.

Diaconis:2000:BRBb

[Dia00b] Persi Diaconis. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1009, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669492.

Diaconis:2000:BRBc

[Dia00c] Persi Diaconis. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1377–1378, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669801.

Diaconis:2003:BRB

[Dia03a] Persi Diaconis. Book review: Sequential Monte Carlo Methods in Practice, by Arnaud Doucet; Nando de Freitas; Neil Gordon. Journal of the American Statistical Association, 98 (462):496–497, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045268.

Diaconis:2003:SMC

[Dia03b] Persi Diaconis. Sequential Monte Carlo methods in practice. Journal of the American Statistical Association, 98(462):496–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s32/p496.

Dias:2007:BRB

[Dia07] Alexandra Dias. Book review: Simulation Techniques in Financial Risk Management. Journal of the American Statistical Association, 102(478):758–759, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:2005:BRB

[Dig05] Peter Diggle. Book review: Applied Spatial Statistics for Public Health Data, by Lance A. Waller; Carol A. Gotway. Journal of the American Statistical Association, 100(470):702–703, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s29/p702;   http://www.
jstor.org/stable/27590588.

Dixon:2006:BRB

[Dix06] Philip Dixon. Book review: Nondetects and Data Analysis: Statistics for Censored Environmental Data, by Dennis R. Helsel. Journal of the American Statistical Association, 101 (473):401, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047486.

Deng:2009:ALT

[DJSW09] Xinwei Deng, V. Roshan Joseph, Agus Sudjianto, and C. F. Jeff Wu. Active learning through sequential design, with applications to detection of money laundering. Journal of the American Statistical Association, 104(487):969–981, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dimatteo:2001:VTD

[DK01] Ilaria Dimatteo and Joseph B. Kadane. Vote tampering in a district justice election in Beaver County, PA. Journal of the American Statistical Association, 96(454):510–518, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670289;   reqidx=/cw/asa/01621459/v96n454/
s15/p510.

Davis:2006:BRBc

[DKC^{+}06] J. Wade Davis, Alan F. Karr, Joseph Cavanaugh, Richard J. Cleary, and Robert Lund. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 101 (475):1318–1320, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590821.

Dikta:2006:BAM

[DKS06] Gerhard Dikta, Marsel Kvesic, and Christian Schmidt. Bootstrap approximations in model checks for binary data. Journal of the American Statistical Association, 101(474):521–530, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590714.

Diaconis:2008:CSA

[DL08] Persi Diaconis and Erich Lehmann. Comment: On Student’s 1908 article “The Probable Error of a Mean”. Journal of the American Statistical Association, 103(481):16–19, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Denby:2004:CEP

[DLM04] L. Denby, J. M. Landwehr, and C. L. Mallows. Comment: The estimation of prediction error: Covariance penalties and cross-validation. Journal of the American Statistical Association, 99(467):633–634, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n467/s11/p633.

Davis:2006:SBE

[DLRY06] Richard A. Davis, Thomas C. M. Lee, and Gabriel A. Rodriguez-Yam. Structural break estimation for nonstationary time series models. Journal of the American Statistical Association, 101(473):223–239, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047452.

Duncan:2000:ODL

[DM00] George T. Duncan and Sumitra Mukherjee. Optimal disclosure limitation strategy in statistical databases: Deterring tracker attacks through additive noise. Journal of the American Statistical Association, 95(451):720–729, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669452.

DiCiccio:2004:IAS

[DM04] Thomas J. DiCiccio and Anna Clara Monti. Inferential aspects of the skew exponential power distribution. Journal of the American Statistical Association, 99(466):439–450, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s13/p439;   http://www.jstor.org/stable/27590399.

Dubin:2005:DCM

[DM05] Joel A. Dubin and Hans-Georg Müller. Dynamical correlation for multivariate longitudinal data. Journal of the American Statistical Association, 100(471):872–881, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590619.

Drignei:2006:EBA

[DM06] Dorin Drignei and Max D. Morris. Empirical Bayesian analysis for computer experiments involving finite-difference codes. Journal of the American Statistical Association, 101 (476):1527–1536, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639770.

Delaigle:2007:NRE

[DM07] Aurore Delaigle and Alexander Meister. Nonparametric regression estimation in the heteroscedastic errors-in-variables problem. Journal of the American Statistical Association, 102 (480):1416–1426, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dmitrienko:2005:BRB

[Dmi05] Alex Dmitrienko. Book review: Multiple Analyses in Clinical Trials: Fundamentals for Investigators, by Lemuel A. Moyé. Journal of the American Statistical Association, 100(472): 1463, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590695.

Dasgupta:2008:SMA

[DMJ^{+}08] Tirthankar Dasgupta, Christopher Ma, V. Roshan Joseph, Z. L. Wang, and C. F. Jeff Wu. Statistical modeling and analysis for robust synthesis of nanostructures. Journal of the American Statistical Association, 103(482):594–603, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DeIorio:2004:AMD

[DMRM04] Maria De Iorio, Peter Müller, Gary L. Rosner, and Steven N. MacEachern. An ANOVA model for dependent random measures. Journal of the American Statistical Association, 99(465):205–215, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s20/p205;   http:
//www.jstor.org/stable/27590366.

Dette:2005:ODG

[DMW05] Holger Dette, Viatcheslav B. Melas, and Weng Kee Wong. Optimal design for goodness-of-fit of the Michaelis–Menten enzyme kinetic function. Journal of the American Statistical Association, 100(472):1370–1381, December 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590678.

Dass:2003:EDS

[DN03] Sarat C. Dass and Vijayan N. Nair. Edge detection, spatial smoothing, and image reconstruction with partially observed multivariate data. Journal of the American Statistical Association, 98(461):77–89, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s9/p77;   http://www.jstor.org/stable/30045196.

Dahl:2007:MHT

[DN07] David B. Dahl and Michael A. Newton. Multiple hypothesis testing by clustering treatment effects. Journal of the American Statistical Association, 102(478):517–526, June 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2005:NNE

[DNP05] Holger Dette, Natalie Neumeyer, and Kay F. Pilz. A note on nonparametric estimation of the effective dose in quantal bioassay. Journal of the American Statistical Association, 100(470):503–510, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s13/p503;   http://www.
jstor.org/stable/27590572.

Dobrow:2000:BRB

[Dob00] R. Dobrow. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1018–1019, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669503.

Dobler:2002:BRB

[Dob02a] Carolyn Pillers Dobler. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1201, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085842.

Dobler:2002:SC

[Dob02b] Carolyn Pillers Dobler. Statisticians of the centuries. Journal of the American Statistical Association, 97(460): 1201–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s23/p1201.

Dobler:2003:BRBa

[Dob03a] Carolyn Pillers Dobler. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 98(464): 1087, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045362.

Dobler:2003:BRBb

[Dob03b] Carolyn Pillers Dobler. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 98(464): 1087, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045363.

Dobler:2004:BRB

[Dob04] Carolyn Pillers Dobler. Book review: Statistical Tables, Explained and Applied, by Louis Laurencelle; François-A. Dupuis. Journal of the American Statistical Association, 99(466):568–569, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s40/p568;   http:
//www.jstor.org/stable/27590426.

Douglas:2006:BRB

[Dou06] Jeffrey Douglas. Book review: Linear Models for Optimal Test Design, by Wim J. Van Der Linden. Journal of the American Statistical Association, 101(476):1727, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639796.

Down:2000:BRB

[Dow00] D. Down. Book review: [untitled]. Journal of the American Statistical Association, 95(449):346–347, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669582.

Dodd:2003:SRA

[DP03] Lori E. Dodd and Margaret Sullivan Pepe. Semiparametric regression for the area under the receiver operating characteristic curve. Journal of the American Statistical Association, 98(462):409–417, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s13/p409;   http://
www.jstor.org/stable/30045250.

Dukic:2005:VEM

[DP05] Vanja M. Dukić and Edsel A. Peña. Variance estimation in a model with Gaussian submodels. Journal of the American Statistical Association, 100(469):296–309, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s33/
p296;   http://www.jstor.org/stable/27590539.

Dodd:2006:CSP

[DP06] Lori E. Dodd and Margaret Sullivan Pepe. Corrections: Semiparametric regression for the area under the receiver operating characteristic curve. Journal of the American Statistical Association, 101(473):407, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047495.

DAgostino:2000:EUP

[DR00] Ralph B. D’Agostino, Jr. and Donald B. Rubin. Estimating and using propensity scores with partially missing data.

Journal of the American Statistical Association, 95(451):749– 759, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669455.

Dorazio:2005:ESC

[DR05] Robert M. Dorazio and J. Andrew Royle. Estimating size and composition of biological communities by modeling the occurrence of species. Journal of the American Statistical Association, 100(470):389–398, June 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s4/p389;   http:
//www.jstor.org/stable/27590563.

Dragalin:2006:BRB

[Dra06] Vladimir Dragalin. Book review: Inference for Change Point and Post Change Means after a CUSUM Test, by Yanhong Wu. Journal of the American Statistical Association, 101 (474):848, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590747.

Datta:2005:RST

[DS05] Somnath Datta and Glen A. Satten. Rank-sum tests for clustered data. Journal of the American Statistical Association, 100(471):908–915, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590622.

Dror:2008:SED

[DS08] Hovav A. Dror and David M. Steinberg. Sequential experimental designs for generalized linear models. Journal of the American Statistical Association, 103(481):288–298, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delouille:2004:SDA

[DSV04] V. Delouille, J. Simoens, and R. Von Sachs. Smooth design-adapted wavelets for nonparametric stochastic regression. Journal of the American Statistical Association, 99(467):643–658, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s15/p643;   http:
//www.jstor.org/stable/27590437.

Duval:2000:NTF

[DT00] Sue Duval and Richard Tweedie. A nonparametric “trim and fill” method of accounting for publication bias in metaanalysis. Journal of the American Statistical Association, 95(449):89–98, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/Duval.htm;   http:
//www.jstor.org/stable/2669529.

Doksum:2008:NVS

[DTT08] Kjell Doksum, Shijie Tang, and Kam-Wah Tsui. Nonparametric variable selection: The EARTH algorithm. Journal of the American Statistical Association, 103(484):1609–1620, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duarte:2001:BRB

[Dua01] James E. Duarte. Book review: [untitled]. Journal of the American Statistical Association, 96(454):781–782, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670324.

Dubin:2001:BRB

[Dub01] Joel Dubin. Book review: [untitled]. Journal of the American Statistical Association, 96(454):777–778, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670313.

Duivesteijn:2016:CJT

[Dui16] Wouter Duivesteijn. Correction to Jin-Ting Zhang’s “Approximate and Asymptotic Distributions of Chi-Squared-Type Mixtures With Applications”. Journal of the American Statistical Association, 111(515):1370–1371, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Zha05b].

Dunson:2003:DLT

[Dun03] David B. Dunson. Dynamic latent trait models for multidimensional longitudinal data. Journal of the American Statistical Association, 98(463):555–563, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s6/p555;   http://www.jstor.org/stable/30045281.

Dunson:2005:BSI

[Dun05] David B. Dunson. Bayesian semiparametric isotonic regression for count data. Journal of the American Statistical Association, 100(470):618–627, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s23/p618;   http://www.
jstor.org/stable/27590582.

Dunson:2008:CBB

[Dun08] David B. Dunson. Comment: Bivariate binomial spatial modeling of Loa loa prevalence in tropical Africa. Journal of the American Statistical Association, 103(481):40–41, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CDR08a, CDR08b, Pac08, Tho08].

Durkalski:2006:BRB

[Dur06] Valerie L. Durkalski. Book review: Analysis of Clinical Trials Using SAS(R): A Practical Guide, by Alex Dmitrienko; Geert Molenberghs; Christy Chuang-Stein; Walter Offen. Journal of the American Statistical Association, 101(474):858, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590761.

Duval:2003:BME

[Duv03a] Sue Duval. Biostatistical methods in epidemiology. Journal of the American Statistical Association, 98(461):251–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s30/p251.

Duval:2003:BRB

[Duv03b] Sue Duval. Book review: Biostatistical Methods in Epidemiology, by Stephen C. Newman. Journal of the American Statistical Association, 98(461):251, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045217.

Dubinin:2003:LBI

[DV03] Thomas Michael Dubinin and Stephen Bruce Vardeman. Likelihood-based inference in some continuous exponential families with unknown threshold parameters. Journal of the American Statistical Association, 98(463):741–749, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045302.

deValpine:2004:MCS

[dV04] Perry de Valpine. Monte Carlo state-space likelihoods by weighted posterior kernel density estimation. Journal of the American Statistical Association, 99(466):523–536, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s21/p523;   http://www.jstor.org/stable/27590407.

Ding:2004:TIB

[DW04] A. Adam Ding and Weijing Wang. Testing independence for bivariate current status data. Journal of the American Statistical Association, 99(465):145–155, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s15/p145;   http://www.jstor.org/stable/27590361.

Dunson:2009:NBM

[DX09] David B. Dunson and Chuanhua Xing. Nonparametric Bayes modeling of multivariate categorical data. Journal of the American Statistical Association, 104(487):1042–1051, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dunson:2008:MSB

[DXC08] David B. Dunson, Ya Xue, and Lawrence Carin. The matrix stick-breaking process. Journal of the American Statistical Association, 103(481):317–327, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dykstra:2000:BRB

[Dyk00] Richard L. Dykstra. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1017–1018, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669502.

Dunson:2000:BMF

[DZ00] David B. Dunson and Haibo Zhou. A Bayesian model for fecundability and sterility. Journal of the American Statistical Association, 95(452):1054–1062, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669742.

DeGooijer:2003:ACQ

[DZ03] Jan G. De Gooijer and Dawit Zerom. On additive conditional quantiles with high dimensional covariates. Journal of the American Statistical Association, 98(461):135–146, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s14/p135;   http://www.jstor.org/stable/30045201.

Eastwood:2003:BRB

[Eas03] Brian J. Eastwood. Book review: Applied Statistics in the Pharmaceutical Industry with Case Studies Using S-PLUS, by Steven P. Millard; Andreas Krause. Journal of the American Statistical Association, 98(461):250, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30045215.

Efron:2004:EPE

[EBD^{+}04] Bradley Efron, Prabir Burman, L. Denby, J. M. Landwehr, C. L. Mallows, Xiaotong Shen, Hsin-Cheng Huang, Jianming Ye, Jimmy Ye, and Chunming Zhang. The estimation of prediction error: Covariance penalties and cross-validation [with comments, rejoinder]. Journal of the American Statistical Association, 99(467):619–632, September 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/s9/
p619;   http://www.jstor.org/stable/27590436.

Edwards:2005:CFD

[Edw05] Don Edwards. Comment: False discovery rate-adjusted multiple confidence intervals for selected parameters. Journal of the American Statistical Association, 100(469):81–82, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s9/
p81.

Edwards:2008:CSA

[Edw08] A. W. F. Edwards. Comment: On Student’s 1908 article “The Probable Error of a Mean”. Journal of the American Statistical Association, 103(481):11–13, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2000:BMS

[Efr00] Bradley Efron. The bootstrap and modern statistics. Journal of the American Statistical Association, 95(452):1293– 1296, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669773.

Efromovich:2001:DEU

[Efr01] Sam Efromovich. Density estimation under random censorship and order restrictions: From asymptotic to small samples. Journal of the American Statistical Association, 96(454):667–684, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670305;   reqidx=/cw/
asa/01621459/v96n454/s26/p667.

Efron:2004:LSS

[Efr04a] Bradley Efron. Large-scale simultaneous hypothesis testing: The choice of a null hypothesis. Journal of the American Statistical Association, 99(465):96–104, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s10/p96;   http://www.jstor.org/stable/27590356.

Efron:2004:REP

[Efr04b] Bradley Efron. Rejoinder: The estimation of prediction error: Covariance penalties and cross-validation. Journal of the American Statistical Association, 99(467):640–642, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s14/p640.

Efromovich:2005:UNR

[Efr05a] Sam Efromovich. Univariate nonparametric regression in the presence of auxiliary covariates. Journal of the American Statistical Association, 100(472):1185–1201, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590664.

Efron:2005:BFS

[Efr05b] Bradley Efron. Bayesians, frequentists, and scientists. Journal of the American Statistical Association, 100(469):1–5, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s1/
p1;   http://www.jstor.org/stable/27590513.

Efron:2007:CLS

[Efr07] Bradley Efron. Correlation and large-scale simultaneous significance testing. Journal of the American Statistical Association, 102(477):93–103, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2009:EBE

[Efr09] Bradley Efron. Empirical Bayes estimates for large-scale prediction problems. Journal of the American Statistical Association, 104(487):1015–1028, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

ElGhouch:2009:LPQ

[EG09] Anouar El Ghouch and Marc G. Genton. Local polynomial quantile regression with parametric features. Journal of the American Statistical Association, 104(488):1416–1429, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

E:2008:FIV

[EHI08] Lidong E, Jan Hannig, and Hari Iyer. Fiducial intervals for variance components in an unbalanced two-component normal mixed linear model. Journal of the American Statistical Association, 103(482):854–865, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Elmore:2006:NDE

[EHT06] Ryan T. Elmore, Peter Hall, and Vladimir S. Troynikov. Nonparametric density estimation from covariate information. Journal of the American Statistical Association, 101(474): 701–711, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590728.

Eisenberg:2000:EML

[Eis00] Theodore Eisenberg. Empirical methods and the law. Journal of the American Statistical Association, 95(450):665–669, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669414.

Elliott:2000:BAC

[EL00] Michael R. Elliott and Roderick J. A. Little. A Bayesian approach to combining information from a census, a coverage measurement survey, and demographic analysis. Journal of the American Statistical Association, 95(450):351–362, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/2000/June/Elliott.htm;   http://www.jstor.org/
stable/2669372.

Ekstrom:2004:SME

[EL04] Magnus Ekström and Sara Sjöstedt-De Luna. Subsampling methods to estimate the variance of sample means based on nonstationary spatial data with varying expected values. Journal of the American Statistical Association, 99(465):82–95, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n465/s9/p82;   http://www.
jstor.org/stable/27590355.

Elliott:2005:BAC

[EL05] Michael R. Elliott and Roderick J. A. Little. A Bayesian approach to 2000 census evaluation using ACE survey data and demographic analysis. Journal of the American Statistical Association, 100(470):380–388, June 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s3/p380;   http:
//www.jstor.org/stable/27590562.

Elbasha:2008:BRB

[Elb08] Elamin H. Elbasha. Book review: Statistical Analysis of Cost-Effectiveness Data. Journal of the American Statistical Association, 103(482):891, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ellenberg:2000:SS

[Ell00a] Jonas H. Ellenberg. Statisticians’ significance. Journal of the American Statistical Association, 95(449):1–8, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/Ellenberg.htm;   http://www.jstor.org/stable/
2669517.

Elliott:2000:SCI

[ELL00b] Michael R. Elliott, Roderick J. A. Little, and Steve Lewitzky. Subsampling callbacks to improve survey efficiency. Journal of the American Statistical Association, 95(451):730– 738, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669453.

Elliott:2001:BRB

[Ell01] Michael Elliott. Book review: [untitled]. Journal of the American Statistical Association, 96(453):351–352, March 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670384.

Einmahl:2009:TEE

[ELL09] John H. J. Einmahl, Jun Li, and Regina Y. Liu. Thresholding events of extreme in simultaneous monitoring of multiple risks. Journal of the American Statistical Association, 104(487):982– 992, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Elsner:2002:ATS

[Els02a] James B. Elsner. Analysis of time series structure: SSA and related techniques. Journal of the American Statistical Association, 97(460):1207–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s34/p1207.

Elsner:2002:BRB

[Els02b] James B. Elsner. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1207–1208, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085853.

Elston:2002:APH

[Els02c] Robert C. Elston. [adjusting for population heterogeneity and misspecified haplotype frequencies when testing nonparametric null hypotheses in statistical genetics]: Comment. Journal of the American Statistical Association, 97(459):751– 753, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085715.

ElBarmi:2005:IUS

[EM05] Hammou El Barmi and Hari Mukerjee. Inferences under a stochastic ordering constraint: The k-sample case. Journal of the American Statistical Association, 100(469):252–261, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s29/
p252;   http://www.jstor.org/stable/27590535.

Einmahl:2008:RAT

[EM08] John H. J. Einmahl and Jan R. Magnus. Records in athletics through extreme-value theory. Journal of the American Statistical Association, 103(484):1382–1391, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Embrechts:2001:BRBa

[Emb01a] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 96(454):783, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670328.

Embrechts:2001:BRBb

[Emb01b] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 96(454):783, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670329.

Embrechts:2001:BRBc

[Emb01c] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085940.

Embrechts:2002:BRBa

[Emb02a] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 97(459):928, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085748.

Embrechts:2002:BRBb

[Emb02b] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1204, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085846.

Embrechts:2002:BRBc

[Emb02c] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1217, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085869.

Embrechts:2003:BRB

[Emb03] Paul Embrechts. Book review: Martingales and Markov Chains: Solved Exercises and Elements of Theory, by Paolo Baldi; Laurent Mazliak; Pierre Priouret. Journal of the American Statistical Association, 98(463):769, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045309.

Ensor:2002:BRB

[Ens02] Katherine Bennett Ensor. Book review: [untitled]. Journal of the American Statistical Association, 97(458):656–657, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085686.

Eide:2002:PRV

[EOU02] Alfhild Lien Eide, Henning Omre, and Bjørn Ursin. Prediction of reservoir variables based on seismic data and well observations. Journal of the American Statistical Association, 97(457):18–28, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085755;   reqidx=/cw/
asa/01621459/v97n457/s3/p18.

Espejo:2006:BRB

[EP06] Mariano Ruiz Espejo and Miguel Delgado Pineda. Book review: Estimation in Surveys with Nonresponse, by Carl-Erik Särndal; Sixten Lundström. Journal of the American Statistical Association, 101(474):854, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590754.

Epifani:2004:BRB

[Epi04] Ilenia Epifani. Book review: Bayesian Nonparametrics, by J. K. Ghosh; R. V. Ramamoorthi. Journal of the American Statistical Association, 99(467):898–899, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s43/p898;   http://www.jstor.org/stable/27590459.

Eppstein:2004:CLS

[Epp04] David Eppstein. Comment: Location-scale depth. Journal of the American Statistical Association, 99(468):976–979, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s9/p976.

Ernst:2000:BRB

[Ern00] Michael D. Ernst. Book review: [untitled]. Journal of the American Statistical Association, 95(449):333, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669566.

Elliott:2002:EDD

[ERS02] Michael R. Elliott, Trivellore E. Raghunathan, and Jean T. Shope. The effect of duration and delay of licensure on risk of crash: a Bayesian analysis of repeated time-to-event measures. Journal of the American Statistical Association, 97(458):420–431, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s6/p420;   http://
www.jstor.org/stable/3085655.

Elliott:2002:LCA

[ES02] Michael R. Elliott and Mary D. Sammel. [latent class analysis of complex sample survey data: Application to dietary data]: Comment. Journal of the American Statistical Association, 97(459):732–734, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085710.

Escanciano:2006:GFT

[Esc06] J. Carlos Escanciano. Goodness-of-fit tests for linear and nonlinear time series models. Journal of the American Statistical Association, 101(474):531–541, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590715.

Efron:2001:EBA

[ETST01] Bradley Efron, Robert Tibshirani, John D. Storey, and Virginia Tusher. Empirical Bayes analysis of a microarray experiment. Journal of the American Statistical Association, 96(456):1151–1160, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085878;   reqidx=/cw/
asa/01621459/v96n456/s2/p1151.

Eubank:2007:BRB

[Eub07] Lane A. Eubank. Book review: Introduction to Linear Models and Statistical Inference. Journal of the American Statistical Association, 102(477):385–386, March 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Evans:2000:ADV

[Eva00a] Michael Evans. Asymptotic distribution of P values in composite null models: Comment. Journal of the American Statistical Association, 95(452):1160–1163, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669752.

Evans:2000:POM

[Eva00b] Michael Evans. On the probability of observing misleading statistical evidence: Comment. Journal of the American Statistical Association, 95(451):768–769, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669457.

Ellis:2008:LCB

[EW08] Byron Ellis and Wing Hung Wong. Learning causal Bayesian network structures from experimental data. Journal of the American Statistical Association, 103(482):778–789, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2000:PNM

[Fan00] Jianqing Fan. Prospects of nonparametric modeling. Journal of the American Statistical Association, 95(452):1296– 1300, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669774.

Faraway:2001:IDF

[Far01] Julian Faraway. Inference for density families using functional principal component analysis: Comment. Journal of the American Statistical Association, 96(454):537, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670293.

Fine:2000:RAR

[FB00] Jason P. Fine and Ronald J. Bosch. Risk assessment via a robust probit model, with application to toxicology. Journal of the American Statistical Association, 95(450): 375–382, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/2000/June/Fine.htm;   http://
www.jstor.org/stable/2669374.

Frangakis:2004:MEP

[FBV^{+}04] Constantine E. Frangakis, Ronald S. Brookmeyer, Ravi Varadhan, Mahboobeh Safaeian, David Vlahov, and Steffanie A. Strathdee. Methodology for evaluating a partially controlled longitudinal treatment using principal stratification, with application to a needle exchange program. Journal of the American Statistical Association, 99(465):239–249, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s23/p239;   http://www.jstor.org/stable/27590369.

Foulkes:2003:CPV

[FD03] A. S. Foulkes and V. De Gruttola. Characterizing the progression of viral mutations over time. Journal of the American Statistical Association, 98(464):859–867, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/s9/
p859;   http://www.jstor.org/stable/30045337.

Fouskakis:2008:CSO

[FD08] Dimitris Fouskakis and David Draper. Comparing stochastic optimization methods for variable selection in binary outcome prediction, with application to health policy. Journal of the American Statistical Association, 103(484):1367–1381, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fried:2007:WRM

[FEG07] Roland Fried, Jochen Einbeck, and Ursula Gather. Weighted repeated median smoothing and filtering. Journal of the American Statistical Association, 102(480):1300–1308, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feldman:2002:BRBa

[Fel02a] Raisa E. Feldman. Book review: [untitled]. Journal of the American Statistical Association, 97(458):658–659, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085691.

Feldman:2002:BRBb

[Fel02b] Raisa E. Feldman. Book review: [untitled]. Journal of the American Statistical Association, 97(459):925–926, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085742.

Feldman:2002:LDS

[Fel02c] Raisa E. Feldman. Limit distributions for sums of independent random vectors. Journal of the American Statistical Association, 97(459):925–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s32/p925.

Fewster:2004:BRB

[Few04] Rachel M. Fewster. Book review: Estimating Animal Abundance: Closed Populations, by David Borchers; Stephen T. Buckland; Walter Zucchini. Journal of the American Statistical Association, 99(466):558, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s24/p558;   http:
//www.jstor.org/stable/27590410.

Fan:2006:CQA

[FF06] Jianqing Fan and Yingying Fan. Comment: Quantile autoregression. Journal of the American Statistical Association, 101 (475):991–994, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2009:CNP

[FF09] Jianqing Fan and Yang Feng. Comment: Nonparametric prediction in measurement error models. Journal of the American Statistical Association, 104(487):1003–1007, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2007:DIT

[FFJ07] Jianqing Fan, Yingying Fan, and Jiancheng Jiang. Dynamic integration of time- and state-domain methods for volatility estimation. Journal of the American Statistical Association, 102(478):618–631, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fischetti:2000:MAO

[FG00] Matteo Fischetti and Juan Jose Salazar Gonzalez. Models and algorithms for optimizing cell suppression in tabular data with linear constraints. Journal of the American Statistical Association, 95(451):916–928, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669474.

Fischer:2004:SME

[FG04] Krista Fischer and Els Goetghebeur. Structural mean effects of noncompliance: Estimating interaction with baseline prognosis and selection effects. Journal of the American Statistical Association, 99(468):918–928, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/s2/
p918;   http://www.jstor.org/stable/27590475.

Faes:2008:FMM

[FGM^{+}08] Christel Faes, Helena Geys, Geert Molenberghs, Marc Aerts, Carmen Cadarso-Suárez, Carlos Acuña, and Mónica Cano. A flexible method to measure synchrony in neuronal firing. Journal of the American Statistical Association, 103(481):149–161, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feuerverger:2000:MDE

[FH00] Andrey Feuerverger and Peter Hall. Methods for density estimation in thick-slice versions of Wicksell’s problem. Journal of the American Statistical Association, 95(450):535–546, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/2000/June/Feuerverger.htm;   http://www.jstor.
org/stable/2669397.

Fan:2001:GFT

[FH01] Jianqing Fan and Li-Shan Huang. Goodness-of-fit tests for parametric regression models. Journal of the American Statistical Association, 96(454):640–652, June 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670303;   reqidx=/cw/asa/01621459/v96n454/s24/p640.

Fan:2007:ALD

[FHL07] Jianqing Fan, Tao Huang, and Runze Li. Analysis of longitudinal data with semiparametric estimation of covariance function. Journal of the American Statistical Association, 102 (478):632–641, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Forni:2005:GDF

[FHLR05] Mario Forni, Marc Hallin, Marco Lippi, and Lucrezia Reichlin. The generalized dynamic factor model: One-sided estimation and forecasting. Journal of the American Statistical Association, 100(471):830–840, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590616.

Fan:2000:GUL

[FHW00] Jianqing Fan, Hui-Nien Hung, and Wing-Hung Wong. Geometric understanding of likelihood ratio statistics. Journal of the American Statistical Association, 95(451):836– 841, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669467.

Fan:2007:HMS

[FHY07] Jianqing Fan, Peter Hall, and Qiwei Yao. To how many simultaneous hypothesis tests can normal, Student’s t or bootstrap calibration be applied? Journal of the American Statistical Association, 102(480):1282–1288, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fairley:2001:CII

[FIC01] William B. Fairley, Alan J. Izenman, and Steven M. Crunk. Combining incomplete information from independent assessment surveys for estimating masonry deterioration. Journal of the American Statistical Association, 96(454):488–499, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/

stable/2670287;   reqidx=/cw/asa/01621459/v96n454/
s13/p488.

Fienberg:2000:CTL

[Fie00] Stephen E. Fienberg. Contingency tables and log-linear models: Basic results and new developments. Journal of the American Statistical Association, 95(450):643–647, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669409.

Fienberg:2006:BRB

[Fie06] Stephen E. Fienberg. Book review: Analysis of Survey Data, by R. L. Chambers; C. J. Skinner. Journal of the American Statistical Association, 101(473):401–402, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047487.

Filipovic:2006:BRB

[Fil06] Damir Filipovic. Book review: Probability: A Graduate Course, by Allan Gut. Journal of the American Statistical Association, 101(474):845, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590742.

Fan:2005:NIA

[FJ05] Jianqing Fan and Jiancheng Jiang. Nonparametric inferences for additive models. Journal of the American Statistical Association, 100(471):890–907, September 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590621.

Fraiman:2008:SVC

[FJS08] Ricardo Fraiman, Ana Justel, and Marcela Svarc. Selection of variables for cluster analysis and classification rules. Journal of the American Statistical Association, 103(483):1294–1303, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Friguet:2009:FMA

[FKC09] Chloé Friguet, Maela Kloareg, and David Causeur. A factor model approach to multiple testing under dependence. Journal of the American Statistical Association, 104(488):1406–1415, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fernandez:2002:MOP

[FKS02] Carmen Fernández, Gary Koop, and Mark F. J. Steel. Multiple-output production with undesirable outputs: An application to nitrogen surplus in agriculture. Journal of the American Statistical Association, 97(458):432–442, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s7/p432;   http://www.jstor.org/stable/3085656.

French:2001:SNM

[FKW01] J. L. French, E. E. Kammann, and M. P. Wand. [semiparametric nonlinear mixed-effects models and their applications]: Comment. Journal of the American Statistical Association, 96 (456):1285–1288, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085897.

Fan:2001:VSN

[FL01] Jianqing Fan and Runze Li. Variable selection via nonconcave penalized likelihood and its oracle properties. Journal of the American Statistical Association, 96(456):1348–1360, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085904;   reqidx=/cw/asa/01621459/v96n456/
s17/p1348.

Fan:2004:NEM

[FL04] Jianqing Fan and Runze Li. New estimation and model selection procedures for semiparametric modeling in longitudinal data analysis. Journal of the American Statistical Association, 99(467):710–723, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s20/p710;   http:
//www.jstor.org/stable/27590442.

Fan:2009:CBM

[FL09] Xiaodan Fan and Jun S. Liu. Comment: a Bayesian model for cross-study differential gene expression. Journal of the American Statistical Association, 104(488):1314–1318, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fligner:2006:NTS

[FM06] Michael A. Fligner and Steven N. MacEachern. Nonparametric two-sample methods for ranked-set sample data. Journal of the American Statistical Association, 101(475):1107– 1118, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590788.

Fan:2009:OPM

[FM09] Jianqing Fan and Loriano Mancini. Option pricing with modelguided nonparametric methods. Journal of the American Statistical Association, 104(488):1351–1372, December 2009. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fryzlewicz:2009:CCN

[FO09] Piotr Fryzlewicz and Hernando Ombao. Consistent classification of nonstationary time series using stochastic wavelet representations. Journal of the American Statistical Association, 104(485):299–312, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frey:2007:NTP

[FOD07] Jesse Frey, Omer Ozturk, and Jayant V. Deshpande. Nonparametric tests for perfect judgment rankings. Journal of the American Statistical Association, 102(478):708–717, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Follmann:2000:ETA

[Fol00] Dean A. Follmann. On the effect of treatment among wouldbe treatment compliers: An analysis of the multiple risk factor intervention trial. Journal of the American Statistical Association, 95(452):1101–1109, December 2000. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669746.

Fan:2005:SHD

[FPH^{+}05] Jianqing Fan, Heng Peng, Tao Huang, Chiara Sabatti, Bruce A. Craig, Jian Huang, Cun-Hui Zhang, Heping Zhang, Michael R. Kosorok, Shuangge Ma, Robert Tibshirani, and Yi Ren. Semilinear high-dimensional model for normalization of microarray data: a theoretical analysis and partial consistency [with comments, rejoinder]. Journal of the American Statistical Association, 100(471):781–796, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590614.

Fan:2005:RSH

[FPHR05] Jianqing Fan, Heng Peng, Tao Huang, and Yi Ren. Rejoinder: Semilinear high-dimensional model for normalization of microarray data. Journal of the American Statistical Association, 100(471):808–813, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fraley:2002:MBC

[FR02] Chris Fraley and Adrian E. Raftery. Model-based clustering, discriminant analysis, and density estimation. Journal of the American Statistical Association, 97(458):611–631, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s22/p611;   http://www.jstor.org/stable/3085676.

Frame:2008:BRB

[Fra08] Samuel J. Frame. Book review: Simulation and Monte Carlo: With Applications in Finance and MCMC. Journal of the American Statistical Association, 103(482):890–891, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frey:2004:BRB

[Fre04] Jesse Frey. Book review: Introduction to Stochastic Search and Optimization: Estimation, Simulation, and Control, by James C. Spall. Journal of the American Statistical Association, 99(468):1204–1205, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s35/p1204;   http://www.jstor.org/stable/27590502.

Friedl:2002:BRB

[Fri02a] Herwig Friedl. Book review: [untitled]. Journal of the American Statistical Association, 97(457):362, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085801.

Fries:2002:RJH

[Fri02b] Arthur Fries. [response to Joel L. Horowitz article]. Journal of the American Statistical Association, 97(458):660, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085695. See

[Hor01a, Hor01b, HK02].

Fricks:2007:BRB

[Fri07] John Fricks. Book review: Stochastic Processes and Models. Journal of the American Statistical Association, 102(477):381, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frolich:2008:STC

[Frö08] Markus Frölich. Statistical treatment choice. Journal of the American Statistical Association, 103(482):547–558, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fernandez:2006:IAI

[FRS06] Miguel A. Fernández, Cristina Rueda, and Bonifacio Salvador. Incorporating additional information to normal linear discriminant rules. Journal of the American Statistical Association, 101(474):569–577, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590718.

Fokianos:2009:PA

[FRT09] Konstantinos Fokianos, Anders Rahbek, and Dag Tjøstheim. Poisson autoregression. Journal of the American Statistical Association, 104(488):1430–1439, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frydman:2005:EMM

[Fry05] Halina Frydman. Estimation in the mixture of Markov chains moving with different speeds. Journal of the American Statistical Association, 100(471):1046–1053, September 2005. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590634.

Fryzlewicz:2007:UHT

[Fry07] Piotr Fryzlewicz. Unbalanced Haar technique for nonparametric function estimation. Journal of the American Statistical Association, 102(480):1318–1327, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Francq:2005:DCA

[FRZ05] Christian Francq, Roch Roy, and Jean-Michel Zakoïan. Diagnostic checking in ARMA models with uncorrelated errors. Journal of the American Statistical Association, 100(470):532–544, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s16/p532;   http://www.
jstor.org/stable/27590575.

Fruhwirth-Schnatter:2000:BRB

[FS00] Sylvia Fruhwirth-Schnatter. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1373– 1374, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669796.

Fruhwirth-Schnatter:2001:MCM

[FS01] Sylvia Frühwirth-Schnatter. Markov chain Monte Carlo estimation of classical and dynamic switching and mixture models. Journal of the American Statistical Association, 96(453):194–209, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670359;   reqidx=/cw/
asa/01621459/v96n453/s15/p194.

Ferreira:2004:TDB

[FS04a] E. Ferreira and W. Stute. Testing for differences between conditional means in a time series context. Journal of the American Statistical Association, 99(465):169–174, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s17/p169;   http://www.jstor.org/stable/27590363.

Foster:2004:VSD

[FS04b] Dean P. Foster and Robert A. Stine. Variable selection in data mining: Building a predictive model for bankruptcy. Journal of the American Statistical Association, 99(466):303–313, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n466/s1/p303;   http://www.
jstor.org/stable/27590387.

Ferreira:2006:CRU

[FS06a] José T. A. S. Ferreira and Mark F. J. Steel. A constructive representation of univariate skewed distributions. Journal of the American Statistical Association, 101(474):823–829, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590738.

Finner:2006:EBS

[FS06b] H. Finner and K. Strassburger. On δ-equivalence with the best in k-sample models. Journal of the American Statistical Association, 101(474):737–746, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590731.

Fan:2006:TCS

[FSL^{+}06] Juanjuan Fan, Xiao-Gang Su, Richard A. Levine, Martha E. Nunn, and Michael LeBlanc. Trees for correlated survival data by goodness of split, with applications to tooth prognosis. Journal of the American Statistical Association, 101 (475):959–967, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590775.

Farewell:2004:IDE

[FTR04] Vern T. Farewell, Brian D. M. Tom, and Patrick Royston. The impact of dichotomization on the efficiency of testing for an interaction effect in exponential family models. Journal of the American Statistical Association, 99(467):822–831, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s30/p822;   http://www.jstor.org/stable/27590452.

Foster:2001:EBC

[FTW01] A. M. Foster, L. Tian, and L. J. Wei. Estimation for the Box–Cox transformation model without assuming parametric error distribution. Journal of the American Statistical Association, 96(455):1097–1101, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670255;   reqidx=/cw/
asa/01621459/v96n455/s27/p1097.

Fuentes:2003:BRB

[Fue03] Montserrat Fuentes. Book review: Predictions in Time Series Using Regression Models, by Frantisek Stulajter. Journal of the American Statistical Association, 98(463):768– 769, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045308.

Fuentes:2007:ALL

[Fue07] Montserrat Fuentes. Approximate likelihood for large irregularly spaced spatial data. Journal of the American Statistical Association, 102(477):321–331, March 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fulkerson:2002:BRB

[Ful02] William F. Fulkerson. Book review: [untitled]. Journal of the American Statistical Association, 97(457):366, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085810.

Frees:2008:HIC

[FV08] Edward W. Frees and Emiliano A. Valdez. Hierarchical insurance claims modeling. Journal of the American Statistical Association, 103(484):1457–1469, December 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fearnhead:2009:BAI

[FV09] Paul Fearnhead and Despina Vasileiou. Bayesian analysis of isochores. Journal of the American Statistical Association, 104 (485):132–141, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2000:PLC

[FW00a] Jianqing Fan and Wing-Hung Wong. On profile likelihood: Comment. Journal of the American Statistical Association, 95 (450):468–471, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669388.

Fang:2000:IVM

[FW00b] Zhide Fang and Douglas P. Wiens. Integer-valued, minimax robust designs for estimation and extrapolation in heteroscedastic, approximately linear models. Journal of the American Statistical Association, 95(451):807–818, September 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669464.

Farrington:2005:CSM

[FW05] C. Paddy Farrington and Heather J. Whitaker. Contact surface models for infectious diseases: Estimation from serologic survey data. Journal of the American Statistical Association, 100(470):370–379, June 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s2/p370;   http:
//www.jstor.org/stable/27590561.

Fan:2007:MSJ

[FW07] Jianqing Fan and Yazhen Wang. Multi-scale jump and volatility analysis for high-frequency financial data. Journal of the American Statistical Association, 102(480):1349–1362, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2008:SEC

[FW08] Jianqing Fan and Yichao Wu. Semiparametric estimation of covariance matrixes for longitudinal data. Journal of the American Statistical Association, 103(484):1520–1533, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2005:FSM

[FWP05] Shibao Feng, Robert A. Wolfe, and Friedrich K. Port. Frailty survival model analysis of the national deceased donor kidney transplant dataset using Poisson variance structures. Journal of the American Statistical Association, 100(471):728– 735, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590610.

Fan:2003:RDE

[FZ03] Jianqing Fan and Chunming Zhang. A reexamination of diffusion estimators with applications to financial model validation. Journal of the American Statistical Association, 98(461):118–134, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s13/p118;   http://
www.jstor.org/stable/30045200.

Francq:2009:TNG

[FZ09] Christian Francq and Jean-Michel Zakoïan. Testing the nullity of GARCH coefficients: Correction of the standard tests and relative efficiency comparisons. Journal of the American Statistical Association, 104(485):313–324, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gueorguieva:2001:CPM

[GA01] Ralitza V. Gueorguieva and Alan Agresti. A correlated probit model for joint modeling of clustered binary and continuous responses. Journal of the American Statistical Association, 96(455):1102–1112, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670256;   reqidx=/cw/
asa/01621459/v96n455/s28/p1102.

Gao:2005:UNA

[GA05] Xin Gao and Mayer Alvo. A unified nonparametric approach for unbalanced factorial designs. Journal of the American Statistical Association, 100(471):926–941, September 2005. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590624.

Gadbury:2000:BRB

[Gad00] Gary Gadbury. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1013–1014, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669497.

Gallagher:2005:BRB

[Gal05] Colin M. Gallagher. Book review: Diagnostic Checks in Time Series, by Wai Keung Li. Journal of the American Statistical Association, 100(471):1091–1092, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590638.

Gallagher:2007:BRB

[Gal07] Colin Gallagher. Book review: Extending the Linear Model With R: Generalized Linear, Mixed Effects and Nonparametric Regression Models. Journal of the American Statistical Association, 102(480):1477, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gani:2002:BRB

[Gan02] Joseph Gani. Book review: [untitled]. Journal of the American Statistical Association, 97(458):659, June 2002. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/3085692.

Gander:2003:BRB

[Gan03a] Walter Gander. Book review: From Elementary Probability to Stochastic Differential Equations with MAPLE, by Sasha Cyganowski; Peter Kloeden; Jerzy Ombach. Journal of the American Statistical Association, 98(461):254–255, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045222.

Gander:2003:EPS

[Gan03b] Walter Gander. From elementary probability to stochastic differential equations with MAPLE. Journal of the American Statistical Association, 98(461):254–??, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s35/p254.

Gandy:2009:SIM

[Gan09] Axel Gandy. Sequential implementation of Monte Carlo tests with uniformly bounded resampling risk. Journal of the American Statistical Association, 104(488):1504–1511, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garbers:2000:BRB

[Gar00a] Hermann Garbers. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1373, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669795.

Garcia:2000:BRB

[Gar00b] Rene Garcia. Book review: [untitled]. Journal of the American Statistical Association, 95(450):688, June 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669430.

Gasser:2003:BRB

[Gas03a] Theo Gasser. Book review: Nonlinear Models in Medical Statistics, by J. K. Lindsey. Journal of the American Statistical Association, 98(461):251–252, March 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30045218.

Gasser:2003:NMM

[Gas03b] Theo Gasser. Nonlinear models in medical statistics. Journal of the American Statistical Association, 98(461):251–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s31/p251.

Gray:2000:MLD

[GB00] Sarah M. Gray and Ron Brookmeyer. Multidimensional longitudinal data: Estimating a treatment effect from continuous, discrete, or time-to-event response variables. Journal of the American Statistical Association, 95(450):396–406, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/2000/June/Gray.htm;   http://www.
jstor.org/stable/2669376.

Goldman:2001:EIM

[GBM^{+}01] Dana P. Goldman, Jayanta Bhattacharya, Daniel F. Mc-Caffrey, Naihua Duan, Arleen A. Leibowitz, Geoffrey F. Joyce, and Sally C. Morton. Effect of insurance on mortality in an HIV-positive population in care. Journal of the American Statistical Association, 96(455):883–894, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670228;   reqidx=/cw/asa/01621459/v96n455/
s10/p883.

Ghosh:2009:BAC

[GBT09] Pulak Ghosh, Sanjib Basu, and Ram C. Tiwari. Bayesian analysis of cancer rates from SEER program using parametric and semiparametric joinpoint regression models. Journal of the American Statistical Association, 104(486):439–452, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:2009:CBM

[GC09a] Debashis Ghosh and Hyungwon Choi. Comment: a Bayesian model for cross-study differential gene expression. Journal of the American Statistical Association, 104(488):1310–1314, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gill:2009:NPO

[GC09b] Jeff Gill and George Casella. Nonparametric priors for ordinal Bayesian social science models: Specification and estimation. Journal of the American Statistical Association, 104(486):453– 454, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gerlach:2000:EBI

[GCK00] Richard Gerlach, Chris Carter, and Robert Kohn. Efficient Bayesian inference for dynamic mixture models. Journal of the American Statistical Association, 95(451):819–828, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669465.

Guo:2009:BHM

[GDH09] Feng Guo, Dipak K. Dey, and Kent E. Holsinger. A Bayesian hierarchical model for analysis of single-nucleotide polymorphisms diversity in multilocus, multipopulation samples. Journal of the American Statistical Association, 104(485):142–154, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2003:SSA

[GDOvS03] Wensheng Guo, Ming Dai, Hernando C. Ombao, and Rainer von Sachs. Smoothing spline ANOVA for time-dependent spectral analysis. Journal of the American Statistical Association, 98(463):643–652, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s16/p643;   http://
www.jstor.org/stable/30045291.

Ghosh-Dastidar:2003:MEM

[GDS03] Bonnie Ghosh-Dastidar and Joseph L. Schafer. Multiple edit/multiple imputation for multivariate continuous data. Journal of the American Statistical Association, 98 (464):807–817, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s4/p807;   http://www.
jstor.org/stable/30045332.

Godsill:2004:MCS

[GDW04] Simon J. Godsill, Arnaud Doucet, and Mike West. Monte Carlo smoothing for nonlinear time series. Journal of the American Statistical Association, 99(465):156–168, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s16/p156;   http://www.jstor.org/stable/27590362.

Garcia-Escudero:2005:GRP

[GEG05] Luis Angel García-Escudero and Alfonso Gordaliza. Generalized radius processes for elliptically contoured distributions. Journal of the American Statistical Association, 100 (471):1036–1045, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590633.

Gelfand:2000:GS

[Gel00] Alan E. Gelfand. Gibbs sampling. Journal of the American Statistical Association, 95(452):1300–1304, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669775.

Gelman:2004:PBM

[Gel04] Andrew Gelman. Parameterization and Bayesian modeling. Journal of the American Statistical Association, 99 (466):537–545, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n466/s22/p537;   http://www.
jstor.org/stable/27590408.

Gelman:2008:REI

[Gel08] Andrew Gelman. Rejoinder: Estimating incumbency advantage and its variation, as an example of a before–after study. Journal of the American Statistical Association, 103(482):449– 451, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geman:2006:BRB

[Gem06] Don Geman. Book review: Modeling and Inverse Problems in Image Analysis, by Bernard Chalmond. Journal of the American Statistical Association, 101(474):847, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590745.

Genovese:2000:BTCa

[Gen00a] Christopher R. Genovese. A Bayesian time-course model for functional magnetic resonance imaging data. Journal of the American Statistical Association, 95(451):691–703, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669445.

Genovese:2000:BTCb

[Gen00b] Christopher R. Genovese. A Bayesian time-course model for functional magnetic resonance imaging data: Rejoinder. Journal of the American Statistical Association, 95(451):716– 719, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669451.

Gentle:2001:BRB

[Gen01] James Gentle. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1136–1137, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670261.

Genest:2002:BRB

[Gen02] Christian Genest. Book review: [untitled]. Journal of the American Statistical Association, 97(458):653–654, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085682.

Gentleman:2004:BRB

[Gen04] Robert Gentleman. Book review: Current Topics in Computational Molecular Biology, by Tao Jiang; Ying Xu; Michael Q. Zhang. Journal of the American Statistical Association, 99(466):560, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s26/p560;   http:
//www.jstor.org/stable/27590412.

Gentleman:2005:BRBb

[Gen05a] Robert Gentleman. Book review: Design and Analysis of DNA Microarray Investigations, by Richard M. Simon; Edward L. Korn; Lisa M. McShane; Michael D. Radmacher; George W. Wright; Yingdong Zhao. Journal of the American Statistical Association, 100(470):711–712, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590600.

Gentleman:2005:BRBa

[Gen05b] Robert Gentleman. Book review: Microarray Quality Control, by Wei Zhang; Ilya Shmulevich; Jaakko Astola. Journal of the American Statistical Association, 100(470):711, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590599.

Genovese:2006:BRB

[Gen06] Christopher Genovese. Book review: Multivariate Bayesian Statistics: Models for Source Separation and Signal Unmixing, by Daniel B. Rowe. Journal of the American Statistical Association, 101(475):1314–1315, September 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590814.

Genest:2007:BRB

[Gen07a] Christian Genest. Book review: Reliability and Risk: A Bayesian Perspective. Journal of the American Statistical Association, 102(480):1475, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gentle:2007:BRB

[Gen07b] James E. Gentle. Book review: Vertical Density Representations and Its Applications. Journal of the American Statistical Association, 102(477):383, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

George:2000:VSP

[Geo00] Edward I. George. The variable selection problem. Journal of the American Statistical Association, 95(452):1304– 1308, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669776.

Geweke:2004:GIR

[Gew04] John Geweke. Getting it right: Joint distribution tests of posterior simulators. Journal of the American Statistical Association, 99(467):799–804, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s27/p799;   http:
//www.jstor.org/stable/27590449.

Geyer:2001:BRB

[Gey01] C. Geyer. Book review: [untitled]. Journal of the American Statistical Association, 96(454):779, June 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670317.

Gelman:2007:ANY

[GFK07] Andrew Gelman, Jeffrey Fagan, and Alex Kiss. An analysis of the New York City Police Department’s “stop-and-frisk” policy in the context of claims of racial bias. Journal of the American Statistical Association, 102(479):813–823, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gonzalez:2001:BHM

[GG01] Enrique González and Josep Ginebra. Bayesian heuristic for multiperiod control. Journal of the American Statistical Association, 96(455):1113–1121, September 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670257;   reqidx=/cw/asa/01621459/v96n455/s29/p1113.

Gluhovsky:2007:SLD

[GG07] Ilya Gluhovsky and Alexander Gluhovsky. Smooth locationdependent bandwidth selection for local polynomial regression. Journal of the American Statistical Association, 102(478):718– 725, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gao:2008:BSN

[GG08] Jiti Gao and Irène Gijbels. Bandwidth selection in nonparametric kernel testing. Journal of the American Statistical Association, 103(484):1584–1594, December 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Genest:2007:RBE

[GGR07] Christian Genest, Kilani Ghoudi, and Bruno Rémillard. Rankbased extensions of the brock, dechert, and scheinkman test. Journal of the American Statistical Association, 102(480): 1363–1376, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gunter:2000:SPP

[GH00] Bert Gunter and Dan Holder. Statistics in preclinical pharmaceutical research and development. Journal of the American Statistical Association, 95(451):998–1001, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669485.

Gelman:2008:EIA

[GH08] Andrew Gelman and Zaiying Huang. Estimating incumbency advantage and its variation, as an example of a before–after study. Journal of the American Statistical Association, 103 (482):437–446, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gheorghe:2001:BRB

[Ghe01] Adrian V. Gheorghe. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1139–1140, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2670265.

Ghosh:2001:BRB

[Gho01] Subir Ghosh. Book review: [untitled]. Journal of the American Statistical Association, 96(453):352, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670385.

Ghosh:2006:BRB

[Gho06] Malay Ghosh. Book review: Bayesian Nonparametrics via Neural Networks, by Herbert K. H. Lee. Journal of the American Statistical Association, 101(475):1313, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590812.

Ghosh:2007:BRB

[Gho07] Anil K. Ghosh. Book review: Applied MANOVA and Discriminant Analysis. Journal of the American Statistical Association, 102(479):1075–1076, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gupta:2007:VSR

[GI07] Mayetri Gupta and Joseph G. Ibrahim. Variable selection in regression mixture modeling for the discovery of gene regulatory networks. Journal of the American Statistical Association, 102(479):867–880, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gianola:2000:SAB

[Gia00] Daniel Gianola. Statistics in animal breeding. Journal of the American Statistical Association, 95(449):296–299, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669551.

Gilbert:2003:DRG

[Gil03] Scott Gilbert. Distribution of rankings for groups exhibiting heteroscedasticity and correlation. Journal of the American Statistical Association, 98(461):147–157, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s15/p147;   http://www.jstor.org/stable/30045202.

Gilbert:2008:BRB

[Gil08a] Gregory E. Gilbert. Book review: Linear Mixed Models: A Practical Guide Using Statistical Software. Journal of the American Statistical Association, 103(481):427–428, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gill:2008:BRB

[Gil08b] Jeff Gill. Book review: Introduction to Applied Bayesian Statistics and Estimation for Social Scientists. Journal of the American Statistical Association, 103(483):1322–1323, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gabrys:2007:PTI

[GK07] Robertas Gabrys and Piotr Kokoszka. Portmanteau test of independence for functional observations. Journal of the American Statistical Association, 102(480):1338–1348, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:2005:BNS

[GKM05] Alan E. Gelfand, Athanasios Kottas, and Steven N. MacEachern. Bayesian nonparametric spatial modeling with Dirichlet process mixing. Journal of the American Statistical Association, 100(471):1021–1035, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590632.

Garthwaite:2005:SME

[GKO05] Paul H. Garthwaite, Joseph B. Kadane, and Anthony O’Hagan. Statistical methods for eliciting probability distributions. Journal of the American Statistical Association, 100(470):680–700, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s28/p680;   http://www.
jstor.org/stable/27590587.

Gelfand:2003:SMS

[GKSB03] Alan E. Gelfand, Hyon-Jung Kim, C. F. Sirmans, and Sudipto Banerjee. Spatial modeling with spatially varying coefficient processes. Journal of the American Statistical Association, 98(462):387–396, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s11/p387;   http://
www.jstor.org/stable/30045248.

Gupta:2003:DCS

[GL03] Mayetri Gupta and Jun S. Liu. Discovery of conserved sequence patterns using a stochastic dictionary model. Journal of the American Statistical Association, 98(461):55–66, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s7/p55;   http://www.jstor.org/stable/30045194.

Griffin:2006:AFT

[GL06] Beth Ann Griffin and Stephen Lagakos. Analysis of failure time data arising from studies with alternating treatment schedules. Journal of the American Statistical Association, 101 (474):510–520, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590713.

Guan:2007:TBB

[GL07] Yongtao Guan and Ji Meng Loh. A thinned block bootstrap variance estimation procedure for inhomogeneous spatial point patterns. Journal of the American Statistical Association, 102 (480):1377–1386, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gramacy:2008:BTG

[GL08] Robert B. Gramacy and Herbert K. H. Lee. Bayesian treed Gaussian process models with an application to computer modeling. Journal of the American Statistical Association, 103 (483):1119–1130, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guha:2008:BHM

[GLN08] Subharup Guha, Yi Li, and Donna Neuberg. Bayesian hidden Markov modeling of array CGH data. Journal of the American Statistical Association, 103(482):485–497, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garratt:2003:FUM

[GLPS03] Anthony Garratt, Kevin Lee, M. Hashem Pesaran, and Yongcheol Shin. Forecast uncertainties in macroeconomic modeling: An application to the U.K. economy. Journal of the American Statistical Association, 98(464):829–838, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/s6/
p829;   http://www.jstor.org/stable/30045334.

Gneiting:2006:CPF

[GLW^{+}06] Tilmann Gneiting, Kristin Larson, Kenneth Westrick, Marc G. Genton, and Eric Aldrich. Calibrated probabilistic forecasting at the Stateline Wind Energy Center: The regime-switching space: Time method. Journal of the American Statistical Association, 101(475):968–979, September 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590776.

Guha:2006:GPI

[GM06] Subharup Guha and Steven N. MacEachern. Generalized poststratification and importance sampling for subsampled Markov chain Monte Carlo estimation. Journal of the American Statistical Association, 101(475):1175–1184, September 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590793.

Gormley:2008:EVB

[GM08] Isobel Claire Gormley and Thomas Brendan Murphy. Exploring voting blocs within the Irish electorate. Journal of the American Statistical Association, 103(483):1014–1027, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gallant:2009:DGS

[GM09] A. Ronald Gallant and Robert E. McCulloch. On the determination of general scientific models with application to asset pricing. Journal of the American Statistical Association, 104 (485):117–131, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:2004:HBN

[GMK^{+}04] Malay Ghosh, Tapabrata Maiti, Dalho Kim, Sounak Chakraborty, and Ashutosh Tewari. Hierarchical Bayesian neural networks: An application to a prostate cancer study. Journal of the American Statistical Association, 99(467):601–608, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/s7/
p601;   http://www.jstor.org/stable/27590434.

Gneiting:2002:NSC

[Gne02] Tilmann Gneiting. Nonseparable, stationary covariance functions for space–time data. Journal of the American Statistical Association, 97(458):590–600, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s20/p590;   http://
www.jstor.org/stable/3085674.

Golan:2002:BRB

[Gol02a] Amos Golan. Book review: [untitled]. Journal of the American Statistical Association, 97(458):655–656, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085685.

Goldsman:2002:BRB

[Gol02b] David Goldsman. Book review: [untitled]. Journal of the American Statistical Association, 97(459):924–925, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085740.

Goldsman:2002:ISR

[Gol02c] David Goldsman. Introductory statistics and random phenomena: Uncertainty, complexity and chaotic behavior in engineering and science. Journal of the American Statistical Association, 97(459):924–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s30/p924.

Goldie:2005:BRB

[Gol05a] Charles M. Goldie. Book review: Distribution Theory of Runs and Patterns and Its Applications: A Finite Markov Chain Imbedding Approach, by James C. Fu; W. Y. Lou. Journal of the American Statistical Association, 100(469):349–350, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s39/p350;   http://www.
jstor.org/stable/27590545.

Goldstein:2005:BRB

[Gol05b] Darlene R. Goldstein. Book review: Analyzing Microarray Gene Expression Data, by Geoffrey J. McLachlan; Kim-Anh Do; Christophe Ambroise. Journal of the American Statistical Association, 100(472):1464–1465, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590698.

Goldie:2007:BRB

[Gol07] Charles M. Goldie. Book review: Probabilistic Symmetries and Invariance Principles. Journal of the American Statistical Association, 102(479):1078–1079, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gore:2008:BRB

[Gor08] Sharad D. Gore. Book review: Sampling for Natural Resource Monitoring. Journal of the American Statistical Association, 103(482):889–890, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Givens:2001:EFG

[GP01] Geof H. Givens and David Poole. Estimation of fetal growth and gestation in bowhead whales: Comment. Journal of the American Statistical Association, 96(455):926–928, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670233.

Gomes:2007:SRB

[GP07] M. Ivette Gomes and Dinis Pestana. A sturdy reduced-bias extreme quantile (VaR) estimator. Journal of the American Statistical Association, 102(477):280–292, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Galeano:2006:ODM

[GPT06] Pedro Galeano, Daniel Peña, and Ruey S. Tsay. Outlier detection in multivariate time series by projection pursuit. Journal of the American Statistical Association, 101(474): 654–669, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590725.

Green:2002:HMM

[GR02] Peter J. Green and Sylvia Richardson. Hidden Markov models and disease mapping. Journal of the American Statistical Association, 97(460):1055–1070, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s11/p1055;   http://www.jstor.org/stable/3085830.

Genton:2003:RII

[GR03] Marc G. Genton and Elvezio Ronchetti. Robust indirect inference. Journal of the American Statistical Association, 98(461):67–76, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s8/p67;   http://www.
jstor.org/stable/30045195.

Goncalves:2005:DMP

[GR05] Paulo Gonçalvès and Rudolf Riedi. Diverging moments and parameter estimation. Journal of the American Statistical Association, 100(472):1382–1393, December 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590679.

Goldstein:2006:BLC

[GR06a] Michael Goldstein and Jonathan Rougier. Bayes linear calibrated prediction for complex systems. Journal of the American Statistical Association, 101(475):1132–1143, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590790.

Grogger:2006:TRP

[GR06b] Jeffrey Grogger and Greg Ridgeway. Testing for racial profiling in traffic stops from behind a veil of darkness. Journal of the American Statistical Association, 101(475):878– 887, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590768.

Gneiting:2007:SPS

[GR07] Tilmann Gneiting and Adrian E. Raftery. Strictly proper scoring rules, prediction, and estimation. Journal of the American Statistical Association, 102(477):359–378, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gray:2002:BRB

[Gra02] Robert J. Gray. Book review: [untitled]. Journal of the American Statistical Association, 97(457):353–354, March 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085789.

Greenland:2000:CAH

[Gre00] Sander Greenland. Causal analysis in the health sciences. Journal of the American Statistical Association, 95(449):286– 289, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669548.

Greenland:2003:IPD

[Gre03] Sander Greenland. The impact of prior distributions for uncontrolled confounding and response bias: a case study of the relation of wire codes and magnetic fields to childhood leukemia. Journal of the American Statistical Association, 98(461):47–54, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s6/p47;   http://www.jstor.org/stable/30045193.

Green:2007:BRB

[Gre07a] Paul F. Green. Book review: Statistics for Fission Track Analysis. Journal of the American Statistical Association, 102(477): 380–381, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Greenwood:2007:BRB

[Gre07b] Mark Greenwood. Book review: Nonparametric Functional Data Analysis: Theory and Practice. Journal of the American Statistical Association, 102(479):1077, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Graves:2003:HMP

[GRF03] Todd Graves, C. Shane Reese, and Mark Fitzgerald. Hierarchical models for permutations: Analysis of auto racing results. Journal of the American Statistical Association, 98(462):282–291, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s3/p282;   http://
www.jstor.org/stable/30045236.

Gel:2004:CPM

[GRG^{+}04] Yulia Gel, Adrian E. Raftery, Tilmann Gneiting, Claudia Tebaldi, Doug Nychka, William Briggs, Mark S. Roulston, and Veronica J. Berrocal. Calibrated probabilistic mesoscale weather field forecasting: The geostatistical output perturbation method [with comments, rejoinder]. Journal of the American Statistical Association, 99(467):575–583, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/s1/
p575;   http://www.jstor.org/stable/27590432.

Gel:2004:RCP

[GRGB04] Yulia Gel, Adrian E. Raftery, Tilmann Gneiting, and Veronica J. Berrocal. Rejoinder: Calibrated probabilistic mesoscale weather field forecasting: The geostatistical output perturbation method. Journal of the American Statistical Association, 99(467):588–590, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n467/s5/p588.

Grimes:2002:BRB

[Gri02a] Carrie Grimes. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1214–1215, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085861.

Grimes:2002:ICA

[Gri02b] Carrie Grimes. Independent component analysis: Principles and practice. Journal of the American Statistical Association, 97(460):1214–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s42/p1214.

Grogger:2001:SDS

[Gro01] Jeff Grogger. [should the DEA’s STRIDE data be used for economic analyses of markets for illegal drugs?]: Comment. Journal of the American Statistical Association, 96(456):1268– 1269, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085892.

Guo:2004:RPM

[GRT04] Wensheng Guo, Sarah J. Ratcliffe, and Thomas R. Ten Have. A random pattern-mixture model for longitudinal data with dropouts. Journal of the American Statistical Association, 99(468):929–937, December 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/s3/
p929;   http://www.jstor.org/stable/27590476.

Gottardo:2006:QCR

[GRYB06] Raphael Gottardo, Adrian E. Raftery, Ka Yee Yeung, and Roger E. Bumgarner. Quality control and robust estimation for cDNA microarrays with replicates. Journal of the American Statistical Association, 101(473):30–40, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30047438.

Gasser:2001:SNM

[GS01] Theo Gasser and Burkhardt Seifert. [semiparametric nonlinear mixed-effects models and their applications]: Comment. Journal of the American Statistical Association, 96(456):1282– 1283, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085895.

Griffin:2006:OBD

[GS06] J. E. Griffin and M. F. J. Steel. Order-based dependent Dirichlet processes. Journal of the American Statistical Association, 101(473):179–194, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047448.

Grigg:2007:SRA

[GS07] Olivia Grigg and David Spiegelhalter. A simple risk-adjusted exponentially weighted moving average. Journal of the American Statistical Association, 102(477):140–152, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2004:NTS

[GSC04] Yongtao Guan, Michael Sherman, and James A. Calvin. A nonparametric test for spatial isotropy using subsampling. Journal of the American Statistical Association, 99(467):810–821, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s29/p810;   http:
//www.jstor.org/stable/27590451.

Greevy:2004:RII

[GSCR04] Robert Greevy, Jeffrey H. Silber, Avital Cnaan, and Paul R. Rosenbaum. Randomization inference with imperfect compliance in the ACE-inhibitor after anthracycline randomized trial. Journal of the American Statistical Association, 99(465):7–15, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s2/p7;   http://
www.jstor.org/stable/27590348.

Ghosh:2007:PUC

[GT07] Kaushik Ghosh and Ram C. Tiwari. Prediction of U.S. cancer mortality counts using semiparametric Bayesian techniques. Journal of the American Statistical Association, 102(477):7– 15, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:2009:ASA

[GT09] Pulak Ghosh and Wanzhu Tu. Assessing sexual attitudes and behaviors of young women: a joint model with nonlinear time effects, time varying covariates, and dropouts. Journal of the American Statistical Association, 104(486):474–485, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2006:CLA

[Gua06] Yongtao Guan. A composite likelihood approach in fitting spatial point process models. Journal of the American Statistical Association, 101(476):1502–1512, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639768.

Guan:2008:CNI

[Gua08] Yongtao Guan. On consistent nonparametric intensity estimation for inhomogeneous spatial point processes. Journal of the American Statistical Association, 103(483):1238–1247, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2009:NVE

[Gua09] Yongtao Guan. On nonparametric variance estimation for second-order statistics of inhomogeneous spatial point processes with a known parametric intensity form. Journal of the American Statistical Association, 104(488):1482–1491, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guerra:2008:BRB

[Gue08] Rudy Guerra. Book review: Likelihood, Bayesian and MCMC Methods in Quantitative Genetics. Journal of the American Statistical Association, 103(481):432, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gunter:2002:BRB

[Gun02] Berton H. Gunter. Book review: [untitled]. Journal of the American Statistical Association, 97(457):353, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085788.

Gustafson:2000:BRM

[Gus00] Paul Gustafson. Bayesian regression modeling with interactions and smooth effects. Journal of the American Statistical Association, 95(451):795–806, September 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669463.

Gustafson:2009:WLP

[Gus09] Paul Gustafson. What are the limits of posterior distributions arising from nonidentified models, and why should we care? Journal of the American Statistical Association, 104 (488):1682–1695, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guttorp:2000:ES

[Gut00] Peter Guttorp. Environmental statistics. Journal of the American Statistical Association, 95(449):289–292, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669549.

Goncalves:2005:BSE

[GW05] Sílvia Gonçalves and Halbert White. Bootstrap standard error estimates for linear regression. Journal of the American Statistical Association, 100(471):970–979, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590627.

Genovese:2006:ECF

[GW06] Christopher R. Genovese and Larry Wasserman. Exceedance control of the false discovery proportion. Journal of the American Statistical Association, 101(476):1408–1417, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27639761.

Guan:2008:SOA

[GWB08] Yongtao Guan, Rasmus Waagepetersen, and Colin M. Beale. Second-order analysis of inhomogeneous spatial point processes with proportional intensity functions. Journal of the American Statistical Association, 103(482):769–777, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gao:2001:SSAa

[GWKK01a] Fangyu Gao, Grace Wahba, Ronald Klein, and Barbara Klein. Smoothing spline ANOVA for multivariate Bernoulli observations, with application to ophthalmology data. Journal of the American Statistical Association, 96(453):127–147, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670350;   reqidx=/cw/asa/01621459/v96n453/
s11/p127.

Gao:2001:SSAb

[GWKK01b] Fangyu Gao, Grace Wahba, Ronald Klein, and Barbara Klein. Smoothing spline ANOVA for multivariate Bernoulli observations, with application to ophthalmology data: Rejoinder. Journal of the American Statistical Association, 96(453): 158–160, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670355.

Gotway:2002:CIS

[GY02] Carol A. Gotway and Linda J. Young. Combining incompatible spatial data. Journal of the American Statistical Association, 97(458):632–648, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&

reqidx=/cw/asa/01621459/v97n458/s23/p632;   http://
www.jstor.org/stable/3085677.

Habing:2003:BRB

[Hab03] Brian Habing. Book review: Introduction to Nonparametric Item Response Theory, by Klaas Sijtsma; Ivo W. Molenaar. Journal of the American Statistical Association, 98(463):774– 775, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045319.

Hade:2005:BRB

[Had05a] Erinn M. Hade. Book review: Quantitative Methods in Population Health: Extensions of Ordinary Regression, by Mari Palta. Journal of the American Statistical Association, 100(470):704–705, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s32/p704;   http://www.
jstor.org/stable/27590591.

Hadlock:2005:BRB

[Had05b] Charles R. Hadlock. Book review: Causality: Models, Reasoning, and Inference, by Judea Pearl. Journal of the American Statistical Association, 100(471):1095–1096, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590643.

Hall:2002:SBV

[Hal02] Peter Hall. [sieve bootstrap with variable-length Markov chains for stationary categorical time series]: Comment. Journal of the American Statistical Association, 97(458):456–457, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085658.

Hall:2003:ALM

[Hal03a] Daniel B. Hall. Advanced linear modeling (second edition). Journal of the American Statistical Association, 98(461):258–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s39/p258.

Hall:2003:BRB

[Hal03b] Daniel B. Hall. Book review: Advanced Linear Modeling, by Ronald Christensen. Journal of the American Statistical Association, 98(461):258, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045226.

Hall:2006:BRBa

[Hal06] Daniel B. Hall. Book review: Statistical Modelling in Glim4, by Murray Aitkin; Brian Francis; John Hinde. Journal of the American Statistical Association, 101(476):1733–1734, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27639807.

Hall:2007:BRB

[Hal07] Daniel B. Hall. Book review: Models for Intensive Longitudinal Data. Journal of the American Statistical Association, 102 (480):1473, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hall:2008:BRB

[Hal08] Daniel B. Hall. Book review: Measurement Error in Nonlinear Models: A Modern Perspective. Journal of the American Statistical Association, 103(481):427, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hansen:2003:RUE

[Han03] Bruce E. Hansen. Recounts from undervotes: Evidence from the 2000 Presidential Election. Journal of the American Statistical Association, 98(462):292–298, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s4/p292;   http://www.jstor.org/stable/30045237.

Hansen:2004:FMO

[Han04] Ben B. Hansen. Full matching in an observational study of coaching for the SAT. Journal of the American Statistical Association, 99(467):609–618, September 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=

1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/s8/
p609;   http://www.jstor.org/stable/27590435.

Hanson:2006:IMF

[Han06] Timothy E. Hanson. Inference for mixtures of finite Polya tree models. Journal of the American Statistical Association, 101 (476):1548–1565, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639772.

Harrington:2000:RCT

[Har00a] David P. Harrington. The randomized clinical trial. Journal of the American Statistical Association, 95(449):312–315, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669555.

Harter:2000:BRB

[Har00b] H. Leon Harter. Book review: [untitled]. Journal of the American Statistical Association, 95(449):331–332, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669564.

Harville:2003:SSC

[Har03] David A. Harville. The selection or seeding of college basketball or football teams for postseason competition. Journal of the American Statistical Association, 98(461):17–27, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s3/p17;   http://www.jstor.org/stable/30045190.

Harvill:2007:BRB

[Har07] Jane L. Harvill. Book review: Analysis of Integrated and Co-Integrated Time Series With R. Journal of the American Statistical Association, 102(477):389–390, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harvill:2008:BRB

[Har08] Jane L. Harvill. Book review: Essentials of Statistical Inference. Journal of the American Statistical Association, 103 (481):425–426, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Havlak:2007:BRB

[Hav07] Paul Havlak. Book review: Computational Genome Analysis: An Introduction. Journal of the American Statistical Association, 102(477):388, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haynor:2000:BTC

[Hay00] David R. Haynor. A Bayesian time-course model for functional magnetic resonance imaging data: Comment. Journal of the American Statistical Association, 95(451):704, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669446.

Hazeghi:2000:BRB

[Haz00] Kasra Hazeghi. Book review: [untitled]. Journal of the American Statistical Association, 95(449):343, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669577.

Hansen:2009:AEC

[HB09] Ben B. Hansen and Jake Bowers. Attributing effects to a cluster-randomized get-out-the-vote campaign. Journal of the American Statistical Association, 104(487):873–885, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hennerfeind:2006:GSM

[HBF06] Andrea Hennerfeind, Andreas Brezger, and Ludwig Fahrmeir. Geoadditive survival models. Journal of the American Statistical Association, 101(475):1065–1075, September 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590784.

Henderson:2009:BEC

[HBK^{+}09] Daniel A. Henderson, Richard J. Boys, Kim J. Krishnan, Conor Lawless, and Darren J. Wilkinson. Bayesian emulation and calibration of a stochastic computer model of mitochondrial DNA deletions in substantia nigra neurons. Journal of the American Statistical Association, 104(485):76–87, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hernan:2001:MSM

[HBR01] Miguel A. Hernán, Babette Brumback, and James M. Robins. Marginal structural models to estimate the joint causal effect of nonrandomized treatments. Journal of the American Statistical Association, 96(454):440–448, June 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670282;   reqidx=/cw/asa/01621459/v96n454/s8/p440.

Han:2001:MCM

[HC01] Cong Han and Bradley P. Carlin. Markov chain Monte Carlo methods for computing Bayes factors: a comparative review. Journal of the American Statistical Association, 96(455):1122–1132, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670258;   reqidx=/cw/
asa/01621459/v96n455/s30/p1122.

Hong:2002:TSC

[HC02] Han Hong and Victor Chernozhukov. Three-step censored quantile regression and extramarital affairs. Journal of the American Statistical Association, 97(459):872–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://giorgio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n459/s18/p872.

Hjort:2003:FMA

[HC03a] Nils Lid Hjort and Gerda Claeskens. Frequentist model average estimators. Journal of the American Statistical Association, 98(464):879–899, December 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v98n464/s11/p879;   http:
//www.jstor.org/stable/30045339.

Hjort:2003:RPB

[HC03b] Nils Lid Hjort and Gerda Claeskens. Rejoinder: Performance of Bayesian model averaging. Journal of the American Statistical Association, 98(464):938–945, December 2003.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s19/p938.

Hjort:2006:FIC

[HC06] Nils Lid Hjort and Gerda Claeskens. Focused information criteria and model averaging for the Cox hazard regression model. Journal of the American Statistical Association, 101 (476):1449–1464, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639764.

Huang:2007:OGM

[HC07] Hsin-Cheng Huang and Chun-Shu Chen. Optimal geostatistical model selection. Journal of the American Statistical Association, 102(479):1009–1024, September 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hitchcock:2006:IED

[HCB06] David B. Hitchcock, George Casella, and James G. Booth. Improved estimation of dissimilarities by presmoothing functional data. Journal of the American Statistical Association, 101(473):211–222, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047451.

Houseman:2006:NNB

[HCS06] E. Andrés Houseman, Brent A. Coull, and James P. Shine. A nonstationary negative binomial time series with timedependent covariates: Enterococcus counts in Boston Harbor. Journal of the American Statistical Association, 101 (476):1365–1376, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639757.

Hjort:2006:PPP

[HDS06] Nils Lid Hjort, Fredrik A. Dahl, and Gunnhildur Högnadóttir Steinbakk. Post-processing posterior predictive p values. Journal of the American Statistical Association, 101(475):1157– 1174, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590792.

Hans:2007:SSS

[HDW07] Chris Hans, Adrian Dobra, and Mike West. Shotgun stochastic search for “large p” regression. Journal of the American Statistical Association, 102(478):507–516, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Helmers:2000:BRB

[Hel00] R. Helmers. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1017, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669501.

Heller:2007:SRR

[Hel07] Glenn Heller. Smoothed rank regression with censored data. Journal of the American Statistical Association, 102(478):552– 559, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Henderson:2002:BRB

[Hen02] Shane Henderson. Book review: [untitled]. Journal of the American Statistical Association, 97(457):355–356, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085792.

Hesterberg:2000:BRB

[Hes00] Tim Hesterberg. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1376–1377, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669800.

Heyen:2008:BRB

[Hey08] Keith Heyen. Book review: The Basel II Risk Parameters: Estimation, Validation, and Stress Testing. Journal of the American Statistical Association, 103(483):1318–1319, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2005:REG

[HFZ05] Xuming He, Wing K. Fung, and Zhongyi Zhu. Robust estimation in generalized partial linear models for clustered data. Journal of the American Statistical Association, 100 (472):1176–1184, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590663.

Hunter:2008:GFS

[HGH08] David R. Hunter, Steven M. Goodreau, and Mark S. Handcock. Goodness of fit of social network models. Journal of the American Statistical Association, 103(481):248–258, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Higdon:2008:CMC

[HGWR08] Dave Higdon, James Gattiker, Brian Williams, and Maria Rightley. Computer model calibration using high-dimensional output. Journal of the American Statistical Association, 103 (482):570–583, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2002:MCM

[HH02] Xuming He and Feifang Hu. Markov chain marginal bootstrap. Journal of the American Statistical Association, 97 (459):783–795, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s11/p783;   http://
www.jstor.org/stable/3085721.

Hudgens:2006:CVE

[HH06] Michael G. Hudgens and M. Elizabeth Halloran. Causal vaccine effects on binary postinfection outcomes. Journal of the American Statistical Association, 101(473):51–64, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047440.

He:2007:CIE

[HH07] Xuming He and Feifang Hu. Comment: Implementation of estimating function-based inference procedures with Markov chain Monte Carlo samplers. Journal of the American Statistical Association, 102(479):889–890, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hudgens:2008:TCI

[HH08] Michael G. Hudgens and M. Elizabeth Halloran. Toward causal inference with interference. Journal of the American Statistical Association, 103(482):832–842, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2009:SVD

[HHCK09] Jianhua Hu, Xuming He, Gilbert J. Cote, and Ralf Krahe. Singular value decomposition-based alternative splicing detection. Journal of the American Statistical Association, 104(487):944– 953, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heard:2006:QSG

[HHS06] Nicholas A. Heard, Christopher C. Holmes, and David A. Stephens. A quantitative study of gene regulation involved in the immune response of anopheline mosquitoes: An application of Bayesian hierarchical clustering of curves. Journal of the American Statistical Association, 101(473):18–29, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047437.

Herring:2001:LBM

[HI01] Amy H. Herring and Joseph G. Ibrahim. Likelihood-based methods for missing covariates in the Cox proportional hazards model. Journal of the American Statistical Association, 96(453):292–302, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670367;   reqidx=/cw/
asa/01621459/v96n453/s23/p292.

Ho:2006:RIN

[HI06] Daniel E. Ho and Kosuke Imai. Randomization inference with natural experiments: An analysis of ballot effects in the 2003 California Recall Election. Journal of the American Statistical Association, 101(475):888–900, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590769.

Hubbard:2009:JMS

[HID09] Rebecca A. Hubbard, Lurdes Y. T. Inoue, and Paula Diehr. Joint modeling of self-rated health and changes in physical functioning. Journal of the American Statistical Association, 104(487):912–928, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Higdon:2000:BTC

[Hig00] Dave Higdon. A Bayesian time-course model for functional magnetic resonance imaging data: Comment. Journal of the American Statistical Association, 95(451):705–706, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669447.

Higdon:2009:CST

[Hig09] D. Higdon. Comment: a spatio-temporal model for mean, anomaly, and trend fields of North Atlantic sea surface temperature. Journal of the American Statistical Association, 104 (485):18–20, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LS09b].

Hill:2008:CCN

[Hil08] Jennifer Hill. Comment: Can nonrandomized experiments yield accurate answers? A randomized experiment comparing random and nonrandom assignments. Journal of the American Statistical Association, 103(484):1346–1350, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hilbe:2009:NMF

[Hil09] Joseph M. Hilbe. Note: Modeling future record performances in athletics. Journal of the American Statistical Association, 104(487):1293–1294, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hinz:2007:BRB

[Hin07] Juri Hinz. Book review: Weather Derivative Valuation: The Meteorological, Statistical, Financial and Mathematical Foundations. Journal of the American Statistical Association, 102 (477):380, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hannig:2006:FGC

[HIP06] Jan Hannig, Hari Iyer, and Paul Patterson. Fiducial generalized confidence intervals. Journal of the American Statistical Association, 101(473):254–269, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047454.

Hanson:2002:MRE

[HJ02] Timothy Hanson and Wesley O. Johnson. Modeling regression error with a mixture of Polya trees. Journal of the American Statistical Association, 97(460):1020–1033, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s8/p1020;   http://www.jstor.org/stable/3085827.

Hu:2009:LLM

[HJJ09] Jianhua Hu, Adarsh Joshi, and Valen E. Johnson. Log-linear models for gene association. Journal of the American Statistical Association, 104(486):597–607, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Horowitz:2002:RLA

[HK02] Joel L. Horowitz and K. Katsouyanni. Response to letter of Arthur Fries. Journal of the American Statistical Association, 97(458):660, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www. jstor.org/stable/3085696. See [Hor01a, Hor01b, Fri02b].

Heagerty:2000:WSE

[HL00] Patrick J. Heagerty and Thomas Lumley. Window subsampling of estimating functions with application to regression models. Journal of the American Statistical Association, 95 (449):197–211, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/Heagerty.htm;   http:/
/www.jstor.org/stable/2669538.

Hu:2004:SFT

[HL04] Chengcheng Hu and D. Y. Lin. Semiparametric failure time regression with replicates of mismeasured covariates. Journal of the American Statistical Association, 99(465):105–118, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s11/p105;   http://www.jstor.org/stable/27590357.

Horowitz:2005:NEA

[HL05] Joel L. Horowitz and Sokbae Lee. Nonparametric estimation of an additive quantile regression model. Journal of the American Statistical Association, 100(472):1238–1249, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590668.

Hafner:2006:CQA

[HL06a] Christian M. Hafner and Oliver B. Linton. Comment: Quantile autoregression. Journal of the American Statistical Association, 101(475):998–1001, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hassan:2006:MMP

[HL06b] Mohamed Yusuf Hassan and Keh-Shin Lii. Modeling marked point processes via bivariate mixture transition distribution models. Journal of the American Statistical Association, 101 (475):1241–1252, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590798.

Hallin:2007:DNF

[HL07] Marc Hallin and Roman Liška. Determining the number of factors in the general dynamic factor model. Journal of the American Statistical Association, 102(478):603–617, June 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hurvich:2005:ELM

[HLS05] Clifford Hurvich, Gabriel Lang, and Philippe Soulier. Estimation of long memory in the presence of a smooth nonparametric trend. Journal of the American Statistical Association, 100(471):853–871, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590618.

Horowirtz:2000:RNA

[HM00a] Joel L. Horowirtz and Charles F. Manski. Rejoinder: Nonparametric analysis of randomized experiments with missing covariate and outcome data. Journal of the American Statistical Association, 95(449):87–??, March 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Horowitz:2000:NARa

[HM00b] Joel L. Horowitz and Charles F. Manski. Nonparametric analysis of randomized experiments with missing covariate and outcome data. Journal of the American Statistical Association, 95(449):77–84, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/Horowitz1.htm;   http:
//www.jstor.org/stable/2669526.

Horowitz:2000:NARb

[HM00c] Joel L. Horowitz and Charles F. Manski. Nonparametric analysis of randomized experiments with missing covariate and outcome data: Rejoinder. Journal of the American Statistical Association, 95(449):87–88, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669528.

Hall:2003:OPN

[HM03a] Peter Hall and Hans-Georg Müller. Order-preserving nonparametric regression, with applications to conditional distribution and quantile function estimation. Journal of the American Statistical Association, 98(463):598–608, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s13/p598;   http://www.jstor.org/stable/30045288.

Holmes:2003:GNM

[HM03b] C. C. Holmes and B. K. Mallick. Generalized nonlinear modeling with multivariate free-knot regression splines. Journal of the American Statistical Association, 98(462):352–368, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n462/s8/p352;   http://www.jstor.org/stable/30045245.

Hannig:2006:ADT

[HM06] J. Hannig and J. S. Marron. Advanced distribution theory for SiZer. Journal of the American Statistical Association, 101 (474):484–499, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590711.

Hettmansperger:2000:RNM

[HMS00] Thomas P. Hettmansperger, Joseph W. McKean, and Simon J. Sheather. Robust nonparametric methods. Journal of the American Statistical Association, 95(452):1308– 1312, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669777.

Hawkins:2002:IRAa

[HO02a] Douglas M. Hawkins and David J. Olive. Inconsistency of resampling algorithms for high-breakdown regression estimators and a new algorithm. Journal of the American Statistical Association, 97(457):136–148, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085765;   reqidx=/cw/
asa/01621459/v97n457/s13/p136.

Hawkins:2002:IRAb

[HO02b] Douglas M. Hawkins and David J. Olive. [inconsistency of resampling algorithms for high-breakdown regression estimators and a new algorithm]: Rejoinder. Journal of the American Statistical Association, 97(457):156–159, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085770.

Hobert:2000:HMC

[Hob00] James P. Hobert. Hierarchical models: a current computational perspective. Journal of the American Statistical Association, 95(452):1312–1316, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669778.

Hodges:2008:CEI

[Hod08] James S. Hodges. Comment: Estimating incumbency advantage and its variation, as an example of a before–after study.

Journal of the American Statistical Association, 103(482):448– 449, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoff:2005:BME

[Hof05] Peter D. Hoff. Bilinear mixed-effects models for dyadic data. Journal of the American Statistical Association, 100 (469):286–295, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s32/p286;   http://www.
jstor.org/stable/27590538.

Hoff:2007:MAD

[Hof07] Peter D. Hoff. Model averaging and dimension selection for the singular value decomposition. Journal of the American Statistical Association, 102(478):674–685, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holmes:2000:BRBb

[Hol00a] Susan Holmes. Book review: [untitled]. Journal of the American Statistical Association, 95(450):688, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669431.

Holmes:2000:BRBa

[Hol00b] Susan P. Holmes. Book review: [untitled]. Journal of the American Statistical Association, 95(449):328–329, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669559.

Holland:2005:CFD

[Hol05] Burt Holland. Comment: False discovery rate-adjusted multiple confidence intervals for selected parameters. Journal of the American Statistical Association, 100(469):89–90, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s13/
p89.

Hooten:2007:BRB

[Hoo07] Mevin B. Hooten. Book review: Statistical Analysis of Environmental Space–time Processes. Journal of the American Statistical Association, 102(480):1477, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Horowitz:2001:SDSa

[Hor01a] Joel L. Horowitz. Should the DEA’s STRIDE data be used for economic analyses of markets for illegal drugs? Journal of the American Statistical Association, 96(456):1254–1262, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085888;   reqidx=/cw/asa/01621459/v96n456/

s12/p1254. See rejoinder [Hor01b], response [Fri02b] and reply [HK02].

Horowitz:2001:SDSb

[Hor01b] Joel L. Horowitz. [should the DEA’s STRIDE data be used for economic analyses of markets for illegal drugs?]: Rejoinder. Journal of the American Statistical Association, 96(456):1269– 1271, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/3085893. See [Hor01a], response [Fri02b], and reply [HK02].

Horne:2002:BRB

[Hor02] A. Dale Horne. Book review: [untitled]. Journal of the American Statistical Association, 97(457):356–357, March 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085794.

Hossjer:2003:DID

[Hös03] Ola Hössjer. Determining inheritance distributions via stochastic penetrances. Journal of the American Statistical Association, 98(464):1035–1051, December 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v98n464/s28/p1035.

Huang:2004:DCN

[HOS04] Hsiao-Yun Huang, Hernando Ombao, and David S. Stoffer. Discrimination and classification of nonstationary time series using the SLEX model. Journal of the American Statistical Association, 99(467):763–774, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s24/p763;   http://www.jstor.org/stable/27590446.

He:2004:CLS

[HP04] Xuming He and Stephen Portnoy. Comment: Locationscale depth. Journal of the American Statistical Association, 99(468):973–976, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s8/p973.

Halloran:2003:EVE

[HPC03] M. Elizabeth Halloran, Marie Pierre Préziosi, and Haitao Chu. Estimating vaccine efficacy from secondary attack rates. Journal of the American Statistical Association, 98(461):38–46, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s5/p38;   http://www.jstor.org/stable/30045192.

Huang:2008:ELB

[HQF08] Chiung-Yu Huang, Jing Qin, and Dean A. Follmann. Empirical likelihood-based estimation of the treatment effect in a pretest–posttest study. Journal of the American Statistical Association, 103(483):1270–1280, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hughes:2000:ART

[HR00] James P. Hughes and Barbra A. Richardson. Analysis of a randomized trial to prevent vertical transmission of HIV-1. Journal of the American Statistical Association, 95(452):1032– 1043, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669740.

Hu:2003:OVP

[HR03] Feifang Hu and William F. Rosenberger. Optimality, variability, power: Evaluating response-adaptive randomization procedures for treatment comparisons. Journal of the American Statistical Association, 98(463):671–678, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-

bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s19/p671;   http://www.jstor.org/stable/30045294.

Hong:2006:EKR

[HR06] Guanglei Hong and Stephen W. Raudenbush. Evaluating kindergarten retention policy: a case study of causal inference for multilevel observational data. Journal of the American Statistical Association, 101(475):901–910, September 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590770.

Houseman:2004:CRA

[HRC04] E. Andrés Houseman, Louise M. Ryan, and Brent A. Coull. Cholesky residuals for assessing normal errors in a linear model with correlated outcomes. Journal of the American Statistical Association, 99(466):383–394, April 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/s8/
p383;   http://www.jstor.org/stable/27590394.

Hoff:2002:LSA

[HRH02] Peter D. Hoff, Adrian E. Raftery, and Mark S. Handcock. Latent space approaches to social network analysis. Journal of the American Statistical Association, 97 (460):1090–1098, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s14/p1090;   http:/
/www.jstor.org/stable/3085833.

Hall:2004:CVE

[HRL04] Peter Hall, Jeff Racine, and Qi Li. Cross-validation and the estimation of conditional probability densities. Journal of the American Statistical Association, 99(468): 1015–1026, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s14/p1015;   http://www.
jstor.org/stable/27590481.

Heller:2009:SSD

[HRS09] Ruth Heller, Paul R. Rosenbaum, and Dylan S. Small. Split samples and design sensitivity in observational studies. Journal of the American Statistical Association, 104(487):1090– 1101, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hubert:2002:IRA

[HRV02] Mia Hubert, Peter J. Rousseeuw, and Stefan Van Aelst. [inconsistency of resampling algorithms for high-breakdown regression estimators and a new algorithm]: Comment. Journal of the American Statistical Association, 97(457):151–153, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085768.

Hubert:2004:CLS

[HRV04] Mia Hubert, Peter J. Rousseeuw, and Karlien Vanden Branden. Comment: Location-scale depth. Journal of the American Statistical Association, 99(468):966–969, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/s6/
p966.

Hall:2002:ECB

[HS02a] Peter Hall and Léopold Simar. Estimating a changepoint, boundary, or frontier in the presence of observation error. Journal of the American Statistical Association, 97(458):523–534, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s14/p523;   http://
www.jstor.org/stable/3085668.

Horowitz:2002:ARO

[HS02b] Joel L. Horowitz and Vladimir G. Spokoiny. An adaptive, rate-optimal test of linearity for median regression models. Journal of the American Statistical Association, 97 (459):822–835, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s14/p822;   http://
www.jstor.org/stable/3085724.

Hallin:2007:OTN

[HS07a] Marc Hallin and Abdessamad Saidi. Optimal tests of noncorrelation between multivariate time series. Journal of the American Statistical Association, 102(479):938–951, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hinich:2007:ENE

[HS07b] Melvin J. Hinich and Apostolos Serletis. Episodic nonlinear event detection in the Canadian exchange rate. Journal of the American Statistical Association, 102(477):68–74, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2009:ATW

[HSB09] Jianhua Z. Huang, Haipeng Shen, and Andreas Buja. The analysis of two-way functional data using two-way regularized singular value decompositions. Journal of the American Statistical Association, 104(488):1609–1620, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Henderson:2002:MSV

[HSG02] Robin Henderson, Silvia Shimakura, and David Gorst. Modeling spatial variation in leukemia survival data. Journal of the American Statistical Association, 97(460):965–972, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n460/s4/p965;   http://www.jstor.org/stable/3085818.

Hardle:2001:STA

[HSS01] Wolfgang Härdle, Stefan Sperlich, and Vladimir Spokoiny. Structural tests in additive regression. Journal of the American Statistical Association, 96(456):1333–1347, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085903;   reqidx=/cw/asa/01621459/v96n456/
s16/p1333.

Hssjer:2003:DID

[Hss03] Ola Hssjer. Determining inheritance distributions via stochastic penetrances. Journal of the American Statistical Association, 98(464):1035–1051, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045349.

Hall:2006:BRBb

[HSSM06] Daniel B. Hall, Debajyoti Sinha, Michael Sherman, and Tapabrata Maiti. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 101(476):1735– 1736, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639810.

Hedayat:2006:OEC

[HSY06] A. S. Hedayat, John Stufken, and Min Yang. Optimal and efficient crossover designs when subject effects are random. Journal of the American Statistical Association, 101(475):1031– 1038, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590780.

Hogan:2004:BFA

[HT04] Joseph W. Hogan and Rusty Tchernis. Bayesian factor analysis for spatially correlated data, with application to summarizing area-level material deprivation from census data. Journal of the American Statistical Association, 99(466):314–324, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/s2/
p314;   http://www.jstor.org/stable/27590388.

Han:2006:RFC

[HT06] Dong Han and Fugee Tsung. A reference-free Cuscore chart for dynamic mean change detection and a unified framework for charting performance comparison. Journal of the American Statistical Association, 101(473):368–386, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047464.

Hoeting:2003:TEP

[HTO03] Jennifer A. Hoeting, Richard L. Tweedie, and Christine S. Olver. Transform estimation of parameters for stage-frequency data. Journal of the American Statistical Association, 98 (463):503–514, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s1/p503;   http://
www.jstor.org/stable/30045276.

Hall:2009:MBC

[HTX09] Peter Hall, D. M. Titterington, and Jing-Hao Xue. Medianbased classifiers for high-dimensional data. Journal of the American Statistical Association, 104(488):1597–1608, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2009:WNS

[HTZ^{+}09] Lan Huang, Ram C. Tiwari, Zhaohui Zou, Martin Kulldorff, and Eric J. Feuer. Weighted normal spatial scan statistic for heterogeneous population data. Journal of the American Statistical Association, 104(487):886–898, September 2009. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2004:BRB

[Hu04] X. Joan Hu. Book review: Applied Longitudinal Data Analysis for Epidemiology: A Practical Guide, by Jos W. R. Twisk. Journal of the American Statistical Association, 99(468):1207–1208, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s39/p1207;   http://www.
jstor.org/stable/27590506.

Huang:2000:TSM

[Hua00] Yijian Huang. Two-sample multistate accelerated sojourn times model. Journal of the American Statistical Association, 95(450):619–627, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/2000/June/Huang.htm;
http://www.jstor.org/stable/2669404.

Huang:2002:CRC

[Hua02] Yijian Huang. Calibration regression of censored lifetime medical cost. Journal of the American Statistical Association, 97(457):318–327, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085785;   reqidx=/cw/
asa/01621459/v97n457/s28/p318.

Hurd:2000:BRB

[Hur00] Harry Hurd. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1372–1373, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669793.

Huang:2000:CRA

[HW00] Yijian Huang and C. Y. Wang. Cox regression with accurate covariates unascertainable: a nonparametric-correction approach. Journal of the American Statistical Association, 95(452):1209–1219, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669761.

Hoover:2001:SNR

[HW01a] Donald R. Hoover and Colin O. Wu. Semiparametric and nonparametric regression analysis of longitudinal data: Comment. Journal of the American Statistical Association, 96 (453):116–119, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670347.

Huang:2001:CFM

[HW01b] Yijian Huang and C. Y. Wang. Consistent functional methods for logistic regression with errors in covariates. Journal of the American Statistical Association, 96(456):1469–1482, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085914;   reqidx=/cw/asa/01621459/v96n456/
s27/p1469.

Huang:2003:FRE

[HW03] Yijian Huang and Mei-Cheng Wang. Frequency of recurrent events at failure time: Modeling and inference. Journal of the American Statistical Association, 98(463): 663–670, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s18/p663;   http://
www.jstor.org/stable/30045293.

Huang:2004:JME

[HW04] Chiung-Yu Huang and Mei-Cheng Wang. Joint modeling and estimation for recurrent event processes and failure time data. Journal of the American Statistical Association, 99(468):1153–1165, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s26/p1153;   http://www.
jstor.org/stable/27590493.

Hallin:2006:CQA

[HW06] Marc Hallin and Bas J. M. Werker. Comment: Quantile autoregression. Journal of the American Statistical Association, 101(475):996–998, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2005:TWS

[HWZ05] Jian Huang, Deli Wang, and Cun-Hui Zhang. A two-way semilinear model for normalization and analysis of cDNA microarray data. Journal of the American Statistical Association, 100(471):814–829, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590615.

Hu:2006:EEI

[HWZ06] Jianhua Hu, Fred A. Wright, and Fei Zou. Estimation of expression indexes for oligonucleotide arrays using the singular value decomposition. Journal of the American Statistical Association, 101(473):41–50, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047439.

Hansen:2001:MSP

[HY01a] Mark H. Hansen and Bin Yu. Model selection and the principle of minimum description length. Journal of the American Statistical Association, 96(454):746–774, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670311;   reqidx=/cw/asa/01621459/v96n454/
s32/p746.

Higdon:2001:EHS

[HY01b] Dave Higdon and Steve Yamamoto. Estimation of the head sensitivity function in scanning magnetoresistance microscopy. Journal of the American Statistical Association, 96 (455):785–793, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670219;   reqidx=/cw/
asa/01621459/v96n455/s1/p785.

Hedayat:2004:UOS

[HY04] A. S. Hedayat and Min Yang. Universal optimality for selected crossover designs. Journal of the American Statistical Association, 99(466):461–466, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s15/p461;   http:
//www.jstor.org/stable/27590401.

Harville:2007:CPF

[HY07] David A. Harville and Emmanuel Yashchin. Cost (or price) forecasting in the face of technological advance. Journal of the American Statistical Association, 102(477):28–43, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2005:EPI

[HYFW05] Shuyuan He, Grace L. Yang, Kai Tai Fang, and John F. Widmann. Estimation of Poisson intensity in the presence of dead time. Journal of the American Statistical Association, 100(470):669–679, June 2005. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s27/p669;   http://www.
jstor.org/stable/27590586.

He:2003:LFT

[HZ03] Xuming He and Li-Xing Zhu. A lack-of-fit test for quantile regression. Journal of the American Statistical Association, 98(464):1013–1022, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s26/p1013;   http://www.
jstor.org/stable/30045347.

Huang:2005:CSH

[HZ05] Jian Huang and Cun-Hui Zhang. Comment: Semilinear highdimensional model for normalization of microarray data. Journal of the American Statistical Association, 100(471):800–804, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hyrien:2008:MMD

[HZ08] Ollivier Hyrien and Martin S. Zand. A mixture model with dependent observations for the analysis of CSFE-labeling experiments. Journal of the American Statistical Association, 103 (481):222–239, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2001:LVM

[HZAG01] Wenzheng Huang, Scott L. Zeger, James C. Anthony, and Elizabeth Garrett. Latent variable model for joint analysis of multiple repeated measures and bivariate event times. Journal of the American Statistical Association, 96 (455):906–914, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670230;   reqidx=/cw/
asa/01621459/v96n455/s12/p906.

Huang:2007:BAF

[HZW07] Hanwen Huang, Fei Zou, and Fred A. Wright. Bayesian analysis of frequency of allelic loss data. Journal of the American Statistical Association, 102(480):1245–1253, December 2007.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hung:2008:BTS

[HZZ^{+}08] Ying Hung, Veronika Zarnitsyna, Yan Zhang, Cheng Zhu, and C. F. Jeff Wu. Binary time series modeling with application to adhesion frequency experiments. Journal of the American Statistical Association, 103(483):1248–1259, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ishwaran:2004:RRF

[IBPL04] Hemant Ishwaran, Eugene H. Blackstone, Claire E. Pothier, and Michael S. Lauer. Relative risk forests for exercise heart rate recovery as a predictor of mortality. Journal of the American Statistical Association, 99(467):591–600, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/s6/
p591;   http://www.jstor.org/stable/27590433.

Iversen:2005:PCG

[IC05] Edwin S. Iversen, Jr. and Sining Chen. Population-calibrated gene characterization: Estimating age at onset distributions associated with cancer genes. Journal of the American Statistical Association, 100(470):399–409, June 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s5/p399;   http:
//www.jstor.org/stable/27590564.

Ibrahim:2002:BMG

[ICG02] Joseph G. Ibrahim, Ming-Hui Chen, and Robert J. Gray. Bayesian models for gene expression with DNA microarray data. Journal of the American Statistical Association, 97(457):88–99, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085761;   reqidx=/cw/
asa/01621459/v97n457/s9/p88.

Ibrahim:2005:MDM

[ICLH05] Joseph G. Ibrahim, Ming-Hui Chen, Stuart R. Lipsitz, and Amy H. Herring. Missing-data methods for generalized linear models: a comparative review. Journal of the American Statistical Association, 100(469):332–346, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s36/
p332;   http://www.jstor.org/stable/27590542.

Ibrahim:2003:OPP

[ICS03] Joseph G. Ibrahim, Ming-Hui Chen, and Debajyoti Sinha. On optimality properties of the power prior. Journal of the American Statistical Association, 98(461):204–213, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s20/p204;   http://www.jstor.org/stable/30045207.

Inoue:2008:MDP

[IEMM08] Lurdes Y. T. Inoue, Ruth Etzioni, Christopher Morrell, and Peter Müller. Modeling disease progression with longitudinal markers. Journal of the American Statistical Association, 103 (481):259–270, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ihaka:2006:BRB

[Iha06] Ross Ihaka. Book review: Developing Statistical Software in Fortran 95, by David R. Lemmon; Joseph L. Schafer. Journal of the American Statistical Association, 101(474):857, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590759.

Ishwaran:2001:GSM

[IJ01] Hemant Ishwaran and Lancelot F. James. Gibbs sampling methods for stick breaking priors. Journal of the American Statistical Association, 96(453):161–173, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670356;   reqidx=/cw/asa/01621459/v96n453/
s12/p161.

Ishwaran:2004:CMM

[IJ04] Hemant Ishwaran and Lancelot F. James. Computational methods for multiplicative intensity models using weighted gamma processes: Proportional hazards, marked point processes, and panel count data. Journal of the American Statistical Association, 99(465):175–190, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s18/p175;   http:
//www.jstor.org/stable/27590364.

Ishwaran:2001:BMS

[IJS01] Hemant Ishwaran, Lancelot F. James, and Jiayang Sun. Bayesian model selection in finite mixtures by marginal density decompositions. Journal of the American Statistical Association, 96(456):1316–1332, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085902;   reqidx=/cw/
asa/01621459/v96n456/s15/p1316.

Inoue:2008:HUB

[IK08] Atsushi Inoue and Lutz Kilian. How useful is bagging in forecasting economic time series? A case study of U.S. consumer price inflation. Journal of the American Statistical Association, 103(482):511–522, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Imhof:2003:BRB

[Imh03a] Lorens A. Imhof. Book review: Matrix Algebra and Its Applications to Statistics and Econometrics, by C. Radhakrishna Rao; M. Bhaskara Rao. Journal of the American Statistical Association, 98(462):494–495, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045265.

Imhof:2003:MAA

[Imh03b] Lorens A. Imhof. Matrix algebra and its applications to statistics and econometrics. Journal of the American Statistical Association, 98(462):494–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s29/p494.

Inglis:2000:BRB

[Ing00] James Inglis. Book review: [untitled]. Journal of the American Statistical Association, 95(449):340–341, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669573.

Inoue:2002:DFT

[IP02] Lurdes Y. T. Inoue and Giovanni Parmigiani. Designing follow-up times. Journal of the American Statistical Association, 97(459):847–858, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s16/p847;   http://
www.jstor.org/stable/3085726.

Iversen:2000:GSS

[IPBS00] Edwin S. Iversen, Jr., Giovanni Parmigiani, Donald A. Berry, and Joellen M. Schildkraut. Genetic susceptibility and survival: Application to breast cancer. Journal of the American Statistical Association, 95(449):28–42, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
Iverson.htm;   http://www.jstor.org/stable/2669520.

Iversen:2008:MME

[IPC08] Edwin S. Iversen, Jr., Giovanni Parmigiani, and Sining Chen. Multiple model evaluation absent the gold standard through model combination. Journal of the American Statistical Association, 103(483):897–909, September 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ishwaran:2003:DDE

[IR03a] Hemant Ishwaran and J. Sunil Rao. Detecting differentially expressed genes in microarrays using Bayesian model selection. Journal of the American Statistical Association, 98(462):438–455, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s16/p438;   http://
www.jstor.org/stable/30045253.

Ishwaran:2003:DII

[IR03b] Hemant Ishwaran and J. Sunil Rao. Discussion: Inference and interpretability considerations in frequentist model averaging and selection. Journal of the American Statistical Association, 98(464):922–925, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s15/p922.

Ishwaran:2005:SSG

[IR05] Hemant Ishwaran and J. Sunil Rao. Spike and slab gene selection for multigroup microarray data. Journal of the American Statistical Association, 100(471):764–780, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590613.

Insua:2009:ARA

[IRB09] David Rios Insua, Jesus Rios, and David Banks. Adversarial risk analysis. Journal of the American Statistical Association, 104(486):841–854, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Irizarry:2001:IPP

[Iri01a] Rafael A. Irizarry. Information and posterior probability criteria for model selection in local likelihood estimation. Journal of the American Statistical Association, 96(453):303–315, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670368;   reqidx=/cw/asa/01621459/v96n453/
s24/p303.

Irizarry:2001:LHE

[Iri01b] Rafael A. Irizarry. Local harmonic estimation in musical sound signals. Journal of the American Statistical Association, 96(454):357–367, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670275;   reqidx=/cw/
asa/01621459/v96n454/s1/p357.

Imai:2007:EDF

[IS07] Kosuke Imai and Samir Soneji. On the estimation of disabilityfree life expectancy. Journal of the American Statistical Association, 102(480):1199–1211, December 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Im:2007:SES

[ISZ07] Hae Kyung Im, Michael L. Stein, and Zhengyuan Zhu. Semiparametric estimation of spectral density with irregular observations. Journal of the American Statistical Association, 102 (478):726–735, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ishwaran:2002:IID

[IT02] Hemant Ishwaran and Glen Takahara. Independent and identically distributed Monte Carlo algorithms for semiparametric linear mixed models. Journal of the American Statistical Association, 97(460):1154–1166, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s19/p1154;   http://www.jstor.org/stable/3085838.

Imai:2004:CIG

[IV04] Kosuke Imai and David A. Van Dyk. Causal inference with general treatment regimes: Generalizing the propensity score. Journal of the American Statistical Association, 99(467):854–866, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s33/p854;   http:
//www.jstor.org/stable/27590455.

Iyer:2004:MCI

[IWM04] Hari K. Iyer, C. M. Jack Wang, and Thomas Mathew. Models and confidence intervals for true values in interlaboratory trials. Journal of the American Statistical Association, 99(468):1060–1071, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s18/p1060;   http://www.
jstor.org/stable/27590485.

Iaci:2008:IMA

[IYSK08] Ross Iaci, Xiangrong Yin, T. N. Sriram, and Christian Peter Klingenberg. An informational measure of association and dimension reduction for multiple sets and groups with applications in morphometric analysis. Journal of the American Statistical Association, 103(483):1166–1176, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ibrahim:2007:CIE

[IZ07] Joseph G. Ibrahim and Hongtu Zhu. Comment: Implementation of estimating function-based inference procedures with Markov chain Monte Carlo samplers. Journal of the American Statistical Association, 102(479):893, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ibrahim:2008:MSC

[IZT08] Joseph G. Ibrahim, Hongtu Zhu, and Niansheng Tang. Model selection criteria for missing-data problems using the EM algorithm. Journal of the American Statistical Association, 103 (484):1648–1658, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jammalamadaka:2000:BRBa

[Jam00a] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 95(449):330– 331, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669562.

Jammalamadaka:2000:BRBb

[Jam00b] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 95(449):331, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669563.

Jammalamadaka:2000:BRBc

[Jam00c] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 95(451): 1008, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669489.

Jammalamadaka:2000:BRBd

[Jam00d] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 95(451): 1016, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669499.

Jammalamadaka:2000:BRBe

[Jam00e] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 95(452): 1376, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669799.

Jammalamadaka:2001:BRBa

[Jam01a] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 96(453):354– 355, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670394.

Jammalamadaka:2001:BRBb

[Jam01b] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 96(454): 779–780, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670319.

Jammalamadaka:2001:BRBc

[Jam01c] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 96(454):783, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2670330.

Jammalamadaka:2001:BRBd

[Jam01d] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 96(456): 1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085939.

Jammalamadaka:2002:BRBa

[Jam02a] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 97(459): 927, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085744.

Jammalamadaka:2002:BRBb

[Jam02b] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 97(459): 927, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085745.

Jammalamadaka:2002:BRBc

[Jam02c] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 97(459):927– 928, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085746.

Jammalamadaka:2002:BRBd

[Jam02d] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 97(460): 1206, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085849.

Jammalamadaka:2002:CTA

[Jam02e] Sreenivasa Rao Jammalamadaka. A contingency table approach to nonparametric testing. Journal of the American Statistical Association, 97(459):927–??, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s34/p927.

Jammalamadaka:2002:IAS

[Jam02f] Sreenivasa Rao Jammalamadaka. An invariant approach to statistical analysis of shapes. Journal of the American Statistical Association, 97(459):927–??, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s36/p927.

Jammalamadaka:2002:MAA

[Jam02g] Sreenivasa Rao Jammalamadaka. Matrix algebra for applied economics, and matrix algebra exercises and solutions. Journal of the American Statistical Association, 97 (460):1206–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s30/p1206.

Jammalamadaka:2002:SS

[Jam02h] Sreenivasa Rao Jammalamadaka. Scan statistics. Journal of the American Statistical Association, 97(459):927–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://giorgio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n459/s35/p927.

Jammalamadaka:2003:BRB

[Jam03a] Sreenivasa Rao Jammalamadaka. Book review: Permutation Methods: A Distance Function Approach, by Paul W. Mielke, Jr.; Kenneth J. Berry. Journal of the American Statistical Association, 98(461):258–259, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045227.

Jammalamadaka:2003:PMD

[Jam03b] Sreenivasa Rao Jammalamadaka. Permutation methods: a distance function approach. Journal of the American Statistical Association, 98(461):258–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s40/p258.

Jansen:2001:EBA

[JB01] Maarten Jansen and Adhemar Bultheel. Empirical Bayes approach to improve wavelet thresholding for image noise reduction. Journal of the American Statistical Association, 96(454):629–639, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670302;   reqidx=/cw/
asa/01621459/v96n454/s23/p629.

Jin:2007:ENP

[JC07] Jiashun Jin and T. Tony Cai. Estimating the null and the proportion of nonnull effects in large-scale multiple comparisons. Journal of the American Statistical Association, 102(478):495– 506, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:2002:BAR

[JDvS02] Valen E. Johnson, Robert O. Deaner, and Carel P. van Schaik. Bayesian analysis of rank data with application to primate intelligence experiments. Journal of the American Statistical Association, 97(457):8–17, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085754;   reqidx=/cw/
asa/01621459/v97n457/s2/p8.

Jensen:2002:BRB

[Jen02a] Uwe Jensen. Book review: [untitled]. Journal of the American Statistical Association, 97(459):925, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085741.

Jensen:2002:PRA

[Jen02b] Uwe Jensen. Probabilistic risk analysis: Foundations and methods. Journal of the American Statistical Association, 97(459):925–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s31/p925.

Jensen:2006:BRB

[Jen06] Willis A. Jensen. Book review: Introduction to Nonparametric Regression, by Kunio Takezawa. Journal of the American Statistical Association, 101(476):1721–1722, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639789.

Jersky:2006:BRB

[Jer06] Brian Jersky. Book review: Exploratory Data Analysis with MATLAB, by Wendy L. Martinez; Angel R. Martinez. Journal of the American Statistical Association, 101(473):399–400, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047483.

Jeske:2003:BRB

[Jes03a] Daniel R. Jeske. Book review: Statistical Inference: An Integrated Approach, by Helio S. Migon; Dani Gamerman. Journal of the American Statistical Association, 98(461):259, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045228.

Jeske:2003:SII

[Jes03b] Daniel R. Jeske. Statistical inference: An integrated approach. Journal of the American Statistical Association, 98(461):259–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s41/p259.

Jin:2007:USF

[JFY07] Chunfang Jin, Jason P. Fine, and Brian S. Yandell. A unified semiparametric framework for quantitative trait loci analyses, with application to spike phenotypes. Journal of the American Statistical Association, 102(477):56–67, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jones:2006:FWO

[JHCN06] Galin L. Jones, Murali Haran, Brian S. Caffo, and Ronald Neath. Fixed-width output analysis for Markov chain Monte Carlo. Journal of the American Statistical Association, 101 (476):1537–1547, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639771. See remarks [BL07].

Johnson:2007:SME

[JHISR07] Brent A. Johnson, Amy H. Herring, Joseph G. Ibrahim, and Anna Maria Siega-Riz. Structured measurement error in nutritional epidemiology. Journal of the American Statistical Association, 102(479):856–866, September 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2005:CMP

[Jia05] Jiming Jiang. Comment: Maximization by parts in likelihood inference. Journal of the American Statistical Association, 100 (472):1158–1159, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jupp:2003:IDS

[JKKW03] Peter E. Jupp, Peter T. Kim, Ja-Yong Koo, and Paul Wiegert. The intrinsic distribution and selection bias of long-period cometary orbits. Journal of the American Statistical Association, 98(463):515–521, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s2/p515;   http://www.jstor.org/stable/30045277.

Jun:2008:SAQ

[JKN08] Mikyoung Jun, Reto Knutti, and Douglas W. Nychka. Spatial analysis to quantify numerical model bias and dependence. Journal of the American Statistical Association, 103(483):934– 947, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2006:EFP

[JL06] Jiming Jiang and P. Lahiri. Estimation of finite population domain means: a model-assisted empirical best prediction approach. Journal of the American Statistical Association, 101(473):301–311, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047458.

Jensen:2008:BCT

[JL08] Shane T. Jensen and Jun S. Liu. Bayesian clustering of transcription factor binding motifs. Journal of the American Statistical Association, 103(481):188–200, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnstone:2009:CSP

[JL09a] Iain M. Johnstone and Arthur Yu Lu. On consistency and sparsity for principal components analysis in high dimensions. Journal of the American Statistical Association, 104(486):682– 693, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnstone:2009:RC

[JL09b] Iain M. Johnstone and Arthur Yu Lu. Rejoinder: Comments. Journal of the American Statistical Association, 104(486):701– 703, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:2008:PEF

[JLZ08] Brent A. Johnson, D. Y. Lin, and Donglin Zeng. Penalized estimating functions and variable selection in semiparametric regression models. Journal of the American Statistical Association, 103(482):672–680, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Judge:2004:SBC

[JM04] George G. Judge and Ron C. Mittelhammer. A semiparametric basis for combining estimation problems under quadratic loss. Journal of the American Statistical Association, 99(466):479–487, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s17/p479;   http:
//www.jstor.org/stable/27590403.

Jensen:2006:ELM

[JM06] Søren Tolver Jensen and Jesper Madsen. Estimation in linear models based on observations with unknown and possibly unequal scaling. Journal of the American Statistical Association, 101(475):1059–1064, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590783.

Johns:2003:ISR

[JNKD03] Craig J. Johns, Douglas Nychka, Timothy G. F. Kittel, and Chris Daly. Infilling sparse records of spatial fields. Journal of the American Statistical Association, 98 (464):796–806, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s3/p796;   http://www.
jstor.org/stable/30045331.

Joe:2003:BRB

[Joe03] Harry Joe. Book review: Multivariate Dispersion, Central Regions and Depth: The Lift Zonoid Approach, by Karl Mosler. Journal of the American Statistical Association, 98(464):1082, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s34/p1082;   http://www.
jstor.org/stable/30045355.

Johnson:2003:DII

[Joh03] Wesley O. Johnson. Discussion: Inference and interpretability considerations in frequentist model averaging and selection. Journal of the American Statistical Association, 98 (464):919–921, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s14/p919.

Jones:2000:BRB

[Jon00] M. C. Jones. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1370–1371, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669791.

Jones:2006:BRB

[Jon06] Galin L. Jones. Book review: Computational Statistics, by Geof H. Givens; Jennifer A. Hoeting. Journal of the American Statistical Association, 101(474):856–857, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590758.

Jones:2005:FSD

[JP05] M. C. Jones and Arthur Pewsey. A family of symmetric distributions on the circle. Journal of the American Statistical Association, 100(472):1422–1428, December 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590682.

Javaras:2007:ULV

[JR07] Kristin N. Javaras and Brian D. Ripley. An “unfolding” latent variable model for Likert attitude data. Journal of the American Statistical Association, 102(478):454–463, June 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jin:2008:PSC

[JR08] Hui Jin and Donald B. Rubin. Principal stratification for causal inference with extended partial compliance. Journal of the American Statistical Association, 103(481):101–111, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jarrow:2004:EIR

[JRY04] Robert Jarrow, David Ruppert, and Yan Yu. Estimating the interest rate term structure of corporate debt with a semiparametric penalized spline model. Journal of the American Statistical Association, 99(465):57–66, March 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/s7/
p57;   http://www.jstor.org/stable/27590353.

James:2003:CSS

[JS03] Gareth M. James and Catherine A. Sugar. Clustering for sparsely sampled functional data. Journal of the American Statistical Association, 98(462):397–408, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s12/p397;   http://www.jstor.org/stable/30045249.

James:2005:FAM

[JS05] Gareth M. James and Bernard W. Silverman. Functional adaptive model estimation. Journal of the American Statistical Association, 100(470):565–576, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/

rpsv/cw/asa/01621459/v100n470/s19/p565;   http://www.
jstor.org/stable/27590578.

Jureckova:2006:BRB

[Jur06] Jana Jurečková. Book review: Quantile Regression, by Roger Koenker. Journal of the American Statistical Association, 101(476):1723, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639791.

Javaras:2003:MII

[JvD03] Kristin N. Javaras and David A. van Dyk. Multiple imputation for incomplete data with semicontinuous variables. Journal of the American Statistical Association, 98(463):703– 715, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045298.

Jing:2009:JEL

[JYZ09] Bing-Yi Jing, Junqing Yuan, and Wang Zhou. Jackknife empirical likelihood. Journal of the American Statistical Association, 104(487):1224–1232, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kachman:2004:BRB

[Kac04] Stephen D. Kachman. Book review: Components of Variance, by D. R. Cox; P. J. Solomon. Journal of the American Statistical Association, 99(467):899, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s44/p899;   http://www.jstor.org/stable/27590460.

Kadane:2007:CSL

[Kad07] Joseph B. Kadane. Comment: Subjective likelihood for the assessment of trends in the ocean’s mixed-layer depth. Journal of the American Statistical Association, 102(479):781–782, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kafadar:2003:BRB

[Kaf03] Karen Kafadar. Book review: Testing for Normality, by Henry C. Thode, Jr. Journal of the American Statistical Association, 98(463):765, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045304.

Kalbfleisch:2000:POM

[Kal00] John D. Kalbfleisch. On the probability of observing misleading statistical evidence: Comment. Journal of the American Statistical Association, 95(451):770–771, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669458.

Karr:2002:BRBa

[Kar02a] Alan F. Karr. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1216, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085865.

Karr:2002:BRBb

[Kar02b] Alan F. Karr. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1216, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085867.

Karr:2002:BRBc

[Kar02c] Alan F. Karr. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1216–1217, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085868.

Karr:2003:BRB

[Kar03a] Alan F. Karr. Book review: Maximum Penalized Likelihood Estimation, Vol. I: Density Estimation, by P. P. B. Eggermont; V. N. LaRiccia. Journal of the American Statistical Association, 98(462):493, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045262.

Karr:2003:MPL

[Kar03b] Alan F. Karr. Maximum penalized likelihood estimation, vol. I: Density estimation. Journal of the American Statistical Association, 98(462):493–??, June 2003. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s26/p493.

Karr:2006:BRBa

[Kar06a] Alan F. Karr. Book review: Evolution of Biological Systems in Random Media: Limit Theorems and Stability, by Anatoly Swishchuk; Jianhong Wu. Journal of the American Statistical Association, 101(473):397, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047478.

Karr:2006:BRBb

[Kar06b] Alan F. Karr. Book review: Exploratory Data Mining and Data Cleaning, by Tamraparni Dasu; Theodore Johnson. Journal of the American Statistical Association, 101(473):399, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047482.

Karr:2006:BRBc

[Kar06c] Alan F. Karr. Book review: Random Graphs for Statistical Pattern Recognition, by David J. Marchette. Journal of the American Statistical Association, 101(475):1317, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590819.

Karr:2007:BRBa

[Kar07a] Alan F. Karr. Book review: Partial Identification of Probability Distributions. Journal of the American Statistical Association, 102(477):383, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Karr:2007:BRBb

[Kar07b] Alan F. Karr. Book review: Statistical Matching: A Frequentist Theory, Practical Applications and Alternative Bayesian Approaches. Journal of the American Statistical Association, 102(477):386, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Katz:2008:CEI

[Kat08] Jonathan N. Katz. Comment: Estimating incumbency advantage and its variation, as an example of a before–after study.

Journal of the American Statistical Association, 103(482):446– 448, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

King:2004:CSC

[KB04] R. King and S. P. Brooks. A classical study of catcheffort models for Hector’s dolphins. Journal of the American Statistical Association, 99(466):325–333, April 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/s3/
p325;   http://www.jstor.org/stable/27590389.

Kauermann:2001:NES

[KC01] Göran Kauermann and Raymond J. Carroll. A note on the efficiency of sandwich covariance matrix estimation. Journal of the American Statistical Association, 96(456): 1387–1396, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085907;   reqidx=/cw/
asa/01621459/v96n456/s20/p1387.

Khalili:2007:VSF

[KC07] Abbas Khalili and Jiahua Chen. Variable selection in finite mixture of regression models. Journal of the American Statistical Association, 102(479):1025–1038, September 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2008:SCA

[KCO08] Yongdai Kim, Hosik Choi, and Hee-Seok Oh. Smoothly clipped absolute deviation on high dimensions. Journal of the American Statistical Association, 103(484):1665–1673, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kou:2002:SGM

[KE02] S. C. Kou and Bradley Efron. Smoothers and the C_{p}, generalized maximum likelihood, and extended exponential criteria: a geometric approach. Journal of the American Statistical Association, 97(459):766–782, September 2002. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s10/p766;   http://
www.jstor.org/stable/3085720.

Kendall:2003:BRB

[Ken03a] Wilfrid S. Kendall. Book review: Statistics of the Galaxy Distribution, by Vicent J. Martínez; Enn Saar. Journal of the American Statistical Association, 98(461):248, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045211.

Kendall:2003:SGD

[Ken03b] Wilfrid S. Kendall. Statistics of the galaxy distribution. Journal of the American Statistical Association, 98(461):248–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s24/p248.

Kerr:2006:BRB

[Ker06] Kathleen F. Kerr. Book review: Design and Analysis of Experiments, Vol. 2: Advanced Experimental Design, by Klaus Hinkelmann; Oscar Kempthorne. Journal of the American Statistical Association, 101(474):853–854, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590752.

Kanaan:2002:EWV

[KF02] Mona N. Kanaan and C. Paddy Farrington. Estimation of waning vaccine efficacy. Journal of the American Statistical Association, 97(458):389–397, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s3/p389;   http://www.jstor.org/stable/3085652.

Kosorok:2007:CIE

[KF07] Michael R. Kosorok and Jason P. Fine. Comment: Implementation of estimating function-based inference procedures with Markov chain Monte Carlo samplers. Journal of the American Statistical Association, 102(479):896–897, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:2001:RML

[KG01a] Roger Koenker and Olga Geling. Reappraising medfly longevity: a quantile regression survival analysis. Journal of the American Statistical Association, 96(454):458–468, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670284;   reqidx=/cw/asa/01621459/v96n454/
s10/p458.

Kottas:2001:BSM

[KG01b] Athanasios Kottas and Alan E. Gelfand. Bayesian semiparametric median regression modeling. Journal of the American Statistical Association, 96(456):1458–1468, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085913;   reqidx=/cw/asa/01621459/v96n456/
s26/p1458.

Karr:2006:BRBd

[KHACM06] Alan F. Karr, Daniel B. Hall, Christine M. Anderson-Cook, and Tapabrata Maiti. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 101(473): 405–406, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047494.

Kiessler:2005:BRB

[Kie05] Peter C. Kiessler. Book review: Comparison Methods for Stochastic Models and Risks, by Alfred Müller; Dietrich Stoyan. Journal of the American Statistical Association, 100(470):704, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s31/p704;   http://www.
jstor.org/stable/27590590.

Kiessler:2006:BRB

[Kie06] Peter C. Kiessler. Book review: Statistical Inference for Ergodic Diffusion Processes, by Yury A. Kutoyants. Journal of the American Statistical Association, 101(474):846, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590744.

Kiessler:2007:BRBa

[Kie07a] Peter C. Kiessler. Book review: Measure Theory and Probability Theory. Journal of the American Statistical Association, 102(479):1078, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kiessler:2007:BRBb

[Kie07b] Peter C. Kiessler. Book review: Probabilistic Applications of Tauberian Theorems. Journal of the American Statistical Association, 102(480):1476, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kiessler:2008:BRBe

[Kie08a] Peter C. Kiessler. Book review: An Introduction to Random Sets; Theory of Random Sets. Journal of the American Statistical Association, 103(482):884, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kiessler:2008:BRBd

[Kie08b] Peter C. Kiessler. Book review: Extreme Value Theory: An Introduction. Journal of the American Statistical Association, 103(482):882–883, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kiessler:2008:BRBc

[Kie08c] Peter C. Kiessler. Book review: Stochastic Switching Systems: Analysis and Design. Journal of the American Statistical Association, 103(481):430, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2000:OSC

[Kim00] Jong Tae Kim. An order selection criterion for testing goodness of fit. Journal of the American Statistical Association, 95(451):829–835, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669466.

Kim:2001:BRB

[Kim01a] Hyune-Ju Kim. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1138–1139, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670264.

Kimber:2001:BRB

[Kim01b] Alan C. Kimber. Book review: [untitled]. Journal of the American Statistical Association, 96(454):780–781, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670322.

Kimmel:2006:BRB

[Kim06] Marek Kimmel. Book review: Statistical Methods in Molecular Evolution, by Rasmus Nielsen. Journal of the American Statistical Association, 101(475):1312–1313, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590811.

Kalbfleisch:2007:CIE

[KJ07] John D. Kalbfleisch and Wenyu Jiang. Comment: Implementation of estimating function-based inference procedures with Markov chain Monte Carlo samplers. Journal of the American Statistical Association, 102(479):890–893, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kjell:2000:LBS

[Kje00] Gunnar Kjell. The level-based stratified sampling plan. Journal of the American Statistical Association, 95(452):1185– 1191, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669759.

Kolaczyk:2005:MMS

[KJG05] Eric D. Kolaczyk, Junchang Ju, and Sucharita Gopal. Multiscale, multigranular statistical image segmentation. Journal of the American Statistical Association, 100(472):1358– 1369, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590677.

Krivobokova:2007:NPS

[KK07] Tatyana Krivobokova and Göran Kauermann. A note on penalized spline smoothing with correlated errors. Journal of the American Statistical Association, 102(480):1328–1337, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kulich:2000:AHR

[KL00] Michal Kulich and D. Y. Lin. Additive hazards regression with covariate measurement error. Journal of the American Statistical Association, 95(449):238–248, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
Kulich.htm;   http://www.jstor.org/stable/2669541.

Kim:2001:CTU

[KL01] Hyunjoong Kim and Wei Yin Loh. Classification trees with unbiased multiway splits. Journal of the American Statistical Association, 96(454):589–604, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670299;   reqidx=/cw/
asa/01621459/v96n454/s20/p589.

Kadane:2004:MCM

[KL04a] Joseph B. Kadane and Nicole A. Lazar. Methods and criteria for model selection. Journal of the American Statistical Association, 99(465):279–290, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s26/p279;   http:
//www.jstor.org/stable/27590372.

Kulich:2004:IER

[KL04b] Michal Kulich and D. Y. Lin. Improving the efficiency of relative-risk estimation in case-cohort studies. Journal of the American Statistical Association, 99(467):832–844, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/

vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s31/p832;   http://www.jstor.org/stable/27590453.

Kraft:2005:BRB

[KL05] Peter Kraft and Christoph Lange. Book review: Statistical Methods in Genetic Epidemiology, by Duncan C. Thomas. Journal of the American Statistical Association, 100(472): 1464, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590697.

Kabaila:2006:LSM

[KL06a] Paul Kabaila and Hannes Leeb. On the large-sample minimal coverage probability of confidence intervals after model selection. Journal of the American Statistical Association, 101 (474):619–629, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590722.

Kuan:2006:RTC

[KL06b] Chung-Ming Kuan and Wei-Ming Lee. Robust M tests without consistent estimation of the asymptotic covariance matrix. Journal of the American Statistical Association, 101 (475):1264–1275, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590800.

Komarek:2008:BAF

[KL08] Arnošt Komárek and Emmanuel Lesaffre. Bayesian accelerated failure time model with multivariate doubly intervalcensored data and flexible distributional assumptions. Journal of the American Statistical Association, 103(482):523–533, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Klein:2006:BRB

[Kle06] John P. Klein. Book review: Flowgraph Models for Multistate Time-to-Event Data, by Aparna V. Huzurbazar. Journal of the American Statistical Association, 101(475):1316– 1317, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590817.

Klaassen:2001:PTI

[KM01] Franc J. G. M. Klaassen and Jan R. Magnus. Are points in tennis independent and identically distributed? evidence from a dynamic binary panel data model. Journal of the American Statistical Association, 96(454):500–509, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670288;   reqidx=/cw/asa/01621459/v96n454/
s14/p500.

Kaplan:2002:APH

[KM02] Norman Kaplan and Richard Morris. [adjusting for population heterogeneity and misspecified haplotype frequencies when testing nonparametric null hypotheses in statistical genetics]: Comment. Journal of the American Statistical Association, 97(459):753–754, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085716.

Kosorok:2005:CSH

[KM05] Michael R. Kosorok and Shuangge Ma. Comment: Semilinear high-dimensional model for normalization of microarray data. Journal of the American Statistical Association, 100(471):805– 807, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keller-McNulty:2007:DP

[KM07] Sallie Keller-McNulty. From data to policy. Journal of the American Statistical Association, 102(478):395–399, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2005:ANS

[KMH05] Hyoung-Moon Kim, Bani K. Mallick, and C. C. Holmes. Analyzing nonstationary spatial data using piecewise Gaussian processes. Journal of the American Statistical Association, 100(470):653–668, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s26/p653;   http://www.
jstor.org/stable/27590585.

Kloke:2009:RBE

[KMR09] John D. Kloke, Joseph W. McKean, and M. Mushfiqur Rashid. Rank-based estimation and associated inferences for linear models with cluster correlated errors. Journal of the American Statistical Association, 104(485):384–390, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2006:RVE

[KNF06] Jae Kwang Kim, Alfredo Navarro, and Wayne A. Fuller. Replication variance estimation for two-phase stratified sampling. Journal of the American Statistical Association, 101(473): 312–320, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047459.

Knight:2006:CQA

[Kni06] Keith Knight. Comment: Quantile autoregression. Journal of the American Statistical Association, 101(475):994–996, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koopman:2007:PSR

[KOC07] Siem Jan Koopman, Marius Ooms, and M. Angeles Carnero. Periodic seasonal Reg-ARFIMA–GARCH models for daily electricity spot prices. Journal of the American Statistical Association, 102(477):16–27, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koehler:2002:AMD

[Koe02a] Kenneth Koehler. Analysis of messy data, vol III: Analysis of covariance. Journal of the American Statistical Association, 97(460):1206–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s31/p1206.

Koehler:2002:BRB

[Koe02b] Kenneth Koehler. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1206–1207, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085850.

Kokoszka:2001:BRB

[Kok01] Piotr Kokoszka. Book review: [untitled]. Journal of the American Statistical Association, 96(453):353–354, March 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670387.

Kolassa:2002:BRB

[Kol02a] John Kolassa. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1204–1205, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085847.

Kolassa:2002:CSS

[Kol02b] John Kolassa. Conditional specification of statistical models. Journal of the American Statistical Association, 97 (460):1204–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s27/p1204;   http:
//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s28/p1204.

Kolaczyk:2003:C

[Kol03a] Eric D. Kolaczyk. Comment. Journal of the American Statistical Association, 98(463):585–??, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s10/p585.

Kolaczyk:2003:WBN

[Kol03b] Eric D. Kolaczyk. Wavelet-based nonparametric modeling of hierarchical functions in colon carcinogenesis [with comment]. Journal of the American Statistical Association, 98(463):585– 588, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045285.

Kong:2006:DRL

[Kon06] Yong Kong. Distribution of runs and longest runs: a new generating function approach. Journal of the American Statistical Association, 101(475):1253–1263, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590799.

Koop:2002:CPB

[Koo02] Gary Koop. Comparing the performance of baseball players: a multiple-output approach. Journal of the American Statistical Association, 97(459):710–720, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s6/p710;   http://www.jstor.org/stable/3085707.

Korn:2004:BRB

[Kor04] Ralf Korn. Book review: An Elementary Introduction to Mathematical Finance: Options and Other Topics, by Sheldon M. Ross. Journal of the American Statistical Association, 99(466):563, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s32/p563;   http:
//www.jstor.org/stable/27590418.

Korn:2005:BRB

[Kor05] Ralf Korn. Book review: An Introduction to Financial Option Valuation: Mathematics, Stochastics and Computation, by Desmond J. Higham. Journal of the American Statistical Association, 100(471):1098, September 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590648.

Kou:2008:SNN

[Kou08] S. C. Kou. Stochastic networks in nanoscale biophysics. Journal of the American Statistical Association, 103(483):961–975, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kvam:2005:ELS

[KP05] Paul H. Kvam and Edsel A. Peña. Estimating load-sharing properties in a dynamic reliability system. Journal of the American Statistical Association, 100(469):262–272, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s30/
p262;   http://www.jstor.org/stable/27590536.

Kreider:2007:DE

[KP07] Brent Kreider and John V. Pepper. Disability and employment. Journal of the American Statistical Association, 102 (478):432–441, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kowalski:2002:NTG

[KPD02] Jeanne Kowalski, Marcello Pagano, and Victor DeGruttola. A nonparametric test of gene region heterogeneity associated with phenotype. Journal of the American Statistical Association, 97(458):398–408, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s4/p398;   http://www.jstor.org/stable/3085653.

Krieger:2003:SSL

[KPY03] Abba M. Krieger, Moshe Pollak, and Benjamin Yakir. Surveillance of a simple linear regression. Journal of the American Statistical Association, 98(462):456–469, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s17/p456;   http://www.jstor.org/stable/30045254.

Kulkarni:2002:NEB

[KR02] H. V. Kulkarni and R. N. Rattihalli. Nonparametric estimation of a bivariate mean residual life function. Journal of the American Statistical Association, 97(459):907–917, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s21/p907;   http://www.jstor.org/stable/3085731.

Kneip:2008:CRF

[KR08] Alois Kneip and James O. Ramsay. Combining registration and fitting for functional models. Journal of the American Statistical Association, 103(483):1155–1165, September 2008.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krauth:2004:BRB

[Kra04] Joachim Krauth. Book review: Configural Frequency Analysis: Methods, Models, and Applications, by Alexander von Eye. Journal of the American Statistical Association, 99(465):292–293, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s28/p292;   http:
//www.jstor.org/stable/27590374.

Krause:2008:BRB

[Kra08] Andreas Krause. Book review: Pharmacometrics: The Science of Quantitative Pharmacology. Journal of the American Statistical Association, 103(481):431–432, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kristal:2005:BRB

[Kri05] Alan R. Kristal. Book review: Statistical Estimation of Epidemiological Risk, by Kung-Jong Lui. Journal of the American Statistical Association, 100(472):1464, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590696.

Kaciroti:2006:BAC

[KRS^{+}06] Niko A. Kaciroti, Trivellore E. Raghunathan, M. Anthony Schork, Noreen M. Clark, and Molly Gong. A Bayesian approach for clustered longitudinal ordinal outcome with nonignorable missing data: Evaluation of an asthma education program. Journal of the American Statistical Association, 101 (474):435–446, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590707.

Kovac:2000:ESW

[KS00] Arne Kovac and Bernard W. Silverman. Extending the scope of wavelet regression methods by coefficient-dependent thresholding. Journal of the American Statistical Association, 95(449):172–183, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/Kovac.htm;   http:
//www.jstor.org/stable/2669536.

Kunert:2002:OCD

[KS02] J. Kunert and J. Stufken. Optimal crossover designs in a model with self and mixed carryover effects. Journal of the American Statistical Association, 97(459):898–906, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s20/p898;   http://www.jstor.org/stable/3085730.

Kaziska:2007:GBH

[KS07a] David Kaziska and Anuj Srivastava. Gait-based human recognition by classification of cyclostationary processes on nonlinear shape manifolds. Journal of the American Statistical Association, 102(480):1114–1124, December 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kaziska:2007:RGB

[KS07b] David M. Kaziska and Anuj Srivastava. Rejoinder: Gaitbased human recognition by classification of cyclostationary processes on nonlinear shape manifolds. Journal of the American Statistical Association, 102(480):1127, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kunert:2008:OCD

[KS08] J. Kunert and J. Stufken. Optimal crossover designs for two treatments in the presence of mixed and self-carryover effects. Journal of the American Statistical Association, 103 (484):1641–1647, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koyama:2005:CDT

[KSG05] Tatsuki Koyama, Allan R. Sampson, and Leon J. Gleser. A calculus for design of two-stage adaptive procedures. Journal of the American Statistical Association, 100(469):197–203, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s23/
p197;   http://www.jstor.org/stable/27590529.

Kaufman:2008:CTL

[KSN08] Cari G. Kaufman, Mark J. Schervish, and Douglas W. Nychka. Covariance tapering for likelihood-based estimation in large spatial data sets. Journal of the American Statistical Association, 103(484):1545–1555, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kimball:2008:IRT

[KSS08] Miles S. Kimball, Claudia R. Sahm, and Matthew D. Shapiro. Imputing risk tolerance from survey responses. Journal of the American Statistical Association, 103(483):1028–1038, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2001:BBM

[KST01] Hoon Kim, Dongchu Sun, and Robert K. Tsutakawa. A bivariate Bayes method for improving the estimates of mortality rates with a twofold conditional autoregressive model. Journal of the American Statistical Association, 96(456): 1506–1521, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085917;   reqidx=/cw/
asa/01621459/v96n456/s30/p1506.

Kibria:2002:BSP

[KSZL02] B. M. Golam Kibria, Li Sun, James V. Zidek, and Nhu D. Le. Bayesian spatial prediction of random space–time fields with application to mapping PM_{2.5} exposure. Journal of the American Statistical Association, 97(457):112–124, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085763;   reqidx=/cw/asa/01621459/v97n457/
s11/p112.

Kumbhakar:2005:JMT

[KT05] Subal C. Kumbhakar and Efthymios G. Tsionas. The joint measurement of technical and allocative inefficiencies: An application of Bayesian inference in nonlinear random-effects models. Journal of the American Statistical Association, 100 (471):736–747, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590611.

Kuk:2009:ETV

[KT09] Anthony Y. C. Kuk and C. C. Tan. Estimating the timevarying rate of transmission of SARS in Singapore and Hong Kong under two environments. Journal of the American Statistical Association, 104(485):88–96, March 2009. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kvam:2000:NBE

[KTZ00] Paul H. Kvam, Ram C. Tiwari, and Jyoti N. Zalkikar. Nonparametric Bayes estimation of contamination levels using observations from the residual distribution. Journal of the American Statistical Association, 95(452):1119–1126, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669748.

Kneip:2001:IDFa

[KU01a] Alois Kneip and Klaus J. Utikal. Inference for density families using functional principal component analysis. Journal of the American Statistical Association, 96(454):519–532, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670290;   reqidx=/cw/asa/01621459/v96n454/
s16/p519.

Kneip:2001:IDFb

[KU01b] Alois Kneip and Klaus J. Utikal. Inference for density families using functional principal component analysis: Rejoinder. Journal of the American Statistical Association, 96(454): 540–542, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670295.

Kuersteiner:2000:BRB

[Kue00] Guido M. Kuersteiner. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1014–1016, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669498.

Kulldorff:2006:TSR

[Kul06] Martin Kulldorff. Tests of spatial randomness adjusted for an inhomogeneity: a general framework. Journal of the American Statistical Association, 101(475):1289–1305, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590802.

Khan:2007:RLM

[KVZ07] Jafar A. Khan, Stefan Van Aelst, and Ruben H. Zamar. Robust linear model selection based on least angle regression. Journal of the American Statistical Association, 102(480): 1289–1299, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ke:2001:SNMa

[KW01a] Chunlei Ke and Yuedong Wang. Semiparametric nonlinear mixed-effects models and their applications. Journal of the American Statistical Association, 96(456):1272–1281, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085894;   reqidx=/cw/asa/01621459/v96n456/
s13/p1272.

Ke:2001:SNMb

[KW01b] Chunlei Ke and Yuedong Wang. [semiparametric nonlinear mixed-effects models and their applications]: Rejoinder. Journal of the American Statistical Association, 96(456):1294– 1298, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085900.

Kim:2001:JST

[KW01c] Tae-Hwan Kim and Halbert White. James–Stein-type estimators in large samples with application to the least absolute deviations estimator. Journal of the American Statistical Association, 96(454):697–705, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670307;   reqidx=/cw/
asa/01621459/v96n454/s28/p697.

Kelsall:2002:MSV

[KW02] Julia Kelsall and Jonathan Wakefield. Modeling spatial variation in disease risk: a geostatistical approach. Journal of the American Statistical Association, 97(459): 692–701, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s4/p692;   http://
www.jstor.org/stable/3085705.

Ke:2004:SSN

[KW04] Chunlei Ke and Yuedong Wang. Smoothing spline nonlinear nonparametric regression models. Journal of the American Statistical Association, 99(468):1166–1175, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s27/p1166;   http://www.jstor.org/stable/27590494.

Koenker:2004:URQ

[KX04] Roger Koenker and Zhijie Xiao. Unit root quantile autoregression inference. Journal of the American Statistical Association, 99(467):775–787, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s25/p775;   http:
//www.jstor.org/stable/27590447.

Koenker:2006:QA

[KX06a] Roger Koenker and Zhijie Xiao. Quantile autoregression. Journal of the American Statistical Association, 101(475):980–990, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:2006:RQA

[KX06b] Roger Koenker and Zhijie Xiao. Rejoinder: Quantile autoregression. Journal of the American Statistical Association, 101 (475):1002–1006, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:2006:QAC

[KXF^{+}06] Roger Koenker, Zhijie Xiao, Jianqing Fan, Yingying Fan, Marc Knight, Marc Hallin, Bas J. M. Werker, Christian M. Hafner, Oliver B. Linton, and P. M. Robinson. Quantile autoregression [with comments, rejoinder]. Journal of the American Statistical Association, 101(475):980–1006, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590777.

Krueger:2003:PSA

[KZ03] Alan B. Krueger and Pei Zhu. Principal stratification approach to broken randomized experiments: a case study of school choice vouchers in New York City [with comment]. Journal of the American Statistical Association, 98(462):314–318, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045240.

Lu:2006:AEI

[LA06] Guobing Lu and A. E. Ades. Assessing evidence inconsistency in mixed treatment comparisons. Journal of the American Statistical Association, 101(474):447–459, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590708.

Liu:2008:AEG

[LA08] Qing Liu and Keaven M. Anderson. On adaptive extensions of group sequential trials for clinical investigations. Journal of the American Statistical Association, 103(484):1621–1630, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lachenbruch:2009:CSD

[Lac09] Peter A. Lachenbruch. Communicating statistics and developing professionals: The 2008 ASA Presidential Address. Journal of the American Statistical Association, 104(485):1–4, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lund:2003:BRBa

[LACPE03] Robert Lund, Christine M. Anderson-Cook, Mario Peruggia, and Paul Embrechts. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 98(463):777– 778, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045322.

Lahiri:2002:SBV

[Lah02] S. N. Lahiri. [sieve bootstrap with variable-length Markov chains for stationary categorical time series]: Comment. Journal of the American Statistical Association, 97(458):460–462, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085660.

Lambert:2000:I

[Lam00] Diane Lambert. [introduction]. Journal of the American Statistical Association, 95(451):971, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669478.

Lang:2000:BRB

[Lan00] Joseph B. Lang. Book review: [untitled]. Journal of the American Statistical Association, 95(450):680–681, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669421.

Lang:2003:BRB

[Lan03] Joseph B. Lang. Book review: An Introduction to Generalized Linear Models (second edition), by Annette J. Dobson. Journal of the American Statistical Association, 98 (464):1086–1087, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s39/p1086;   http://www.
jstor.org/stable/30045360.

Lang:2005:HLP

[Lan05] Joseph B. Lang. Homogeneous linear predictor models for contingency tables. Journal of the American Statistical Association, 100(469):121–134, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s17/p121;   http://www.
jstor.org/stable/27590523.

Larsen:2002:BRB

[Lar02] Michael D. Larsen. Book review: [untitled]. Journal of the American Statistical Association, 97(457):358–359, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085796.

Larsen:2006:BRB

[Lar06] Michael D. Larsen. Book review: Longitudinal and Panel Data: Analysis and Applications in the Social Sciences, by Edward W. Frees. Journal of the American Statistical Association, 101(473):402, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047489.

Last:2006:BRB

[Las06] M. Last. Book review: Kernel Methods for Pattern Analysis, by John Shawe-Taylor; Nello Cristianini. Journal of the American Statistical Association, 101(476):1730, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639801.

Last:2007:BRBb

[Las07a] M. Last. Book review: Analysis of Phylogenetics and Evolution With R. Emmanuel Paradis. Journal of the American Statistical Association, 102(480):1477–1478, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Last:2007:BRBa

[Las07b] M. Last. Book review: Pattern Recognition Algorithms for Data Mining. Journal of the American Statistical Association, 102(478):759, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Last:2008:BRB

[Las08] M. Last. Book review: Applying Statistics in the Courtroom: A New Approach for Attorneys and Expert Witnesses; The Expert: The Statistical Analyst in Litigation. Journal of the American Statistical Association, 103(482):880–881, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2003:DRM

[LASA03] Ker-Chau Li, Yve Aragon, Kerby Shedden, and C. Thomas Agnan. Dimension reduction for multivariate response data. Journal of the American Statistical Association, 98(461):99–109, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s11/p99;   http://
www.jstor.org/stable/30045198.

Laviolette:2004:CMF

[Lav04] Michael Laviolette. Comment: Membership functions and probability measures of fuzzy sets. Journal of the American Statistical Association, 99(467):879–880, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s36/p879.

Lawless:2000:SR

[Law00] Jerry Lawless. Statistics in reliability. Journal of the American Statistical Association, 95(451):989–992, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669483.

Lazar:2007:CSA

[Laz07] Nicole A. Lazar. Comment: Statistical analysis of diffusion tensors in diffusion-weighted magnetic resonance imaging data. Journal of the American Statistical Association, 102 (480):1105, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2001:SBA

[LB01] Jaeyong Lee and James O. Berger. Semiparametric Bayesian analysis of selection models. Journal of the American Statistical Association, 96(456):1397–1409, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085908;   reqidx=/cw/asa/01621459/v96n456/
s21/p1397.

Lu:2006:ACC

[LBS06] Wilson W. Lu, J. Michael Brick, and Randy R. Sitter. Algorithms for constructing combined strata variance estimators. Journal of the American Statistical Association, 101 (476):1680–1692, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639782.

Lin:2000:NFE

[LC00] Xihong Lin and Raymond J. Carroll. Nonparametric function estimation for clustered data when the predictor is measured without/with error. Journal of the American Statistical Association, 95(450):520–534, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/2000/June/Lin.htm;
 http://www.jstor.org/stable/2669396.

Lin:2001:SNRb

[LC01a] Xihong Lin and Raymond J. Carroll. Semiparametric and nonparametric regression analysis of longitudinal data: Comment. Journal of the American Statistical Association, 96 (453):114–116, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670346.

Lin:2001:SRC

[LC01b] Xihong Lin and Raymond J. Carroll. Semiparametric regression for clustered data using generalized estimating equations. Journal of the American Statistical Association, 96 (455):1045–1056, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670250;   reqidx=/cw/
asa/01621459/v96n455/s22/p1045.

Love:2009:RMD

[LC09] Tanzy Love and Alicia Carriquiry. Repeated measurements on distinct scales with censoring — a Bayesian approach applied to microarray analysis of maize. Journal of the American Statistical Association, 104(486):524–540, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leman:2009:MS

[LCL09] Scotland C. Leman, Yuguo Chen, and Michael Lavine. The multiset sampler. Journal of the American Statistical Association, 104(487):1029–1041, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2008:ASC

[LCLC08] Sheng Luo, Ciprian M. Crainiceanu, Thomas A. Louis, and Nilanjan Chatterjee. Analysis of smoking cessation patterns using a stochastic mixed-effects model with a latent cured state. Journal of the American Statistical Association, 103 (483):1002–1013, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lou:2005:GSF

[LCT^{+}05] Xiang-Yang Lou, George Casella, Rory J. Todhunter, Mark C. K. Yang, and Rongling Wu. A general statistical framework for unifying interval and linkage disequilibrium mapping: Toward high-resolution mapping of quantitative traits. Journal of the American Statistical Association, 100(469):158–171, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s20/p158;   http://www.
jstor.org/stable/27590526.

Liu:2009:JMA

[LDM09] Xuefeng Liu, Michael J. Daniels, and Bess Marcus. Joint models for the association of longitudinal binary and continuous processes with application to a smoking cessation trial. Journal of the American Statistical Association, 104(486):429–438, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lennox:2009:DEP

[LDVT09] Kristin P. Lennox, David B. Dahl, Marina Vannucci, and Jerry W. Tsai. Density estimation for protein conformation angles using a bivariate von Mises distribution and Bayesian nonparametrics. Journal of the American Statistical Association, 104(486):586–596, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2000:MDL

[Lee00] Thomas C. M. Lee. A minimum description length–based image segmentation procedure, and its comparison with a cross-validation–based segmentation procedure. Journal of the American Statistical Association, 95(449):259– 270, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/Lee.htm;   http://www.jstor.
org/stable/2669543.

Lee:2003:BRB

[Lee03] Herbert K. H. Lee. Book review: Elements of Computational Statistics, by James E. Gentle. Journal of the American Statistical Association, 98(463):772–773, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30045315.

Lee:2006:BRB

[Lee06] Yoonkyung Lee. Book review: Semiparametric Regression, by David Ruppert; M. P. Wand; R. J. Carroll. Journal of the American Statistical Association, 101(476):1722–1723, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27639790.

Lee:2008:BRB

[Lee08] Thomas C. M. Lee. Book review: Information and Complexity in Statistical Modeling. Journal of the American Statistical Association, 103(483):1321–1322, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lehtonen:2005:BRB

[Leh05] Risto Lehtonen. Book review: Multilevel Statistical Models, third edition, by Harvey Goldstein. Journal of the American Statistical Association, 100(469):353–354, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s45/
p354;   http://www.jstor.org/stable/27590551.

Leippold:2005:BRB

[Lei05] Markus Leippold. Book review: Statistics, Econometrics and Forecasting, by Arnold Zellner. Journal of the American Statistical Association, 100(472):1458, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590687.

Lenth:2002:BRBa

[Len02a] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 97(458):654, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085683.

Lenth:2002:BRBb

[Len02b] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 97(458):654–655, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085684.

Lenth:2002:BRBc

[Len02c] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 97(459):924, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085738.

Lenth:2002:BRBd

[Len02d] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 97(459):924, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085739.

Lenth:2002:BRBe

[Len02e] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1217, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085871.

Lenth:2002:EQS

[Len02f] Russell V. Lenth. Experimental quality: a strategic approach to achieve and improve quality. Journal of the American Statistical Association, 97(459):924–??, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s29/p924.

Lenth:2002:TDE

[Len02g] Russell V. Lenth. The theory of the design of experiments. Journal of the American Statistical Association, 97 (459):924–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s28/p924.

Lenth:2003:BDH

[Len03a] Russell V. Lenth. Binomial distribution handbook for scientists and engineers. Journal of the American Statistical Association, 98(462):492–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s25/p492.

Lenth:2003:BRB

[Len03b] Russell V. Lenth. Book review: Binomial Distribution Handbook for Scientists and Engineers, by Elart von Collanist; Klaus Drager. Journal of the American Statistical Association, 98(462):492–493, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045261.

Lenth:2004:BRBb

[Len04a] Russell V. Lenth. Book review: Power Analysis for Experimental Research: A Practical Guide for the Biological, Medical and Social Sciences, by R. Barker Bausell; Yu-Fang Lid; Statistical Power Analysis: A Simple and General Model for Traditional and Modern Hypothesis Tests (second edition), by Kevin R. Murphy; Brett Myors. Journal of the American Statistical Association, 99(468):1206–1207, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s38/p1206;   http://www.jstor.org/stable/27590505.

Lenth:2004:BRBa

[Len04b] Russell V. Lenth. Book review: The Cognitive Style of PowerPoint, by Edward R. Tufte. Journal of the American Statistical Association, 99(466):569, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s41/p569;   http:
//www.jstor.org/stable/27590427.

Lenth:2005:BRBc

[Len05] Russell V. Lenth. Book review: Recurrent Events Data Analysis for Product Repairs, Disease Recurrences, and Other Applications, by Wayne B. Nelson. Journal of the American Statistical Association, 100(470):713, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s44/p713;   http://www.
jstor.org/stable/27590603.

Lenth:2006:BRB

[Len06] Russell V. Lenth. Book review: Applied Adaptive Statistical Methods: Tests of Significance and Confidence Intervals, by Thomas W. O’Gorman. Journal of the American Statistical Association, 101(476):1727–1728, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639797.

Levine:2000:BRB

[Lev00] Richard A. Levine. Book review: [untitled]. Journal of the American Statistical Association, 95(449):344–345, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669579.

Lopez-Fidalgo:2004:OED

[LFGR04] Jesús López-Fidalgo and Sandra A. Garcet-Rodríguez. Optimal experimental designs when some independent variables are not subject to control. Journal of the American Statistical Association, 99(468):1190–1199, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s29/p1190;   http://www.jstor.org/stable/27590496.

Losinger:2003:MOP

[LFKS03] Willard C. Losinger, Carmen Fernández, Gary Koop, and Mark F. J. Steel. Multiple-output production with undesirable outputs: An application to nitrogen surplus in agriculture (with response). Journal of the American Statistical Association, 98(461):261, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045230.

Lin:2008:AMS

[LFY^{+}08] Haiqun Lin, Ziding Feng, Yan Yu, Yingye Zheng, Narayan Shivapurkar, and Adi F. Gazdar. Application of multidimensional selective item response regression model for studying multiple gene methylation in SV40 oncogenic pathways. Journal of the American Statistical Association, 103(481):201–211, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2007:NAP

[LGS07] Bo Li, Marc G. Genton, and Michael Sherman. A nonparametric assessment of properties of space–time covariance functions. Journal of the American Statistical Association, 102(478):736– 744, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2006:D

[LH06] Hongzhe Li and Fangxin Hong. Discussion. Journal of the American Statistical Association, 101(476):1332–1334, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2007:MAC

[LH07] W. Liu and A. J. Hayter. Minimum area confidence set optimality for confidence bands in simple linear regression. Journal of the American Statistical Association, 102(477):181–190, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2009:TEP

[LH09] Yi Liu and Jason Hsu. Testing for efficacy in primary and secondary endpoints by partitioning decision paths. Journal of the American Statistical Association, 104(488):1661–1670, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Logan:2008:TSE

[LHN08] John Allen Logan, Peter D. Hoff, and Michael A. Newton. Two-sided estimation of mate preferences for similarities in age, education, and religion. Journal of the American Statistical Association, 103(482):559–569, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2008:SSC

[LHNM08] Yufeng Liu, David Neil Hayes, Andrew Nobel, and J. S. Marron. Statistical significance of clustering for high-dimension, low-sample size data. Journal of the American Statistical Association, 103(483):1281–1293, September 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2002:SMA

[LHSY02] Lawrence Lin, A. S. Hedayat, Bikas Sinha, and Min Yang. Statistical methods in assessing agreement: Models, issues, and tools. Journal of the American Statistical Association, 97(457):257–270, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085780;   reqidx=/cw/
asa/01621459/v97n457/s23/p257.

Li:2000:PLC

[Li00] Bing Li. On profile likelihood: Comment. Journal of the American Statistical Association, 95(450):472–474, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669389.

Li:2005:HFM

[Li05] Ta-Hsin Li. A hierarchical framework for modeling and forecasting Web server workload. Journal of the American Statistical Association, 100(471):748–763, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590612.

Li:2006:CLB

[Li06] Hongzhe Li. Comment: Likelihood-based inference on haplotype effects in genetic association studies. Journal of the American Statistical Association, 101(473):114–115, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2007:SFO

[Li07] Ta-Hsin Li. A statistical framework of optimal workload consolidation with application to capacity planning for on-demand computing. Journal of the American Statistical Association, 102(479):841–855, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2008:BRB

[Li08a] Bo Li. Book review: Model-Based Geostatistics. Journal of the American Statistical Association, 103(483):1325–1326, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2008:LPT

[Li08b] Ta-Hsin Li. Laplace periodogram for time series analysis. Journal of the American Statistical Association, 103(482):757–768, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2002:DWI

[Lia02] Faming Liang. Dynamically weighted importance sampling in Monte Carlo computation. Journal of the American Statistical Association, 97(459):807–821, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s13/p807;   http://www.jstor.org/stable/3085723.

Liang:2005:GWL

[Lia05] Faming Liang. A generalized Wang–Landau algorithm for Monte Carlo computation. Journal of the American Statistical Association, 100(472):1311–1327, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590674.

Lindsey:2003:BRB

[Lin03a] Jane Lindsey. Book review: Clinical Trials, by Lelia Duley; Barbara Farrell. Journal of the American Statistical Association, 98(463):774, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045318.

Ling:2003:AET

[Lin03b] Shiqing Ling. Adaptive estimators and tests of stationary and nonstationary short- and long-memory ARFIMA-GARCH models. Journal of the American Statistical Association, 98(464):955–967, December 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v98n464/s21/p955;   http:
//www.jstor.org/stable/30045342.

Lin:2004:BRB

[Lin04a] X. Sheldon Lin. Book review: Generalized Poisson Models and Their Applications in Insurance and Finance, by Vladimir E. Bening; Victor Yu Korolev. Journal of the American Statistical Association, 99(466):563, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s31/p563;   http://www.jstor.org/stable/27590417.

Linder:2004:BRB

[Lin04b] R. Scott Linder. Book review: Parametric and Nonparametric Inference from Record-Breaking Data, by Sneh Gulati; William J. Padgett. Journal of the American Statistical Association, 99(465):297, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s33/p297;   http:
//www.jstor.org/stable/27590379.

Lindley:2004:CMF

[Lin04c] D. V. Lindley. Comment: Membership functions and probability measures of fuzzy sets. Journal of the American Statistical Association, 99(467):877–879, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s35/p877.

Lindsey:2004:BRB

[Lin04d] Jane C. Lindsey. Book review: Health Science Research: A Handbook of Quantitative Methods, by Jennifer Peat. Journal of the American Statistical Association, 99 (468):1208, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s41/p1208;   http://www.
jstor.org/stable/27590508.

Lindskog:2007:BRB

[Lin07] Filip Lindskog. Book review: Measures, Integrals and Martingales. Journal of the American Statistical Association, 102 (478):766, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Littell:2000:BRB

[Lit00] Ramon C. Littell. Book review: [untitled]. Journal of the American Statistical Association, 95(449):345, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669580.

Little:2004:MMC

[Lit04] Roderick J. Little. To model or not to model? competing modes of inference for finite population sampling. Journal of the American Statistical Association, 99(466):546–556, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s23/p546;   http://www.jstor.org/stable/27590409.

Little:2007:CMA

[Lit07] Roderick J. Little. Comment: Model-assisted estimation of forest resources with generalized additive models. Journal of the American Statistical Association, 102(478):412–415, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2000:EDD

[Liu00] Chuanhai Liu. Estimation of discrete distributions with a class of simplex constraints. Journal of the American Statistical Association, 95(449):109–120, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
Liu2.htm;   http://www.jstor.org/stable/2669531.

Liu:2003:ASS

[Liu03] Chuanhai Liu. Alternating subspace-spanning resampling to accelerate Markov chain Monte Carlo simulation. Journal of the American Statistical Association, 98(461):110–117, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.

com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s12/p110;   http://www.jstor.org/stable/30045199.

Lloyd:2000:NDE

[LJ00] Chris J. Lloyd and M. C. Jones. Nonparametric density estimation from biased data with unknown biasing function. Journal of the American Statistical Association, 95(451):865– 876, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669470.

Lin:2006:RFA

[LJ06] Yi Lin and Yongho Jeon. Random forests and adaptive nearest neighbors. Journal of the American Statistical Association, 101(474):578–590, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590719.

Liu:2004:MCS

[LJZ04] Wei Liu, Mortaza Jamshidian, and Ying Zhang. Multiple comparison of several linear regression models. Journal of the American Statistical Association, 99(466):395–403, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/s9/
p395;   http://www.jstor.org/stable/27590395.

Lee:2005:PS

[LKF05] Bee Leng Lee, Michael R. Kosorok, and Jason P. Fine. The profile sampler. Journal of the American Statistical Association, 100(471):960–969, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590626.

Layton:2003:HMD

[LL03a] David F. Layton and Richard A. Levine. How much does the far future matter? A hierarchical Bayesian analysis of the public’s willingness to mitigate ecological impacts of climate change. Journal of the American Statistical Association, 98(463):533–544, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-

bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s4/p533;   http://www.jstor.org/stable/30045279.

Li:2003:FIF

[LL03b] Yi Li and Xihong Lin. Functional inference in frailty measurement error models for clustered survival data using the SIMEX approach. Journal of the American Statistical Association, 98(461):191–203, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s19/p191;   http://
www.jstor.org/stable/30045206.

Larocque:2004:CDF

[LL04] Denis Larocque and Mélanie Labarre. A conditionally distribution-free multivariate sign test for one-sided alternatives. Journal of the American Statistical Association, 99(466):499–509, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s19/p499;   http:
//www.jstor.org/stable/27590405.

Lahiri:2005:RAL

[LL05a] P. Lahiri and Michael D. Larsen. Regression analysis with linked data. Journal of the American Statistical Association, 100(469):222–230, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s26/p222;   http://www.
jstor.org/stable/27590532.

Lai:2005:OAP

[LL05b] P. Y. Lai and Stephen M. S. Lee. An overview of asymptotic properties of L_{p} regression under general classes of error distributions. Journal of the American Statistical Association, 100(470):446–458, June 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s8/p446;   http:
//www.jstor.org/stable/27590567.

Lambert:2006:AT

[LL06a] Diane Lambert and Chuanhai Liu. Adaptive thresholds. Journal of the American Statistical Association, 101(473):78–88, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:2006:ATM

[LL06b] Diane Lambert and Chuanhai Liu. Adaptive thresholds: Monitoring streams of network counts. Journal of the American Statistical Association, 101(473):78–88, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30047442.

Li:2006:SNT

[LL06c] Yi Li and Xihong Lin. Semiparametric normal transformation models for spatially correlated survival data. Journal of the American Statistical Association, 101(474):591–603, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590720.

Li:2007:RMB

[LL07] Yan Li and P. Lahiri. Robust model-based and model-assisted predictors of the finite population total. Journal of the American Statistical Association, 102(478):664–673, June 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2008:SDR

[LL08] Lexin Li and Wenbin Lu. Sufficient dimension reduction with missing predictors. Journal of the American Statistical Association, 103(482):822–831, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2009:VSP

[LL09] Hua Liang and Runze Li. Variable selection for partially linear models with measurement errors. Journal of the American Statistical Association, 104(485):234–248, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2000:ITS

[LLC00] Ker-Chau Li, Heng-Hui Lue, and Chun-Houh Chen. Interactive tree-structured regression via principal Hessian directions. Journal of the American Statistical Association, 95(450):547–560, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.amstat.org/publications/jasa/2000/June/Li.htm;
 http://www.jstor.org/stable/2669398.

Liang:2007:SAM

[LLC07] Faming Liang, Chuanhai Liu, and Raymond J. Carroll. Stochastic approximation in Monte Carlo computation. Journal of the American Statistical Association, 102(477):305–320, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:2008:CCN

[LLL08a] Roderick J. Little, Qi Long, and Xihong Lin. Comment: Can nonrandomized experiments yield accurate answers? A randomized experiment comparing random and nonrandom assignments. Journal of the American Statistical Association, 103(484):1344–1346, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Long:2008:CIH

[LLL08b] Qi Long, Roderick J. Little, and Xihong Lin. Causal inference in hybrid intervention trials involving treatment choice. Journal of the American Statistical Association, 103(482):474– 484, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2000:MTM

[LLW00] Jun S. Liu, Faming Liang, and Wing Hung Wong. The multiple-try method and local optimization in Metropolis sampling. Journal of the American Statistical Association, 95(449):121–134, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/Liu.htm;   http://www.
jstor.org/stable/2669532.

Liu:2001:TDW

[LLW01] Jun S. Liu, Faming Liang, and Wing Hung Wong. A theory for dynamic weighting in Monte Carlo computation. Journal of the American Statistical Association, 96(454):561–573, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/

cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670297;   reqidx=/cw/asa/01621459/v96n454/
s18/p561.

Lee:2004:MSV

[LLW04] Yoonkyung Lee, Yi Lin, and Grace Wahba. Multicategory support vector machines: Theory and application to the classification of microarray data and satellite radiance data. Journal of the American Statistical Association, 99(465):67–81, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/s8/
p67;   http://www.jstor.org/stable/27590354.

Li:2007:QRR

[LLZ07] Youjuan Li, Yufeng Liu, and Ji Zhu. Quantile regression in reproducing kernel Hilbert spaces. Journal of the American Statistical Association, 102(477):255–268, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lynn:2000:UPC

[LM00] Henry S. Lynn and Charles E. McCulloch. Using principal component analysis and correspondence analysis for estimation in latent variable models. Journal of the American Statistical Association, 95(450):561–572, June 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/2000/June/
Lynn.htm;   http://www.jstor.org/stable/2669399.

Lele:2002:IIM

[LM02] Subhash R. Lele and Charles E. McCulloch. Invariance, identifiability, and morphometrics. Journal of the American Statistical Association, 97(459):796–806, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s12/p796;   http://www.jstor.org/stable/3085722.

Liu:2004:FCA

[LM04] Xueli Liu and Hans-Georg Müller. Functional convex averaging and synchronization for time-warped random curves. Journal of the American Statistical Association, 99(467):687–699, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s18/p687;   http:
//www.jstor.org/stable/27590440.

Lindquist:2009:LRB

[LM09a] Martin A. Lindquist and Ian W. McKeague. Logistic regression with Brownian-like predictors. Journal of the American Statistical Association, 104(488):1575–1585, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2009:EDS

[LM09b] Bitao Liu and Hans-Georg Müller. Estimating derivatives for samples of sparsely observed functions, with application to online auction dynamics. Journal of the American Statistical Association, 104(486):704–717, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lijoi:2005:HMM

[LMP05] Antonio Lijoi, Ramsés H. Mena, and Igor Prünster. Hierarchical mixture modeling with normalized inverse-gaussian priors. Journal of the American Statistical Association, 100 (472):1278–1291, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590671.

Landrum:2003:SRM

[LNR03] Mary Beth Landrum, Sharon-Lise T. Normand, and Robert A. Rosenheck. Selection of related multivariate means: Monitoring psychiatric care in the Department of Veterans Affairs. Journal of the American Statistical Association, 98(461):7–16, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s2/p7;   http://www.jstor.org/stable/30045189.

Lo:2000:FSS

[Lo00] Andrew W. Lo. Finance: a selective survey. Journal of the American Statistical Association, 95(450):629–635, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669406.

Liu:2006:EMD

[LO06] Ruixue Liu and Art B. Owen. Estimating mean dimensionality of analysis of variance decompositions. Journal of the American Statistical Association, 101(474):712–721, June 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590729.

Lobato:2001:TDP

[Lob01] Ignacio N. Lobato. Testing that a dependent process is uncorrelated. Journal of the American Statistical Association, 96(455):1066–1076, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670252;   reqidx=/cw/
asa/01621459/v96n455/s24/p1066.

Lock:2002:BRB

[Loc02a] Robin H. Lock. Book review: [untitled]. Journal of the American Statistical Association, 97(459):919–920, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085733.

Lock:2002:MSM

[Loc02b] Robin H. Lock. Mathematics and sports (mathematical world, vol. 3). Journal of the American Statistical Association, 97(459):919–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s23/p919.

Lockwood:2004:BRB

[Loc04] J. R. Lockwood. Book review: Risk Analysis in Engineering and Economics, by Bilal M. Ayyub. Journal of the American Statistical Association, 99(466):563–564, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s33/p563;   http://www.jstor.org/stable/27590419.

Louis:2005:CMP

[Lou05] Thomas A. Louis. Comment: Maximization by parts in likelihood inference. Journal of the American Statistical Association, 100(472):1159–1160, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2005:PCD

[LP05] Qing Liu and Gordon W. Pledger. Phase 2 and 3 combination designs to accelerate drug development. Journal of the American Statistical Association, 100(470):493–502, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s12/
p493;   http://www.jstor.org/stable/27590571.

Lee:2006:MBN

[LP06] Stephen M. S. Lee and M. C. Pun. On m out of n bootstrapping for nonstandard M-estimation with nuisance parameters. Journal of the American Statistical Association, 101 (475):1185–1197, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590794.

Li:2000:PTC

[LPD00] Shuying Li, Dennis K. Pearl, and Hani Doss. Phylogenetic tree construction using Markov chain Monte Carlo. Journal of the American Statistical Association, 95(450): 493–508, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/2000/June/Li2.htm;   http://
www.jstor.org/stable/2669394.

Lee:2006:SID

[LPL06] Chu-In Charles Lee, Jianan Peng, and Lin Liu. Statistical inference for the difference between the best treatment mean and a control mean. Journal of the American Statistical Association, 101(475):1050–1058, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590782.

Liang:2008:MGP

[LPM^{+}08] Feng Liang, Rui Paulo, German Molina, Merlise A. Clyde, and Jim O. Berger. Mixtures of g priors for Bayesian variable selection. Journal of the American Statistical Association, 103 (481):410–423, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2002:UTT

[LPP02] Qing Liu, Michael A. Proschan, and Gordon W. Pledger. A unified theory of two-stage adaptive designs. Journal of the American Statistical Association, 97(460):1034–1041, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n460/s9/p1034;   http://www.jstor.org/stable/3085828.

Li:2001:BRS

[LPR01] Yunfei Paul Li, Kathleen J. Propert, and Paul R. Rosenbaum. Balanced risk set matching. Journal of the American Statistical Association, 96(455):870–882, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670227;   reqidx=/cw/asa/01621459/v96n455/s9/p870.

Lopez-Pintado:2009:CDF

[LPR09] Sara López-Pintado and Juan Romo. On the concept of depth for functional data. Journal of the American Statistical Association, 104(486):718–734, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:2001:EMD

[LPS01] Diane Lambert, José C. Pinheiro, and Don X. Sun. Estimating millions of dynamic timing patterns in real time. Journal of the American Statistical Association, 96(453):316–330, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670369;   reqidx=/cw/asa/01621459/v96n453/
s25/p316.

Lijoi:2005:CNN

[LPW05] Antonio Lijoi, Igor Prünster, and Stephen G. Walker. On consistency of nonparametric normal mixtures for Bayesian density estimation. Journal of the American Statistical Association, 100(472):1292–1296, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590672.

Liao:2005:CMP

[LQ05] J. G. Liao and Bahjat F. Qaqish. Comment: Maximization by parts in likelihood inference. Journal of the American Statistical Association, 100(472):1160–1161, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lohr:2000:IDF

[LR00] Sharon L. Lohr and J. N. K. Rao. Inference from dual frame surveys. Journal of the American Statistical Association, 95(449):271–280, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.amstat.org/publications/jasa/Lohr.htm;   http:
//www.jstor.org/stable/2669544.

Larsen:2001:IAR

[LR01] Michael D. Larsen and Donald B. Rubin. Iterative automated record linkage using mixture models. Journal of the American Statistical Association, 96(453):32–41, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670338;   reqidx=/cw/asa/01621459/v96n453/s4/p32.

Li:2004:SAH

[LR04] Yi Li and Louise Ryan. Survival analysis with heterogeneous covariate measurement error. Journal of the American Statistical Association, 99(467):724–735, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s21/p724;   http://www.jstor.org/stable/27590443.

Lohr:2006:EMF

[LR06] Sharon Lohr and J. N. K. Rao. Estimation in multiple-frame surveys. Journal of the American Statistical Association, 101 (475):1019–1030, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590779.

Li:2001:MCD

[LS01] Ker-Chau Li and Kerby Shedden. Monte Carlo deconvolution of digital signals guided by the inverse filter. Journal of the American Statistical Association, 96(455):1014–1021, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670247;   reqidx=/cw/asa/01621459/v96n455/
s19/p1014.

Li:2002:ISC

[LS02] Ker-Chau Li and Kerby Shedden. Identification of shared components in large ensembles of time series using dimension reduction. Journal of the American Statistical Association, 97(459):759–765, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s9/p759;   http://
www.jstor.org/stable/3085719.

Liu:2006:ML

[LS06] Yufeng Liu and Xiaotong Shen. Multicategory ψ-learning. Journal of the American Statistical Association, 101(474): 500–509, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590712.

Lemos:2009:RST

[LS09a] R. T. Lemos and B. Sansó. Rejoinder: a spatio-temporal model for mean, anomaly, and trend fields of North Atlantic sea surface temperature. Journal of the American Statistical Association, 104(485):23–25, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LS09b].

Lemos:2009:STM

[LS09b] Ricardo T. Lemos and Bruno Sansó. A spatio-temporal model for mean, anomaly, and trend fields of North Atlantic sea surface temperature. Journal of the American Statistical Association, 104(485):5–18, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Hig09, Men09, WM09b, LS09a].

Liu:2009:SRC

[LS09c] Hao Liu and Yu Shen. A semiparametric regression cure model for interval-censored data. Journal of the American Statistical Association, 104(487):1168–1178, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2004:MLS

[LSC04] Yonghee Lee, Jun Shao, and Shein-Chung Chow. Modified large-sample confidence intervals for linear combinations of variance components: Extension, theory, and application. Journal of the American Statistical Association, 99(466):467–478, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s16/p467;   http:
//www.jstor.org/stable/27590402.

Lockwood:2001:CAO

[LSGS01] J. R. Lockwood, Mark J. Schervish, Patrick Gurian, and Mitchell J. Small. Characterization of arsenic occurrence in source waters of U.S. community water systems. Journal of the American Statistical Association, 96(456):1184–1193, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085881;   reqidx=/cw/asa/01621459/v96n456/
s5/p1184.

Lockwood:2004:ACC

[LSGS04] J. R. Lockwood, Mark J. Schervish, Patrick L. Gurian, and Mitchell J. Small. Analysis of contaminant co-occurrence in community water systems. Journal of the American Statistical Association, 99(465):45–56, March 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/s6/
p45;   http://www.jstor.org/stable/27590352.

Loh:2003:ELS

[LSQ03] J. M. Loh, M. L. Stein, and J. M. Quashnock. Estimating the large-scale structure of the universe using quasi-stellar object carbon IV absorbers. Journal of the American Statistical Association, 98(463):522–532, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-

bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s3/p522;   http://www.jstor.org/stable/30045278.

Lee:2008:IPA

[LSZH08] Herbert K. H. Lee, Bruno Sansó, Weining Zhou, and David M. Higdon. Inference for a proton accelerator using convolution models. Journal of the American Statistical Association, 103 (482):604–613, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2008:LNC

[LTE08] Julia Y. Lin, Thomas R. Ten Have, and Michael R. Elliott. Longitudinal nested compliance class model in the presence of time-varying noncompliance. Journal of the American Statistical Association, 103(482):462–473, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2002:LCM

[LTMS02] Haiqun Lin, Bruce W. Turnbull, Charles E. McCulloch, and Elizabeth H. Slate. Latent class models for joint analysis of longitudinal biomarker and event process data: Application to longitudinal prostate-specific antigen readings and prostate cancer. Journal of the American Statistical Association, 97(457):53–65, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085758;   reqidx=/cw/
asa/01621459/v97n457/s6/p53.

Luethi:2005:BRB

[Lue05] Hans-Jakob Luethi. Book review: Convex Optimization, by Stephen Boyd; Lieven Vandenberghe. Journal of the American Statistical Association, 100(471):1097, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590646.

Lund:2000:BRBa

[Lun00a] Robert B. Lund. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1019, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669504.

Lund:2000:BRBb

[Lun00b] Robert B. Lund. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1375–1376, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669798.

Lund:2001:BRB

[Lun01] Robert B. Lund. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085935.

Lund:2002:BRB

[Lun02] Robert Lund. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1216, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085866.

Lund:2003:BRBb

[Lun03] Robert B. Lund. Book review: Elements of Applied Stochastic Processes (3rd ed.), by U. Narayan Bhat; Gregory K. Miller. Journal of the American Statistical Association, 98 (464):1085–1086, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s38/p1085;   http://www.
jstor.org/stable/30045359.

Lund:2005:BRB

[Lun05] Robert Lund. Book review: Periodic Time Series Models, by Philip Hans Franses; Richard Paap. Journal of the American Statistical Association, 100(472):1458–1459, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590688.

Lund:2006:BRB

[Lun06] Ulric J. Lund. Book review: An R and S-Plus(R) Companion to Multivariate Analysis, by Brian S. Everitt. Journal of the American Statistical Association, 101(474):857–858, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590760.

Lund:2007:BRBa

[Lun07a] Robert Lund. Book review: Statistical Analysis of Stochastic Processes in Time. Journal of the American Statistical Association, 102(477):382, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lund:2007:BRBb

[Lun07b] Robert Lund. Book review: Time Series Analysis and Its Applications: With R Examples. Journal of the American Statistical Association, 102(479):1079, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luthi:2003:BRB

[Lüt03] Hans-Jakob Lüthi. Book review: Algorithms for Worst-Case Design and Applications to Risk Management, by Berç Rustem; Melendres Howe. Journal of the American Statistical Association, 98(463):770, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045311.

Lai:2001:SNN

[LW01a] Tze Leung Lai and Samuel Po-Shing Wong. Stochastic neural networks with applications to nonlinear time series. Journal of the American Statistical Association, 96 (455):968–981, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670243;   reqidx=/cw/
asa/01621459/v96n455/s15/p968.

Liang:2001:RPE

[LW01b] Faming Liang and Wing Hung Wong. Real-parameter evolutionary Monte Carlo with applications to Bayesian mixture models. Journal of the American Statistical Association, 96(454):653–666, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670304;   reqidx=/cw/
asa/01621459/v96n454/s25/p653.

Li:2007:DRD

[LW07] Bing Li and Shaoli Wang. On directional regression for dimension reduction. Journal of the American Statistical Association, 102(479):997–1008, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2008:PED

[LW08] Hua Liang and Hulin Wu. Parameter estimation for differential equation models using a framework of measurement error in regression models. Journal of the American Statistical Association, 103(484):1570–1583, December 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2004:EPL

[LWRC04] Hua Liang, Suojin Wang, James M. Robins, and Raymond J. Carroll. Estimation in partially linear models with missing covariates. Journal of the American Statistical Association, 99(466):357–367, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n466/s6/p357;   http://www.
jstor.org/stable/27590392.

Lin:2001:STM

[LWY01] D. Y. Lin, L. J. Wei, and Z. Ying. Semiparametric transformation models for point processes. Journal of the American Statistical Association, 96(454):620–628, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670301;   reqidx=/cw/asa/01621459/v96n454/
s22/p620.

Li:2008:PRM

[LWZ08] Bing Li, Songqiao Wen, and Lixing Zhu. On a projective resampling method for dimension reduction with multivariate responses. Journal of the American Statistical Association, 103 (483):1177–1186, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2001:SNRa

[LY01a] D. Y. Lin and Z. Ying. Semiparametric and nonparametric regression analysis of longitudinal data. Journal of the American Statistical Association, 96(453):103–113, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/

cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670344;   reqidx=/cw/asa/01621459/v96n453/
s10/p103.

Lin:2001:SNRc

[LY01b] D. Y. Lin and Z. Ying. Semiparametric and nonparametric regression analysis of longitudinal data: Rejoinder. Journal of the American Statistical Association, 96(453):123–126, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670349.

Lyles:2006:BRB

[Lyl06] Robert H. Lyles. Book review: Regression Methods in Biostatistics: Linear, Logistic, Survival, and Repeated Measures Models, by Eric Vittinghoff; David V. Glidden; Stephen C. Shiboski; Charles E. Mcculloch. Journal of the American Statistical Association, 101(473):403–404, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047491.

Lin:2001:SNM

[LZ01] Xihong Lin and Daowen Zhang. [semiparametric nonlinear mixed-effects models and their applications]: Comment. Journal of the American Statistical Association, 96(456):1288– 1291, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085898.

Lin:2006:RLB

[LZ06] D. Y. Lin and D. Zeng. Rejoinder: Likelihood-based inference on haplotype effects in genetic association studies. Journal of the American Statistical Association, 101(473):116–118, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2008:UAN

[LZ08] Gang Li and Kefei Zhou. A unified approach to nonparametric comparison of receiver operating characteristic curves for longitudinal and clustered data. Journal of the American Statistical Association, 103(482):705–713, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2005:IPF

[LZCC05] Ming T. Lin, Junni L. Zhang, Qiansheng Cheng, and Rong Chen. Independent particle filters. Journal of the American Statistical Association, 100(472):1412–1421, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590681.

Lu:2009:SEM

[LZH09] Minggen Lu, Ying Zhang, and Jian Huang. Semiparametric estimation methods for panel count data using monotone Bsplines. Journal of the American Statistical Association, 104 (487):1060–1070, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2001:MDO

[LZHR01] Bo Lu, Elaine Zanutto, Robert Hornik, and Paul R. Rosenbaum. Matching with doses in an observational study of a media campaign against drug abuse. Journal of the American Statistical Association, 96(456):1245–1253, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085887;   reqidx=/cw/asa/01621459/v96n456/
s11/p1245.

Lin:2006:LBI

[LZS^{+}06] D. Y. Lin, D. Zeng, Chiara Sabatti, Glen A. Satten, Andrew S. Allen, Michael P. Epstein, Nilanjan Chatterjee, Christine Spinka, Jinbo Chen, Raymond J. Carroll, Jung-Ying Tzeng, Kathryn Roeder, Hongzhe Li, D. Y. Lin, and D. Zeng. Likelihood-based inference on haplotype effects in genetic association studies [with comment]. Journal of the American Statistical Association, 101(473):89–104, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30047443.

MacCall:2001:BRB

[Mac01] Alec D. MacCall. Book review: [untitled]. Journal of the American Statistical Association, 96(454):781, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670323.

MacEachern:2006:BRB

[Mac06a] Steven N. MacEachern. Book review: Introduction to Statistical Throught, by Michael Lavine. Journal of the American Statistical Association, 101(476):1719, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639785.

Morris:2006:UWB

[MAC^{+}06b] Jeffrey S. Morris, Cassandra Arroyo, Brent A. Coull, Louise M. Ryan, Richard Herrick, and Steven L. Gortmaker. Using wavelet-based functional mixed models to characterize population heterogeneity in accelerometer profiles: a case study. Journal of the American Statistical Association, 101(476): 1352–1364, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639756.

Madigan:2001:BRB

[Mad01] David Madigan. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1524, December 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085919.

Magee:2001:SDS

[Mag01] Lonnie Magee. [should the DEA’s STRIDE data be used for economic analyses of markets for illegal drugs?]: Comment. Journal of the American Statistical Association, 96 (456):1266–1268, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085891.

Malevergne:2005:BRB

[Mal05] Yannick Malevergne. Book review: Preparing for the Worst: Incorporating Downside Risk in Stock Market Investments, by Hrishikesh D. Vinod; Derrick P. Reagle. Journal of the American Statistical Association, 100(472):1459–1460, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590689.

Manly:2007:BRB

[Man07] Bryan Manly. Book review: Using Statistical Methods for Water Quality Management: Issues, Problems, and Solutions.

Journal of the American Statistical Association, 102(479): 1080, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mao:2004:PCP

[Mao04] Chang Xuan Mao. Predicting the conditional probability of discovering a new class. Journal of the American Statistical Association, 99(468):1108–1118, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s22/p1108;   http://www.jstor.org/stable/27590489.

Mao:2006:INS

[Mao06] Chang Xuan Mao. Inference on the number of species through geometric lower bounds. Journal of the American Statistical Association, 101(476):1663–1670, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639780.

Marden:2000:ADV

[Mar00a] John I. Marden. Asymptotic distribution of P values in composite null models: Comment. Journal of the American Statistical Association, 95(452):1164–1166, December 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669754.

Marden:2000:HTV

[Mar00b] John I. Marden. Hypothesis testing: From p values to Bayes factors. Journal of the American Statistical Association, 95 (452):1316–1320, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669779.

Marron:2001:IDF

[Mar01] J. S. Marron. Inference for density families using functional principal component analysis: Comment. Journal of the American Statistical Association, 96(454):532–533, June 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670291.

Marasinghe:2003:BRB

[Mar03a] Mervyn Marasinghe. Book review: Numerical Methods of Statistics, by John F. Monahan. Journal of the American Statistical Association, 98(462):497–498, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30045269.

Marasinghe:2003:NMS

[Mar03b] Mervyn Marasinghe. Numerical methods of statistics. Journal of the American Statistical Association, 98(462):497–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n462/s33/p497.

Marcoulides:2005:BRB

[Mar05] George A. Marcoulides. Book review: Discovering Knowledge in Data: An Introduction to Data Mining, by Daniel T. Larose. Journal of the American Statistical Association, 100(472):1465, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590699.

Mardia:2007:CGB

[Mar07] Kanti V. Mardia. Comment: Gait-based human recognition by classification of cyclostationary processes on nonlinear shape manifolds. Journal of the American Statistical Association, 102(480):1124, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mason:2004:DSP

[Mas04] Robert L. Mason. Does the statistics profession have an identity crisis? Journal of the American Statistical Association, 99(465):1–6, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s1/p1;   http://
www.jstor.org/stable/27590347.

Matache:2005:BRB

[Mat05] Ana-Maria Matache. Book review: C++ Design Patterns and Derivatives Pricing, by Mark Joshi. Journal of the American Statistical Association, 100(471):1098, September 2005.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590647.

Marshall:2002:UAC

[MB02] Pablo Marshall and Eric T. Bradlow. A unified approach to conjoint analysis models. Journal of the American Statistical Association, 97(459):674–682, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s2/p674;   http://www.jstor.org/stable/3085703.

Martin:2005:TQI

[MB05] Emily C. Martin and Rebecca A. Betensky. Testing quasiindependence of failure and truncation times via conditional Kendall’s tau. Journal of the American Statistical Association, 100(470):484–492, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s11/p484;   http://www.
jstor.org/stable/27590570.

Murphy:2009:SED

[MB09] S. A. Murphy and D. Bingham. Screening experiments for developing dynamic treatment regimes. Journal of the American Statistical Association, 104(485):391–408, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2006:LEE

[MC06] Yanyuan Ma and Raymond J. Carroll. Locally efficient estimators for semiparametric models with measurement error. Journal of the American Statistical Association, 101(476):1465– 1474, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639765.

McBride:2002:MPP

[McB02] Sandra J. McBride. A marked point process model for the source proximity effect in the indoor environment. Journal of the American Statistical Association, 97(459): 683–691, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s3/p683;   http://
www.jstor.org/stable/3085704.

McCulloch:2000:GLM

[McC00] Charles E. McCulloch. Generalized linear models. Journal of the American Statistical Association, 95(452):1320– 1324, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669780.

McCullough:2002:BRM

[McC02a] B. D. McCullough. Book review: Mathematical Statistics With Mathematica, by Colin Rose and Murray Smith. New York: Springer-Verlag, 2002. ISBN 0-387-95234-9. xiii + 481 pp. (+2 CDs). $79.97. Journal of the American Statistical Association, 97(460):1202–1203, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s25/p1202.

McCullough:2002:BRB

[McC02b] B. D. McCullough. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1202–1203, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085844.

McCullagh:2004:CLS

[McC04] Peter McCullagh. Comment: Location-scale depth. Journal of the American Statistical Association, 99(468):980–981, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s10/p980.

McCullough:2005:BRB

[McC05] B. D. McCullough. Book review: Measurement Theory and Practice: The World through Quantification, by David J. Hand. Journal of the American Statistical Association, 100 (472):1462–1463, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590694.

Mugglin:2000:FMB

[MCG00] Andrew S. Mugglin, Bradley P. Carlin, and Alan E. Gelfand. Fully model-based approaches for spatially misaligned data. Journal of the American Statistical Association, 95(451):877– 887, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669471.

McGee:2002:BRB

[McG02] Monnie McGee. Book review: [untitled]. Journal of the American Statistical Association, 97(457):363–364, March 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085803.

McKeague:2005:SMS

[McK05] Ian W. McKeague. A statistical model for signature verification. Journal of the American Statistical Association, 100(469):231–241, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s27/p231;   http://www.
jstor.org/stable/27590533.

Maiti:2005:BRBb

[MCL05] Tapabrata Maiti, Richard J. Cleary, and Russell V. Lenth. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 100(470):713–715, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590604.

McNeil:2001:BRB

[McN01] Alexander J. McNeil. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1525–1526, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085921.

McNeil:2002:BRB

[McN02a] Alexander J. McNeil. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1210–1211, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085856.

McNeil:2002:LDG

[McN02b] Alexander J. McNeil. The Laplace distribution and generalizations: a revisit with applications to communications, economics, engineering, and finance. Journal of the American Statistical Association, 97(460):1210–??, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s37/p1210.

McNeil:2004:BRB

[McN04] Alexander J. McNeil. Book review: Modeling Financial Time Series with S-Plus, by Eric Zivot; Jiahui Wang. Journal of the American Statistical Association, 99(466):564–565, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s34/p564;   http://www.jstor.org/stable/27590420.

McNeil:2006:BRB

[McN06] Alexander J. McNeil. Book review: Multivariate t Distributions and Their Applications, by Samuel Kotz; Saralees Nadarajah. Journal of the American Statistical Association, 101(473):390–391, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047468.

McNeil:2007:BRB

[McN07] Alexander J. McNeil. Book review: Latent Curve Models: A Structural Equation Approach. Journal of the American Statistical Association, 102(480):1479–1481, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maitra:2001:PPP

[MD01] Ranjan Maitra and Siddhartha R. Dalal. Pay phones, parking meters, vending machines, and optimal Bayesian decisions on collection times. Journal of the American Statistical Association, 96(454):476–487, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670286;   reqidx=/cw/
asa/01621459/v96n454/s12/p476.

Maiti:2005:BRBa

[MDACL05] Tapabrata Maiti, Carolyn Pillers Dobler, Christine M. Anderson-Cook, and Russell V. Lenth. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 100(469):357, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590556.

Meeden:2000:DTA

[Mee00] Glen Meeden. A decision theoretic approach to imputation in finite population sampling. Journal of the American Statistical Association, 95(450):586–595, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.amstat.org/publications/jasa/2000/June/Meeden.
htm;   http://www.jstor.org/stable/2669401.

Meeden:2003:BSS

[Mee03] Glen Meeden. A Bayesian solution for a statistical auditing problem. Journal of the American Statistical Association, 98 (463):735–740, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045301.

Meiring:2007:OTT

[Mei07] Wendy Meiring. Oscillations and time trends in stratospheric ozone levels. Journal of the American Statistical Association, 102(479):788–802, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meng:2000:MDD

[Men00] Xiao-Li Meng. Missing data: Dial M for ??? Journal of the American Statistical Association, 95(452):1325–1330, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669781.

Mendelssohn:2009:CST

[Men09] Roy Mendelssohn. Comment: a spatio-temporal model for mean, anomaly, and trend fields of North Atlantic sea surface temperature. Journal of the American Statistical Association, 104(485):20–22, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LS09b].

Merkouris:2004:CIR

[Mer04] Takis Merkouris. Combining independent regression estimators from multiple surveys. Journal of the American Statistical Association, 99(468):1131–1139, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s24/p1131;   http://www.jstor.org/stable/27590491.

Meyer:2001:BRB

[Mey01] Mary C. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1140–1141, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670266.

Midthune:2005:MRD

[MFCF05] Douglas N. Midthune, Michael P. Fay, Limin X. Clegg, and Eric J. Feuer. Modeling reporting delays and reporting corrections in cancer registry data. Journal of the American Statistical Association, 100(469):61–70, March 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n469/s7/p61;   http:/
/www.jstor.org/stable/27590519.

Maitra:2000:BTC

[MG00] Ranjan Maitra and Rao P. Gullapalli. A Bayesian time-course model for functional magnetic resonance imaging data: Comment. Journal of the American Statistical Association, 95 (451):707–708, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669448.

Mandel:2007:ETE

[MGG^{+}07] Micha Mandel, Susan A. Gauthier, Charles R. G. Guttmann, Howard L. Weiner, and Rebecca A. Betensky. Estimating time to event from longitudinal categorical data. Journal of the American Statistical Association, 102(480):1254–1266, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2005:LES

[MGT05] Yanyuan Ma, Marc G. Genton, and Anastasios A. Tsiatis. Locally efficient semiparametric estimators for generalized skewelliptical distributions. Journal of the American Statistical Association, 100(471):980–989, September 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590628.

McLachlan:2002:NNV

[MH02] Geoffrey J. McLachlan and Karyn L. Hamaty. [nearestneighbor variance estimation (NNVE): Robust covariance estimation via nearest-neighbor cleaning]: Comment. Journal of the American Statistical Association, 97(460):1009– 1011, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085823.

Marchev:2004:GEV

[MH04] Dobrin Marchev and James P. Hobert. Geometric ergodicity of van Dyk and Meng’s algorithm for the multivariate Student’s t model. Journal of the American Statistical Association, 99(465):228–238, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s22/p228;   http:
//www.jstor.org/stable/27590368.

Mu:2007:PTT

[MH07] Yunming Mu and Xuming He. Power transformation toward a linear regression quantile. Journal of the American Statistical Association, 102(477):269–279, March 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Micheas:2006:BRB

[Mic06] Athanasios Micheas. Book review: Constrained Statistical Inference: Inequality, Order, and Shape Restrictions, by Mervyn J. Silvapulle; Pranab K. Sen. Journal of the American Statistical Association, 101(473):391–392, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047470.

Morales:2006:BMA

[MICR06] Knashawn H. Morales, Joseph G. Ibrahim, Chien-Jen Chen, and Louise M. Ryan. Bayesian model averaging with applications to benchmark dose estimation for arsenic in drinking water. Journal of the American Statistical Association, 101 (473):9–17, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047436.

Miller:2006:BRB

[Mil06] Webb Miller. Book review: An Introduction to Bioinformatics Algorithms, by Neil C. Jones; Pavel A. Pevzner. Journal of the American Statistical Association, 101(474):855, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590756.

Miller:2007:CSL

[Mil07] Robert N. Miller. Comment: Subjective likelihood for the assessment of trends in the ocean’s mixed-layer depth. Journal of the American Statistical Association, 102(479):780–781, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mira:2005:BRB

[Mir05] Antonietta Mira. Book review: Highly Structured Stochastic Systems, by Peter J. Green; Nils Lid Hjort; Sylvia Richardson. Journal of the American Statistical Association, 100(469):351–352, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s41/p351;   http://www.
jstor.org/stable/27590547.

Miyata:2004:FEL

[Miy04] Yoichi Miyata. Fully exponential Laplace approximations using asymptotic modes. Journal of the American Statistical Association, 99(468):1037–1049, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s16/p1037;   http://www.jstor.org/stable/27590483.

Muthen:2003:PSA

[MJB03] Bengt Muthén, Booil Jo, and C. Hendricks Brown. Principal stratification approach to broken randomized experiments: a case study of school choice vouchers in New York City [with comment]. Journal of the American Statistical Association, 98 (462):311–314, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045239.

Mittelhammer:2002:CBE

[MJvAC02] Ron Mittelhammer, George Judge, Marco van Akkeren, and N. Scott Cardell. Coordinate based empirical likelihood-like estimation in ill-conditioned inverse problems. Journal of the American Statistical Association, 97(460):1108–1121, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n460/s16/p1108;   http://www.jstor.org/stable/3085835.

MLan:2006:BSS

[MJW06] Cyr Emile M’Lan, Lawrence Joseph, and David B. Wolfson. Bayesian sample size determination for case-control studies. Journal of the American Statistical Association, 101(474): 760–772, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590733.

Mark:2006:SIN

[MK06] Steven D. Mark and Hormuzd A. Katki. Specifying and implementing nonparametric and semiparametric survival estimators in two-stage (nested) cohort studies with missing case data. Journal of the American Statistical Association, 101 (474):460–471, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590709.

Meier:2004:PKM

[MKCX04] Paul Meier, Theodore Karrison, Rick Chappell, and Hui Xie. The price of Kaplan–Meier. Journal of the American Statistical Association, 99(467):890–896, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s40/p890;   http://www.jstor.org/stable/27590457.

Meyer:2002:PVS

[ML02] Mary C. Meyer and Purushottam W. Laud. Predictive variable selection in generalized linear models. Journal of the American Statistical Association, 97(459):859–871, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s17/p859;   http://www.jstor.org/stable/3085727.

Mizera:2004:RLS

[MM04] Ivan Mizera and Christine H. Müller. Rejoinder: Locationscale depth. Journal of the American Statistical Association, 99(468):981–989, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s11/p981.

Meinshausen:2009:VHD

[MMB09] Nicolai Meinshausen, Lukas Meier, and Peter Bühlmann. pvalues for high-dimensional regression. Journal of the American Statistical Association, 104(488):1671–1681, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maity:2007:EEP

[MMC07] Arnab Maity, Yanyuan Ma, and Raymond J. Carroll. Efficient estimation of population-level summaries in general semiparametric regression models. Journal of the American Statistical Association, 102(477):123–139, March 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mizera:2004:LSD

[MMH^{+}04] Ivan Mizera, Christine H. Müller, Mia Hubert, Peter J. Rousseeuw, Karlien Vanden Branden, Robert Serfling, Xuming He, Stephen Portnoy, David Eppstein, and Peter Mc-Cullagh. Location-scale depth [with comments, rejoinder]. Journal of the American Statistical Association, 99 (468):949–966, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s5/p949;   http://www.
jstor.org/stable/27590478.

Miglioretti:2002:CIA

[MMZ02] Diana L. Miglioretti, Colin McCulloch, and Scott L. Zeger. Combining images across multiple subjects: a study of direct cortical electrical interference. Journal of the American Statistical Association, 97(457):125–135, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085764;   reqidx=/cw/asa/01621459/v97n457/
s12/p125.

Moeschberger:2003:BRB

[Moe03a] Melvin L. Moeschberger. Book review: Statistical Methods for the Analysis of Repeated Measurements, by Charles S. Davis. Journal of the American Statistical Association, 98(461):248–249, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045212.

Moeschberger:2003:SMA

[Moe03b] Melvin L. Moeschberger. Statistical methods for the analysis of repeated measurements. Journal of the American Statistical Association, 98(461):248–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s25/p248.

Maydeu-Olivares:2005:LFI

[MOJ05] Albert Maydeu-Olivares and Harry Joe. Limited- and fullinformation estimation and goodness-of-fit testing in 2^{n} contingency tables: a unified framework. Journal of the American Statistical Association, 100(471):1009–1020, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590631.

Molenberghs:2005:BRB

[Mol05] Geert Molenberghs. Book review: Applied Longitudinal Analysis, by Garrett M. Fitzmaurice; Nan M. Laird; James H. Ware. Journal of the American Statistical Association, 100(470):709–710, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s38/p709;   http://www.
jstor.org/stable/27590597.

Morgan:2003:BRB

[Mor03] J. P. Morgan. Book review: A First Course in Linear Model Theory, by Nalini Ravishanker; Dipak K. Dey. Journal of the American Statistical Association, 98(463):766–767, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045306.

Morris:2004:BRB

[Mor04] Max D. Morris. Book review: The Design and Analysis of Computer Experiments, by Thomas J. Santner; Brian J. Williams; William I. Notz. Journal of the American Statistical Association, 99(468):1203–1204, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s33/p1203;   http://www.jstor.org/stable/27590500.

Morgan:2005:BRB

[Mor05] J. P. Morgan. Book review: Association Schemes: Designed Experiments, Algebra and Combinatorics, by Rosemary Bailey. Journal of the American Statistical Association, 100 (471):1092–1093, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590639.

Morgan:2007:OIB

[Mor07] J. P. Morgan. Optimal incomplete block designs. Journal of the American Statistical Association, 102(478):655–663, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moulin:2001:RWA

[Mou01] Pierre Moulin. Regularization of wavelet approximations: Discussion. Journal of the American Statistical Association, 96 (455):959–960, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670239.

Merl:2008:AES

[MPE08] Daniel Merl, Raquel Prado, and Ananías A. Escalante. Assessing the effect of selection at the amino acid level in malaria antigen sequences through Bayesian generalized linear models. Journal of the American Statistical Association, 103(484): 1496–1507, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moodie:2009:ERM

[MPK09] Erica E. M. Moodie, Robert W. Platt, and Michael S. Kramer. Estimating response-maximized decision rules with applications to breastfeeding. Journal of the American Statistical Association, 104(485):155–165, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maiti:2004:BRBb

[MPL^{+}04] Tapabrata Maiti, Mario Peruggia, Robert Lund, Mary Kathryn Cowles, Michael D. Larsen, Paul A. L. Embrechts, and Christine M. Anderson-Cook. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 99(467):905– 908, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590469.

Muller:2004:OSS

[MPRR04] Peter Müller, Giovanni Parmigiani, Christian Robert, and Judith Rousseau. Optimal sample size for multiple testing: The case of gene expression microarrays. Journal of the American Statistical Association, 99(468):990–1001, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s12/p990;   http://www.jstor.org/stable/27590479.

Mehta:2000:EMC

[MPS00] Cyrus R. Mehta, Nitin R. Patel, and Pralay Senchaudhuri. Efficient Monte Carlo methods for conditional logistic regression. Journal of the American Statistical Association, 95(449):99–108, March 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/Mehta.
htm;   http://www.jstor.org/stable/2669530.

Montanari:2005:NMC

[MR05] Giorgio E. Montanari and M. Giovanna Ranalli. Nonparametric model calibration estimation in survey sampling. Journal of the American Statistical Association, 100(472):1429– 1442, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590683.

Muller:2001:BMD

[MRID01] Peter Müller, Gary L. Rosner, Lurdes Y. T. Inoue, and Mark W. Dewhirst. A Bayesian model for detecting acute change in nonlinear profiles. Journal of the American Statistical Association, 96(456):1215–1222, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085884;   reqidx=/cw/asa/01621459/v96n456/
s8/p1215.

Mach:2006:OEO

[MRŞK06] Lenka Mach, Philip T. Reiss, and Ioana Şchiopu-Kratina. Optimizing the expected overlap of survey samples via the northwest corner rule. Journal of the American Statistical Association, 101(476):1671–1679, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27639781.

Mancini:2005:OCU

[MRT05] Loriano Mancini, Elvezio Ronchetti, and Fabio Trojani. Optimal conditionally unbiased bounded-influence inference in dynamic location and scale models. Journal of the American Statistical Association, 100(470):628–641, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s24/
p628;   http://www.jstor.org/stable/27590583.

MacEachern:2007:RLC

[MRW07] Steven N. MacEachern, Youlan Rao, and Chunjie Wu. A robust-likelihood cumulative sum chart. Journal of the American Statistical Association, 102(480):1440–1447, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Machado:2005:QC

[MS05] José A. F. Machado and J. M. C. Santos Silva. Quantiles for counts. Journal of the American Statistical Association, 100 (472):1226–1237, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590667.

Muller:2004:OBD

[MSD04] Peter Müller, Bruno Sansó, and Maria De Iorio. Optimal Bayesian design by inhomogeneous Markov chain simulation. Journal of the American Statistical Association, 99(467):788–798, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s26/p788;   http:
//www.jstor.org/stable/27590448.

Merrick:2005:MPR

[MSM05] Jason R. W. Merrick, Refik Soyer, and Thomas A. Mazzuchi. Are maintenance practices for railroad tracks effective? Journal of the American Statistical Association, 100(469):17–25, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s3/p17;   http://www.
jstor.org/stable/27590515.

Muller:2006:FVP

[MSY06] Hans-Georg Müller, Ulrich Stadtmüller, and Fang Yao. Functional variance processes. Journal of the American Statistical Association, 101(475):1007–1018, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590778.

Maller:2002:EET

[MSZ02] Ross A. Maller, Liuquan Sun, and Xian Zhou. Estimating the expected total number of events in a process. Journal of the American Statistical Association, 97(458):577–589, June 2002.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s19/p577;   http://www.jstor.org/stable/3085673.

Marron:2007:DWD

[MTA07] J. S. Marron, Michael J. Todd, and Jeongyoun Ahn. Distanceweighted discrimination. Journal of the American Statistical Association, 102(480):1267–1271, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miller:2001:MMA

[MTR^{+}01] Michael E. Miller, Thomas R. Ten Have, Beth A. Reboussin, Kurt K. Lohman, and W. Jack Rejeski. A marginal model for analyzing discrete outcomes from longitudinal surveys with outcomes subject to multiple-cause nonresponse. Journal of the American Statistical Association, 96(455):844–857, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670225;   reqidx=/cw/
asa/01621459/v96n455/s7/p844.

MTW:2000:BRBa

[MTW00a] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(449):349, March 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669587.

MTW:2000:BRBb

[MTW00b] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(449):349, March 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669588.

MTW:2000:BRBc

[MTW00c] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(449):349, March 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669589.

MTW:2000:BRBd

[MTW00d] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(449):349–350, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669592.

MTW:2000:BRBe

[MTW00e] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(449):350, March 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669594.

MTW:2000:BRBf

[MTW00f] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(450):689, June 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669433.

MTW:2000:BRBg

[MTW00g] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(450):689, June 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669435.

MTW:2000:BRBh

[MTW00h] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(450):690, June 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669441.

MTW:2000:BRBi

[MTW00i] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(450):690, June 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669442.

MTW:2000:BRBj

[MTW00j] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1019, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669506.

MTW:2000:BRBk

[MTW00k] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1019–1020, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669507.

MTW:2000:BRBl

[MTW00l] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1020, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669509.

MTW:2000:BRBm

[MTW00m] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1381, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669807.

MTW:2000:BRBn

[MTW00n] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1381, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669808.

MTW:2000:BRBo

[MTW00o] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1382, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669811.

MTW:2000:BRBp

[MTW00p] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1382, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669813.

MTW:2000:BRBq

[MTW00q] MTW. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1382, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669815.

Muller:2005:BRB

[Mül05] Peter Müller. Book review: Applied Bayesian Modelling, by Peter Congdon. Journal of the American Statistical Association, 100(469):355–356, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s48/p356;   http://www.
jstor.org/stable/27590554.

Munk:2004:BRB

[Mun04] Axel Munk. Book review: Testing Statistical Hypotheses of Equivalence, by Stefan Wellek. Journal of the American Statistical Association, 99(465):293, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s29/p293;   http:
//www.jstor.org/stable/27590375.

Munnich:2008:BRB

[Mün08] Ralf Münnich. Book review: Sampling Algorithms. Journal of the American Statistical Association, 103(481):428–429, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Murphy:2000:PL

[MV00] S. A. Murphy and A. W. Van Der Vaart. On profile likelihood. Journal of the American Statistical Association, 95 (450):449–465, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/2000/June/Murphy.htm;   http:
//www.jstor.org/stable/2669386.

Morris:2003:R

[MVBC03a] Jeffrey S. Morris, Marina Vannucci, Philip J. Brown, and Raymond J. Carroll. Rejoinder. Journal of the American Statistical Association, 98(463):591–??, September 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s12/p591.

Morris:2003:WBNa

[MVBC03b] Jeffrey S. Morris, Marina Vannucci, Philip J. Brown, and Raymond J. Carroll. Wavelet-based nonparametric modeling of hierarchical functions in colon carcinogenesis. Journal of the American Statistical Association, 98 (463):573–583, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s8/p573;   http://
www.jstor.org/stable/30045283.

Morris:2003:WBNb

[MVBC03c] Jeffrey S. Morris, Marina Vannucci, Philip J. Brown, and Raymond J. Carroll. Wavelet-based nonparametric modeling of hierarchical functions in colon carcinogenesis [with rejoinder]. Journal of the American Statistical Association, 98(463):591– 597, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045287.

Murphy:2001:MMM

[MvdLRC01] S. A. Murphy, M. J. van der Laan, J. M. Robins, and Conduct Problems Prevention Research Group. Marginal mean models for dynamic regimes. Journal of the American Statistical Association, 96(456):1410–1423, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085909;   reqidx=/cw/asa/01621459/v96n456/
s22/p1410.

Murphy:2000:PLR

[MvdV00] S. A. Murphy and A. W. van der Vaart. On profile likelihood: Rejoinder. Journal of the American Statistical Association, 95(450):482–485, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669392.

Moustaki:2006:BIR

[MVF06] Irini Moustaki and Maria-Pia Victoria-Feser. Boundedinfluence robust estimation in generalized linear latent variable models. Journal of the American Statistical Association, 101(474):644–653, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590724.

Marazzi:2009:RRT

[MVY09] Alfio Marazzi, Ana J. Villar, and Victor J. Yohai. Robust response transformations based on optimal prediction. Journal of the American Statistical Association, 104(485):360–370, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:2005:ORM

[MW05] Samuel Müller and A. H. Welsh. Outlier robust model selection in linear regression. Journal of the American Statistical Association, 100(472):1297–1310, December 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590673.

Micheas:2009:BHN

[MW09] Athanasios C. Micheas and Christopher K. Wikle. A Bayesian hierarchical nonoverlapping random disc growth model. Journal of the American Statistical Association, 104(485):274–283, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maiti:2004:BRBa

[MWE^{+}04] Tapabrata Maiti, Jon Wakefield, Paul A. L. Embrechts, Vincent J. Carey, Mario Peruggia, Robert Gentleman, Phil Spector, Roger Ghanem, Mary Kathryn Cowles, David Hunter, and Russell V. Lenth. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 99(466):570–573, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590428.

Morris:2001:PNM

[MWL^{+}01] Jeffrey S. Morris, Naisyin Wang, Joanne R. Lupton, Robert S. Chapkin, Nancy D. Turner, Mee Young Hong, and Raymond J. Carroll. Parametric and nonparametric methods for understanding the relationship between carcinogen-induced DNA adduct levels in distal and proximal regions of the colon. Journal of the American Statistical Association, 96 (455):816–826, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670222;   reqidx=/cw/
asa/01621459/v96n455/s4/p816.

Maronna:2002:IRA

[MY02] Ricardo A. Maronna and Victor J. Yohai. [inconsistency of resampling algorithms for high-breakdown regression estimators and a new algorithm]: Comment. Journal of the American Statistical Association, 97(457):154–155, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085769.

Ma:2008:CRM

[MY08a] Yanyuan Ma and Guosheng Yin. Cure rate model with mismeasured covariates under transformation. Journal of the American Statistical Association, 103(482):743–756, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:2008:FAM

[MY08b] Hans-Georg Müller and Fang Yao. Functional additive models. Journal of the American Statistical Association, 103(484): 1534–1544, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Myers:2007:BRB

[Mye07] Kary L. Myers. Book review: A Kalman Filter Primer. Journal of the American Statistical Association, 102(477):384, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2008:PCL

[MZ08] Ping Ma and Wenxuan Zhong. Penalized clustering of largescale functional data with multiple covariates. Journal of the American Statistical Association, 103(482):625–636, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nadler:2009:D

[Nad09] Boaz Nadler. Discussion. Journal of the American Statistical Association, 104(486):694–697, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nandram:2002:HBN

[NC02] Balgobin Nandram and Jai Won Choi. Hierarchical Bayesian nonresponse models for binary data from small areas with uncertainty about ignorability. Journal of the American Statistical Association, 97(458):381–388, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s2/p381;   http://www.jstor.org/stable/3085651.

Nordman:2008:PIE

[NC08] Daniel J. Nordman and Petruţa C. Caragea. Point and interval estimation of variogram models using spatial empirical likelihood. Journal of the American Statistical Association, 103 (481):350–361, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nott:2001:SCD

[NDKW01] David J. Nott, William T. M. Dunsmuir, Robert Kohn, and Frank Woodcock. Statistical correction of a deterministic numerical weather prediction model. Journal of the American Statistical Association, 96(455):794–804, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670220;   reqidx=/cw/asa/01621459/v96n455/s2/p794.

Neath:2006:BRB

[Nea06] Andrew A. Neath. Book review: Testing Statistical Hypotheses, by E. L. Lehmann; Joseph P. Romano. Journal of the American Statistical Association, 101(474):847–848, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590746.

Neddermeyer:2009:CEN

[Ned09] Jan C. Neddermeyer. Computationally efficient nonparametric importance sampling. Journal of the American Statistical Association, 104(486):788–802, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nembhard:2004:BRB

[Nem04] Harriet Black Nembhard. Book review: Statistical Process Adjustment Methods for Quality Control, by Enrique del Castillo. Journal of the American Statistical Association, 99(466):567–568, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s38/p567;   http:
//www.jstor.org/stable/27590424.

Neslehova:2005:BRB

[Nes05] Johanna Neslehova. Book review: Weibull Models, by Prabhakar D. N. Murthy; Min Xie; Renyan Jiang. Journal of the American Statistical Association, 100(471):1094, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590641.

Neslehova:2007:BRB

[Nes07] Johanna Neslehova. Book review: Fractal-Based Point Processes. Journal of the American Statistical Association, 102 (477):382–383, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nettleton:2006:D

[Net06] Dan Nettleton. Discussion. Journal of the American Statistical Association, 101(476):1334–1338, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nettleton:2009:TSM

[Net09] Dan Nettleton. Testing for the supremacy of a multinomial cell probability. Journal of the American Statistical Association, 104(487):1052–1059, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Neuhaus:2006:BRB

[Neu06] John M. Neuhaus. Book review: Fixed Effects Regression Methods for Longitudinal Data Using SAS, by Paul D. Allison. Journal of the American Statistical Association, 101(475): 1308, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590804.

Newton:2002:DCG

[New02] Michael A. Newton. Discovering combinations of genomic aberrations associated with cancer. Journal of the American Statistical Association, 97(460):931–942, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s1/p931;   http://www.jstor.org/stable/3085815.

Nair:2000:SAM

[NHS00] Vijay Nair, Mark Hansen, and Jan Shi. Statistics in advanced manufacturing. Journal of the American Statistical Association, 95(451):1002–1005, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669486.

Natarajan:2000:RBM

[NK00] Ranjini Natarajan and Robert E. Kass. Reference Bayesian methods for generalized linear mixed models. Journal of the American Statistical Association, 95(449):227–237, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/Natarajan.htm;   http://www.jstor.org/stable/
2669540.

Noh:2007:RMI

[NL07] Maengseok Noh and Youngjo Lee. Robust modeling for inference from generalized linear model classes. Journal of the American Statistical Association, 102(479):1059–1072, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nan:2006:PCC

[NLLH06] Bin Nan, Xihong Lin, Lynda D. Lisabeth, and Siobán D. Harlow. Piecewise constant cross-ratio estimation for association of age at a marker event and age at menopause. Journal of the American Statistical Association, 101(473):65–77, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047441.

Nygren:2006:LSD

[NN06] Kjell Nygren and Lan Ma Nygren. Likelihood subgradient densities. Journal of the American Statistical Association, 101 (475):1144–1156, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590791.

Norton:2009:IAS

[NN09] Jonathan D. Norton and Xu-Feng Niu. Intrinsically autoregressive spatiotemporal models with application to aggregated birth outcomes. Journal of the American Statistical Association, 104(486):638–649, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nolan:2003:BRB

[Nol03a] Deborah Nolan. Book review: Statistics in the 21st Century, by Adrian E. Raftery; Martin A. Tanner; Martin T. Wells. Journal of the American Statistical Association, 98 (462):491, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045258.

Nolan:2003:SC

[Nol03b] Deborah Nolan. Statistics in the 21st Century. Journal of the American Statistical Association, 98(462):491–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n462/s22/p491.

Notz:2003:BRB

[Not03a] William I. Notz. Book review: Topics in Optimal Design, by Erkki P. Liski; Nripes Mandal; Kirti R. Shah; Bikas Siniha. Journal of the American Statistical Association, 98 (461):249–250, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045214.

Notz:2003:TOD

[Not03b] William I. Notz. Topics in optimal design. Journal of the American Statistical Association, 98(461):249–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s27/p249.

Novikov:2005:BRB

[Nov05] Alexander Novikov. Book review: Lévy Processes and Stochastic Calculus, by David Applebaum. Journal of the American Statistical Association, 100(472):1461, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590691.

Newman:2002:MSC

[NR02] Ken B. Newman and John Rice. Modeling the survival of Chinook salmon smolts outmigrating through the Lower Sacramento River system. Journal of the American Statistical Association, 97(460):983–993, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s6/p983;   http://www.jstor.org/stable/3085820.

Nowicki:2001:EPS

[NS01] Krzysztof Nowicki and Tom A. B. Snijders. Estimation and prediction for stochastic blockstructures. Journal of the American Statistical Association, 96(455):1077–1087, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670253;   reqidx=/cw/asa/01621459/v96n455/
s25/p1077.

Novick:2002:CSE

[NS02] Steven J. Novick and Leonard A. Stefanski. Corrected score estimation via complex variable simulation extrapolation. Journal of the American Statistical Association, 97(458):472–481, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s9/p472;   http://
www.jstor.org/stable/3085663.

Nandram:2000:BAM

[NSP00] Balgobin Nandram, J. Sedransk, and Linda Williams Pickle. Bayesian analysis and mapping of mortality rates for chronic obstructive pulmonary disease. Journal of the American Statistical Association, 95(452):1110–1118, December 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669747.

Naik:2007:EAI

[NST07] Prasad A. Naik, Peide Shi, and Chih-Ling Tsai. Extending the Akaike information criterion to mixture regression models. Journal of the American Statistical Association, 102(477):244– 254, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Naik:2005:CIR

[NT05] Prasad A. Naik and Chih-Ling Tsai. Constrained inverse regression for incorporating prior information. Journal of the American Statistical Association, 100(469):204–211, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s24/
p204;   http://www.jstor.org/stable/27590530.

Ntasin:2007:BRB

[Nta07] Louis N. Ntasin. Book review: Operational Risks: Modeling Analytics. Journal of the American Statistical Association, 102 (480):1475–1476, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nusser:2006:BRB

[Nus06] Sarah M. Nusser. Book review: Survey Methodology, by Robert M. Groves; Floyd J. Fowler, Jr.; Mick P. Couper; James M. Lepkowski; Eleanor Singer; Roger Tourangeau. Journal of the American Statistical Association, 101(475):1310– 1311, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590808.

Neocleous:2006:CCR

[NVP06] Tereza Neocleous, Karlien Vanden Branden, and Stephen Portnoy. Correction to Censored Regression Quantiles by S. Portnoy, 98 (2003), 1001–1012. Journal of the American Statistical Association, 101(474):860–861, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590763. See [Por03].

Nychka:2000:CUA

[Nyc00] Doug Nychka. Challenges in understanding the atmosphere. Journal of the American Statistical Association, 95(451):972– 975, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669479.

Oakes:2000:SA

[Oak00] David Oakes. Survival analysis. Journal of the American Statistical Association, 95(449):282–285, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669547.

Opsomer:2007:MAE

[OBMK07a] Jean D. Opsomer, F. Jay Breidt, Gretchen G. Moisen, and Göran Kauermann. Model-assisted estimation of forest resources with generalized additive models. Journal of the American Statistical Association, 102(478):400–409, June 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Opsomer:2007:RMA

[OBMK07b] Jean D. Opsomer, F. Jay Breidt, Gretchen G. Moisen, and Göran Kauermann. Rejoinder: Model-assisted estimation of forest resources with generalized additive models. Journal of the American Statistical Association, 102(478):415–416, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OSullivan:2006:SMR

[OC06] Finbarr O’Sullivan and Kingshuk Roy Choudhury. A statistical measure of regularity for the study of wind-generated wave field images. Journal of the American Statistical Association, 101(475):1119–1131, September 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590789.

OFallon:2001:SMV

[O’F01] W. Michael O’Fallon. Statistical meetings: View of past, vision of future. Journal of the American Statistical Association, 96(456):1143–1150, December 2001. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085877;   reqidx=/cw/
asa/01621459/v96n456/s1/p1143.

Ogden:2000:BRBa

[Ogd00a] R. T. Ogden. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1007, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669487.

Ogden:2000:BRBb

[Ogd00b] R. T. Ogden. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1007–1008, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669488.

Ogden:2002:BRB

[Ogd02] Todd Ogden. Book review: [untitled]. Journal of the American Statistical Association, 97(457):362–363, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085802.

Olsen:2002:BRB

[Ols02] Chris Olsen. Book review: [untitled]. Journal of the American Statistical Association, 97(458):650, June 2002. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/3085678.

Othus:2009:CSM

[OLT09] Megan Othus, Yi Li, and Ram C. Tiwari. A class of semiparametric mixture cure survival models with dependent censoring. Journal of the American Statistical Association, 104 (487):1241–1250, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OSullivan:2009:NRA

[OMS^{+}09] Finbarr O’Sullivan, Mark Muzi, Alexander M. Spence, David M. Mankoff, Janet N. O’Sullivan, Niall Fitzgerald, George C. Newman, and Kenneth A. Krohn. Nonparametric residue analysis of dynamic PET data with application to cerebral FDG studies in normals. Journal of the American Statistical Association, 104(486):556–571, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ollila:2003:ERC

[OOK03] Esa Ollila, Hannu Oja, and Visa Koivunen. Estimates of regression coefficients based on lift rank covariance matrix. Journal of the American Statistical Association, 98(461):90–98, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s10/p90;   http://
www.jstor.org/stable/30045197.

Oh:2001:BMS

[OR01] Man-Suk Oh and Adrian E. Raftery. Bayesian multidimensional scaling and choice of dimension. Journal of the American Statistical Association, 96(455):1031–1044, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670249;   reqidx=/cw/asa/01621459/v96n455/
s21/p1031.

Ombao:2001:ASA

[ORvSM01] Hernando C. Ombao, Jonathan A. Raz, Rainer von Sachs, and Beth A. Malow. Automatic statistical analysis of bivariate nonstationary time series. in memory of Jonathan A. Raz. Journal of the American Statistical Association, 96(454):543–560, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670296;   reqidx=/cw/
asa/01621459/v96n454/s17/p543.

Osborne:2000:IEO

[OS00] Jason A. Osborne and Thomas A. Severini. Inference for exponential order statistic models based on an integrated likelihood function. Journal of the American Statistical Association, 95 (452):1220–1228, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669762.

Olsen:2001:TPR

[OS01] Maren K. Olsen and Joseph L. Schafer. A two-part randomeffects model for semicontinuous longitudinal data. Journal of the American Statistical Association, 96(454):730–745, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670310;   reqidx=/cw/asa/01621459/v96n454/
s31/p730.

Otter:2006:BRB

[Ott06] Thomas Otter. Book review: Contemporary Bayesian Econometrics and Statistics, by John Geweke. Journal of the American Statistical Association, 101(475):1313–1314, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590813.

Ounpraseuth:2008:BRB

[Oun08] Songthip T. Ounpraseuth. Book review: Gaussian Processes for Machine Learning. Journal of the American Statistical Association, 103(481):429, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ombao:2005:SAM

[OvSG05] Hernando Ombao, Rainer von Sachs, and Wensheng Guo. SLEX analysis of multivariate nonstationary time series. Journal of the American Statistical Association, 100(470):519–531, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/
s15/p519;   http://www.jstor.org/stable/27590574.

Owen:2000:SEI

[OZ00] Art Owen and Yi Zhou. Safe and effective importance sampling. Journal of the American Statistical Association, 95(449):135–143, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/Owen.htm;   http:/
/www.jstor.org/stable/2669533.

OMalley:2008:DLC

[OZ08] A. James O’Malley and Alan M. Zaslavsky. Domain-level covariance analysis for multilevel survey data with structured nonresponse. Journal of the American Statistical Association, 103(484):1405–1418, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ozturk:2002:RRR

[Ozt02] Omer Ozturk. Rank regression in ranked-set samples. Journal of the American Statistical Association, 97(460): 1180–1191, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s21/p1180;   http:/
/www.jstor.org/stable/3085840.

Ozturk:2005:BRB

[Ozt05] Omer Ozturk. Book review: Ranked Set Sampling: Theory and Applications, by Zehua Chen; Zhidong Bai; Bimal K. Sinha. Journal of the American Statistical Association, 100 (471):1093–1094, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590640.

Paciorek:2008:CBB

[Pac08] Christopher J. Paciorek. Comment: Bivariate binomial spatial modeling of Loa loa prevalence in tropical Africa. Journal of the American Statistical Association, 103(481):37–40, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CDR08a, CDR08b, Dun08, Tho08].

Peruggia:2003:BRBb

[PACP^{+}03] Mario Peruggia, Christine M. Anderson-Cook, Mario Peruggia, Alan F. Karr, and Russell V. Lenth. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 98(462):499–500, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045270.

Padgett:2000:BRB

[Pad00] William J. Padgett. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1374–1375, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669797.

Paindaveine:2009:MRT

[Pai09] Davy Paindaveine. On multivariate runs tests for randomness. Journal of the American Statistical Association, 104 (488):1525–1538, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Palesch:2008:BRB

[Pal08] Yuko Palesch. Book review: Adaptive Design Methods in Clinical Trials. Journal of the American Statistical Association, 103 (481):431, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paruolo:2000:BRB

[Par00] Paolo Paruolo. Book review: [untitled]. Journal of the American Statistical Association, 95(450):683–685, June 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669425.

Parzen:2002:BRB

[Par02a] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1207, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085852.

Parzen:2002:FTS

[Par02b] Emanuel Parzen. Foundations of time series analysis and prediction theory. Journal of the American Statistical Association, 97(460):1207–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s33/p1207.

Paruggia:2006:BRB

[Par06] Mario Paruggia. Book review: Sensitivity Analysis in Practice: A Guide to Assessing Scientific Models. Journal of the American Statistical Association, 101(473):398–399, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2007:BRB

[Par07] Do-Hwan Park. Book review: The Statistical Analysis of Interval-Censored Failure Time Data. Journal of the American Statistical Association, 102(480):1473–1474, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Parmigiani:2003:CCS

[PAS^{+}03] Giovanni Parmigiani, Heidi W. Ashih, Gregory P. Samsa, Pamela W. Duncan, Sue Min Lai, and David B. Matchar. Cross-calibration of stroke disability measures: Bayesian analysis of longitudinal ordinal categorical data using negative dependence. Journal of the American Statistical Association, 98(462):273–281, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s2/p273;   http://
www.jstor.org/stable/30045235.

Presnell:2004:ITM

[PB04] Brett Presnell and Dennis D. Boos. The IOS test for model misspecification. Journal of the American Statistical Association, 99(465):216–227, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s21/p216;   http:
//www.jstor.org/stable/27590367.

Putt:2000:RAC

[PC00] Mary E. Putt and Vernon M. Chinchilli. A robust analysis of crossover designs using multisample generalized L-statistics. Journal of the American Statistical Association, 95(452):1256– 1262, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669765.

Peruggia:2006:BRBb

[PC06] Mario Peruggia and Joseph Cavanaugh. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 101(474):859, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590762.

Park:2008:BL

[PC08] Trevor Park and George Casella. The Bayesian lasso. Journal of the American Statistical Association, 103(482):681–686, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2006:FCA

[PCK06] Yousung Park, Jai Won Choi, and Hee-Young Kim. Forecasting cause-age specific mortality using two random processes. Journal of the American Statistical Association, 101 (474):472–483, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590710.

Patterson:2002:LCAa

[PDG02a] Blossom H. Patterson, C. Mitchell Dayton, and Barry I. Graubard. Latent class analysis of complex sample survey data: Application to dietary data. Journal of the American Statistical Association, 97(459):721–729, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s7/p721;   http://www.jstor.org/stable/3085708.

Patterson:2002:LCAb

[PDG02b] Blossom H. Patterson, C. Mitchell Dayton, and Barry I. Graubard. [latent class analysis of complex sample survey data: Application to dietary data]: Rejoinder. Journal of the American Statistical Association, 97(459):738–741, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085713.

Peddada:2007:IHC

[PDK07] Shyamal D. Peddada, Gregg E. Dinse, and Grace E. Kissling. Incorporating historical control data when comparing tumor incidence rates. Journal of the American Statistical Association, 102(480):1212–1220, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

PE:2000:BRBa

[PE00a] PE. Book review: [untitled]. Journal of the American Statistical Association, 95(449):349, March 2000. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669590.

PE:2000:BRBb

[PE00b] PE. Book review: [untitled]. Journal of the American Statistical Association, 95(449):349, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669591.

PE:2000:BRBc

[PE00c] PE. Book review: [untitled]. Journal of the American Statistical Association, 95(449):350, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669593.

PE:2000:BRBd

[PE00d] PE. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1020, September 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669508.

PE:2000:BRBe

[PE00e] PE. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1020, September 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669511.

Pearl:2000:CIC

[Pea00] Judea Pearl. Causal inference without counterfactuals: Comment. Journal of the American Statistical Association, 95 (450):428–431, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669380.

Pea:2002:BRB

[Pea02] Edsel A. Pea. Book review: [untitled]. Journal of the American Statistical Association, 97(458):652–653, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085681.

Peruggia:2004:BRB

[PEACL04] Mario Peruggia, Paul A. L. Embrechts, Christine M. Anderson-Cook, and Russell V. Lenth. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 99(465):300–301, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590383.

Pedroni:2001:BRB

[Ped01] Peter Pedroni. Book review: [untitled]. Journal of the American Statistical Association, 96(453):345–346, March 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670378.

Pena:2000:CMS

[Pen00a] Edsel A. Pena. Comment: Model selection and semiparametric inference for bivariate failure-time data. Journal of the American Statistical Association, 95(449):73–??, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pena:2000:MSS

[Pen00b] Edsel A. Pena. Model selection and semiparametric inference for bivariate failure-time data: Comment. Journal of the American Statistical Association, 95(449):73–75, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669524.

Pena:2006:BRB

[Peñ06] Edsel A. Peña. Book review: Nonparametric Statistical Methods for Complete and Censored Data, by M. M. Desu; D. Raghavarao. Journal of the American Statistical Association, 101(474):852–853, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590751.

Pepe:2000:ROC

[Pep00] Margaret Sullivan Pepe. Receiver operating characteristic methodology. Journal of the American Statistical Association, 95(449):308–311, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669554.

Peruggia:2002:BRBa

[Per02a] Mario Peruggia. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1215, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085862.

Peruggia:2002:BRBb

[Per02b] Mario Peruggia. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1217, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085870.

Peruggia:2006:BRBa

[Per06] Mario Peruggia. Book review: Sensitivity Analysis in Practice: A Guide to Assessing Scientific Models, by Andrea Saltelli; Stefano Tarantola; Francesca Campolongo; Marco Ratto. Journal of the American Statistical Association, 101 (473):398–399, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047481.

Peruggia:2003:BRBa

[PESJ03] Mario Peruggia, Paul Embrechts, Eric P. Smith, and Sreenivasa Rao Jammalamadaka. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 98 (461):259–260, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045229.

Peng:2006:REA

[PF06] Limin Peng and Jason P. Fine. Rank estimation of accelerated lifetime models with dependent censoring. Journal of the American Statistical Association, 101(475):1085– 1093, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590786.

Posch:2008:UIB

[PF08] Martin Posch and Andreas Futschik. A uniform improvement of Bonferroni-type tests by sequential tests. Journal of the American Statistical Association, 103(481):299–308, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peng:2009:CRQ

[PF09] Limin Peng and Jason P. Fine. Competing risks quantile regression. Journal of the American Statistical Association, 104 (488):1440–1453, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2001:MRM

[PGH01] Eun Sug Park, Peter Guttorp, and Ronald C. Henry. Multivariate receptor modeling for temporally correlated data by using MCMC. Journal of the American Statistical Association, 96(456):1171–1183, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085880;   reqidx=/cw/
asa/01621459/v96n456/s4/p1171.

Preisser:2000:AST

[PGLW00] John S. Preisser, Andrzej T. Galecki, Kurt K. Lohman, and Lynne E. Wagenknecht. Analysis of smoking trends with incomplete longitudinal binary responses. Journal of the American Statistical Association, 95(452):1021–1031, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669739.

Pacifico:2004:FDC

[PGVW04] M. Perone Pacifico, C. Genovese, I. Verdinelli, and L. Wasserman. False discovery control for random fields. Journal of the American Statistical Association, 99(468):1002–1014, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s13/p1002;   http://www.jstor.org/stable/27590480.

Portnoy:2000:RJN

[PH00] Stephen Portnoy and Xuming He. A robust journey in the new millennium. Journal of the American Statistical Association, 95(452):1331–1335, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669782.

Peng:2008:SAQ

[PH08] Limin Peng and Yijian Huang. Survival analysis with quantile regression models. Journal of the American Statistical Association, 103(482):637–649, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Piccinato:2000:ADV

[Pic00] Ludovico Piccinato. Asymptotic distribution of P values in composite null models: Comment. Journal of the American Statistical Association, 95(452):1166–1167, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669755.

Picard:2005:BRB

[Pic05] Richard R. Picard. Book review: Introduction to Rare Event Simulation, by James Antonio Bucklew. Journal of the American Statistical Association, 100(471):1091, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590637.

Park:2001:OLR

[PKC01] You Sung Park, Kee Whan Kim, and Jai Won Choi. One-level rotation design balanced on time in monthly sample and in rotation group. Journal of the American Statistical Association, 96(456):1483–1496, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085915;   reqidx=/cw/
asa/01621459/v96n456/s28/p1483.

Peruggia:2005:BRB

[PLL^{+}05] Mario Peruggia, Russell V. Lenth, R. V. L., Carolyn Pillers Dobler, Robert Lund, Tapabrata Maiti, C. P. D., Robert Gentleman, and R. L. Book review: Telegraphic Reviews. Journal of the American Statistical Association, 100(472):1466– 1469, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590701.

Pavia:2008:EFU

[PLM08] Jose Manuel Pavía, Beatriz Larraz, and Jose Marí Montero. Election forecasts using spatiotemporal models. Journal of the American Statistical Association, 103(483):1050–1059, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pavia-Miralles:2005:FNS

[PM05] Jose Manuel Pavía-Miralles. Forecasts from nonrandom samples: The election night case. Journal of the American Statistical Association, 100(472):1113–1122, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590658.

Paciorek:2009:MAF

[PM09] Christopher J. Paciorek and Jason S. McLachlan. Mapping ancient forests: Bayesian inference for spatio-temporal trends in forest composition using the fossil pollen proxy record. Journal of the American Statistical Association, 104(486):608–622, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2009:TSM

[PMHB09] Byeong U. Park, Enno Mammen, Wolfgang Härdle, and Szymon Borak. Time series modelling with semiparametric factor dynamics. Journal of the American Statistical Association, 104(485):284–298, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pollock:2000:CRM

[Pol00] Kenneth H. Pollock. Capture-recapture models. Journal of the American Statistical Association, 95(449):293–296, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669550.

Politis:2002:SBV

[Pol02] Dimitris N. Politis. [sieve bootstrap with variable-length Markov chains for stationary categorical time series]: Comment. Journal of the American Statistical Association, 97 (458):463–465, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085661.

Pooler:2004:BRB

[Poo04] Penelope S. Pooler. Book review: Statistical Tools for Environmental Quality Measurement, by Michael E. Ginevan; Douglas E. Splitstone. Journal of the American Statistical Association, 99(467):900–901, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s46/p900;   http:
//www.jstor.org/stable/27590462.

Pooler:2006:BRB

[Poo06] Penelope S. Pooler. Book review: Handling Missing Data: Applications to Environmental Analysis, by G. Latini; G. Passerini. Journal of the American Statistical Association, 101(473):400–401, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047485.

Portnoy:2002:IRA

[Por02] Stephen Portnoy. [inconsistency of resampling algorithms for high-breakdown regression estimators and a new algorithm]: Comment. Journal of the American Statistical Association, 97(457):149–150, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085767.

Portnoy:2003:CRQ

[Por03] Stephen Portnoy. Censored regression quantiles. Journal of the American Statistical Association, 98(464):1001–1012, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s25/p1001;   http://www.jstor.org/stable/30045346. See

correction [NVP06].

Pouet:2004:BRB

[Pou04] Christophe Pouet. Book review: Nonparametric Goodnessof-Fit Testing under Gaussian Models, by Yu I. Ingster; Irina A. Suslina. Journal of the American Statistical Association, 99(466):561–562, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s29/p561;   http:
//www.jstor.org/stable/27590415.

Pena:2001:CIU

[PP01] Daniel Peña and Francisco J. Prieto. Cluster identification using projections. Journal of the American Statistical Association, 96(456):1433–1445, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085911;   reqidx=/cw/
asa/01621459/v96n456/s24/p1433.

Paparoditis:2005:BUR

[PP05] Efstathios Paparoditis and Dimitris N. Politis. Bootstrapping unit root tests for autoregressive time series. Journal of the American Statistical Association, 100(470):545–553, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s17/
p545;   http://www.jstor.org/stable/27590576.

Perin:2009:SEE

[PPR09] Jamie Perin, John S. Preisser, and Paul J. Rathouz. Semiparametric efficient estimation for incomplete longitudinal binary data, with application to smoking trends. Journal of the American Statistical Association, 104(488):1373–1384, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Poole:2000:IDS

[PR00] David Poole and Adrian E. Raftery. Inference for deterministic simulation models: The Bayesian melding approach. Journal of the American Statistical Association, 95(452):1244– 1255, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669764.

Pena:2002:PPT

[PR02] Daniel Peña and Julio Rodríguez. A powerful portmanteau test of lack of fit for time series. Journal of the American Statistical Association, 97(458):601–610, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s21/p601;   http://
www.jstor.org/stable/3085675.

Pratelli:2000:BRB

[Pra00] M. Pratelli. Book review: [untitled]. Journal of the American Statistical Association, 95(450):688–689, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669432.

Presnell:2000:BRB

[Pre00] Brett Presnell. Book review: [untitled]. Journal of the American Statistical Association, 95(449):333–334, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669567.

Presnell:2002:BRB

[Pre02] Brett D. Presnell. Book review: [untitled]. Journal of the American Statistical Association, 97(457):355, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085791.

Prewitt:2003:BRB

[Pre03] Kathryn A. Prewitt. Book review: A Distribution-Free Theory of Nonparametric Regression, by Laszlo Gyrfi; Michael Kohler; Adam Krzyzak; Harro Walk. Journal of the American Statistical Association, 98(464):1084, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s36/p1084;   http://www.jstor.org/stable/30045357.

Pena:2005:MPV

[PS05] Daniel Peña and Ismael Sánchez. Multifold predictive validation in ARMAX time series models. Journal of the American Statistical Association, 100(469):135–146, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s18/
p135;   http://www.jstor.org/stable/27590524.

Palacios:2006:NGB

[PS06a] M. Blanca Palacios and Mark F. J. Steel. Non-Gaussian Bayesian geostatistical modeling. Journal of the American Statistical Association, 101(474):604–618, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590721.

Pena:2006:GVL

[PS06b] Edsel A. Peña and Elizabeth H. Slate. Global validation of linear model assumptions. Journal of the American Statistical Association, 101(473):341–354, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047462.

Pratesi:2006:BRB

[PS06c] Monica Pratesi and Nicola Salvati. Book review: Missing Data and Small-Area Estimation: Modern Analytical Equipment for the Survey Statistician, by Nicholas T. Longford. Journal of the American Statistical Association, 101(476):1729– 1730, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639800.

Pfeffermann:2007:SAE

[PS07] Danny Pfeffermann and Michail Sverchkov. Small-area estimation under informative probability sampling of areas and within the selected areas. Journal of the American Statistical Association, 102(480):1427–1439, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pena:2001:NER

[PSH01] E. A. Peña, R. L. Strawderman, and M. Hollander. Nonparametric estimation with recurrent event data. Journal of the American Statistical Association, 96(456):1299–1315, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085901;   reqidx=/cw/asa/01621459/v96n456/
s14/p1299.

Peng:2005:STC

[PSW05] Roger D. Peng, Frederic Paik Schoenberg, and James A. Woods. A space–time conditional intensity model for evaluating a wildfire hazard index. Journal of the American Statistical Association, 100(469):26–35, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/

rpsv/cw/asa/01621459/v100n469/s4/p26;   http://www.
jstor.org/stable/27590516.

Pfeffermann:2006:SAE

[PT06] Danny Pfeffermann and Richard Tiller. Small-area estimation with state–space models subject to benchmark constraints. Journal of the American Statistical Association, 101 (476):1387–1397, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639759.

Patton:2007:TFO

[PT07] Andrew J. Patton and Allan Timmermann. Testing forecast optimality under unknown loss. Journal of the American Statistical Association, 102(480):1172–1184, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pesaran:2009:TDA

[PT09] M. Hashem Pesaran and Allan Timmermann. Testing dependence among serially correlated multicategory variables. Journal of the American Statistical Association, 104(485):325–337, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Puig:2003:CAC

[Pui03] Pedro Puig. Characterizing additively closed discrete models by a property of their maximum likelihood estimators, with an application to generalized Hermite distributions. Journal of the American Statistical Association, 98 (463):687–692, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s21/p687;   http://
www.jstor.org/stable/30045296.

Puig:2006:CDD

[PV06] Pedro Puig and Jordi Valero. Count data distributions: Some characterizations with applications. Journal of the American Statistical Association, 101(473):332–340, March 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047461.

Pitt:2005:CST

[PW05] Michael K. Pitt and Stephen G. Walker. Constructing stationary time series models using auxiliary variables with applications. Journal of the American Statistical Association, 100(470):554–564, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s18/p554;   http://www.
jstor.org/stable/27590577.

Peel:2001:FMK

[PWM01] David Peel, William J. Whiten, and Geoffrey J. McLachlan. Fitting mixtures of Kent distributions to aid in joint set identification. Journal of the American Statistical Association, 96(453):56–63, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670340;   reqidx=/cw/
asa/01621459/v96n453/s6/p56.

Peng:2009:PCE

[PWZZ09] Jie Peng, Pei Wang, Nengfeng Zhou, and Ji Zhu. Partial correlation estimation by joint sparse regression models. Journal of the American Statistical Association, 104(486):735–746, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Polonik:2000:CMV

[PY00] Wolfgang Polonik and Qiwei Yao. Conditional minimum volume predictive regions for stochastic processes. Journal of the American Statistical Association, 95(450):509–519, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/2000/June/Polonik.htm;   http://www.jstor.org/
stable/2669395.

Posch:2009:HSF

[PZB09] Martin Posch, Sonja Zehetmayer, and Peter Bauer. Hunting for significance with the false discovery rate. Journal of the American Statistical Association, 104(486):832–840, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Queen:2009:ICF

[QA09] Catriona M. Queen and Casper J. Albers. Intervention and causality: Forecasting traffic flows using a dynamic Bayesian network. Journal of the American Statistical Association, 104 (486):669–681, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2005:BRB

[Qiu05] Peihua Qiu. Book review: The Statistical Evaluation of Medical Tests for Classification and Prediction, by Margaret Sullivan Pepe. Journal of the American Statistical Association, 100(470):705, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s33/p705;   http://www.
jstor.org/stable/27590592.

Qiu:2007:JSE

[Qiu07] Peihua Qiu. Jump surface estimation, edge detection, and image restoration. Journal of the American Statistical Association, 102(478):745–756, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qin:2002:ESD

[QLS02] Jing Qin, Denis Leung, and Jun Shao. Estimation with survey data under nonignorable nonresponse or informative sampling. Journal of the American Statistical Association, 97(457):193–200, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085774;   reqidx=/cw/
asa/01621459/v97n457/s17/p193.

Quinn:2001:EFG

[QM01] Terrance J. Quinn, II and Sally A. Mizroch. Estimation of fetal growth and gestation in bowhead whales: Comment. Journal of the American Statistical Association, 96(455):928– 933, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670234.

Qiu:2007:LSI

[QS07] Peihua Qiu and Jingran Sun. Local smoothing image segmentation for spotted microarray images. Journal of the American Statistical Association, 102(480):1129–1144, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qin:2008:EDR

[QSZ08] Jing Qin, Jun Shao, and Biao Zhang. Efficient and doubly robust imputation for covariate-dependent missing responses. Journal of the American Statistical Association, 103(482):797– 810, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Quale:2006:LEE

[QvdLR06] Christopher M. Quale, Mark J. van der Laan, and James R. Robins. Locally efficient estimation with bivariate rightcensored data. Journal of the American Statistical Association, 101(475):1076–1084, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590785.

Qi:2005:WEP

[QWP05] Lihong Qi, C. Y. Wang, and Ross L. Prentice. Weighted estimators for proportional hazards regression with missing covariates. Journal of the American Statistical Association, 100 (472):1250–1263, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590669.

Qin:2009:ELM

[QZL09] Jing Qin, Biao Zhang, and Denis H. Y. Leung. Empirical likelihood in missing data problems. Journal of the American Statistical Association, 104(488):1492–1503, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rossi:2000:SM

[RA00] Peter E. Rossi and Greg M. Allenby. Statistics and marketing. Journal of the American Statistical Association, 95(450): 635–638, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669407.

Rabinowitz:2002:APHa

[Rab02a]  Daniel Rabinowitz.  Adjusting for population hetero-

geneity and misspecified haplotype frequencies when testing nonparametric null hypotheses in statistical genetics. Journal of the American Statistical Association, 97(459): 742–751, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s8/p742;   http://
www.jstor.org/stable/3085714.

Rabinowitz:2002:APHb

[Rab02b] Daniel Rabinowitz. [adjusting for population heterogeneity and misspecified haplotype frequencies when testing nonparametric null hypotheses in statistical genetics]: Rejoinder. Journal of the American Statistical Association, 97(459):757– 758, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085718.

Racine:2001:BRB

[Rac01] Jeffrey S. Racine. Book review: [untitled]. Journal of the American Statistical Association, 96(453):343–345, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670376.

Raftery:2000:SS

[Raf00] Adrian E. Raftery. Statistics in sociology, 1950–2000. Journal of the American Statistical Association, 95(450):654–661, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669412.

Raghunathan:2000:CNA

[Rag00a] T. E. Raghunathan. Comment: Nonparametric analysis of randomized experiments with missing covariate and outcome data. Journal of the American Statistical Association, 95 (449):85–??, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raghunathan:2000:NAR

[Rag00b] T. E. Raghunathan. Nonparametric analysis of randomized experiments with missing covariate and outcome data: Comment. Journal of the American Statistical Association, 95 (449):85–87, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669527.

Roy:2006:CIM

[RAHT06] Jason Roy, Don Alderson, Joseph W. Hogan, and Karen T. Tashima. Conditional inference methods for incomplete Poisson data with endogenous time-varying covariates: Emergency department use among HIV-infected women. Journal of the American Statistical Association, 101(474):424–434, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590706.

Ramsay:2000:FCV

[Ram00] J. O. Ramsay. Functional components of variation in handwriting. Journal of the American Statistical Association, 95 (449):9–15, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/Ramsay.htm;   http://
www.jstor.org/stable/2669518.

Ramsay:2001:IDF

[Ram01a] J. O. Ramsay. Inference for density families using functional principal component analysis: Comment. Journal of the American Statistical Association, 96(454):538–539, June 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670294.

Ramsay:2001:SNM

[Ram01b] J. O. Ramsay. [semiparametric nonlinear mixed-effects models and their applications]: Comment. Journal of the American Statistical Association, 96(456):1283–1284, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085896.

Ramoni:2005:BRB

[Ram05] Marco F. Ramoni. Book review: Bayesian Artificial Intelligence, by Keven B. Korb; Ann E. Nicholson. Journal of the American Statistical Association, 100(471):1096– 1097, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590644.

Ramsay:2009:C

[Ram09] James O. Ramsay. Comments. Journal of the American Statistical Association, 104(486):700, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Randles:2000:SAI

[Ran00] Ronald H. Randles. A simpler, affine-invariant, multivariate, distribution-free sign test. Journal of the American Statistical Association, 95(452):1263–1268, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669766.

Rao:2000:BRB

[Rao00] P. V. Rao. Book review: [untitled]. Journal of the American Statistical Association, 95(450):681–682, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669422.

Rao:2003:BRB

[Rao03a] Sunil J. Rao. Book review: Regression Modeling Strategies: With Applications to Linear Models, Logistic Regression, and Survival Analysis, by Frank E. Harrell, Jr. Journal of the American Statistical Association, 98(461):257–258, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045225.

Rao:2003:RMS

[Rao03b] Sunil J. Rao. Regression modeling strategies: With applications to linear models, logistic regression, and survival analysis. Journal of the American Statistical Association, 98(461):257–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s38/p257.

Rathbun:2000:BRB

[Rat00] Stephen L. Rathbun. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1010–1011, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669494.

Raz:2000:BTC

[Raz00] Jonathan Raz. A Bayesian time-course model for functional magnetic resonance imaging data: Comment. Journal of the American Statistical Association, 95(451):709–711, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669449.

Rivest:2008:DME

[RBP08] Louis-Paul Rivest, Sophie Baillargeon, and Michael Pierrynowski. A directional model for the estimation of the rotation axes of the ankle joint. Journal of the American Statistical Association, 103(483):1060–1069, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reese:2001:EFGa

[RCGT01a] C. Shane Reese, James A. Calvin, John C. George, and Raymond J. Tarpley. Estimation of fetal growth and gestation in bowhead whales. Journal of the American Statistical Association, 96(455):915–923, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670231;   reqidx=/cw/
asa/01621459/v96n455/s13/p915.

Reese:2001:EFGb

[RCGT01b] C. Shane Reese, James A. Calvin, John C. George, and Raymond J. Tarpley. Estimation of fetal growth and gestation in bowhead whales: Rejoinder. Journal of the American Statistical Association, 96(455):935–938, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670236.

Raftery:2006:VSM

[RD06] Adrian E. Raftery and Nema Dean. Variable selection for model-based clustering. Journal of the American Statistical Association, 101(473):168–178, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30047447.

Rodriguez:2008:NDP

[RDG08] Abel Rodríguez, David B. Dunson, and Alan E. Gelfand. The nested Dirichlet process. Journal of the American Statistical Association, 103(483):1131–1154, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reese:2008:BRB

[Ree08] C. Shane Reese. Book review: An Introduction to Bayesian Analysis: Theory and Methods. Journal of the American Statistical Association, 103(482):883–884, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reid:2000:L

[Rei00a] N. Reid. Likelihood. Journal of the American Statistical Association, 95(452):1335–1340, December 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669783.

Reimers:2000:BRB

[Rei00b] Hans Eggert Reimers. Book review: [untitled]. Journal of the American Statistical Association, 95(450):686–687, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669428.

Reiter:2002:BRB

[Rei02] Jerome P. Reiter. Book review: [untitled]. Journal of the American Statistical Association, 97(457):357–358, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085795.

Reiter:2005:ERI

[Rei05] Jerome P. Reiter. Estimating risks of identification disclosure in microdata. Journal of the American Statistical Association, 100(472):1103–1112, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590657.

Reiter:2008:BRB

[Rei08] Jerome Reiter. Book review: Data Quality and Record Linkage Techniques. Journal of the American Statistical Association, 103(482):881, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roy:2004:TTP

[RFF04] Anindya Roy, Barry Falk, and Wayne A. Fuller. Testing for trend in the presence of autoregressive error. Journal of the American Statistical Association, 99(468): 1082–1091, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s20/p1082;   http://www.
jstor.org/stable/27590487.

Rueda:2009:EPS

[RFP09] Cristina Rueda, Miguel A. Fernández, and Shyamal Das Peddada. Estimation of parameters subject to order restrictions on a circle with application to estimation of phase angles of cell cycle genes. Journal of the American Statistical Association, 104(485):338–347, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rotnitzky:2006:DRE

[RFS06] Andrea Rotnitzky, David Faraggi, and Enrique Schisterman. Doubly robust estimation of the area under the receiveroperating characteristic curve in the presence of verification bias. Journal of the American Statistical Association, 101 (475):1276–1288, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590801.

Robins:2000:CIC

[RG00] James M. Robins and Sander Greenland. Causal inference without counterfactuals: Comment. Journal of the American Statistical Association, 95(450):431–435, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669381.

Rossi:2001:OSU

[RGA01] Peter E. Rossi, Zvi Gilula, and Greg M. Allenby. Overcoming scale usage heterogeneity: a Bayesian hierarchical approach. Journal of the American Statistical Association, 96(453):20–31, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670337;   reqidx=/cw/
asa/01621459/v96n453/s3/p20.

Reilly:2001:PPL

[RGK01] Cavan Reilly, Andrew Gelman, and Jonathan Katz. Poststratification without population level information on the poststratifying variable with application to political polling. Journal of the American Statistical Association, 96(453):1–11, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670335;   reqidx=/cw/asa/01621459/v96n453/s1/p1.

Reich:2007:SAP

[RHC07] Brian J. Reich, James S. Hodges, and Bradley P. Carlin. Spatial analyses of periodontal data using conditionally autoregressive priors having two classes of neighbor relations. Journal of the American Statistical Association, 102(477):44–55, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rheinlaender:2002:BRB

[Rhe02] Thorsten Rheinlaender. Book review: [untitled]. Journal of the American Statistical Association, 97(457):364–365, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085805.

Rheinlander:2003:BRB

[Rhe03a] Thorsten Rheinländer. Book review: Risk Management: Value at Risk and Beyond, by M. A. H. Dempster. Journal of the American Statistical Association, 98(462):494, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045264.

Rheinlander:2003:RMV

[Rhe03b] Thorsten Rheinländer. Risk management: Value at risk and beyond. Journal of the American Statistical Association, 98(462):494–??, June 2003. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s28/p494.

Rheinlander:2007:BRB

[Rhe07] Thorsten Rheinlander. Book review: Introductory Stochastic Analysis for Finance and Insurance. Journal of the American Statistical Association, 102(478):765–766, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rheinlander:2008:BRB

[Rhe08] Thorsten Rheinländer. Book review: Introduction to Stochastic Integration. Journal of the American Statistical Association, 103(483):1324, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rice:2004:EBC

[Ric04] Kenneth M. Rice. Equivalence between conditional and mixture approaches to the Rasch model and matched casecontrol studies, with applications. Journal of the American Statistical Association, 99(466):510–522, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s20/p510;   http://www.jstor.org/stable/27590406.

Rice:2008:EBC

[Ric08] Kenneth Rice. Equivalence between conditional and randomeffects likelihoods for pair-matched case-control studies. Journal of the American Statistical Association, 103(481):385–396, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ridgeway:2007:BRB

[Rid07] Greg Ridgeway. Book review: All of Nonparametric Statistics. Journal of the American Statistical Association, 102(477):389, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Riedi:2006:BRB

[Rie06] Rudolf H. Riedi. Book review: An Introduction to Statistical Signal Processing, by Robert M. Gray; Lee D. Davisson. Journal of the American Statistical Association, 101(475):

1317, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590818.

Rilstone:2002:BRB

[Ril02] Paul Rilstone. Book review: [untitled]. Journal of the American Statistical Association, 97(457):361–362, March 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085800.

Radchenko:2008:VIS

[RJ08] Peter Radchenko and Gareth M. James. Variable inclusion and shrinkage algorithms. Journal of the American Statistical Association, 103(483):1304–1315, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rhodes:2001:SDS

[RK01] William Rhodes and Ryan Kling. [should the DEA’s STRIDE data be used for economic analyses of markets for illegal drugs?]: Comment. Journal of the American Statistical Association, 96(456):1264–1266, December 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/3085890.

Rathouz:2001:LAQ

[RL01] Paul J. Rathouz and Kung-Yee Liang. Locally ancillary quasiscore models for errors-in-covariates. Journal of the American Statistical Association, 96(455):1004–1013, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670246;   reqidx=/cw/asa/01621459/v96n455/
s18/p1004.

Roy:2002:AML

[RL02] Jason Roy and Xihong Lin. Analysis of multivariate longitudinal outcomes with nonignorable dropouts and missing covariates: Changes in methadone treatment practices. Journal of the American Statistical Association, 97(457):40–52, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
3085757;   reqidx=/cw/asa/01621459/v97n457/s5/p40.

Rappold:2007:RSL

[RLL07a] Ana Grohovac Rappold, Michael Lavine, and Susan Lozier. Rejoinder: Subjective likelihood for the assessment of trends in the ocean’s mixed-layer depth. Journal of the American Statistical Association, 102(479):785–787, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rappold:2007:SLA

[RLL07b] Ana Grohovac Rappold, Michael Lavine, and Susan Lozier. Subjective likelihood for the assessment of trends in the ocean’s mixed-layer depth. Journal of the American Statistical Association, 102(479):771–780, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

RLS:2000:BRBa

[RLS00a] RLS. Book review: [untitled]. Journal of the American Statistical Association, 95(450):689, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669434.

RLS:2000:BRBb

[RLS00b] RLS. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1020, September 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669512.

RLS:2000:BRBc

[RLS00c] RLS. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1382, December 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669814.

Rothman:2009:GTL

[RLZ09] Adam J. Rothman, Elizaveta Levina, and Ji Zhu. Generalized thresholding of large covariance matrices. Journal of the American Statistical Association, 104(485):177–186, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ridgeway:2009:DRI

[RM09] Greg Ridgeway and John M. MacDonald. Doubly robust internal benchmarking and false discovery rates for detecting racial bias in police stops. Journal of the American Statistical Association, 104(486):661–668, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rutter:2009:BCM

[RMS09] Carolyn M. Rutter, Diana L. Miglioretti, and James E. Savarino. Bayesian calibration of microsimulation models. Journal of the American Statistical Association, 104(488): 1338–1350, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reiss:2007:FPC

[RO07] Philip T. Reiss and R. Todd Ogden. Functional principal component regression and functional partial least squares. Journal of the American Statistical Association, 102(479):984–996, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robert:2000:BRB

[Rob00] Christian P. Robert. Book review: [untitled]. Journal of the American Statistical Association, 95(449):337–339, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669570.

Robinson:2003:BRB

[Rob03a] Timothy J. Robinson. Book review: Multilevel Analysis: Techniques and Applications, by Joop Hox. Journal of the American Statistical Association, 98(462):496, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30045267.

Robinson:2003:MAT

[Rob03b] Timothy J. Robinson. Multilevel analysis: Techniques and applications. Journal of the American Statistical Association, 98(462):496–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s31/p496.

Robinson:2004:BRB

[Rob04] Timothy J. Robinson. Book review: Optimal Design of Blocked and Split-Plot Experiments, by Peter Goos. Journal of the American Statistical Association, 99(465):296–297, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s32/p296;   http://www.jstor.org/stable/27590378.

Robinson:2006:CQA

[Rob06] P. M. Robinson. Comment: Quantile autoregression. Journal of the American Statistical Association, 101(475):1001–1002, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robinson:2008:BRB

[Rob08] Timothy J. Robinson. Book review: Models for Ecological Data: An Introduction. Journal of the American Statistical Association, 103(483):1326, September 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rocke:2000:BRB

[Roc00] D. M. Rocke. Book review: [untitled]. Journal of the American Statistical Association, 95(449):347–348, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669583.

Rogness:2002:AIS

[Rog02a] Neal Rogness. Applying and interpreting statistics, a comprehensive guide (second edition). Journal of the American Statistical Association, 97(459):926–??, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s33/p926.

Rogness:2002:BRB

[Rog02b] Neal Rogness. Book review: [untitled]. Journal of the American Statistical Association, 97(459):926–927, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085743.

Rohlf:2005:BRB

[Roh05] F. James Rohlf. Book review: Procrustes Problems, by John C. Gower; Garmt B. Dijksterhuis. Journal of the American Statistical Association, 100(471):1097, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590645.

Rolski:2000:BRB

[Rol00] Tomasz Rolski. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1379–1380, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669804.

Rolph:2001:BRB

[Rol01] John E. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 96(453):341–342, March 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670372.

Rolski:2002:BRB

[Rol02] Tomasz Rolski. Book review: [untitled]. Journal of the American Statistical Association, 97(457):360–361, March 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085799.

Ronchetti:2000:BRB

[Ron00] Elvezio Ronchetti. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1008–1009, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669491.

Ronchetti:2005:BRB

[Ron05] Elvezio Ronchetti. Book review: Finite-Sample Econometrics: Advanced Texts in Econometrics, by Aman Ullah. Journal of the American Statistical Association, 100(471):1098– 1099, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590649.

Roos:2002:BRB

[Roo02a] Małgorzata Roos. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1205, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085848.

Roos:2002:RSA

[Roo02b] Małgorzata Roos. Runs and scans with applications. Journal of the American Statistical Association, 97(460): 1205–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s29/p1205.

Rosenbaum:2001:SAT

[Ros01] Paul R. Rosenbaum. Stability in the absence of treatment. Journal of the American Statistical Association, 96 (453):210–219, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670360;   reqidx=/cw/
asa/01621459/v96n453/s16/p210.

Rosenbaum:2002:AET

[Ros02] Paul R. Rosenbaum. Attributing effects to treatment in matched observational studies. Journal of the American Statistical Association, 97(457):183–192, March 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
3085773;   reqidx=/cw/asa/01621459/v97n457/s16/p183.

Rosenthal:2003:AVC

[Ros03] Jeffrey S. Rosenthal. Asymptotic variance and convergence rates of nearly-periodic Markov chain Monte Carlo algorithms. Journal of the American Statistical Association, 98(461):169–177, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s17/p169;   http://
www.jstor.org/stable/30045204.

Rosenberg:2004:BRB

[Ros04] Noah A. Rosenberg. Book review: Probability Models for DNA Sequence Evolution, by Rick Durrett. Journal of the American Statistical Association, 99(466):560–561, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s27/p560;   http://www.jstor.org/stable/27590413.

Rosenbaum:2005:ENI

[Ros05] Paul R. Rosenbaum. Exact, nonparametric inference when doses are measured with random errors. Journal of the American Statistical Association, 100(470):511–518, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s14/
p511;   http://www.jstor.org/stable/27590573.

Rosenberger:2006:BRB

[Ros06] William F. Rosenberger. Book review: Statistical Concepts and Applications in Clinical Medicine, by John Aitchison; Jim W. Kay; Ian J. Lauder. Journal of the American Statistical Association, 101(473):404, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047492.

Rosenbaum:2007:IBU

[Ros07] Paul R. Rosenbaum. Interference between units in randomized experiments. Journal of the American Statistical Association, 102(477):191–200, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roulston:2004:CCP

[Rou04] Mark S. Roulston. Comment: Calibrated probabilistic mesoscale weather field forecasting: The geostatistical output perturbation method. Journal of the American Statistical Association, 99(467):587–588, September 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/s4/
p587.

Rousseau:2006:BRB

[Rou06] Judith Rousseau. Book review: Probability Matching Priors: Higher Order Asymptotics, by Gauri Sankar Datta; Rahul Mukerjee. Journal of the American Statistical Association, 101 (473):391, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047469.

Royall:2000:POMa

[Roy00a] Richard Royall. On the probability of observing misleading statistical evidence. Journal of the American Statistical Association, 95(451):760–768, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669456.

Royall:2000:POMb

[Roy00b] Richard Royall. On the probability of observing misleading statistical evidence: Rejoinder. Journal of the American Statistical Association, 95(451):773–780, September 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669461.

Reiter:2007:MAM

[RR07] Jerome P. Reiter and Trivellore E. Raghunathan. The multiple adaptations of multiple imputation. Journal of the American Statistical Association, 102(480):1462–1471, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2007:MDM

[RRS07] Paul R. Rosenbaum, Richard N. Ross, and Jeffrey H. Silber. Minimum distance matched sampling with fine balance in an observational study of treatment for ovarian cancer. Journal of the American Statistical Association, 102(477):75–83, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robins:2000:PLC

[RRvdL00] James M. Robins, Andrea Rotnitzky, and Mark van der Laan. On profile likelihood: Comment. Journal of the American Statistical Association, 95(450):477–482, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669391.

Rosenbaum:2008:AET

[RS08] Paul R. Rosenbaum and Jeffrey H. Silber. Aberrant effects of treatment. Journal of the American Statistical Association, 103(481):240–247, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2009:ASA

[RS09a] Paul R. Rosenbaum and Jeffrey H. Silber. Amplification of sensitivity analysis in matched observational studies. Journal of the American Statistical Association, 104(488):1398–1405, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2009:SAE

[RS09b] Paul R. Rosenbaum and Jeffrey H. Silber. Sensitivity analysis for equivalence and difference in an observational study of neonatal intensive care units. Journal of the American Statistical Association, 104(486):501–511, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reilly:2002:CIS

[RSH^{+}02] Cavan Reilly, Timothy Schacker, Ashley T. Haase, Steve Wietgrefe, and David Krason. The clustering of infected SIV cells in lymphatic tissue. Journal of the American Statistical Association, 97(460):943–954, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s2/p943;   http://www.jstor.org/stable/3085816.

Ryu:2007:LSO

[RSM^{+}07] Duchwan Ryu, Debajyoti Sinha, Bani Mallick, Stuart R. Lipsitz, and Steven E. Lipshultz. Longitudinal studies with outcome-dependent follow-up. Journal of the American Statistical Association, 102(479):952–961, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosen:2009:LSA

[RSW09] Ori Rosen, David S. Stoffer, and Sally Wood. Local spectral analysis via a Bayesian mixture of smoothing splines. Journal of the American Statistical Association, 104(485):249–262, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rubin:2000:CPS

[RT00] Donald B. Rubin and Neal Thomas. Combining propensity score matching with additional adjustments for prognostic covariates. Journal of the American Statistical Association, 95 (450):573–585, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/2000/June/Rubin.htm;   http:/
/www.jstor.org/stable/2669400.

Raz:2001:IRW

[RTD01] Jonathan Raz, Bruce I. Turetsky, and Linda W. Dickerson. Inference for a random wavelet packet model of single-channel event-related potentials. Journal of the American Statistical Association, 96(454):409–420, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670279;   reqidx=/cw/
asa/01621459/v96n454/s5/p409.

Raftery:2000:I

[RTW00a] Adrian E. Raftery, Martin A. Tanner, and Martin T. Wells. [introduction]. Journal of the American Statistical Association, 95(449):281, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669545.

Raftery:2000:SYV

[RTW00b] Adrian E. Raftery, Martin A. Tanner, and Martin T. Wells. Statistics in the year 2000: Vignettes. Journal of the American Statistical Association, 95(449):281–??, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rubin:2000:CIC

[Rub00] Donald B. Rubin. Causal inference without counterfactuals: Comment. Journal of the American Statistical Association, 95(450):435–438, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669382.

Rubin:2005:CIU

[Rub05] Donald B. Rubin. Causal inference using potential outcomes: Design, modeling, decisions. Journal of the American Statistical Association, 100(469):322–331, March 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n469/s35/p322;   http:
//www.jstor.org/stable/27590541.

Rubin:2008:CDA

[Rub08] Donald B. Rubin. Comment: The design and analysis of gold standard randomized experiments. Journal of the American Statistical Association, 103(484):1350–1353, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rue:2004:BRB

[Rue04] Håvard Rue. Book review: 2D Object Detection and Recognition: Models, Algorithms, and Networks, by Yali Amit. Journal of the American Statistical Association, 99(467):902–903, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s49/p902;   http:
//www.jstor.org/stable/27590465.

Rukhin:2003:BRB

[Ruk03a] Andrew L. Rukhin. Book review: Matrix Variate Distributions, by A. K. Gupta; D. K. Nagar. Journal of the American Statistical Association, 98(462):495–496, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/30045266.

Rukhin:2003:MVD

[Ruk03b] Andrew L. Rukhin. Matrix variate distributions. Journal of the American Statistical Association, 98(462):495–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n462/s30/p495.

Ruppert:2001:BRB

[Rup01] David Ruppert. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1523–1524, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085918.

Ruppert:2002:BRB

[Rup02a] David Ruppert. Book review: [untitled]. Journal of the American Statistical Association, 97(458):658, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085690.

Ruppert:2002:IRA

[Rup02b] David Ruppert. [inconsistency of resampling algorithms for high-breakdown regression estimators and a new algorithm]: Comment. Journal of the American Statistical Association, 97(457):148–149, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/3085766.

Ruppert:2004:BRB

[Rup04] David Ruppert. Book review: The Elements of Statistical Learning: Data Mining, Inference, and Prediction, by Trevor Hastie; Robert Tibshirani; Jerome Friedman. Journal of the American Statistical Association, 99(466):567, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s37/p567;   http://www.jstor.org/stable/27590423.

Ruppert:2005:CMP

[Rup05] David Ruppert. Comment: Maximization by parts in likelihood inference. Journal of the American Statistical Association, 100(472):1161–1163, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ruppert:2007:CMA

[Rup07] David Ruppert. Comment: Model-assisted estimation of forest resources with generalized additive models. Journal of the American Statistical Association, 102(478):409–411, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rust:2005:BRB

[Rus05] Philip F. Rust. Book review: Statistical Methods of Analysis, by Chin Long Chiang. Journal of the American Statistical Association, 100(471):1094–1095, September 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590642.

Rutkowski:2000:BRB

[Rut00] Marek Rutkowski. Book review: [untitled]. Journal of the American Statistical Association, 95(450):685–686, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669426.

Robins:2000:ADVa

[RvdVV00a] James M. Robins, Aad van der Vaart, and Valerie Ventura. Asymptotic distribution of P values in composite null models. Journal of the American Statistical Association, 95(452):1143– 1156, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669750.

Robins:2000:ADVb

[RvdVV00b] James M. Robins, Aad van der Vaart, and Valerie Ventura. Asymptotic distribution of P values in composite null models: Rejoinder. Journal of the American Statistical Association, 95 (452):1171–1172, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669757.

RVL:2000:BRB

[RVL00] RVL. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1381–1382, December 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669809.

Robins:2000:CLC

[RW00] James Robins and Larry Wasserman. Conditioning, likelihood, and coherence: a review of some foundational concepts. Journal of the American Statistical Association, 95 (452):1340–1346, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669784.

Romano:2005:EAS

[RW05] Joseph P. Romano and Michael Wolf. Exact and approximate stepdown methods for multiple hypothesis testing. Journal of the American Statistical Association, 100 (469):94–108, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s15/p94;   http://www.
jstor.org/stable/27590521.

Reilly:2003:MNM

[RWR03] Cavan Reilly, Changchun Wang, and Mark Rutherford. A method for normalizing microarrays using genes that are not differentially expressed. Journal of the American Statistical Association, 98(464):868–878, December 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v98n464/s10/p868;   http:
//www.jstor.org/stable/30045338.

Raghunathan:2007:CIT

[RXS^{+}07] Trivellore E. Raghunathan, Dawei Xie, Nathaniel Schenker, Van L. Parsons, William W. Davis, Kevin W. Dodd, and Eric J. Feuer. Combining information from two surveys to estimate county-level prevalence rates of cancer risk factors and screening. Journal of the American Statistical Association, 102(478):474–486, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rissanen:2000:CCH

[RY00] Jorma Rissanen and Bin Yu. Coding and compression: a happy union of theory and practice. Journal of the American Statistical Association, 95(451):986–989, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669482.

Ryan:2000:SIT

[Rya00] Louise M. Ryan. Statistical issues in toxicology. Journal of the American Statistical Association, 95(449):304–308, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669553.

Ryan:2005:BRB

[Rya05] Sarah M. Ryan. Book review: Stochastic Models in Queueing Theory (second edition), by J. Medhi. Journal of the American Statistical Association, 100(469):350–351, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s40/
p350;   http://www.jstor.org/stable/27590546.

Raftery:2003:DPB

[RZ03] Adrian E. Raftery and Yingye Zheng. Discussion: Performance of Bayesian model averaging. Journal of the American Statistical Association, 98(464):931–938, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s18/p931.

Robinson:2006:HMT

[RZF06] John W. Robinson, Scott L. Zeger, and Christopher B. Forrest. A hierarchical multivariate two-part model for profiling providers’ effects on health care charges. Journal of the American Statistical Association, 101(475):911–923, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590771.

Sabatti:2005:CSH

[Sab05] Chiara Sabatti. Comment: Semilinear high-dimensional model for normalization of microarray data. Journal of the American Statistical Association, 100(471):796–798, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sabatti:2006:CLB

[Sab06] Chiara Sabatti. Comment: Likelihood-based inference on haplotype effects in genetic association studies. Journal of the American Statistical Association, 101(473):104–106, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sackrowitz:2003:BRB

[Sac03a] Harold Sackrowitz. Book review: Statistical Decision Theory, by Simon French; David Rios Insua. Journal of the American Statistical Association, 98(462):492, June 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/30045260.

Sackrowitz:2003:SDT

[Sac03b] Harold Sackrowitz. Statistical decision theory. Journal of the American Statistical Association, 98(462):492–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n462/s24/p492.

Satten:2006:CLB

[SAE06] Glen A. Satten, Andrew S. Allen, and Michael P. Epstein. Comment: Likelihood-based inference on haplotype effects in genetic association studies. Journal of the American Statistical Association, 101(473):107–108, March 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sain:2006:BRB

[Sai06] Stephan R. Sain. Book review: Analysis and Modelling of Spatial Environmental Data, by Mikhail Kanevski; Michel Maignan. Journal of the American Statistical Association, 101(475): 1312, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590810.

Sanil:2003:BRB

[San03a] Ashish P. Sanil. Book review: Principles of Data Mining, by David Hand; Heikki Mannila; Padhraic Smyth. Journal of the American Statistical Association, 98(461):252–253, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045219.

Sanil:2003:PDM

[San03b] Ashish P. Sanil. Principles of data mining. Journal of the American Statistical Association, 98(461):252–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s32/p252.

Singpurwalla:2004:RMF

[SB04] Nozer D. Singpurwalla and Jane M. Booker. Rejoinder: Membership functions and probability measures of fuzzy sets. Journal of the American Statistical Association, 99 (467):884–889, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n467/s39/p884.

Staudenmayer:2005:MEL

[SB05] John Staudenmayer and John P. Buonaccorsi. Measurement error in linear autoregressive models. Journal of the American Statistical Association, 100(471):841–852, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590617.

Sanchez:2005:SEM

[SBJRH05] Brisa N. Sánchez, Esben Budtz-Jørgensen, Louise M. Ryan, and Howard Hu. Structural equation models: a review with applications to environmental epidemiology. Journal of the American Statistical Association, 100(472):1443–1455, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590684.

Singpurwalla:2004:MFP

[SBL^{+}04] Nozer D. Singpurwalla, Jane M. Booker, D. V. Lindley, Michael Laviolette, Lotfi A. Zadeh, and A. P. Dempster. Membership functions and probability measures of fuzzy sets [with comments, rejoinder]. Journal of the American Statistical Association, 99(467):867–877, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s34/p867;   http://www.jstor.org/stable/27590456.

Salibian-Barrera:2006:PCA

[SBVW06] Matías Salibián-Barrera, Stefan Van Aelst, and Gert Willems. Principal components analysis based on multivariate MM estimators with fast and robust bootstrap. Journal of the American Statistical Association, 101(475):1198–1211, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590795.

Saloff-Coste:2000:BRB

[SC00] Laurent Saloff-Coste. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1378, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669802.

Shen:2003:SSR

[SC03] Yu Shen and Jianwen Cai. Sample size reestimation for clinical trials with censored survival data. Journal of the American Statistical Association, 98(462):418–426, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n462/
s14/p418;   http://www.jstor.org/stable/30045251.

Sun:2007:OAC

[SC07] Wenguang Sun and T. Tony Cai. Oracle and adaptive compound decision rules for false discovery rate control. Journal of the American Statistical Association, 102(479):901–912, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Swartz:2006:IDT

[SCC^{+}06] Richard J. Swartz, Dennis D. Cox, Scott B. Cantor, Kalatu Davies, and Michele Follen. Inverse decision theory: Characterizing losses for a decision rule with applications in cervical cancer screening. Journal of the American Statistical Association, 101(473):1–8, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047435.

Spence:2007:ASD

[SCG^{+}07] Jeffrey S. Spence, Patrick S. Carmack, Richard F. Gunst, William R. Schucany, Wayne A. Woodward, and Robert W. Haley. Accounting for spatial dependence in the analysis of SPECT brain imaging data. Journal of the American Statistical Association, 102(478):464–473, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmock:2000:BRB

[Sch00a] Uwe Schmock. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1380–1381, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669805.

Schork:2000:BRB

[Sch00b] M. Anthony Schork. Book review: [untitled]. Journal of the American Statistical Association, 95(449):339–340, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669572.

Schabenberger:2001:BRB

[Sch01] O. Schabenberger. Book review: [untitled]. Journal of the American Statistical Association, 96(453):348–349, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670382.

Schaid:2002:APH

[Sch02a] Daniel J. Schaid. [adjusting for population heterogeneity and misspecified haplotype frequencies when testing nonparametric null hypotheses in statistical genetics]: Comment. Journal of the American Statistical Association, 97(459):754– 756, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085717.

Scheaffer:2002:SB

[Sch02b] Richard L. Scheaffer. Statistical bridges. Journal of the American Statistical Association, 97(457):1–7, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
3085753;   reqidx=/cw/asa/01621459/v97n457/s1/p1.

Schell:2002:BRB

[Sch02c] Michael J. Schell. Book review: [untitled]. Journal of the American Statistical Association, 97(459):919, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085732.

Schell:2002:CBB

[Sch02d] Michael J. Schell. Curve ball: Baseball, statistics, and the role of chance in the game. Journal of the American Statistical Association, 97(459):919–??, September 2002. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s22/p919.

Schott:2002:BRB

[Sch02e] James R. Schott. Book review: [untitled]. Journal of the American Statistical Association, 97(458):657–658, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085688.

Schenker:2003:AVD

[Sch03a] Nathaniel Schenker. Assessing variability due to race bridging: Application to census counts and vital rates for the year 2000. Journal of the American Statistical Association, 98(464):818–828, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s5/p818;   http://www.
jstor.org/stable/30045333.

Schliekelman:2003:BRB

[Sch03b] Paul Schliekelman. Book review: Statistical Methods in Bioinformatics: An Introduction, by Warren J. Ewens; Gregory R. Grant. Journal of the American Statistical Association, 98 (462):490–491, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045257.

Schliekelman:2003:SMB

[Sch03c] Paul Schliekelman. Statistical methods in bioinformatics: An introduction. Journal of the American Statistical Association, 98(462):490–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s21/p490.

Schoenberg:2003:MRA

[Sch03d] Frederic Paik Schoenberg. Multidimensional residual analysis of point process models for earthquake occurrences. Journal of the American Statistical Association, 98(464): 789–795, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/

rpsv/cw/asa/01621459/v98n464/s2/p789;   http://www.
jstor.org/stable/30045330.

Schonbucher:2004:BRB

[Sch04a] Philipp Schönbucher. Book review: Applied Computational Economics and Finance. Mario J. Miranda and Paul L. Fackler. Journal of the American Statistical Association, 99(466):565–566, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s35/p565;   http:
//www.jstor.org/stable/27590421.

Schuster:2004:BRB

[Sch04b] Christof Schuster. Book review: Measures of Interobserver Agreement, by Mohamed M. Shoukri. Journal of the American Statistical Association, 99(468):1204–1205, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s34/p1204;   http://www.jstor.org/stable/27590501.

Schoenberg:2005:BRB

[Sch05] Frederic P. Schoenberg. Book review: Statistical inference and Simulation for Spatial Point Processes, by Jesper Møller; Rasmus P. Waagepetersen. Journal of the American Statistical Association, 100(469):349–350, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s38/p349;   http://www.
jstor.org/stable/27590544.

Schmidli:2006:BRB

[Sch06] Hanspeter Schmidli. Book review: Insurance Risk and Ruin, by David C. M. Dickson. Journal of the American Statistical Association, 101(475):1316, September 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590816.

Scheuren:2007:BRB

[Sch07a] Fritz Scheuren. Book review: Statistical Matching: Theory and Practice. Journal of the American Statistical Association, 102 (479):1076–1077, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scheuren:2007:PBS

[Sch07b] Fritz Scheuren. The pro bono statistician. Journal of the American Statistical Association, 102(477):1–6, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schlather:2007:BRB

[Sch07c] Martin Schlather. Book review: Point Process Theory and Applications: Marked Point and Piecewise Deterministic Processes. Journal of the American Statistical Association, 102 (479):1077–1078, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwartzman:2007:CSA

[Sch07d] Armin Schwartzman. Comment: Statistical analysis of diffusion tensors in diffusion-weighted magnetic resonance imaging data. Journal of the American Statistical Association, 102 (480):1102, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schweder:2007:BRB

[Sch07e] Tore Schweder. Book review: Advanced Distance Sampling: Estimating Abundance of Biological Populations. Journal of the American Statistical Association, 102(478):763–764, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schennach:2009:CNP

[Sch09] Susanne M. Schennach. Comment: Nonparametric prediction in measurement error models. Journal of the American Statistical Association, 104(487):1007–1010, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scott:2002:BMH

[Sco02] Steven L. Scott. Bayesian methods for hidden Markov models: Recursive computing in the 21st Century. Journal of the American Statistical Association, 97(457):337–351, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085787;   reqidx=/cw/asa/01621459/v97n457/
s30/p337.

Shadish:2008:CNE

[SCS08a] William R. Shadish, M. H. Clark, and Peter M. Steiner. Can nonrandomized experiments yield accurate answers? A randomized experiment comparing random and nonrandom assignments. Journal of the American Statistical Association, 103(484):1334–1344, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shadish:2008:RCD

[SCS08b] William R. Shadish, M. H. Clark, and Peter M. Steiner. Rejoinder: Comment: The design and analysis of gold standard randomized experiments. Journal of the American Statistical Association, 103(484):1353–1356, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Soofi:2002:NNV

[SD02] Ehsan S. Soofi and Ali Dadpay. [nearest-neighbor variance estimation (NNVE): Robust covariance estimation via nearestneighbor cleaning]: Comment. Journal of the American Statistical Association, 97(460):1012–1014, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085824.

Shen:2003:DII

[SD03] Xiaotong Shen and Daniel P. Dougherty. Discussion: Inference and interpretability considerations in frequentist model averaging and selection. Journal of the American Statistical Association, 98(464):917–919, December 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v98n464/s13/p917.

Seastrom:2002:LCA

[Sea02] Marilyn M. Seastrom. [latent class analysis of complex sample survey data: Application to dietary data]: Comment. Journal of the American Statistical Association, 97(459):735– 736, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085711.

Sengupta:2001:BRB

[Sen01] Debasis Sengupta. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1138, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670263.

Seneta:2008:CSA

[Sen08a] Eugene Seneta. Comment: On Student’s 1908 article “The Probable Error of a Mean”. Journal of the American Statistical Association, 103(481):13–15, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Senter:2008:ALS

[Sen08b] Herman F. Senter. Applied linear statistical models. Journal of the American Statistical Association, 103(482):880, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Serfling:2004:CLS

[Ser04] Robert Serfling. Comment: Location-scale depth. Journal of the American Statistical Association, 99(468):970–973, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s7/p970.

Seymour:2005:BRB

[Sey05] Lynne Seymour. Book review: Spatial Data Analysis: Theory and Practice, by Robert Haining. Journal of the American Statistical Association, 100(469):353, March 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n469/s44/p353;   http:
//www.jstor.org/stable/27590550.

Smith:2007:SBV

[SF07] Michael Smith and Ludwig Fahrmeir. Spatial Bayesian variable selection with application to functional magnetic resonance imaging. Journal of the American Statistical Association, 102(478):417–431, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2005:RMP

[SFK05a] Peter X.-K. Song, Yanqin Fan, and John D. Kalbfleisch. Rejoinder: Maximization by parts in likelihood inference. Journal of the American Statistical Association, 100(472):1164–1167, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2005:MPL

[SFK^{+}05b] Peter X.-K. Song, Yanqin Fan, John D. Kalbfleisch, Jiming Jiang, Thomas A. Louis, J. G. Liao, Bahjat F. Qaqish, and David Ruppert. Maximization by parts in likelihood inference [with comments, rejoinder]. Journal of the American Statistical Association, 100(472):1145–1158, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590661.

Suarez-Farinas:2004:LGN

[SFPM04] Mayte Suárez-Fariñas, Carlos E. Pedreira, and Marcelo C. Medeiros. Local global neural networks: a new approach for nonlinear time series modeling. Journal of the American Statistical Association, 99(468):1092–1107, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s21/p1092;   http://www.jstor.org/stable/27590488.

Sanso:2000:NMM

[SG00] Bruno Sanso and Lelys Guenni. A nonstationary multisite model for rainfall. Journal of the American Statistical Association, 95(452):1089–1100, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669745.

Sim:2005:OLB

[SGC05] C. H. Sim, F. F. Gan, and T. C. Chang. Outlier labeling with boxplot procedures. Journal of the American Statistical Association, 100(470):642–652, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/

rpsv/cw/asa/01621459/v100n470/s25/p642;   http://www.
jstor.org/stable/27590584.

Sahu:2007:HRS

[SGH07] Sujit K. Sahu, Alan E. Gelfand, and David M. Holland. Highresolution space–time ozone modeling for assessing trends. Journal of the American Statistical Association, 102(480): 1221–1234, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shepherd:2007:SAC

[SGL07] Bryan E. Shepherd, Peter B. Gilbert, and Thomas Lumley. Sensitivity analyses comparing time-to-event outcomes existing only in a subset selected postrandomization. Journal of the American Statistical Association, 102(478):573–582, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2006:OMA

[SH06] Xiaotong Shen and Hsin-Cheng Huang. Optimal model assessment, selection, and combination. Journal of the American Statistical Association, 101(474):554–568, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590717.

Shafer:2000:CIC

[Sha00] Glenn Shafer. Causal inference without counterfactuals: Comment. Journal of the American Statistical Association, 95 (450):438–442, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669383.

Shaffer:2005:CFD

[Sha05] Juliet Popper Shaffer. Comment: False discovery rateadjusted multiple confidence intervals for selected parameters. Journal of the American Statistical Association, 100(469):82–84, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s10/p82.

Shafer:2007:BRB

[Sha07] Glenn Shafer. Book review: Scientific Reasoning: The Bayesian Approach. Journal of the American Statistical Association, 102(477):387, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2002:NHT

[SHC02] Xiaotong Shen, Hsin-Cheng Huang, and Noel Cressie. Nonparametric hypothesis testing for a spatial signal. Journal of the American Statistical Association, 97(460): 1122–1140, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s17/p1122;   http:/
/www.jstor.org/stable/3085836.

Shen:2000:LRC

[She00a] Xiaotong Shen. Linear regression with current status data. Journal of the American Statistical Association, 95(451):842– 852, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669468.

Shen:2000:PLC

[She00b] Xiaotong Shen. On profile likelihood: Comment. Journal of the American Statistical Association, 95(450):474–477, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669390.

Shen:2002:ANS

[She02a] Xiaotong Shen. Asymptotic normality of semiparametric and nonparametric posterior distributions. Journal of the American Statistical Association, 97(457):222–235, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085777;   reqidx=/cw/asa/01621459/v97n457/
s20/p222.

Shepp:2002:BRB

[She02b] Lawrence A. Shepp. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1213, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085859.

Shepp:2002:PE

[She02c] Lawrence A. Shepp. Probability essentials. Journal of the American Statistical Association, 97(460):1213–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n460/s40/p1213.

Sherman:2002:SBV

[She02d] Michael Sherman. [sieve bootstrap with variable-length Markov chains for stationary categorical time series]: Comment. Journal of the American Statistical Association, 97 (458):457–460, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085659.

Stromberg:2000:LTD

[SHH00] Arnold J. Stromberg, Ola Hossjer, and Douglas M. Hawkins. The least trimmed differences regression estimator and alternatives. Journal of the American Statistical Association, 95 (451):853–864, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669469.

Shiu:2001:BRB

[Shi01] Elias Shiu. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1527, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085924.

Short:2005:BRB

[Sho05] Thomas H. Short. Book review: Subjective and Objective Bayesian Statistics: Principles, Models, and Applications, by S. James. Journal of the American Statistical Association, 100(469):355–356, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s47/p355;   http://www.
jstor.org/stable/27590553.

Shen:2004:CEP

[SHY04a] Xiaotong Shen, Hsin-Cheng Huang, and Jianming Ye. Comment: The estimation of prediction error: Covariance penalties and cross-validation. Journal of the American Statistical Association, 99(467):634–637, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s12/p634.

Shen:2004:IAM

[SHY04b] Xiaotong Shen, Hsin-Cheng Huang, and Jimmy Ye. Inference after model selection. Journal of the American Statistical Association, 99(467):751–762, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s23/p751;   http://www.jstor.org/stable/27590445.

Scott:2002:EBI

[SI02] Steven L. Scott and Edward H. Ip. Empirical Bayes and item-clustering effects in a latent variable hierarchical model: a case study from the National Assessment of Educational Progress. Journal of the American Statistical Association, 97(458):409–419, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s5/p409;   http://
www.jstor.org/stable/3085654.

Sigman:2000:BRB

[Sig00] K. Sigman. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1378–1379, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669803.

Simpson:2002:NNV

[Sim02] Douglas G. Simpson. [nearest-neighbor variance estimation (NNVE): Robust covariance estimation via nearest-neighbor cleaning]: Comment. Journal of the American Statistical Association, 97(460):1015, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/3085825.

Sinha:2004:RAG

[Sin04] Sanjoy K. Sinha. Robust analysis of generalized linear mixed models. Journal of the American Statistical Association, 99(466):451–460, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s14/p451;   http:
//www.jstor.org/stable/27590400.

Singpurwalla:2006:HP

[Sin06a] Nozer D. Singpurwalla. The hazard potential. Journal of the American Statistical Association, 101(476):1705–1717, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Singpurwalla:2006:HPI

[Sin06b] Nozer D. Singpurwalla. The hazard potential: Introduction and overview. Journal of the American Statistical Association, 101(476):1705–1717, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639784.

Sinha:2006:BRB

[Sin06c] Debajyoti Sinha. Book review: Statistical Monitoring of Clinical Trials: Fundamentals for Investigators, by Lemuel A. Moyé. Journal of the American Statistical Association, 101(476):1734, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639808.

Sinha:2007:BRBa

[Sin07a] Debajyoti Sinha. Book review: Biostatistics: a Bayesian Introduction. Journal of the American Statistical Association, 102(478):767–768, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sinha:2007:BRBb

[Sin07b] Debajyoti Sinha. Book review: Dynamic Regression Models for Survival Data. Journal of the American Statistical Association, 102(480):1474, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sirakaya:2006:RSI

[Sir06] Sibel Sirakaya. Recidivism and social interactions. Journal of the American Statistical Association, 101(475):863– 877, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590767.

Sisson:2005:TMC

[Sis05] Scott A. Sisson. Transdimensional Markov chains: a decade of progress and future perspectives. Journal of the American Statistical Association, 100(471):1077–1089, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590636.

Sugar:2003:FNC

[SJ03] Catherine A. Sugar and Gareth M. James. Finding the number of clusters in a dataset: An information-theoretic approach. Journal of the American Statistical Association, 98 (463):750–763, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045303.

Signorini:2004:KEU

[SJ04] D. F. Signorini and M. C. Jones. Kernel estimators for univariate binary regression. Journal of the American Statistical Association, 99(465):119–126, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s12/p119;   http:
//www.jstor.org/stable/27590358.

Su:2006:LSJ

[SJ06] Chun-Lung Su and Wesley O. Johnson. Large-sample joint posterior approximations when full conditionals are approximately normal: Application to generalized linear mixed models. Journal of the American Statistical Association, 101(474): 795–811, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590736.

Sun:2001:ETI

[SjKH^{+}01] Jianguo Sun, Yang jin Kim, John Hewett, Jane C. Johnson, Janet Farmer, and Michelle Gibler. Evaluation of traffic injury prevention programs using counting process approaches. Journal of the American Statistical Association, 96(454):469–475, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670285;   reqidx=/cw/asa/01621459/v96n454/
s11/p469.

Scott:2005:HMM

[SJS05] Steven L. Scott, Gareth M. James, and Catherine A. Sugar. Hidden Markov models for longitudinal comparisons. Journal of the American Statistical Association, 100(470):359–369, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s1/p359;   http:
//www.jstor.org/stable/27590560.

Smith:2002:PCM

[SK02] Michael Smith and Robert Kohn. Parsimonious covariance matrix estimation for longitudinal data. Journal of the American Statistical Association, 97(460):1141–1153, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s18/p1141;   http://www.jstor.org/stable/3085837.

Song:2006:CBC

[SL06] Kai-Sheng Song and Ta-Hsin Li. On convergence and bias correction of a joint estimation algorithm for multiple sinusoidal frequencies. Journal of the American Statistical Association, 101(474):830–842, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590739.

Sabatti:2008:BGM

[SL08] Chiara Sabatti and Kenneth Lange. Bayesian Gaussian mixture models for high-density genotyping arrays. Journal of the American Statistical Association, 103(481):89–100, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Small:2007:SAI

[Sma07] Dylan S. Small. Sensitivity analysis for instrumental variables regression with overidentifying restrictions. Journal of the American Statistical Association, 102(479):1049–1058, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spitzner:2003:MMF

[SME03] Dan J. Spitzner, J. S. Marron, and G. K. Essick. Mixedmodel functional ANOVA for studying human tactile perception. Journal of the American Statistical Association, 98(462):263–272, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s1/p263;   http://
www.jstor.org/stable/30045234.

Sinha:2005:SBA

[SMG^{+}05] Samiran Sinha, Bhramar Mukherjee, Malay Ghosh, Bani K. Mallick, and Raymond J. Carroll. Semiparametric Bayesian analysis of matched case-control studies with missing exposure. Journal of the American Statistical Association, 100(470):591–601, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s21/p591;   http://www.
jstor.org/stable/27590580.

Sinha:2008:CMR

[SMIO08] Debajyoti Sinha, Tapabrata Maiti, Joseph G. Ibrahim, and Bichun Ouyang. Current methods for recurrent events data with dependent termination. Journal of the American Statistical Association, 103(482):866–878, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stroud:2003:NSS

[SMP03] Jonathan R. Stroud, Peter Müller, and Nicholas G. Polson. Nonlinear state-space models with state-dependent variances. Journal of the American Statistical Association, 98(462):377–386, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&

reqidx=/cw/asa/01621459/v98n462/s10/p377;   http://
www.jstor.org/stable/30045247.

Shen:2009:ALB

[SNQ09] Yu Shen, Jing Ning, and Jing Qin. Analyzing length-biased data with semiparametric transformation and accelerated failure time models. Journal of the American Statistical Association, 104(487):1192–1202, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stevens:2004:SBS

[SO04] Don L. Stevens, Jr. and Anthony R. Olsen. Spatially balanced sampling of natural resources. Journal of the American Statistical Association, 99(465):262–278, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s25/p262;   http://www.jstor.org/stable/27590371.

Sobel:2000:CIS

[Sob00] Michael E. Sobel. Causal inference in the social sciences. Journal of the American Statistical Association, 95(450):647–651, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669410.

Sobel:2006:WDR

[Sob06] Michael E. Sobel. What do randomized studies of housing mobility demonstrate?: Causal inference in the face of interference. Journal of the American Statistical Association, 101 (476):1398–1407, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639760.

Solo:2000:ETS

[Sol00] V. Solo. The end of time series. Journal of the American Statistical Association, 95(452):1346–1349, December 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669785.

Solo:2001:RWA

[Sol01] V. Solo. Regularization of wavelet approximations: Discussion. Journal of the American Statistical Association, 96(455):

963–964, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670241.

Soofi:2000:PIT

[Soo00] Ehsan S. Soofi. Principal information theoretic approaches. Journal of the American Statistical Association, 95(452):1349– 1353, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669786.

Sorensen:2000:BRB

[Sor00] Michael Sorensen. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1016–1017, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669500.

Spanos:2002:BRB

[Spa02a] Aris Spanos. Book review: [untitled]. Journal of the American Statistical Association, 97(459):921–923, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085736.

Spanos:2002:PET

[Spa02b] Aris Spanos. Probability, econometrics and truth: The methodology of econometrics. Journal of the American Statistical Association, 97(459):921–??, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s26/p921.

Speckman:2006:BRB

[Spe06a] Paul Speckman. Book review: Nonparametric and Semiparametric Models, by Wolfgang Härdle; Marlene Müller; Stefan Sperlich; Axel Werwatz. Journal of the American Statistical Association, 101(475):1308–1309, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590805.

Speed:2006:BRB

[Spe06b] Terry Speed. Book review: Statistics for Experimenters: Design, Innovation, and Discovery, by George E. P. Box; J. Stuart Hunter; William G. Hunter. Journal of the American Statistical Association, 101(476):1720–1721, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639787.

Spiegelman:2006:BRB

[Spi06] Donna Spiegelman. Book review: Measurement Error and Misclassification in Statistics and Epidemiology, by Paul Gustafson. Journal of the American Statistical Association, 101(473):392–393, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30047471.

Skates:2001:SBR

[SPJ01] Steven J. Skates, Donna K. Pauler, and Ian J. Jacobs. Screening based on the risk of cancer calculation from Bayesian hierarchical changepoint and mixture models of longitudinal markers. Journal of the American Statistical Association, 96(454):429–439, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670281;   reqidx=/cw/
asa/01621459/v96n454/s7/p429.

Sun:2005:SRA

[SPSZ05] Jianguo Sun, Do-Hwan Park, Liuquan Sun, and Xingqiu Zhao. Semiparametric regression analysis of longitudinal data with informative observation times. Journal of the American Statistical Association, 100(471):882–889, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590620.

Shen:2007:ITE

[SQC07] Yu Shen, Jing Qin, and Joseph P. Costantino. Inference of Tamoxifen’s effects on prevention of breast cancer from a randomized controlled trial. Journal of the American Statistical Association, 102(480):1235–1244, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sebastiani:2001:BSD

[SR01] Paola Sebastiani and Marco Ramoni. Bayesian selection of decomposable models with incomplete data. Journal of the American Statistical Association, 96(456):1375–1386, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085906;   reqidx=/cw/asa/01621459/v96n456/
s19/p1375.

Small:2008:WW

[SR08] Dylan S. Small and Paul R. Rosenbaum. War and wages. Journal of the American Statistical Association, 103(483):924– 933, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shepherd:2008:DFA

[SRA08] Bryan E. Shepherd, Mary W. Redman, and Donna P. Ankerst. Does finasteride affect the severity of prostate cancer? A causal sensitivity analysis. Journal of the American Statistical Association, 103(484):1392–1404, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Staudenmayer:2008:DEP

[SRB08] John Staudenmayer, David Ruppert, and John P. Buonaccorsi. Density estimation in the presence of heteroscedastic measurement error. Journal of the American Statistical Association, 103(482):726–736, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schenker:2006:MIM

[SRC^{+}06] Nathaniel Schenker, Trivellore E. Raghunathan, Pei-Lu Chiu, Diane M. Makuc, Guangyu Zhang, and Alan J. Cohen. Multiple imputation of missing income data in the national health interview survey. Journal of the American Statistical Association, 101(475):924–933, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590772.

Sriram:2002:BRB

[Sri02] T. N. Sriram. Book review: [untitled]. Journal of the American Statistical Association, 97(457):356, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085793.

SRJ:2000:BRBa

[SRJ00a] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1020, September 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669510.

SRJ:2000:BRBb

[SRJ00b] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1381, December 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669806.

SRJ:2000:BRBc

[SRJ00c] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1382, December 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669810.

SRJ:2000:BRBd

[SRJ00d] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1382, December 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669812.

Spiegelman:2000:EIL

[SRL00] Donna Spiegelman, Bernard Rosner, and Roger Logan. Estimation and inference for logistic regression with covariate misclassification and measurement error in main study/ validation study designs. Journal of the American Statistical Association, 95(449):51–61, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.amstat.org/publications/jasa/Spiegelman.htm;
http://www.jstor.org/stable/2669522.

Stoumbos:2000:SSP

[SRRW00] Zachary G. Stoumbos, Marion R. Reynolds, Jr., Thomas P. Ryan, and William H. Woodall. The state of statistical process control as we proceed into the 21st century. Journal of the American Statistical Association, 95(451):992–998, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669484.

Schafer:2000:IIC

[SS00] Joseph L. Schafer and Nathaniel Schenker. Inference with imputed conditional means. Journal of the American Statistical Association, 95(449):144–154, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
Schafer.htm;   http://www.jstor.org/stable/2669534.

Stanek:2004:PRE

[SS04] Edward J. Stanek, III and Julio M. Singer. Predicting random effects from finite population clustered samples with response error. Journal of the American Statistical Association, 99(468):1119–1130, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s23/p1119;   http://www.jstor.org/stable/27590490.

Skinner:2008:AIR

[SS08] Chris Skinner and Natalie Shlomo. Assessing identification risk in survey microdata using log-linear models. Journal of the American Statistical Association, 103(483):989–1001, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schafer:2009:CCR

[SS09] Chad M. Schafer and Philip B. Stark. Constructing confidence regions of optimal expected size. Journal of the American Statistical Association, 104(487):1080–1089, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shaked:2003:CRE

[SSL03] Moshe Shaked and Alfonso Suarez-Llorens. On the comparison of reliability experiments based on the convolution order. Journal of the American Statistical Association, 98(463):693–702, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&

reqidx=/cw/asa/01621459/v98n463/s22/p693;   http://
www.jstor.org/stable/30045297.

Sun:2007:RAL

[SSL07] Jianguo Sun, Liuquan Sun, and Dandan Liu. Regression analysis of longitudinal data in the presence of informative observation and censoring times. Journal of the American Statistical Association, 102(480):1397–1406, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samarov:2005:CIE

[SSV05] Alexander Samarov, Vladimir Spokoiny, and Celine Vial. Component identification and estimation in nonlinear highdimensional regression models by structural adaptation. Journal of the American Statistical Association, 100(470):429–445, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/
s7/p429;   http://www.jstor.org/stable/27590566.

Sangalli:2009:CSE

[SSVV09] Laura M. Sangalli, Piercesare Secchi, Simone Vantini, and Alessandro Veneziani. A case study in exploratory functional data analysis: Geometrical features of the internal carotid artery. Journal of the American Statistical Association, 104 (485):37–48, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2007:ISE

[SSY07] Jianxin Shi, David Siegmund, and Benny Yakir. Importance sampling for estimating p values in linkage analysis. Journal of the American Statistical Association, 102(479):929–937, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Saltelli:2002:RII

[ST02] Andrea Saltelli and Stefano Tarantola. On the relative importance of input factors in mathematical models: Safety assessment for nuclear waste disposal. Journal of the American Statistical Association, 97(459):702–709, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-

bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s5/p702;   http://www.jstor.org/stable/3085706.

Sardy:2004:SAS

[ST04] Sylvain Sardy and Paul Tseng. On the statistical analysis of smoothing by maximizing dirty Markov random field posterior distributions. Journal of the American Statistical Association, 99(465):191–204, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s19/p191;   http:
//www.jstor.org/stable/27590365.

Starica:2001:BRB

[Sta01] Catalin Starica. Book review: [untitled]. Journal of the American Statistical Association, 96(454):780, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670320.

Stangl:2008:BRB

[Sta08] Dalene K. Stangl. Book review: Encyclopedia of Statistics in Behavioral Science. Journal of the American Statistical Association, 103(482):881–882, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Staudenmayer:2009:CNP

[Sta09] John Staudenmayer. Comment: Nonparametric prediction in measurement error models. Journal of the American Statistical Association, 104(487):1010–1012, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stefanski:2000:MEM

[Ste00a] L. A. Stefanski. Measurement error models. Journal of the American Statistical Association, 95(452):1353–1358, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669787.

Stein:2000:BRB

[Ste00b] Michael L. Stein. Book review: [untitled]. Journal of the American Statistical Association, 95(449):335–337, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669569.

Stern:2000:ADV

[Ste00c] Hal S. Stern. Asymptotic distribution of P values in composite null models: Comment. Journal of the American Statistical Association, 95(452):1157–1159, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669751.

Steiger:2001:DFR

[Ste01] James H. Steiger. Driving fast in reverse: The relationship between software development, theory, and education in structural equation modeling. Journal of the American Statistical Association, 96(453):331–338, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670370;   reqidx=/cw/
asa/01621459/v96n453/s26/p331.

Stephens:2004:BRB

[Ste04] David A. Stephens. Book review: Bayesian Field Theory, by Jörg C. Lemm. Journal of the American Statistical Association, 99(468):1202, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s31/p1202;   http://www.
jstor.org/stable/27590498.

Steele:2005:BRB

[Ste05a] Russell Steele. Book review: Optimization, by Kenneth Lange. Journal of the American Statistical Association, 100(472): 1457, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590685.

Stein:2005:STC

[Ste05b] Michael L. Stein. Space: Time covariance functions. Journal of the American Statistical Association, 100(469):310–321, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s34/
p310;   http://www.jstor.org/stable/27590540.

Steinley:2006:BRB

[Ste06] Douglas Steinley. Book review: Categorical Data Analysis with SAS and SPSS Applications, by Bayo Lawal; Categorical Data Analysis Using the SAS System, by Maura E. Stokes; Charles S. Davis; Gary G. Koch. Journal of the American Statistical Association, 101(473):400, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047484.

Stigler:2008:CSA

[Sti08] Stephen M. Stigler. Comment: On Student’s 1908 article “The Probable Error of a Mean”. Journal of the American Statistical Association, 103(481):7–8, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sen:2007:HDL

[STJ07] Pranab K. Sen, Ming-Tien Tsai, and Yuh-Shan Jou. Highdimension, low-sample size perspectives in constrained statistical inference. Journal of the American Statistical Association, 102(478):686–694, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:2009:BMU

[STNM09] Richard L. Smith, Claudia Tebaldi, Doug Nychka, and Linda O. Mearns. Bayesian modeling of uncertainty in ensembles of climate models. Journal of the American Statistical Association, 104(485):97–116, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stoffer:2000:BRB

[Sto00] David S. Stoffer. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1373, December 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669794.

Stoner:2004:BRB

[Sto04] Julie A. Stoner. Book review: Cross-Over Trials in Clinical Research (second edition), by Stephen Senn. Journal of the American Statistical Association, 99(468):1208, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/

vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s40/p1208;   http://www.jstor.org/stable/27590507.

Stokes:2008:BRB

[Sto08] Lynne Stokes. Book review: Introduction to Variance Estimation. Journal of the American Statistical Association, 103 (483):1324–1325, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scharpf:2009:BMC

[STPN09a] Robert B. Scharpf, Håkon Tjelmeland, Giovanni Parmigiani, and Andrew B. Nobel. A Bayesian model for cross-study differential gene expression. Journal of the American Statistical Association, 104(488):1295–1310, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scharpf:2009:RBM

[STPN09b] Robert B. Scharpf, Håkon Tjelmeland, Giovanni Parmigiani, and Andrew B. Nobel. Rejoinder: a Bayesian model for crossstudy differential gene expression. Journal of the American Statistical Association, 104(488):1318–1323, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Strawderman:2000:BRBc

[Str00a] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(449):341, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669574.

Strawderman:2000:BRBd

[Str00b] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(449):342, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669575.

Strawderman:2000:BRBe

[Str00c] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(449):346, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669581.

Strawderman:2000:BRBg

[Str00d] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(449):349, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669586.

Strawderman:2000:BRBh

[Str00e] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1019, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669505.

Strawderman:2000:EMI

[Str00f] Robert L. Strawderman. Estimating the mean of an increasing stochastic process at a censored stopping time. Journal of the American Statistical Association, 95(452):1192– 1208, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669760.

Strawderman:2000:HOA

[Str00g] Robert L. Strawderman. Higher-order asymptotic approximation: Laplace, saddlepoint, and related methods. Journal of the American Statistical Association, 95(452):1358– 1364, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669788.

Strawderman:2000:BRBa

[Str00h] William E. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(449):329, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669560.

Strawderman:2000:BRBb

[Str00i] William E. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(449):329–330, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669561.

Strawderman:2000:BRBf

[Str00j] William E. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 95(449):348, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669585.

Strawderman:2000:M

[Str00k] William E. Strawderman. Minimaxity. Journal of the American Statistical Association, 95(452):1364–1368, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669789.

Street:2000:BRB

[Str00l] Deborah J. Street. Book review: [untitled]. Journal of the American Statistical Association, 95(450):677–678, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669418.

Stramer:2001:BRB

[Str01a] Osnat Stramer. Book review: [untitled]. Journal of the American Statistical Association, 96(453):342–343, March 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670374.

Strawderman:2001:BRBa

[Str01b] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 96(454):778, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670314.

Strawderman:2001:BRBb

[Str01c] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 96(454):782–783, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670325.

Strawderman:2001:BRBc

[Str01d] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1141, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670269.

Strawderman:2001:BRBd

[Str01e] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1530, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085929.

Strawderman:2001:BRBe

[Str01f] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1530, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085930.

Strawderman:2001:BRBf

[Str01g] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085937.

Strawderman:2002:BRBb

[Str02a] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 97(458):659, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085694.

Strawderman:2002:BRBa

[Str02b] William E. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 97(457):366, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085809.

Stromberg:2002:BRB

[Str02c] Arnold J. Stromberg. Book review: [untitled]. Journal of the American Statistical Association, 97(457):365–366, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085807.

Straf:2003:SNG

[Str03a] Miron L. Straf. Statistics: The next generation. Journal of the American Statistical Association, 98(461):1–6, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-

bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n461/
s1/p1;   http://www.jstor.org/stable/30045188.

Strawderman:2003:BRB

[Str03b] William E. Strawderman. Book review: Annotated Readings in the History of Statistics, by H. A. David; A. W. F. Edwards. Journal of the American Statistical Association, 98(464):1080, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s32/p1080;   http://www.
jstor.org/stable/30045353.

Strawderman:2007:BRB

[Str07] William E. Strawderman. Book review: Theory of Preliminary Test and Stein-Type Estimation With Applications. Journal of the American Statistical Association, 102(478):762, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Small:2008:RIG

[STR08] Dylan S. Small, Thomas R. Ten Have, and Paul R. Rosenbaum. Randomization inference in a group-randomized trial of treatments for depression. Journal of the American Statistical Association, 103(481):271–279, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stufken:2000:BRB

[Stu00] John Stufken. Book review: [untitled]. Journal of the American Statistical Association, 95(450):679, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669419.

Sturm:2006:BRB

[Stu06] Anja Sturm. Book review: The Theory of Measures and Integration, by Eric M. Vestrup. Journal of the American Statistical Association, 101(473):388, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047465.

Shen:2003:L

[STZW03] Xiaotong Shen, George C. Tseng, Xuegong Zhang, and Wing Hung Wong. On ψ-learning. Journal of the American Statistical Association, 98(463):724–734, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045300.

Sungur:2000:BRB

[Sun00] Engin A. Sungur. Book review: [untitled]. Journal of the American Statistical Association, 95(449):334–335, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669568.

Sundaram:2001:BRB

[Sun01] Rajeshwari Sundaram. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1528, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085926.

Sun:2002:BRB

[Sun02] Yanqing Sun. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1217, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085872.

Sundberg:2008:BRB

[Sun08] Rolf Sundberg. Book review: Applied Chemometrics for Scientists. Journal of the American Statistical Association, 103 (483):1317–1318, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Surles:2004:BRB

[Sur04] James Surles. Book review: The Stress-Strength Model and Its Generalizations: Theory and Applications, by Samuel Kotz; Yan Lumelskii; Marianna Pensky. Journal of the American Statistical Association, 99(467):903–904, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s50/p903;   http://www.jstor.org/stable/27590466.

Segal:2002:CTB

[SW02a] Mark R. Segal and Joseph L. Wiemels. Clustering of translocation breakpoints. Journal of the American Statistical Association, 97(457):66–76, March 2002. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085759;   reqidx=/cw/
asa/01621459/v97n457/s7/p66.

Shao:2002:SCC

[SW02b] Jun Shao and Hansheng Wang. Sample correlation coefficients based on survey data under regression imputation. Journal of the American Statistical Association, 97(458):544–552, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s16/p544;   http://
www.jstor.org/stable/3085670.

Sitter:2002:EEQ

[SW02c] Randy R. Sitter and Changbao Wu. Efficient estimation of quadratic finite population functions in the presence of auxiliary information. Journal of the American Statistical Association, 97(458):535–543, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s15/p535;   http://
www.jstor.org/stable/3085669.

Stock:2002:FUP

[SW02d] James H. Stock and Mark W. Watson. Forecasting using principal components from a large number of predictors. Journal of the American Statistical Association, 97 (460):1167–1179, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s20/p1167;   http:/
/www.jstor.org/stable/3085839.

Serban:2005:C

[SW05a] Nicoleta Serban and Larry Wasserman. CATS. Journal of the American Statistical Association, 100(471):990–999, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Serban:2005:CCA

[SW05b] Nicoleta Serban and Larry Wasserman. CATS: Clustering after transformation and smoothing. Journal of the American Statistical Association, 100(471):990–999, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590629.

Sun:2009:CNP

[SW09] Jiayang Sun and Xiaofeng Wang. Comment: Nonparametric prediction in measurement error models. Journal of the American Statistical Association, 104(487):1012–1013, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Swanson:2000:BRB

[Swa00] Norman R. Swanson. Book review: [untitled]. Journal of the American Statistical Association, 95(450):687–688, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669429.

Suchard:2003:ISP

[SWDS03] Marc A. Suchard, Robert E. Weiss, Karin S. Dorman, and Janet S. Sinsheimer. Inferring spatial phylogenetic variation along nucleotide sequences: a multiple changepoint model. Journal of the American Statistical Association, 98(462):427–437, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s15/p427;   http://
www.jstor.org/stable/30045252.

Serroukh:2000:SPU

[SWP00] A. Serroukh, A. T. Walden, and D. B. Percival. Statistical properties and uses of the wavelet variance estimator for the scale analysis of time series. Journal of the American Statistical Association, 95(449):184–196, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/Serroukh.htm;   http://www.jstor.
org/stable/2669537.

Suchard:2003:ESA

[SWS^{+}03] Marc A. Suchard, Robert E. Weiss, Janet S. Sinsheimer, Karin S. Dorman, Megha Patel, and Edward R. B. McCabe. Evolutionary similarity among genes. Journal of the American Statistical Association, 98(463):653–662, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v98n463/
s17/p653;   http://www.jstor.org/stable/30045292.

Schaubel:2009:EET

[SWSM09] Douglas E. Schaubel, Robert A. Wolfe, Camelia S. Sima, and Robert M. Merion. Estimating the effect of a time-dependent treatment by levels of an internal time-dependent covariate: Application to the contrast between liver wait-list and posttransplant mortality. Journal of the American Statistical Association, 104(485):49–59, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Swygert:2002:BRB

[Swy02a] Kimberly Swygert. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1213–1214, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085860.

Swygert:2002:PCC

[Swy02b] Kimberly Swygert. Practical considerations in computerbased testing. Journal of the American Statistical Association, 97(460):1213–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s41/p1213.

Stute:2007:ELI

[SXZ07] Winfried Stute, Liugen Xue, and Lixing Zhu. Empirical likelihood inference in nonlinear errors-in-covariables models with validation data. Journal of the American Statistical Association, 102(477):332–346, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2002:AMS

[SY02] Xiaotong Shen and Jianming Ye. Adaptive model selection. Journal of the American Statistical Association, 97 (457):210–221, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085776;   reqidx=/cw/
asa/01621459/v97n457/s19/p210.

Su:2006:NLF

[SY06] Zhaohui Su and Shie-Shien Yang. A note on lack-of-fit tests for linear models without replication. Journal of the American Statistical Association, 101(473):205–210, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30047450.

Shi:2008:DAC

[SYCB08] Tao Shi, Bin Yu, Eugene E. Clothiaux, and Amy J. Braverman. Daytime Arctic cloud detection based on multi-angle satellite data with case studies. Journal of the American Statistical Association, 103(482):584–593, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2009:CTM

[SZ09] Liuquan Sun and Zhigang Zhang. A class of transformed mean residual life models with censored survival data. Journal of the American Statistical Association, 104(486):803–815, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Szkutnik:2003:DSE

[Szk03] Zbigniew Szkutnik. Doubly smoothed EM algorithm for statistical inverse problems. Journal of the American Statistical Association, 98(461):178–190, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n461/s18/p178;   http://
www.jstor.org/stable/30045205.

Taksar:2000:BRB

[Tak00] Michael Taksar. Book review: [untitled]. Journal of the American Statistical Association, 95(449):343–344, March 2000. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669578.

Tamhane:2005:CFD

[Tam05] Ajit C. Tamhane. Comment: False discovery rate-adjusted multiple confidence intervals for selected parameters. Journal of the American Statistical Association, 100(469):84–85, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/
s11/p84.

Tan:2004:LAM

[Tan04] Zhiqiang Tan. On a likelihood approach for Monte Carlo integration. Journal of the American Statistical Association, 99(468):1027–1036, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s15/p1027;   http://www.
jstor.org/stable/27590482.

Tan:2006:DAC

[Tan06a] Zhiqiang Tan. A distributional approach for causal inference using propensity scores. Journal of the American Statistical Association, 101(476):1619–1637, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639777.

Tan:2006:RWM

[Tan06b] Zhiqiang Tan. Regression and weighting methods for causal inference using instrumental variables. Journal of the American Statistical Association, 101(476):1607–1618, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639776.

Taylor:2000:BRB

[Tay00] Robert L. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 95(449):348, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669584.

Taylor:2003:BRB

[Tay03a] Robert Taylor. Book review: A User’s Guide to Measure-Theoretic Probability, by David Pollard. Journal of the American Statistical Association, 98(462):493–494, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045263.

Taylor:2003:UGM

[Tay03b] Robert Taylor. A user’s guide to measure-theoretic probability. Journal of the American Statistical Association, 98(462):493–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s27/p493.

Tzeng:2003:ODF

[TBD^{+}03] Jung-Ying Tzeng, William Byerley, B. Devlin, Kathryn Roeder, and Larry Wasserman. Outlier detection and false discovery rates for whole-genome DNA matching. Journal of the American Statistical Association, 98(461):236–246, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s23/p236;   http://www.jstor.org/stable/30045210.

Tu:2008:EFD

[TBJ08] Yufeng Tu, Michael O. Ball, and Wolfgang S. Jank. Estimating flight departure delay distributions — a statistical approach with long-term trend and short-term pattern. Journal of the American Statistical Association, 103(481):112–125, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tamhane:2001:MTP

[TDGW01] Ajit C. Tamhane, Charles W. Dunnett, John W. Green, and Jeffrey D. Wetherington. Multiple test procedures for identifying the maximum safe dose. Journal of the American Statistical Association, 96(455):835–843, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670224;   reqidx=/cw/asa/01621459/v96n455/s6/p835.

Trenkel:2000:FPD

[TEB00] Verena M. Trenkel, David A. Elston, and Stephen T. Buckland. Fitting population dynamics models to count and cull data using sequential importance sampling. Journal of the American Statistical Association, 95(450):363–374, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/2000/June/Trenkel.htm;   http://www.jstor.org/
stable/2669373.

TenHave:2004:CMR

[TEJ^{+}04] Thomas R. Ten Have, Michael R. Elliott, Marshall Joffe, Elaine Zanutto, and Catherine Datto. Causal models for randomized physician encouragement trials in treating primary care depression. Journal of the American Statistical Association, 99(465):16–25, March 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n465/s3/p16;   http://
www.jstor.org/stable/27590349.

Terrell:2004:BRB

[Ter04] George R. Terrell. Book review: Applied Probability, by Kenneth Lange. Journal of the American Statistical Association, 99(466):562, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s30/p562;   http:
//www.jstor.org/stable/27590416.

Teugels:2003:BRB

[Teu03] Jozef L. Teugels. Book review: The Politics of Large Numbers: a History of Statistical Reasoning, by Alain Desrosieres; translated By Camille Naish. Journal of the American Statistical Association, 98(464):1080, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s31/p1080;   http://www.jstor.org/stable/30045352.

Teugels:2008:BRB

[Teu08] Jozef L. Teugels. Book review: Markov Chains: Models, Algorithms and Applications. Journal of the American Statistical Association, 103(483):1325, September 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Teng:2009:SFI

[TH09] Siew Leng Teng and Haiyan Huang. A statistical framework to infer functional gene relationships from biologically interrelated microarray experiments. Journal of the American Statistical Association, 104(486):465–473, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tzeng:2005:FOS

[THC05] ShengLi Tzeng, Hsin-Cheng Huang, and Noel Cressie. A fast, optimal spatial-prediction method for massive datasets. Journal of the American Statistical Association, 100(472):1343– 1357, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590676.

Therneau:2000:BRB

[The00] Terry M. Therneau. Book review: [untitled]. Journal of the American Statistical Association, 95(450):676–677, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669417.

Thomas:2000:SCS

[Tho00] Duncan C. Thomas. Some contributions of statistics to environmental epidemiology. Journal of the American Statistical Association, 95(449):315–319, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669556.

Thompson:2002:BRB

[Tho02] James Thompson. Book review: [untitled]. Journal of the American Statistical Association, 97(458):658, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085689.

Thomson:2008:CBB

[Tho08] Madeleine C. Thomson. Comment: Bivariate binomial spatial modeling of Loa loa prevalence in tropical Africa. Journal of the American Statistical Association, 103(481):42–43, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CDR08a, CDR08b, Pac08, Dun08].

Telesca:2008:BHC

[TI08] Donatello Telesca and Lurdes Y. T. Inoue. Bayesian hierarchical curve registration. Journal of the American Statistical Association, 103(481):328–339, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tibshirani:2005:CSH

[Tib05] Robert Tibshirani. Comment: Semilinear high-dimensional model for normalization of microarray data. Journal of the American Statistical Association, 100(471):807–808, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Timmerman:2003:BRB

[Tim03] Marieke E. Timmerman. Book review: Principal Component Analysis (second edition), by I. T. Jolliffe. Journal of the American Statistical Association, 98(464):1082–1083, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s35/p1082;   http://www.jstor.org/stable/30045356.

Tiwari:2000:BRB

[Tiw00] Ram Tiwari. Book review: [untitled]. Journal of the American Statistical Association, 95(452):1371–1372, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669792.

Tsodikov:2003:ECR

[TIY03] A. D. Tsodikov, J. G. Ibrahim, and A. Y. Yakovlev. Estimating cure rates from survival data: An alternative to two-component mixture models. Journal of the American Statistical Association, 98(464):1063–1078, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s30/p1063;   http://www.jstor.org/stable/30045351.

Teh:2006:HDP

[TJBB06] Yee Whye Teh, Michael I. Jordan, Matthew J. Beal, and David M. Blei. Hierarchical Dirichlet processes. Journal of the American Statistical Association, 101(476):1566–1581, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27639773.

TJV:2000:BRBa

[TJV00a] TJV. Book review: [untitled]. Journal of the American Statistical Association, 95(450):689, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669436.

TJV:2000:BRBb

[TJV00b] TJV. Book review: [untitled]. Journal of the American Statistical Association, 95(450):689, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669437.

TJV:2000:BRBc

[TJV00c] TJV. Book review: [untitled]. Journal of the American Statistical Association, 95(450):689, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669438.

TJV:2000:BRBd

[TJV00d] TJV. Book review: [untitled]. Journal of the American Statistical Association, 95(450):689–690, June 2000. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669439.

TJV:2000:BRBe

[TJV00e] TJV. Book review: [untitled]. Journal of the American Statistical Association, 95(450):690, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669440.

Tamhane:2002:MTP

[TL02] Ajit C. Tamhane and Brent R. Logan. Multiple test procedures for identifying the minimum effective and maximum safe doses of a drug. Journal of the American Statistical Association, 97(457):293–301, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;

http://www.jstor.org/stable/3085783;   reqidx=/cw/
asa/01621459/v97n457/s26/p293.

Tian:2007:RIE

[TLW07a] Lu Tian, Jun Liu, and L. J. Wei. Rejoinder: Implementation of estimating function-based inference procedures with Markov chain Monte Carlo samplers. Journal of the American Statistical Association, 102(479):897–900, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2007:IEF

[TLW07b] Lu Tian, Jun S. Liu, and L. J. Wei. Implementation of estimating function-based inference procedures with Markov chain Monte Carlo samplers. Journal of the American Statistical Association, 102(479):881–888, September 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tebaldi:2004:CCP

[TN04] Claudia Tebaldi and Doug Nychka. Comment: Calibrated probabilistic mesoscale weather field forecasting: The geostatistical output perturbation method. Journal of the American Statistical Association, 99(467):583–586, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/s2/
p583.

Taskinen:2005:MNT

[TOR05] Sara Taskinen, Hannu Oja, and Ronald H. Randles. Multivariate nonparametric tests of independence. Journal of the American Statistical Association, 100(471):916–925, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590623.

Tarpey:2003:PPR

[TPO03] Thaddeus Tarpey, Eva Petkova, and R. Todd Ogden. Profiling placebo responders by self-consistent partitioning of functional data. Journal of the American Statistical Association, 98(464):850–858, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/

rpsv/cw/asa/01621459/v98n464/s8/p850;   http://www.
jstor.org/stable/30045336.

Tzeng:2006:CLB

[TR06] Jung-Ying Tzeng and Kathryn Roeder. Comment: Likelihoodbased inference on haplotype effects in genetic association studies. Journal of the American Statistical Association, 101 (473):111–114, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Traat:2005:BRB

[Tra05] Imbi Traat. Book review: Sample Survey Theory: Some Pythagorean Perspectives, by Paul Knottnerus. Journal of the American Statistical Association, 100(469):356, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s49/
p356;   http://www.jstor.org/stable/27590555.

Tressou:2006:NML

[Tre06] Jessica Tressou. Nonparametric modeling of the left censorship of analytical data in food risk assessment. Journal of the American Statistical Association, 101(476):1377– 1386, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639758.

Tymofyeyev:2007:IOA

[TRH07] Yevgen Tymofyeyev, William F. Rosenberger, and Feifang Hu. Implementing optimal allocation in sequential binary response experiments. Journal of the American Statistical Association, 102(477):224–234, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Trindade:2002:BRB

[Tri02a] A. Alexandre Trindade. Book review: [untitled]. Journal of the American Statistical Association, 97(459):920, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085734.

Trindade:2002:TSF

[Tri02b] A. Alexandre Trindade. Time-series forecasting. Journal of the American Statistical Association, 97(459):920–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://giorgio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v97n459/s24/p920.

Trojani:2006:BRB

[Tro06] Fabio Trojani. Book review: Semiparametric Regression for the Applied Econometrician, by Adonis Yatchew. Journal of the American Statistical Association, 101(473):397–398, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047479.

Tsay:2000:TSF

[Tsa00] Ruey S. Tsay. Time series and forecasting: Brief history and future research. Journal of the American Statistical Association, 95(450):638–643, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669408.

Tsai:2003:DIC

[Tsa03] Chih-Ling Tsai. Discussion: Improved, corrected, and profiled focused information criteria. Journal of the American Statistical Association, 98(464):928–930, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v98n464/
s17/p928.

Thall:2002:STS

[TSE02] Peter F. Thall, Hsi-Guang Sung, and Elihu H. Estey. Selecting therapeutic strategies based on efficacy and death in multicourse clinical trials. Journal of the American Statistical Association, 97(457):29–39, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085756;   reqidx=/cw/
asa/01621459/v97n457/s4/p29.

Tadesse:2005:BVS

[TSV05] Mahlet G. Tadesse, Naijun Sha, and Marina Vannucci. Bayesian variable selection in clustering high-dimensional data. Journal of the American Statistical Association, 100(470):602–617, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s22/p602;   http://www.
jstor.org/stable/27590581.

Taylor:2007:DSS

[TW07a] Jonathan E. Taylor and Keith J. Worsley. Detecting sparse signals in random fields, with an application to brain mapping. Journal of the American Statistical Association, 102(479):913– 928, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tong:2007:OSE

[TW07b] Tiejun Tong and Yuedong Wang. Optimal shrinkage estimation of variances with applications to microarray data analysis. Journal of the American Statistical Association, 102(477):113– 122, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tyler:2007:BRB

[Tyl07] David E. Tyler. Book review: Robust Statistical Methods with R. Journal of the American Statistical Association, 102 (478):759–760, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tyler:2008:BRB

[Tyl08] David E. Tyler. Book review: Robust Statistics: Theory and Methods. Journal of the American Statistical Association, 103 (482):888–889, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2005:CMT

[TZW05] Lu Tian, David Zucker, and L. J. Wei. On the Cox model with time-varying regression coefficients. Journal of the American Statistical Association, 100(469):172–183, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s21/
p172;   http://www.jstor.org/stable/27590527.

Uno:2007:EPR

[UCTW07] Hajime Uno, Tianxi Cai, Lu Tian, and L. J. Wei. Evaluating prediction rules for t-year survivors with censored regression models. Journal of the American Statistical Association, 102 (478):527–537, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

VanZandt:2005:BRB

[Van05] T. Van Zandt. Book review: Numerical Issues in Statistical Computing for the Social Scientist, by Micah Altman; Jeff Gill; Michael P. McDonald. Journal of the American Statistical Association, 100(470):707–708, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s35/p707;   http://www.
jstor.org/stable/27590594.

Ventura:2009:PDP

[VCR09] Laura Ventura, Stefano Cabras, and Walter Racugno. Prior distributions from pseudo-likelihoods in the presence of nuisance parameters. Journal of the American Statistical Association, 104(486):768–774, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

vanderLaan:2002:LEE

[vdLHR02] Mark J. van der Laan, Alan E. Hubbard, and James M. Robins. Locally efficient estimation of a multivariate survival function in longitudinal studies. Journal of the American Statistical Association, 97(458):494–507, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s11/p494;   http://www.jstor.org/stable/3085665.

vanDyk:2008:PCG

[vDP08] David A. van Dyk and Taeyoung Park. Partially collapsed Gibbs samplers. Journal of the American Statistical Association, 103(482):790–796, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

vanderVaart:2001:BRB

[vdV01] A. W. van der Vaart. Book review: [untitled]. Journal of the American Statistical Association, 96(453):352–353, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670386.

vandeWiel:2002:EDM

[vdW02] Mark A. van de Wiel. Exact distributions of multiple comparisons rank statistics. Journal of the American Statistical Association, 97(460):1081–1089, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s13/p1081;   http://www.jstor.org/stable/3085832.

Vermunt:2002:LCA

[Ver02] Jeroen K. Vermunt. [latent class analysis of complex sample survey data: Application to dietary data]: Comment. Journal of the American Statistical Association, 97(459):736– 737, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085712.

VanKeilegom:2002:IME

[VH02] Ingrid Van Keilegom and Thomas P. Hettmansperger. Inference on multivariate M estimators based on bivariate censored data. Journal of the American Statistical Association, 97(457):328–336, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//rosina.catchword.com/cgi-bin/linker?ini=asa&amp%
3Breqidx=/cw/asa/01621459/v97n457/s29/p328;   http:
//www.jstor.org/stable/3085786.

Vidakovic:2001:RWA

[Vid01] Brani Vidakovic. Regularization of wavelet approximations: Discussion. Journal of the American Statistical Association, 96(455):956–958, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670238.

Viens:2002:BRB

[Vie02a] Frederi Viens. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1209–1210, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085855.

Viens:2002:SPP

[Vie02b] Frederi Viens. Stochastic processes: From physics to finance. Journal of the American Statistical Association, 97 (460):1209–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s36/p1209.

Vere-Jones:2000:SSV

[VJ00] David Vere-Jones. Seismology — a statistical vignette. Journal of the American Statistical Association, 95(451):975– 978, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669480.

Vlasiou:2008:BRB

[Vla08] Maria Vlasiou. Book review: A Course on Queueing Models. Journal of the American Statistical Association, 103(481):429– 430, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vogelsang:2001:BRBa

[Vog01a] Timothy J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 96(453):345, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670377.

Vogelsang:2001:BRBb

[Vog01b] Timothy J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 96(453):346–347, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670379.

Vogelsang:2001:BRBc

[Vog01c] Timothy J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 96(453):354, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670388.

Vogelsang:2001:BRBd

[Vog01d] Timothy J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 96(453):354, March 2001.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670389.

Vogelsang:2002:BRBa

[Vog02a] T. J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 97(458):657, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085687.

Vogelsang:2002:BRBb

[Vog02b] T. J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 97(458):659, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085693.

Vogelsang:2002:BRBd

[Vog02c] T. J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 97(459):928, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085749.

Vogelsang:2002:ATE

[Vog02d] Timothy J. Vogelsang. Asymptotic theory for econometricians (rev. ed.). Journal of the American Statistical Association, 97(459):921–??, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n459/s25/p921.

Vogelsang:2002:BRBc

[Vog02e] Timothy J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 97(459):921, September 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085735.

Vonesh:2006:BRB

[Von06] Edward F. Vonesh. Book review: Mixed Models: Theory and Applications, by Eugene Demidenko. Journal of the American Statistical Association, 101(476):1724–1726, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639794.

VanDerLans:2008:EMA

[VPW08] Ralf Van Der Lans, Rik Pieters, and Michel Wedel. Eyemovement analysis of search effectiveness. Journal of the American Statistical Association, 103(482):452–461, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Velasco:2000:WPM

[VR00] Carlos Velasco and Peter M. Robinson. Whittle pseudomaximum likelihood estimation for nonstationary time series. Journal of the American Statistical Association, 95(452):1229– 1243, December 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669763.

Veen:2008:EST

[VS08] Alejandro Veen and Frederic P. Schoenberg. Estimation of space–time branching process models in seismology using an EM–type algorithm. Journal of the American Statistical Association, 103(482):614–624, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

vanSoest:2008:TAY

[vSH08] Arthur van Soest and Michael Hurd. A test for anchoring and yea-saying in experimental consumption data. Journal of the American Statistical Association, 103(481):126–136, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vansteelandt:2008:MRI

[VVTR08] Stijn Vansteelandt, Tyler J. VanderWeele, Eric J. Tchetgen, and James M. Robins. Multiply robust inference for statistical interactions. Journal of the American Statistical Association, 103(484):1693–1704, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vonesh:2001:GLS

[VWM01] Edward F. Vonesh, Hao Wang, and Dibyen Majumdar. Generalized least squares, Taylor series linearization and Fisher’s scoring in multivariate nonlinear regression. Journal of the American Statistical Association, 96(453):282–291, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670366;   reqidx=/cw/asa/01621459/v96n453/
s22/p282.

Vonesh:2002:CSO

[VWNM02] Edward F. Vonesh, Hao Wang, Lei Nie, and Dibyen Majumdar. Conditional second-order generalized estimating equations for generalized linear and nonlinear mixed-effects models. Journal of the American Statistical Association, 97(457):271–283, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/3085781;   reqidx=/cw/
asa/01621459/v97n457/s24/p271.

vanWieringen:2008:TPD

[vWvdWvdV08] Wessel N. van Wieringen, Mark A. van de Wiel, and Aad W. van der Vaart. A test for partial differential expression. Journal of the American Statistical Association, 103(483):1039–1049, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wall:2000:EPS

[WA00] Melanie M. Wall and Yasuo Amemiya. Estimation for polynomial structural equation models. Journal of the American Statistical Association, 95(451):929–940, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669475.

Wang:2006:TCE

[WA06] Lan Wang and Michael G. Akritas. Testing for covariate effects in the fully nonparametric analysis of covariance model. Journal of the American Statistical Association, 101(474): 722–736, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590730.

Walther:2002:DPM

[Wal02] Guenther Walther. Detecting the presence of mixing with multiscale maximum likelihood. Journal of the American Statistical Association, 97(458):508–513, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s12/p508;   http://www.jstor.org/stable/3085666.

Walker:2007:BRB

[Wal07] John H. Walker. Book review: Introduction to Regression Modeling. Journal of the American Statistical Association, 102 (477):384–385, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wand:2001:BRB

[Wan01] M. P. Wand. Book review: [untitled]. Journal of the American Statistical Association, 96(453):343, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670375.

Wang:2002:BRB

[Wan02] Yuedong Wang. Book review: [untitled]. Journal of the American Statistical Association, 97(457):366, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/3085808.

Wang:2006:BRB

[Wan06] Yazhen Wang. Book review: Image Processing and Jump Regression Analysis, by Peihua Qiu. Journal of the American Statistical Association, 101(476):1724, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639792.

Wang:2009:FRU

[Wan09] Hansheng Wang. Forward regression for ultra-high dimensional variable screening. Journal of the American Statistical Association, 104(488):1512–1524, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Warton:2008:PNL

[War08] David I. Warton. Penalized normal likelihood and ridge regularization of correlation and covariance matrices. Journal of the American Statistical Association, 103(481):340–349, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wasserman:2000:CIC

[Was00] Larry Wasserman. Causal inference without counterfactuals: Comment. Journal of the American Statistical Association, 95(450):442–443, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669384.

Wasserman:2001:BRB

[Was01] Stanley Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/3085938.

Watnik:2007:BRB

[Wat07] Mitchell Watnik. Book review: Advanced Statistics From an Elementary Point of View. Journal of the American Statistical Association, 102(478):766–767, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2005:RDG

[WBBR05] Chen-Pin Wang, C. Hendricks Brown, and Karen Bandeen-Roche. Residual diagnostics for growth mixture models: Examining the impact of a preventive intervention on multiple trajectories of aggressive behavior. Journal of the American Statistical Association, 100(471):1054–1076, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590635.

Weinberg:2007:BFI

[WBS07a] Jonathan Weinberg, Lawrence D. Brown, and Jonathan R. Stroud. Bayesian forecasting of an inhomogeneous Poisson process with applications to call center data. Journal of the American Statistical Association, 102(480):1185–1198, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2007:CVS

[WBS07b] Yujun Wu, Dennis D. Boos, and Leonard A. Stefanski. Controlling variable selection by the addition of pseudovariables. Journal of the American Statistical Association, 102(477):235– 243, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2001:SNRa

[WC01] Mei-Cheng Wang and Ying-Qing Chen. Semiparametric and nonparametric regression analysis of longitudinal data: Comment. Journal of the American Statistical Association, 96 (453):113–114, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670345.

Wang:2004:UEE

[WC04] You-Gan Wang and Vincent J. Carey. Unbiased estimating equations from working correlation models for irregularly timed repeated measures. Journal of the American Statistical Association, 99(467):845–853, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s32/p845;   http://www.jstor.org/stable/27590454.

Wei:2009:QRM

[WC09] Ying Wei and Raymond J. Carroll. Quantile regression with measurement error. Journal of the American Statistical Association, 104(487):1129–1143, September 2009. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2005:ESM

[WCL05] Naisyin Wang, Raymond J. Carroll, and Xihong Lin. Efficient semiparametric marginal estimation for longitudinal/ clustered data. Journal of the American Statistical Association, 100(469):147–157, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s19/p147;   http://www.
jstor.org/stable/27590525.

Weinberg:2000:SIA

[WD00] Clarice R. Weinberg and David B. Dunson. Some issues in assessing human fertility. Journal of the American Statistical Association, 95(449):300–303, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2669552.

Walker:2004:PKL

[WDL04] Stephen Walker, Paul Damien, and Peter Lenk. On priors with a Kullback–Leibler property. Journal of the American Statistical Association, 99(466):404–408, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s10/p404;   http://www.jstor.org/stable/27590396.

Wegman:2000:BRB

[Weg00] Edward J. Wegman. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1009–1010, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669493.

Weir:2000:CFS

[Wei00] B. S. Weir. Challenges facing statistical genetics. Journal of the American Statistical Association, 95(449):319–322, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669557.

Wei:2008:AMC

[Wei08] Ying Wei. An approach to multivariate covariate-dependent quantile contours with application to bivariate conditional growth charts. Journal of the American Statistical Association, 103(481):397–409, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wells:2000:BRB

[Wel00] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1008, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669490.

Wells:2001:BRBf

[Wel01a] Martin Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(454):778, June 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670315.

Wells:2001:BRBg

[Wel01b] Martin Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(454):783, June 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670326.

Wells:2001:BRBh

[Wel01c] Martin Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(454):783, June 2001. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670327.

Wells:2001:BRBa

[Wel01d] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(453):342, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670373.

Wells:2001:BRBb

[Wel01e] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(453):354, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670390.

Wells:2001:BRBc

[Wel01f] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(453):354, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670391.

Wells:2001:BRBd

[Wel01g] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(453):354, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670392.

Wells:2001:BRBe

[Wel01h] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(453):354, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670393.

Wells:2001:BRBi

[Wel01i] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1137–1138, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670262.

Wells:2001:BRBj

[Wel01j] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1141, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670267.

Wells:2001:BRBk

[Wel01k] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1141, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670268.

Wells:2001:BRBl

[Wel01l] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1141, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670270.

Wells:2001:BRBm

[Wel01m] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1141, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670271.

Wells:2001:BRBn

[Wel01n] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(455):1141, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670272.

Wells:2001:BRBo

[Wel01o] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1526, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085923.

Wells:2001:BRBp

[Wel01p] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1530–1531, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085931.

Wells:2001:BRBq

[Wel01q] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1531–1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085933.

Wells:2001:BRBr

[Wel01r] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085934.

Wells:2001:BRBs

[Wel01s] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1532, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085936.

Wells:2001:BRBt

[Wel01t] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 96(456):1533, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085941.

Westfall:2005:CFD

[Wes05]  Peter H. Westfall.  Comment: False discovery rate-

adjusted multiple confidence intervals for selected parameters. Journal of the American Statistical Association, 100(469):85–89, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s12/p85.

West:2007:BRB

[Wes07] Webster West. Book review: Applied Statistics for Engineers and Scientists. Journal of the American Statistical Association, 102(478):767, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2003:MSE

[WF03] Junyuan Wang and Wayne A. Fuller. The mean squared error of small area predictors constructed with estimated area variances. Journal of the American Statistical Association, 98(463):716–723, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045299.

Wu:2003:C

[WH03a] Colin O. Wu and Jianhua Z. Huang. Comment. Journal of the American Statistical Association, 98(463):588–??, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n463/s11/p588.

Wu:2003:WBN

[WH03b] Colin O. Wu and Jianhua Z. Huang. Wavelet-based nonparametric modeling of hierarchical functions in colon carcinogenesis [with comment]. Journal of the American Statistical Association, 98(463):588–591, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30045286.

Wang:2007:DDE

[WH07] Huixia Wang and Xuming He. Detecting differential expressions in GeneChip microarray studies. Journal of the American Statistical Association, 102(477):104–112, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Whittaker:2001:SSA

[Whi01] Joe Whittaker. Smoothing spline ANOVA for multivariate Bernoulli observations, with application to ophthalmology data: Comment. Journal of the American Statistical Association, 96(453):157–158, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670354.

Witten:2009:D

[WHT09] Daniela M. Witten, Trevor Hastie, and Robert Tibshirani. Discussion. Journal of the American Statistical Association, 104 (486):698–699, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wiegand:2006:BRB

[Wie06] Ryan E. Wiegand. Book review: Introductory Econometrics: Using Monte Carlo Simulation With Microsoft Excel, by Humberto Barreto; Frank M. Howland. Journal of the American Statistical Association, 101(476):1730–1731, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27639802.

Wu:2004:MBB

[WIG^{+}04] Zhijin Wu, Rafael A. Irizarry, Robert Gentleman, Francisco Martinez-Murillo, and Forrest Spencer. A modelbased background adjustment for oligonucleotide expression arrays. Journal of the American Statistical Association, 99(468):909–917, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s1/p909;   http://www.
jstor.org/stable/27590474.

Wikle:2004:BRB

[Wik04] Christopher K. Wikle. Book review: Computational Statistics Handbook with MATLAB, by Wendy Martinez; Angel R. Martinez. Journal of the American Statistical Association, 99(466):566, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s36/p566;   http:
//www.jstor.org/stable/27590422.

Wikle:2006:BRB

[Wik06] Christopher K. Wikle. Book review: Environmental Statistics: Methods and Applications, by Vic Barnett. Journal of the American Statistical Association, 101(475):1311– 1312, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590809.

Williams:2003:BRB

[Wil03a] Christopher K. I. Williams. Book review: Learning with Kernels: Support Vector Machines, Regularization, Optimization, and Beyond, by Bernhard Schölkopf; Alexander Smola; Learning Kernel Classifiers, by Ralf Herbrich. Journal of the American Statistical Association, 98(462):489–490, June 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/30045256.

Williams:2003:LKC

[Wil03b] Christopher K. I. Williams. Learning kernel classifiers. Journal of the American Statistical Association, 98(462):489–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n462/s20/p489.

Williams:2003:LKS

[Wil03c] Christopher K. I. Williams. Learning with kernels: Support vector machines, regularization, optimization, and beyond. Journal of the American Statistical Association, 98(462):489–??, June 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s19/p489.

Williams:2004:BRBa

[Wil04a] Brian J. Williams. Book review: Visualizing Statistical Models and Concepts, by R. W. Farebrother. Journal of the American Statistical Association, 99(465):292, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s27/p292;   http://www.jstor.org/stable/27590373.

Williams:2004:BRBb

[Wil04b] James D. Williams. Book review: Generalized Estimating Equations, by James W. Hardin; Joseph M. Hilbe. Journal of the American Statistical Association, 99(465):297–298, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s34/p297;   http://www.jstor.org/stable/27590380.

Williamson:2006:BRBb

[Wil06a] John Williamson. Book review: Models for Discrete Longitudinal Data, by Geert Molenberghs; Geert Verbeke. Journal of the American Statistical Association, 101(475):1307, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590803.

Williamson:2006:BRBa

[Wil06b] Patricia Pepple Williamson. Book review: Mathematical Statistics with Applications, by Asha Seth Kapadia; Wenyaw Chan; Lemuel Moyé. Journal of the American Statistical Association, 101(473):394–395, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/30047474.

Winship:2003:BRB

[Win03] Christopher Winship. Book review: Logit and Probit: Ordered and Multinomial Models, by Vani K. Borooah. Journal of the American Statistical Association, 98(463):775–776, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30045320.

Wang:2008:MPD

[WJSS08] Shanshan Wang, Wolfgang Jank, Galit Shmueli, and Paul Smith. Modeling price dynamics in eBay auctions using differential equations. Journal of the American Statistical Association, 103(483):1100–1118, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2009:LRI

[WKL09] Lan Wang, Bo Kai, and Runze Li. Local rank inference for varying coefficient models. Journal of the American Statistical Association, 104(488):1631–1645, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weinberg:2000:ABL

[WL00] Janice M. Weinberg and Stephen W. Lagakos. Asymptotic behavior of linear permutation tests under general alternatives, with application to test selection and study design. Journal of the American Statistical Association, 95(450):596– 607, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/2000/June/Weinberg.htm;   http://
www.jstor.org/stable/2669402.

Wong:2001:MAC

[WL01] Chun Shan Wong and Wai Keung Li. On a mixture autoregressive conditional heteroscedastic model. Journal of the American Statistical Association, 96(455):982–995, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670244;   reqidx=/cw/asa/01621459/v96n455/
s16/p982.

Wang:2005:PNM

[WL05] Ji-Ping Z. Wang and Bruce G. Lindsay. A penalized nonparametric maximum likelihood approach to species richness estimation. Journal of the American Statistical Association, 100(471):942–959, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590625.

Wang:2007:ULE

[WL07a] Hansheng Wang and Chenlei Leng. Unified LASSO estimation by least squares approximation. Journal of the American Statistical Association, 102(479):1039–1048, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2007:RTH

[WL07b] Yichao Wu and Yufeng Liu. Robust truncated hinge loss support vector machines. Journal of the American Statistical Association, 102(479):974–983, September 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Welsh:2002:MLN

[WLC02] Alan H. Welsh, Xihong Lin, and Raymond J. Carroll. Marginal longitudinal nonparametric regression: Locality and efficiency of spline and kernel methods. Journal of the American Statistical Association, 97(458):482–493, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n458/
s10/p482;   http://www.jstor.org/stable/3085664.

Wang:2004:SRA

[WLH04] Qihua Wang, Oliver Linton, and Wolfgang Härdle. Semiparametric regression analysis with missing response at random. Journal of the American Statistical Association, 99(466):334–345, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n466/s4/p334;   http://www.
jstor.org/stable/27590390.

Wang:2008:VSN

[WLH08] Lifeng Wang, Hongzhe Li, and Jianhua Z. Huang. Variable selection in nonparametric varying-coefficient models for analysis of repeated measurements. Journal of the American Statistical Association, 103(484):1556–1569, December 2008. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wakefield:2001:BMD

[WM01] Jon C. Wakefield and Sara E. Morris. The Bayesian modeling of disease risk in relation to a point source. Journal of the American Statistical Association, 96(453):77– 91, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://rosina.

catchword.com/cgi-bin/linker?ini=asa&amp%3Breqidx=/
cw/asa/01621459/v96n453/s8/p77;   http://www.jstor.
org/stable/2670342.

Wang:2009:IDQ

[WM09a] Zuoheng Wang and Mary Sara McPeek. An incomplete-data quasi-likelihood approach to haplotype-based genetic association studies on related individuals. Journal of the American Statistical Association, 104(487):1251–1260, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wikle:2009:CST

[WM09b] C. K. Wikle and R. F. Milliff. Comment: a spatio-temporal model for mean, anomaly, and trend fields of North Atlantic sea surface temperature. Journal of the American Statistical Association, 104(485):22–23, March 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LS09b].

Wikle:2001:SHB

[WMNB01] Christopher K. Wikle, Ralph F. Milliff, Doug Nychka, and L. Mark Berliner. Spatiotemporal hierarchical Bayesian modeling tropical ocean surface winds. Journal of the American Statistical Association, 96(454):382–397, June 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670277;   reqidx=/cw/asa/01621459/v96n454/s3/p382.

Wolpert:2000:POM

[Wol00] Robert L. Wolpert. On the probability of observing misleading statistical evidence: Comment. Journal of the American Statistical Association, 95(451):771–772, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669459.

Wollan:2002:BRB

[Wol02a] Peter C. Wollan. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1207, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085851.

Wollan:2002:OSS

[Wol02b] Peter C. Wollan. Observational studies (second edition). Journal of the American Statistical Association, 97(460): 1207–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s32/p1207.

Wolpert:2007:CSL

[Wol07] Robert L. Wolpert. Comment: Subjective likelihood for the assessment of trends in the ocean’s mixed-layer depth. Journal of the American Statistical Association, 102(479):782–784, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolf:2008:BRB

[Wol08] Michael Wolf. Book review: Nonparametric Econometrics: Theory and Practice. Journal of the American Statistical Association, 103(482):885–886, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:2000:POM

[Won00a] A. C. M. Wong. On the probability of observing misleading statistical evidence: Comment. Journal of the American Statistical Association, 95(451):773, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669460.

Wong:2000:CMB

[Won00b] Wing Hung Wong. Computational molecular biology. Journal of the American Statistical Association, 95(449):322–326, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
2669558.

Wood:2001:SSA

[Woo01] Sally Wood. Smoothing spline ANOVA for multivariate Bernoulli observations, with application to ophthalmology data: Comment. Journal of the American Statistical Association, 96(453):151–155, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670352.

Wood:2004:SEM

[Woo04a] Simon N. Wood. Stable and efficient multiple smoothing parameter estimation for generalized additive models. Journal of the American Statistical Association, 99(467):673–686, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s17/p673;   http:
//www.jstor.org/stable/27590439.

Woodworth:2004:BRB

[Woo04b] George Woodworth. Book review: Bayesian Reasoning in Data Analysis: A Critical Introduction, by Giulio D’Agostini. Journal of the American Statistical Association, 99(468): 1201–1202, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s30/p1201;   http://www.
jstor.org/stable/27590497.

Woodworth:2007:BRB

[Woo07] George Woodworth. Book review: Modeling in Medical Decision Making: A Bayesian Approach. Journal of the American Statistical Association, 102(477):387–388, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Worsley:2000:BTC

[Wor00] K. J. Worsley. A Bayesian time-course model for functional magnetic resonance imaging data: Comment. Journal of the American Statistical Association, 95(451):711–716, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669450.

Wang:2007:RTR

[WQ07] Lan Wang and Annie Qu. Robust tests in regression models with omnibus alternatives and bounded influence. Journal of the American Statistical Association, 102(477):347–358, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2001:ARE

[WQC01] Mei-Cheng Wang, Jing Qin, and Chin-Tsang Chiang. Analyzing recurrent event data with informative censoring. Journal of the American Statistical Association, 96(455): 1057–1065, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;
http://www.jstor.org/stable/2670251;   reqidx=/cw/
asa/01621459/v96n455/s23/p1057.

Wang:2002:NNVa

[WR02a] Naisyin Wang and Adrian E. Raftery. Nearest-neighbor variance estimation (NNVE): Robust covariance estimation via nearest-neighbor cleaning. Journal of the American Statistical Association, 97(460):994–1006, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s7/p994;   http://www.jstor.org/stable/3085821.

Wang:2002:NNVb

[WR02b] Naisyin Wang and Adrian E. Raftery. [nearest-neighbor variance estimation (NNVE): Robust covariance estimation via nearest-neighbor cleaning]: Rejoinder. Journal of the American Statistical Association, 97(460):1016–1019, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
3085826.

Wang:2007:BCC

[WRM07] Xiaohui Wang, Shubhankar Ray, and Bani K. Mallick. Bayesian curve classification using wavelets. Journal of the American Statistical Association, 102(479):962–973, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westfall:2001:UPI

[WS01a] Peter H. Westfall and Keith A. Soper. Using priors to improve multiple animal carcinogenicity tests. Journal of the American Statistical Association, 96(455):827–834, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670223;   reqidx=/cw/asa/01621459/v96n455/s5/p827.

Wu:2001:MCA

[WS01b] Changbao Wu and Randy R. Sitter. A model-calibration approach to using complete auxiliary information from survey data. Journal of the American Statistical Association, 96(453):185–193, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/rosina.catchword.com/cgi-bin/linker?ini=asa&amp;

http://www.jstor.org/stable/2670358;   reqidx=/cw/
asa/01621459/v96n453/s14/p185.

Winnett:2002:ANA

[WS02] Angela Winnett and Peter Sasieni. Adjusted Nelson-Aalen estimates with retrospective matching. Journal of the American Statistical Association, 97(457):245–256, March 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085779;   reqidx=/cw/asa/01621459/v97n457/
s22/p245.

Wolbers:2005:LUM

[WS05] Marcel Wolbers and Werner Stahel. Linear unmixing of multivariate observations: a structural model. Journal of the American Statistical Association, 100(472):1328–1342, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590675.

Woo:2006:REM

[WS06] Mi-Ja Woo and T. N. Sriram. Robust estimation of mixture complexity. Journal of the American Statistical Association, 101(476):1475–1486, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639766.

Wang:2007:NMS

[WS07] Lifeng Wang and Xiaotong Shen. On L_{1}-norm multiclass support vector machines. Journal of the American Statistical Association, 102(478):583–594, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2006:CMO

[WSLC06] Xinlei Wang, Lynne Stokes, Johan Lim, and Min Chen. Concomitants of multivariate order statistics with application to judgment poststratification. Journal of the American Statistical Association, 101(476):1693–1704, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27639783.

Wang:2009:LMH

[WSP09] Junhui Wang, Xiaotong Shen, and Wei Pan. On large margin hierarchical classification with multiple paths. Journal of the American Statistical Association, 104(487):1213–1223, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2001:JML

[WT01] Yan Wang and Jeremy M. G. Taylor. Jointly modeling longitudinal and event time data with application to acquired immunodeficiency syndrome. Journal of the American Statistical Association, 96(455):895–905, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670229;   reqidx=/cw/asa/01621459/v96n455/
s11/p895.

Westfall:2007:MTG

[WT07] Peter H. Westfall and Randall D. Tobias. Multiple testing of general contrasts. Journal of the American Statistical Association, 102(478):487–494, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2009:TIR

[WT09] Hansheng Wang and Chih-Ling Tsai. Tail index regression. Journal of the American Statistical Association, 104(487): 1233–1240, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2002:JMN

[Wu02] Lang Wu. A joint model for nonlinear mixed-effects models with censoring and covariates measured with error, with application to AIDS studies. Journal of the American Statistical Association, 97(460):955–964, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s3/p955;   http://www.jstor.org/stable/3085817.

Wu:2004:EAI

[Wu04] Lang Wu. Exact and approximate inferences for nonlinear mixed-effects models with missing covariates. Journal of the American Statistical Association, 99(467):700–709, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s19/p700;   http://www.jstor.org/stable/27590441.

Wang:2000:MSSa

[WW00a] Weijing Wang and Martin T. Wells. Model selection and semiparametric inference for bivariate failure-time data. Journal of the American Statistical Association, 95(449):62–72, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/Wang.htm;   http://www.jstor.org/
stable/2669523.

Wang:2000:MSSb

[WW00b] Weijing Wang and Martin T. Wells. Model selection and semiparametric inference for bivariate failure-time data: Rejoinder. Journal of the American Statistical Association, 95 (449):75–76, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669525.

Wang:2000:RMS

[WW00c] Weijing Wang and Martin T. Wells. Rejoinder: Model selection and semiparametric inference for bivariate failure-time data. Journal of the American Statistical Association, 95 (449):75–??, March 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2001:SNRb

[WW01] Jane-Ling Wang and Wei Wang. Semiparametric and nonparametric regression analysis of longitudinal data: Comment. Journal of the American Statistical Association, 96 (453):119–123, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670348.

Wang:2009:LWC

[WW09] Huixia Judy Wang and Lan Wang. Locally weighted censored quantile regression. Journal of the American Statistical Association, 104(487):1117–1128, September 2009. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2008:MIE

[WWP^{+}08] Xiao Wang, Matthew Walker, Jayanta Pal, Michael Woodroofe, and Mario Mateo. Model-independent estimates of dark matter distributions. Journal of the American Statistical Association, 103(483):1070–1084, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2008:SRD

[WX08] Hansheng Wang and Yingcun Xia. Sliced regression for dimension reduction. Journal of the American Statistical Association, 103(482):811–821, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2009:SEV

[WX09] Hansheng Wang and Yingcun Xia. Shrinkage estimation of the varying coefficient model. Journal of the American Statistical Association, 104(486):747–757, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2002:LPM

[WZ02] Hulin Wu and Jin-Ting Zhang. Local polynomial mixed-effects models for longitudinal data. Journal of the American Statistical Association, 97(459):883–897, September 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://giorgio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n459/
s19/p883;   http://www.jstor.org/stable/3085729.

Weaver:2005:ELM

[WZ05] Mark A. Weaver and Haibo Zhou. An estimated likelihood method for continuous outcome regression models with outcome-dependent sampling. Journal of the American Statistical Association, 100(470):459–469, June 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v100n470/s9/p459;   http:
//www.jstor.org/stable/27590568.

Xia:2008:MIM

[Xia08] Yingcun Xia. A multiple-index model and dimension reduction. Journal of the American Statistical Association, 103 (484):1631–1640, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2000:DPP

[Xie00] Yu Xie. Demography: Past, present, and future. Journal of the American Statistical Association, 95(450):670–673, June 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669415.

Xie:2003:BRB

[Xie03a] Minge Xie. Book review: Design and Analysis of Bioavailability and Bioequivalence Studies, by Shein-Chung Chow; Jen-Pei Liu. Journal of the American Statistical Association, 98 (461):251, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30045216.

Xie:2003:DAB

[Xie03b] Minge Xie. Design and analysis of bioavailability and bioequivalence studies (second edition, revised and expanded). Journal of the American Statistical Association, 98(461):251–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://fidelio.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/01621459/
v98n461/s29/p251.

Xiao:2009:CQE

[XK09] Zhijie Xiao and Roger Koenker. Conditional quantile estimation for generalized autoregressive conditional heteroscedasticity models. Journal of the American Statistical Association, 104(488):1696–1712, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2004:BID

[XLB04] Jun Xie, Ker-Chau Li, and Minou Bina. A Bayesian insertion/deletion algorithm for distant protein motif searching via entropy filtering. Journal of the American Statistical Association, 99(466):409–420, April 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n466/s11/p409;   http:
//www.jstor.org/stable/27590397.

Xiao:2003:MEL

[XLCM03] Zhijie Xiao, Oliver B. Linton, Raymond J. Carroll, and Enno Mammen. More efficient local polynomial estimation in nonparametric regression with autocorrelated errors. Journal of the American Statistical Association, 98 (464):980–992, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s23/p980;   http://www.
jstor.org/stable/30045344.

Xue:2004:SML

[XLL04] Hongqi Xue, K. F. Lam, and Guoying Li. Sieve maximum likelihood estimator for semiparametric regression models with current status data. Journal of the American Statistical Association, 99(466):346–356, April 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/s5/
p346;   http://www.jstor.org/stable/27590391.

Bowei:2006:ENL

[XMN06] Bowei Xi, George Michailidis, and Vijayan N. Nair. Estimating network loss rates using active tomography. Journal of the American Statistical Association, 101(476):1430– 1448, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639763.

Xu:2009:REC

[XPLT09] Qiang Xu, Myunghee Cho Paik, Xiaodong Luo, and Wei-Yann Tsai. Reweighting estimators for Cox regression with missing covariates. Journal of the American Statistical Association, 104(487):1155–1167, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2008:SAH

[XSC08] Minge Xie, Douglas G. Simpson, and Raymond J. Carroll. Semiparametric analysis of heterogeneous data using varyingscale generalized linear models. Journal of the American Statistical Association, 103(482):650–660, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2009:CIP

[XSZ09] Minge Xie, Kesar Singh, and Cun-Hui Zhang. Confidence intervals for population ranks in the presence of ties and near ties. Journal of the American Statistical Association, 104 (486):775–788, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2001:GTB

[XTSY01] Minge Xie, Kay Tatsuoka, Jerome Sacks, and S. Stanley Young. Group testing with blockers and synergism. Journal of the American Statistical Association, 96(453):92–102, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670343;   reqidx=/cw/asa/01621459/v96n453/s9/p92.

Xu:2005:KBS

[XWF05] Ke Xu, Christopher K. Wikle, and Neil I. Fox. A kernel-based spatio-temporal dynamical model for nowcasting weather radar reflectivities. Journal of the American Statistical Association, 100(472):1133–1144, December 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590660.

Xue:2007:ELV

[XZ07] Liugen Xue and Lixing Zhu. Empirical likelihood for a varying coefficient model with longitudinal data. Journal of the American Statistical Association, 102(478):642–654, June 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2001:ARM

[Yan01] Yuhong Yang. Adaptive regression by mixing. Journal of the American Statistical Association, 96(454):574–588, June 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670298;   reqidx=/cw/asa/01621459/v96n454/
s19/p574.

Yan:2004:BRBb

[Yan04a] Jun Yan. Book review: Bayesian Survival Analysis, by Joseph G. Ibrahim; Ming-Hui Chen; Debajyoti Sinha.

Journal of the American Statistical Association, 99(468): 1202–1203, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v99n468/s32/p1202;   http://www.
jstor.org/stable/27590499.

Yan:2004:BRBa

[Yan04b] Jun Yan. Book review: Survival Analysis: Techniques for Censored and Truncated Data, second edition, by John P. Klein; Melvin L. Moeschberger. Journal of the American Statistical Association, 99(467):900, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s45/p900;   http://www.jstor.org/stable/27590461.

Yang:2004:BRB

[Yan04c] Yuhong Yang. Book review: Nonlinear Estimation and Classification, by David D. Denison; Mark H. Hansen; Christopher C. Holmes; Bani Mallick; Bin Yu. Journal of the American Statistical Association, 99(466):561, April 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n466/
s28/p561;   http://www.jstor.org/stable/27590414.

Yan:2005:BRB

[Yan05a] Jun Yan. Book review: Analysis of Multivariate Survival Data, by Phillip Hougaard. Journal of the American Statistical Association, 100(469):354–355, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s46/p355;   http://www.
jstor.org/stable/27590552.

Yang:2005:BRB

[Yan05b] Yuhong Yang. Book review: Information Theory, Inference, and Learning Algorithms, by David J. C. MacKay. Journal of the American Statistical Association, 100(472):1461– 1462, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590692.

Yan:2006:BRB

[Yan06] Jun Yan. Book review: Gaussian Markov Random Fields: Theory and Applications, by Harvard Rue; Leonhard Held. Journal of the American Statistical Association, 101(473): 388–389, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047466.

Yan:2008:BRB

[Yan08a] Xin Yan. Book review: Linear Models: An Integrated Approach. Journal of the American Statistical Association, 103 (482):884–885, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2008:BRB

[Yan08b] Yuhong Yang. Book review: Elements of Information Theory. Journal of the American Statistical Association, 103(481):429, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yatracos:2005:AAS

[Yat05] Yannis G. Yatracos. Artificially augmented samples, shrinkage, and mean squared error reduction. Journal of the American Statistical Association, 100(472):1168–1175, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590662.

Yi:2002:MMI

[YC02] Grace Y. Yi and Richard J. Cook. Marginal methods for incomplete longitudinal data arising in clusters. Journal of the American Statistical Association, 97(460): 1071–1080, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s12/p1071;   http:/
/www.jstor.org/stable/3085831.

Yekutieli:2008:HFD

[Yek08] Daniel Yekutieli. Hierarchical false discovery rate–controlling methodology. Journal of the American Statistical Association, 103(481):309–316, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yan:2005:FAM

[YF05] Jun Yan and Jason P. Fine. Functional association models for multivariate survival processes. Journal of the American Statistical Association, 100(469):184–196, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s22/
p184;   http://www.jstor.org/stable/27590528.

Yan:2008:AED

[YF08] Jun Yan and Jason P. Fine. Analysis of episodic data with application to recurrent pulmonary exacerbations in cystic fibrosis patients. Journal of the American Statistical Association, 103(482):498–510, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yin:2004:BRB

[Yin04] George Yin. Book review: Stochastic Approximation and Its Applications, by Han-Fu Chen. Journal of the American Statistical Association, 99(465):295–296, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s31/p295;   http://www.jstor.org/stable/27590377.

Yu:2004:LBL

[YJ04] K. Yu and M. C. Jones. Likelihood-based local linear estimation of the conditional variance function. Journal of the American Statistical Association, 99(465):139–144, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s14/p139;   http://www.jstor.org/stable/27590360.

Yang:2004:SAM

[YK04] Bo Yang and John E. Kolassa. Smooth and accurate multivariate confidence regions. Journal of the American Statistical Association, 99(468):1072–1081, December 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n468/
s19/p1072;   http://www.jstor.org/stable/27590486.

Yuan:2006:RD

[YK06] Ming Yuan and Christina Kendziorski. Rejoinder: Discussion. Journal of the American Statistical Association, 101 (476):1338–1340, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yuan:2006:HMM

[YKL^{+}06] Ming Yuan, Christina Kendziorski, Hongzhe Li, Fangxin Hong, and Dan Nettleton. Hidden Markov models for microarray time course data in multiple biological conditions [with discussions, rejoinder]. Journal of the American Statistical Association, 101(476):1323–1332, December 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27639754.

Yau:2001:ICA

[YL01] Linda H. Y. Yau and Roderick J. Little. Inference for the complier-average causal effect from longitudinal data subject to noncompliance and missing data, with application to a job training assessment for the unemployed. Journal of the American Statistical Association, 96(456):1232–1244, December 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/3085886;   reqidx=/cw/asa/01621459/v96n456/
s10/p1232.

Yuan:2005:EEB

[YL05] Ming Yuan and Yi Lin. Efficient empirical Bayes variable selection and estimation in linear models. Journal of the American Statistical Association, 100(472):1215–1225, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590666.

Yao:2009:BML

[YL09] Weixin Yao and Bruce G. Lindsay. Bayesian mixture labeling by highest posterior density. Journal of the American Statistical Association, 104(486):758–767, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ylvisaker:2001:BRB

[Ylv01] Don Ylvisaker. Book review: [untitled]. Journal of the American Statistical Association, 96(453):340–341, March 2001. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670371.

Yin:2008:PLA

[YLZ08] Guosheng Yin, Hui Li, and Donglin Zeng. Partially linear additive hazards regression with varying coefficients. Journal of the American Statistical Association, 103(483):1200–1213, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yao:2005:FDA

[YMW05] Fang Yao, Hans-Georg Müller, and Jane-Ling Wang. Functional data analysis for sparse longitudinal data. Journal of the American Statistical Association, 100(470):577–590, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n470/s20/
p577;   http://www.jstor.org/stable/27590579.

Yor:2006:BRB

[Yor06] Marc Yor. Book review: Stochastic Inequalities and Applications, by Evariste Giné; Christian Houdré; David Nualart. Journal of the American Statistical Association, 101 (474):845–846, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590743.

Yang:2006:ETV

[YPXH06] Lijian Yang, Byeong U. Park, Lan Xue, and Wolfgang Härdle. Estimation and testing for varying coefficients in additive models with marginal integration. Journal of the American Statistical Association, 101(475):1212–1227, September 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590796.

Yung:2000:JVE

[YR00] W. Yung and J. N. K. Rao. Jackknife variance estimation under imputation for estimators using poststratification information. Journal of the American Statistical Association, 95(451):903–915, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669473.

Yu:2002:PSE

[YR02] Yan Yu and David Ruppert. Penalized spline estimation for partially linear single-index models. Journal of the American Statistical Association, 97(460):1042–1054, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/01621459/v97n460/
s10/p1042;   http://www.jstor.org/stable/3085829.

Yakovlev:2008:BPM

[YSY08] Andrei Y. Yakovlev, Vessela K. Stoimenova, and Nikolay M. Yanev. Branching processes as models of progenitor cell populations and estimation of the offspring distributions. Journal of the American Statistical Association, 103(484):1357–1366, December 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2008:IPP

[YTS08] Menggang Yu, Jeremy M. G. Taylor, and Howard M. Sandler. Individual prediction in prostate cancer studies using a joint longitudinal survival–cure model. Journal of the American Statistical Association, 103(481):178–187, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yee:2001:SSA

[YW01] T. W. Yee and C. J. Wild. Smoothing spline ANOVA for multivariate Bernoulli observations, with application to ophthalmology data: Comment. Journal of the American Statistical Association, 96(453):147–151, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2670351.

Ye:2003:UBS

[YW03] Zhishen Ye and Robert E. Weiss. Using the bootstrap to select one of a new class of dimension reduction methods. Journal of the American Statistical Association, 98(464):968–979, December 2003. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v98n464/s22/p968;   http:
//www.jstor.org/stable/30045343.

Yuan:2005:CLR

[YY05] Zheng Yuan and Yuhong Yang. Combining linear regression models: When and how? Journal of the American Statistical Association, 100(472):1202–1214, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/27590665.

Yin:2009:BMA

[YY09] Guosheng Yin and Ying Yuan. Bayesian model averaging continual reassessment method in Phase I clinical trials. Journal of the American Statistical Association, 104(487):954–968, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yucel:2005:IBT

[YZ05] Recai M. Yucel and Alan M. Zaslavsky. Imputation of binary treatment variables with measurement error in administrative data. Journal of the American Statistical Association, 100 (472):1123–1132, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590659.

Yuan:2009:EGA

[YZ09] Ming Yuan and Hui Zou. Efficient global approximation of generalized nonlinear ℓ_{1}-regularized solution paths and its applications. Journal of the American Statistical Association, 104(488):1562–1574, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yin:2008:PTL

[YZL08] Guosheng Yin, Donglin Zeng, and Hui Li. Power-transformed linear quantile regression with censored data. Journal of the American Statistical Association, 103(483):1214–1224, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zabell:2008:SAP

[Zab08a] S. L. Zabell. On Student’s 1908 article “The Probable Error of a Mean”. Journal of the American Statistical Association, 103 (481):1–7, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zabell:2008:RSA

[Zab08b] S. L. Zabell. Rejoinder: On Student’s 1908 article “The Probable Error of a Mean”. Journal of the American Statistical Association, 103(481):19–20, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zadeh:2004:CMF

[Zad04] Lotfi A. Zadeh. Comment: Membership functions and probability measures of fuzzy sets. Journal of the American Statistical Association, 99(467):880–881, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s37/p880.

Zalkikar:2002:BRB

[Zal02] Jyoti Zalkikar. Book review: [untitled]. Journal of the American Statistical Association, 97(458):650–651, June 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/3085679.

Zeng:2007:TMM

[ZC07] Leilei Zeng and Richard J. Cook. Transition models for multivariate longitudinal binary data. Journal of the American Statistical Association, 102(477):211–223, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2009:IRM

[ZCI^{+}09] Hongtu Zhu, Yasheng Chen, Joseph G. Ibrahim, Yimei Li, Colin Hall, and Weili Lin. Intrinsic regression models for positive-definite matrices with applications to diffusion tensor imaging. Journal of the American Statistical Association, 104 (487):1203–1212, September 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zheng:2008:TDP

[ZCPL08] Yingye Zheng, Tianxi Cai, Margaret S. Pepe, and Wayne C. Levy. Time-dependent predictive values of prognostic biomarkers with failure time outcome. Journal of the American Statistical Association, 103(481):362–368, March 2008.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeh:2001:EFG

[Zeh01] Judith E. Zeh. Estimation of fetal growth and gestation in bowhead whales: Comment. Journal of the American Statistical Association, 96(455):933–935, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670235.

Zhang:2003:LVM

[ZFZ03] Heping Zhang, Rui Feng, and Hongtu Zhu. A latent variable model of segregation analysis for ordinal traits. Journal of the American Statistical Association, 98(464): 1023–1034, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/s27/p1023;   http://www.
jstor.org/stable/30045348.

Zhang:2003:CDF

[Zha03] Chunming Zhang. Calibrating the degrees of freedom for automatic data smoothing and effective curve checking. Journal of the American Statistical Association, 98 (463):609–628, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n463/s14/p609;   http://
www.jstor.org/stable/30045289.

Zhang:2004:CEP

[Zha04a] Chunming Zhang. Comment: The estimation of prediction error: Covariance penalties and cross-validation. Journal of the American Statistical Association, 99(467):637–640, September 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n467/
s13/p637.

Zhang:2004:IEA

[Zha04b] Hao Zhang. Inconsistent estimation and asymptotically equal interpolations in model-based geostatistics. Journal of the American Statistical Association, 99(465):250–261, March 2004. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/
vl=1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v99n465/
s24/p250;   http://www.jstor.org/stable/27590370.

Zhang:2005:CSH

[Zha05a] Heping Zhang. Comment: Semilinear high-dimensional model for normalization of microarray data. Journal of the American Statistical Association, 100(471):804, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2005:AAD

[Zha05b] Jin-Ting Zhang. Approximate and asymptotic distributions of chi-squared-type mixtures with applications. Journal of the American Statistical Association, 100(469):273–285, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s31/
p273;   http://www.jstor.org/stable/27590537.  See

correction [Dui16].

Zhang:2006:BRB

[Zha06] Zhigang Zhang. Book review: Modeling Longitudinal Data, by Robert E. Weiss. Journal of the American Statistical Association, 101(476):1726–1727, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27639795.

Zheng:2006:BRB

[Zhe06] W. Jim Zheng. Book review: Statistical Methods in Bioinformatics: An Introduction, by Warren J. Ewens; Gregory Grant. Journal of the American Statistical Association, 101 (473):393, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/30047472.

Zhu:2006:BRB

[Zhu06] Jun Zhu. Book review: Statistical Methods for Spatial Data Analysis, by Oliver Schabenberger; Carol A. Gotway. Journal of the American Statistical Association, 101(473):389–390, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047467.

Zimmerman:2003:BRB

[Zim03] Dale Zimmerman. Book review: Geographically Weighted Regression, by A. Stewart Fotheringham; Chris Brunsdon; Martin Charlton. Journal of the American Statistical Association, 98(463):765–766, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045305.

Zhou:2005:CMM

[ZIP05] Xi Zhou, Edwin S. Iversen, Jr., and Giovanni Parmigiani. Classification of missense mutations of disease genes. Journal of the American Statistical Association, 100(469):51–60, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n469/s6/p51;   http://www.
jstor.org/stable/27590518.

Zivot:2006:BRB

[Ziv06] Eric Zivot. Book review: Statistics and Finance: An Introduction, by David Ruppert. Journal of the American Statistical Association, 101(474):849–850, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/27590749.

Zeng:2007:EEA

[ZL07a] Donglin Zeng and D. Y. Lin. Efficient estimation for the accelerated failure time model. Journal of the American Statistical Association, 102(480):1387–1396, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2007:STM

[ZL07b] Donglin Zeng and D. Y. Lin. Semiparametric transformation models with random effects for recurrent events. Journal of the American Statistical Association, 102(477):167–180, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2009:RMI

[ZLI^{+}09] Hongtu Zhu, Yimei Li, Joseph G. Ibrahim, Xiaoyan Shi, Hongyu An, Yashen Chen, Wei Gao, Weili Lin, Daniel B. Rowe, and Bradley S. Peterson. Regression models for identifying noise sources in magnetic resonance images. Journal of the American Statistical Association, 104(486):623–637, June 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2005:MLE

[ZLY05] Donglin Zeng, D. Y. Lin, and Guosheng Yin. Maximum likelihood estimation for the proportional odds model with random effects. Journal of the American Statistical Association, 100(470):470–483, June 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v100n470/s10/p470;   http://www.
jstor.org/stable/27590569.

Zhang:2005:TTT

[ZMAS05] Lan Zhang, Per A. Mykland, and Yacine Aït-Sahalia. A tale of two time scales: Determining integrated volatility with noisy high-frequency data. Journal of the American Statistical Association, 100(472):1394–1411, December 2005. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/27590680.

Zhu:2006:SIR

[ZMP06] Lixing Zhu, Baiqi Miao, and Heng Peng. On sliced inverse regression with high-dimensional covariates. Journal of the American Statistical Association, 101(474):630–643, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590723.

Zou:2006:ALO

[Zou06] Hui Zou. The adaptive lasso and its oracle properties. Journal of the American Statistical Association, 101(476):1418– 1429, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27639762.

Zhuang:2002:SDS

[ZOVJ02] Jiancang Zhuang, Yosihiko Ogata, and David Vere-Jones. Stochastic declustering of space–time earthquake occurrences. Journal of the American Statistical Association, 97(458):369–380, June 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//giorgio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n458/s1/p369;   http://
www.jstor.org/stable/3085650.

Zou:2009:MSP

[ZQ09] Changliang Zou and Peihua Qiu. Multivariate statistical process control using LASSO. Journal of the American Statistical Association, 104(488):1586–1596, December 2009. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2008:STM

[ZRM^{+}08] Jun Zhu, Jakob G. Rasmussen, Jesper Møller, Brian H. Aukema, and Kenneth F. Raffa. Spatial-temporal modeling of forest gaps generated by colonization from below- and aboveground bark beetle species. Journal of the American Statistical Association, 103(481):162–177, March 2008. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2009:LBA

[ZRM09] Junni L. Zhang, Donald B. Rubin, and Fabrizia Mealli. Likelihood-based analysis of causal effects of job-training programs using principal stratification. Journal of the American Statistical Association, 104(485):166–176, March 2009. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2001:SAR

[ZS01] Shanggang Zhou and Xiaotong Shen. Spatially adaptive regression splines and accurate knot selection schemes. Journal of the American Statistical Association, 96(453):247–259, March 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://rosina.catchword.com/
cgi-bin/linker?ini=asa&amp;   http://www.jstor.org/
stable/2670363;   reqidx=/cw/asa/01621459/v96n453/
s19/p247.

Zhao:2002:BRB

[ZS02a] Yonggang Zhao and Xiaotong Shen. Book review: [untitled]. Journal of the American Statistical Association, 97(460):1203– 1204, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/3085845.

Zhao:2002:EL

[ZS02b] Yonggang Zhao and Xiaotong Shen. Empirical likelihood. Journal of the American Statistical Association, 97 (460):1203–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v97n460/s26/p1203.

Zaslavsky:2001:DIC

[ZSB01] Alan M. Zaslavsky, Nathaniel Schenker, and Thomas R. Belin. Downweighting influential clusters in surveys: Application to the 1990 post enumeration survey. Journal of the American Statistical Association, 96(455):858–869, September 2001. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://rosina.catchword.com/cgi-bin/
linker?ini=asa&amp;   http://www.jstor.org/stable/
2670226;   reqidx=/cw/asa/01621459/v96n455/s8/p858.

Zhang:2006:BIT

[ZSCW06] Min Zhang, Robert L. Strawderman, Mark E. Cowen, and Martin T. Wells. Bayesian inference for a two-part hierarchical model: An application to profiling providers in managed health care. Journal of the American Statistical Association, 101(475):934–945, September 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590773.

Zheng:2006:HMP

[ZSG06] Tian Zheng, Matthew J. Salganik, and Andrew Gelman. How many people do you know in prison?: Using overdispersion in count data to estimate social structure in networks. Journal of the American Statistical Association, 101(474):409–423, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
27590705.

Zucker:2005:PPL

[Zuc05] David M. Zucker. A pseudo–partial likelihood method for semiparametric survival regression with covariate errors. Journal of the American Statistical Association, 100(472):1264– 1277, December 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590670.

Zhao:2008:WBN

[ZW08] Wei Zhao and Rongling Wu. Wavelet-based nonparametric functional mapping of longitudinal curves. Journal of the American Statistical Association, 103(482):714–725, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zwick:2000:BRB

[Zwi00] Rebecca Zwick. Book review: [untitled]. Journal of the American Statistical Association, 95(451):1012–1013, September 2000. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669496.

Zhang:2004:VSM

[ZWL^{+}04] Hao Helen Zhang, Grace Wahba, Yi Lin, Meta Voelker, Michael Ferris, Ronald Klein, and Barbara Klein. Variable selection and model building via likelihood basis pursuit. Journal of the American Statistical Association, 99(467):659–672, September 2004. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://miranda.asa.catchword.org/vl=1515603/cl=66/
nw=1/rpsv/cw/asa/01621459/v99n467/s16/p659;   http:
//www.jstor.org/stable/27590438.

Zhang:2006:CCD

[ZWS06] Peng Zhang, Xiaogang Wang, and Peter X.-K. Song. Clustering categorical data based on distance vectors. Journal of the American Statistical Association, 101(473):355–367, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/
30047463.

Zhang:2005:HGM

[ZWT^{+}05] Dabao Zhang, Martin T. Wells, Bruce W. Turnbull, David Sparrow, and Patricia A. Cassano. Hierarchical graphical models: An application to pulmonary function and cholesterol levels in the normative aging study. Journal of the American Statistical Association, 100(471):719–727, September 2005. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/27590609.

Zhou:2008:EEI

[ZWW08] Yong Zhou, Alan T. K. Wan, and Xiaojing Wang. Estimating equations inference with missing data. Journal of the American Statistical Association, 103(483):1187–1199, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2008:JMT

[ZYDS08] Heping Zhang, Yuanqing Ye, Peter Diggle, and Jian Shi. Joint modeling of time series measures and recurrent events and analysis of the effects of air quality on respiratory symptoms. Journal of the American Statistical Association, 103(481):48– 60, March 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2005:ICT

[ZYI05] Donglin Zeng, Guosheng Yin, and Joseph G. Ibrahim. Inference for a class of transformed hazards models. Journal of the American Statistical Association, 100(471):1000– 1008, September 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/27590630.

Zeng:2006:STM

[ZYI06] Donglin Zeng, Guosheng Yin, and Joseph G. Ibrahim. Semiparametric transformation models for survival data with a cure fraction. Journal of the American Statistical Association, 101(474):670–684, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27590726.

Zhang:2003:ILD

[ZYZS03] Heping Zhang, Chang-Yung Yu, Hongtu Zhu, and Jian Shi. Identification of linear directions in multivariate adaptive spline models. Journal of the American Statistical Association, 98(462):369–376, June 2003. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//fidelio.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/01621459/v98n462/s9/p369;   http://
www.jstor.org/stable/30045246.

Zhu:2006:FME

[ZZ06] Yu Zhu and Peng Zeng. Fourier methods for estimating the central subspace and the central mean subspace in regression. Journal of the American Statistical Association, 101 (476):1638–1651, December 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/27639778.

Zhu:2007:RSA

[ZZIP07a] Hongtu Zhu, Heping Zhang, Joseph G. Ibrahim, and Bradley S. Peterson. Rejoinder: Statistical analysis of diffusion tensors in diffusion-weighted magnetic resonance imaging data. Journal of the American Statistical Association, 102 (480):1110, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2007:SAD

[ZZIP07b] Hongtu Zhu, Heping Zhang, Joseph G. Ibrahim, and Bradley S. Peterson. Statistical analysis of diffusion tensors in diffusion-weighted magnetic resonance imaging data. Journal of the American Statistical Association, 102(480):1085–1102, December 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).